A Complete Bibliography of Publications in Technology and Culture: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.01

Title word cross-reference

$5 [Hou03].

0The [McG00c].

1000 [Poo03]. 1477 [Bac06a]. 1500 [Phi03]. 1520 [Jaf03]. 1540 [Ung05]. 1550 [Ben01c]. 1613 [San03]. 1683 [Gor06]. 1700 [Bro06b, Smi05a]. 1750 [Gle03]. 1760 [Ben01d]. 1763 [Dic05]. 1770 [War03]. 1787 [Low04b]. 1791 [Ben02]. 1800 [Hac05b, Ben01a, Ten05a, Tim09]. 1803 [Bow00a]. 1805 [Wak02]. 1825 [Fin00]. 1830 [Haf09, Pan07, Sew06, Sha09b]. 1832 [Kel05, Pet04c]. 1840 [Gle08]. 1840s [Con00a]. 1844 [Mac02]. 1845 [Jaf05, Pfa04]. 1849 [Isr02]. 1850 [Big02, CK01, Gei00, Lit03, Pri04, Rab04, Roc02]. 1851 [Ink06]. 1853 [Isr02]. 1860 [Gou00, Gud07, Moh08, See07, Sha09a]. 1864 [Buc03b, Twe05]. 1865 [Col07c, Smi06]. 1868 [Bek04b]. 1869 [Hof02]. 1870 [Clo05, War09b]. 1870s [Och02]. 1870s-1990s [Och02]. 1875 [Fin03]. 1877 [Mol00b]. 1878 [Rot01]. 1880 [HP00, Kie09, Mar09]. 1880s [Nis07]. 1880s-1950s [Nis07]. 1881 [Pre05]. 1883 [Coo01a, Pre08a, Sch02c]. 1895 [Sat04]. 1900 [Epp02a, Gor05, Har00a, Lan00, Law00, Poh00, Pur08, Spa02, Wea02]. 1901 [Bla04a]. 1910 [Luc04]. 1911 [Mac00]. 1912 [Bro00b]. 1914 [Nei02, Pet07b, Bur02, Hac05a, Joh00b, Le 05, McB02, Pet01c, Vee03, Wen01, Wer02]. 1915 [O’G04a]. 1916 [Kno06a]. 1917 [Lan07b, Wer05b]. 1918 [Nei02, Tod02]. 1919 [Hor06, Lau02]. 1920 [Gor02a, Her00, Hin08, Mac09, Vol09b, Wos02a]. 1920s [Bra07b, Pat03, Slo08]. 1922 [McB06]. 1925 [Lip00a]. 1926 [Les04b]. 1929 [And09, Sim03]. 1930 [Mom04, Ben00, Bry07, Jac05b, Mar00b, McN02, Per00, See03, Sim09, Wit07, Yal08]. 1930s [Nic02, Ste08b]. 1932 [Zel01]. 1933 [Arn03, Mom02, Wer07a]. 1934 [Cas07, Cha01, Sid04]. 1935 [Pet03]. 1937 [Dou09, Rei05]. 1938 [Jac05a]. 1939 [Lew00, Mul01b, Pal02, Vee01, Zem02]. 1940 [Bor08, Div03a, Mal01, Mei02, Pri02, Sla09, Slo02a, Sol03, Ste08a, War05a, Wol08]. 1940s [Coc03]. 1941 [Ang04, Gui00, Hig03, Mul07]. 1944 [Cas08]. 1945 [Bug01, Doe08b, Feh04, Gis00, Git01, Has06, Hol01, Moy06, MW07, Sch06c, Sch01a, Waj00, Zel02]. 1947 [Con02, Smu05]. 1949 [Bey08, Gam02, Gom00]. 1950 [Car02, Cro08b, Dav01, Gor01b, Hoc03, Low04a, Win05]. 1950s [Fri01, Nis07]. 1951 [Hor07, Gor01c, Har05]. 1952 [Har08b, Rom08]. 1956 [Tea09, Wan07]. 1957 [Lig06]. 1958 [Jeo08, Zac03]. 1959 [Mei07, Neu02, Wat00]. 1960 [Bla00a, Wes03, Whi02]. 1960s [Sie09, Wis09b, Ste01a]. 1960s-1990s [Ste01a]. 1961 [Ric08b]. 1963 [Bey07, Ham07]. 1964 [Pow02]. 1965 [Bud09, Luc07, Pet00]. 1966 [Ber06]. 1968 [Sve03]. 1969 [Sto05]. 1970 [Ben03, Hac07, Och01, Sto03c]. 1970er [Die01]. 1970s [Coc03]. 1971 [Sau05]. 1972 [Wes00]. 1973 [Fox00]. 1974 [Bai01, Wes04]. 1975 [Ell05, Hac09, Wen02]. 1976 [CK04]. 1978 [Bre01]. 1980 [Aik03, Coo02b, Tei05]. 1980s [Sie09]. 1981 [Fan01]. 1983 [Gif07, Zei03]. 1987 [Kra01]. 1990 [McG00c, Bor09a, Div00, Elk02, Mid05a, Sch05b, Smi03, Zac01]. 1990s [Wes07, Bur00a, Och02, Ste01a]. 1991 [Sat08]. 1993 [Sei03]. 1994 [Div03b]. 1995 [Sha00c]. 1998 [Sto01b]. 1999 [Ano00h, Daw06]. 19e [Pan03b]. 19th [Fro05].

2 [Buc05b]. 2.0 [Bes08]. 20 [Wac05]. 2000 [Ano01e, Cro06b, Dis05, Gie02, Hår08a, Smi05b, Ste01c, Wak05, Wis09a, Wor01]. 2001 [Ano02i, Wos02b, McC08a, Mis02, Tri03]. 2003 [Ano04a, Con07]. 2004 [Ano05a, Esb05, HW05, Ree05a, Sat07]. 2005 [Ano06e, Cor06, Nad07, Bas09c, Her08b, Hin07a, The08]. 2006 [Ano07c, GC07]. 2007 [Ano08d, Sch07a, Jon09]. 2008 [AKP09, Ano09e, Bro09c]. 200th [Ano09d]. 20e [Lev01]. 20th [Bas05, Ede09, Kel09]. 24-28 [Sch06b]. 2nd [Mil04c].

3-D [Rom08].

6-9 [Cor06].

850 [Jaf03]. 850-1520 [Jaf03]. 8F [WW04].

9XM [Doe08a].

A-12 [Hou03]. A.M [Gre04c]. Abandoned [Bro01]. Abbey [Kle03a]. Abbeys [Her01b]. Abbott [Mol05]. aboard [McB00]. Abolition [Olw04]. Abraham [Eri07]. Abroad [Pre08a]. Absenkung [Gis01]. absolute [Ash00]. Abwasserentsorgung [Ber06]. abyme [Mic08]. Abyss [Hoh03]. Academic [Tei01]. Academy [Dym08]. Acceptance [WW00]. Accidents [Luc07, Ten00]. Accounting [WW00]. Accounts [Sch02e]. Accuracy [Tod08]. Accursed [Isr05]. Aces [Nei05]. Acoustics [Arn03]. acque [Lon05b, Squ00]. Acre [War07a]. Across [Neu01, Poh07, Vog04, Dic05, Mil03, Mie09]. Act [Sha05a]. Action [Lip02, Par08]. Active [Smu05]. Activism [Smu05]. Actors [Div00]. acueducto [Por05]. Adams [Tap02]. Adaptive [Pet03]. Addiction [Bar00b]. Admiral [Vin06]. Adolf [Wer06]. Advance [Wri06]. Advancement [Wal01c]. Advancing [Gan00]. Advantage [Kir00, Sto06a]. Advent [Mor00, Sch01d, Tak00]. Adventure [De 06a, Dur06]. Adventurers [Wer05a]. Adventures [Jon04b, Kru04, Sid06a, Vol07a]. Advertising [De 03a, Lai03b, Mil09, Smu05, Str01]. Advice [Ham07]. Advisory [Hec09]. Advocating [Lai03a]. Aerial [Cro07, Dou04, Tri04, Zim04]. Aerodynamic [Mid03]. Aerodynamics [Eck05, Lev05]. aeronauta [Lap06]. Aeronautics [Lev03b]. Aérotrain [Gui08]. Aesthetic [Bla05]. Aesthetics [HV06, Jon06, Rob05a, Tho00, Kat03]. Affinity [Gou04a]. Affluence [Cro08b]. Africa [Cro06b, Sto04b, Dav00]. African [Ako07, All07b, Fou06, Gre05, Ham03, Sto03b, Kin09a, Dob04a]. African-American [Gre05]. after [Eck05, How08a, Pic01, Poo03, Pos04, Rot04a, vD03, Les04a]. Afterlife [Nad04b]. Aftermath [Bey08]. Again [Epp06]. against [Vol03a, Wal01b]. Agatha [Cut02, GS02]. Age [Ade08, Ald00, Ben08, Bla04b, Bro09c, CK01, Cer04, Coo05a, Cor09, Dou04, Epp09, Eri08b, Feh04, Gol00a, Gre08a, Joh06, Kat05, Kou01b, Lip06, Mae09, Maz04a, McG03, McM00, Mol04, Mor01, Oli07, Pan07, Pan02b, Pie06, Poh08, Pop07, Pre04c, Rus00, San02, San06, Sho02, Smi09, Tri04, Vol09b, Wal06b, You09, Bar04a, Bel04a, Cas03, Ham06a, Kin09a, Lip03, Nor07, O’H00, Pop01b, Wel07, Wes01, Cer06, Cor02a, Dug09, Eel09, Ens07, Flo00, Jeo09, McB01, Mil09, Mol00a, Mor03, New06, Olw09, Tho09b, Vol05c, War01a]. Agency [Bro06a, Dob07, Kra01, Ram05, Wri04]. Agent [How08a]. Ages [Bla08, Jaf03, Kid01, Ros06, vD03, HPV06, TeB02, Hur02]. Agrarmodernisierung [Zel04]. Agricola [Lon03a]. agricoltura [Squ00]. Agricultural [Cas07, Dal06b]. Agriculture [And09, Hur04, Lan07a, Sol03, Sto03a, And06a, Boy03, Dan03, Lan03a, Lon03b]. AHA [OS05]. Ahead [Hin07a]. Aichi [Nad07]. Aid [Sch05a]. Aiken [Wil00b]. Aiming [War01a]. Ain’t [Gom00]. Air [All04, Bed03, Bed05a, Bug04, Con01a, Dou04, Fri02b, Gor01c, Hig03, Hol01, Low04a, Mor05b, Nei07, Pal00, Sid08, Vai00, Wil00c, Ber04, Bla00a, Kri00a]. Air-Conditioning [Ber04, Bla00a]. Air-Mindedness [Pal00]. Airbus [Mod08]. Aircraft [Cas04b, Gui00, Joh04, Sin05a]. Airfield [Con02]. Airfields [Hol01]. Airline [Tom04]. Airlines [Rot08, Bed03]. Airplane [Bug01, Lev05]. Airports [Bed05b, Con02]. Airpower [Gui00]. Airship [Gli00, Pal02]. Airspace [Dou09]. Aitken [Kan07]. al [Lon02a, Ang04]. Al-Qaeda [Ang04]. Alan [Kid06]. Alarms [Sau05]. Albear [Por05]. Alchemical [Sha00a]. Alchemy [Buc05a, Har07a, Jen05, Pow09]. Alcohol [Ash01]. Alcoy [PP03]. Alfred [Dic04, Hug07, Gif07]. Alienation [Kno03]. Aliens [Mor00]. Aligning [Bas09b]. All-Consuming [Jac02]. All-Inclusive [Rob08b]. Alley [Bas09c]. Alliance [Car03]. Allied [Bro00a]. Allies [Chu07]. Allow [Les08b]. Alltag [Osi00]. Allure [Iri03]. Ally [Ang05a]. Alms [Kha08]. Alpen [Vee01]. Alpes [Bel04b]. Alternative [Sch04a, Wri00, Cep08]. Alternatives [Iri06a]. Altitude [Pri03]. Alva [Pre08b]. am [Zet08]. Amalgamated [Zem02]. Amateur [Fro05]. ambientali [Gag07]. Ambiguity [Sla01]. Ambition [Tat05]. Ambitions [Har07a]. Ambrose [Pet05]. America [Arn03, Bek05, Ber04, Bla09a, Bla00a, Bor08, Bor09a, Bra03a, Bro02, Bur00a, Coa02, Coo07a, Coo04b, Day04a, Day06, Edg08, Fan01, Gam07, Gor01c, Guc08, Hok07, Jos04, Kie09, Luc02, Mac09, Mea05, Mei07, Mel06, Mor07b, Mul07, Pos03, Qui01, Rot04a, Rus08, Sad06, Sha05a, Sil00, Sin06, Sli09, Smi06, Sta01, Sto03b, Tei01, Waj00, Wal01c, Wer05a, Win05, Wor01, dlP07, Cas04a, Lau08, Les04a, Pat03, Sch06a, Smi09, Sti03, Wer06, Wis09b, Ade08, Bed05a, Ber02, Bra01a, Bra08a, Bro05a, Car09, Coc04, Con02, Coo03b, Cor08, End01, Eri08a, Fit02, Fra08, Fra09, Fri01, Fro09, Gin07, Gla01, Gre08a, HM01, Hec09, Hoo04, How08b, Hum09, Iri00, Jac02, Jac06, Jac07, Jag00, Kle06b, Kno07a, Kru04, Lai07, Lak05, Lar03, Las06, Mag00, Mar06a]. America [Nye06, Per06b, Pyn09, Rii03, Rob08a, Ros07, Sch05c, Sch04c, Smi02, Sol04, Ten05b, Tho06, Tho08, Vie08, Wal05]. American [Ake05, And09, Ang06, Bar03c, Bij03, Biv07, Bla01, Bla04b, Ble04, Bro06c, Bro09b, Bry07, Bug01, Bug07, Cas04a, Col08, Con00a, Coo00, Coo02c, Cru03, Cut06, Dan03, Dir07, Don07, Epp09, Fer05, Fos04, Fra03, Gan05, Git01, Gle03, Glo02, Gom00, Goo09, Gor02a, Gor04b, Gre05, Gre02, Hac09, Hal07, Ham06b, Hea02, Hin08, Hov04, Hug09b, Iri01, Iri06b, Iri06c, Iri08a, Jac05b, Jac00, Jac04b, Jeo08, Kee03, Kha01, Kle08, Kno05a, Kol07, Kra07a, Lai03b, Law03, Lev08, Lev03b, Lev05, Lic02, Lip00a, Lon03b, Low04a, Luc07, Mac05b, Mac05a, Mad06, Mag05b, Mar00a, Mar00b, McE06, McE07, McS02, Mel01, Min03, Moh08, Mol08, Mor05b, Mou08, Mul04, Nei02, Nei05, Nie07a, O’H05, O’H07, Ott08, Pet07a, Pfa06, Pra01, Rau01, Rom03, Rot08, Rot04b, Ryd01]. American [Sat07, Sch02d, Sch05d, See03, Seg00, Sew06, Sha09a, Sla09, Spa02, Ste07b, Sto03c, Str01, Tea05, Tea09, Tei05, Tes09, Uss03, Vol07b, Wal01c, Wan02, War06b, Web08, Wei03, Wes04, Win08, Wis09a, Wol08, You09, Ang01, Boy01, Bud05, Bur00b, Clo07b, Flo03, Her08a, Kry04, Lew01, MK01, Ree05b, Van02, Wel07, Cha01, Gro08, Gui00, Ham08a, Hur04, Kin09b, Kou03, Law00, Mai04, New06, Rus00, War05a]. Americanization [Sto01a]. Americans [All07b, Bar04b, Fou06, Goo06, Kin09a]. Americas [Ham03]. Ames [Ryd01, Gol00b]. Ami [Bow00a]. Among [Kru04, Vas01, Bor00, Muk06]. Amos [Coo02c]. Amsterdam [Ano05a]. Analog [Bra03b]. Analogy [Egl07, Eis09]. Analyse [dS09]. Analysis [Day01, MB06, Sch08b, For08]. Anatomy [Nei07]. ancienne [Kah01]. Ancient [DO03, KO08, Lie01, Luc05a, Maz07, McG00a, Rol05, Whe00]. Andreas [Pow09]. Andrew [Nag08, Hin07b]. Androids [O’G05]. Andy [Coo02c]. Anesthesia [Bac02]. Anfange [Mos06]. Angeles [You01, Abb05, Jud05, Sma05, Spo02]. Angewandte [Har08b]. Anglo [Con00a, Fis03]. Anglo-American [Con00a]. Anglo-Saxon [Fis03]. Anhang [Pet01a]. Animal [Rus07]. Animals [For05]. Anime [Nad08]. Annaherung [Rie05]. Annales [Lon05a]. Annihilation [All05]. Anniversary [Bij09, Pos08a]. Annotated [Kar02]. Annual [Cor06, Esb05]. Anspruch [Gal00]. Answer [Cla02c, Epp02b]. Answers [BP02]. Ant [Per06b]. Antarctic [Fal06a]. Antebellum [Coc04, Eel09, Kno07a]. Anthology [Ker06]. Anthracite [Lon07]. Anthropological [Ada02]. Anthropologist [Bor00]. Anthropology [Sha08, Wat02]. Anti [Sta05]. Anti-Typhus [Sta05]. Antibiotics [Bud05]. Anticipating [Nei02]. Anticommunism [Rus00]. Antiquity [Bla08, KO09, Tri04, Wak04a, Hac06]. Antoine [War03]. antologi [Hyl00]. Antonio [Twa00]. Anxieties [Bud02]. Anxiety [Mae09, Rus00]. Any [Ree05b]. Anytime [Slo04]. Anywhere [Slo04]. Apocalypse [Gor04b, Kol06]. Apocalyptic [Jos03]. Apollo [Bai01, Bea02, Eze02, Her09a, Mac05a, Her08a, Joh01, Hun00]. Apollonian [Sch08a]. Appalachia [Ste08b]. Appalachian [Her00]. Apparatus [Rob04]. appareils [War03]. Apparent [Kri00a]. Appetite [Moo02b, Pas06]. Applaud [PB03]. Application [For09, Bek03]. Applied [Gie02, Smi03]. appliquée [Le 07]. apports [Bel04b]. Approaches [HPV06, McG06]. approches [Dav09]. Appropriation [Lew00]. April [Bro09c, Pre08a, Ano00k, Ano01g, Ano02l, Ano02o, Ano03f, Ano04h, Ano05i, Ano06i, Ano07f, Ano08g, Ano09h]. Aquarium [Rit07]. aquis [KO02]. Arabian [Ten05a]. Arcadian [Rob04]. Archaeologist [Col03]. Archaeology [Bek04c, DeV03, Gli00, McC00b, McG02, Mar02, Fal06a]. Archeology [Wer07b]. Arches [Mil03]. Archigram [Pic06]. Archimedean [SD09]. Archimedes [Wal06b, DO03]. Archipelago [Tag05]. Architect [Wer06, Bro09a]. Architects [Ald00, All05]. Architectural [Arn03, Ben00, McG03, Wat02]. Architecture [AST08, Bra08b, Cla07, Dor00, Hex07, How09b, Hyd00, Iri00, Iri05, Iri06a, Kim04, MS08, Mil02, Mil08, Mol02, Pet03, Pic06, Sam05, Sew06, Ste05a, Wer07a, Win08, Iri08b, Abe02, Ber03b, Big07, Der04a, Lit03, McG03, Pic06, San02, San06]. Architectures [Spi01]. Architektur [Pet04c]. Archives [Sch04d]. Arco [Gis07]. Arctic [Guc08, Lon00]. Argand [Bow00a]. Argentina [Cas08]. Argentine [Cul09]. Arguments [DvdE03]. Arizona [Ell05]. Armageddon [Bar04b, Pow02]. arme [Wer02]. Armed [Sho02]. armèe [Pan03a]. Armillary [SS02]. Arms [Gor07, Smi07, Hac05a]. Armstrong [Hac05a, Neu06a]. Army [Mor03, Moy06, Ang01]. Arnold [Bek07]. Arrested [LW05]. Arrows [Rol05]. Ars [Wal04]. Arsenal [Kan07]. Art [Bac05, Bro09c, BW04, Coo00, Fis03, Gan05, Lin05, Lon01b, Low04c, McG02, Pet04b, Pet09, Rey02a, Smi07, Spa02, Tra05, Wal06b, Wri06, Bro09c, Goo09, Rap01, Rob06c, War05b]. Arte [Lon05b]. Artefacts [Kel04]. Artefactual [Kel04]. Artefakter [Har06a]. Artifacts [Bla00b, Sha06d, Col03, Gro08]. Artificial [Pic00, Der04a, Lon04b, Sin00]. Artillery [Bac06a]. Artisan [Lin05, Mil06]. Artisans [Con00a]. Artist [Wos02a]. Arts [Ben01b, Jon00, Tho00, Wak04a]. Asbestos [Kno07b]. Ascendant [Moy05]. Asia [Hac05b, Vol08a, Tag05, Vol01]. Asian [Vol04a, Vol08b]. Aspekte [Sto01b]. Asphalt [Hol03]. Assault [Orr02]. Assembled [Tak01]. Assembling [Gan05]. Assessing [RK08]. Asset [Zel01, Zel02]. Associates [Lig06]. Association [Cor06, Esb05, Wal01c, Gou04c]. Associations [Yar07]. Astrolabes [SS00]. Astronomers [Rot04b]. Astronomical [Tat05]. Astronomy [Dal06a]. Athena [Wil03b]. Atkins [Goo09]. Atlanta [Ano04a]. Atlantic [Dic05, Hub09]. Atlantropa [Gis01]. Atlantropa-Projekt [Gis01]. ATLAS [Mod08]. Atlases [Clo05]. Atmosphere [Kri03]. Atom [Jen01]. Atomic [Che04, Dev00a, Hec08, KM04, Kru04, Mol06, O’H01a, Sil00, Kin09a, Jos03]. Atoms [Wan04, Ham06a]. Atomzeitalters [Hes04]. Attached [Ree05b]. Attack [For08]. Attacks [Wei03]. Attempted [Moy02]. Attempting [Sch04a]. Attention [Dug09, Pop07]. Attic [Pos01b]. aube [Lau02]. Audible [Bij04]. Audience [Dob08c]. Audiences [Mul01b]. Audion [Bow03]. Aufbau [Neu02]. Aufstieg [dS09]. August [Ano01e, Bro09c, Dor08, GC07]. Auschwitz [All05, Moy02]. Austin [Cep08]. Australia [Biv07]. Authenticity [Rob05a]. Author [Ano00a, Ano02a]. Authoritative [Hai03]. Authority [Ash00, Kle09, Zac06]. Authors [How08b]. Authorship [Sch01d, Wak04a]. Auto [Bre00, Bor09b, Cor08]. Autobahnen [Zel06]. Automatic [Kid06]. Automation [Pet04d, Ham06a]. Automobilbau [Mom02]. Automobile [Fra08, Lic02, Luc09, McI00, McS02, Vol05c, Wel08, Bar04a, Bor06b, Lun09]. Automobiles [Fos04]. Automobilitat [Mos06]. Automobils [Mom04]. Automotive [Jon02a, Lug03, Tho09b]. Autophobia [Tho09b]. Autos [Mom04]. Aux [Gou00]. avancement [Gou04c]. Avant [Lar08, Gan05]. Avant-Garde [Lar08, Gan05]. Avatars [Dug08]. Aviation [Bed03, Boc09, Bye04, De 06a, Goo04, Her07b, Jeo08, Sid08, Con03]. Aviator [Bed05a]. Avoiding [Wri06, SHW03]. Awaiting [Bar04b]. Awards [Ano00b, Ano01a, Ano02b, Ano04b, Ano05b, Ano06a, Ano07a, Ano08a, Ano09a]. Away [For09, For09]. Axes [Rab04]. Axis [Fer05, Van02]. Ayanz [San03]. Aztec [Tow02].

B [Bro05b, Mil03, Rot04a]. Babel [New06]. Baby [Cow03]. Back [Ano08d, Gam07]. Backcountry [BS09]. baco [Lon02a]. Bacon [dSL04, McC07a, Eis09, McC09a]. Bacteria [Sha05b]. Bad [War08]. badischer [Pet04c]. Baghdad [Bek04a]. Bahn [Sch05b]. Bahnen [Vee00, Vee01]. Balance [Rau01]. Balancing [Pri00]. Bald [Rob00]. Balloons [Bow00a]. Baltic [Bra08b]. Baltimore [Zem02, Gwy04]. Bamboo [vF05]. Ban [Ham07]. Banana [Lan07a]. Banham [Kat02]. Banjo [HM01]. Banknotes [Rob05a]. Bankrupting [Nis08]. Bark [Coo06]. Barnes [Zai08]. Baroque [San02, HV06]. barrages [Bel04b]. Barrel [Tei01]. Barrier [Bij02]. Barton [Sin09]. Based [Dyb09]. Bases [Hol01]. BASF [Mer05]. Basin [HÖ08]. Bathroom [Ben05a]. Bats [Kel06b]. Battle [Abb03, Hac06, McB05, McB07, Sha01a, Sin06, Wil04]. Bau [Cie01]. Baumeister [Pet04c]. Baustatik [Pet04a]. Bay [Gam07, Sha01a]. Be [Ben05b, Ham05, Sch07e]. Beaumont [Bla04a, San03]. Beautiful [War08]. Beauty [Big07]. Became [Dor08, Gou04b, Pic00, dlP07]. Becoming [Bal00]. Bedini [Pos08b]. Beeching [Cro08a]. Beef [For03]. Been [Moy02]. Beer [Ros06]. Before [Wil00a, Bek05, Bul03a, Gei00, Nis08, Owe03, Sch06a, Ter00]. Beginnings [Con03, Lon04a, vF05, Mar06a]. Begins [Mol07]. Begrenzte [Zel06]. Behavior [Bla00b]. beherrschen [MW07]. behind [Hol06, Tom03]. Being [Cro08b, Dob04b]. Beings [Bla00b]. Beispiel [Zet08]. Beitrag [Mar03]. Beitrage [Heß04]. Belief [Tow02, War01b]. Beliefs [Con99]. believe [Lin00]. Bell [Coo02b, Luc09, Fin09]. Belli [Wal04]. Below [Kai02, Shu09a]. Ben [Ell01]. Benjamin [Lev08, Mol04, Rob06c, Sch04b, Ste07a]. Benvenuto [Pet05]. Berge [Vee01]. Berkner [Rau01]. Berlin [Hut00, Cav08a]. Berube [See08]. besser [MW07]. Best [Rot08]. Bethe [Olw01]. beton [Wer02]. Better [Joh02c, Low01a, Mei07, Coo05b]. between [Cas05, Cas07, Clo07b, Rud09, Ste06, Ash01, Owe03]. Beyond [Ano08d, Osl07, Vol05b, Nor09, Ala09, Arn00, Bea02, Daw07a, Mar05a, Pay05, Sin05b]. Bias [Mul09]. Bias/Time [Mul09]. Bibliography [Ano00c, Ano02c, Hov04, Low00b, Low02b]. Bicycle [Epp02a, Epp03, Vol05a]. Bier [Mar03]. Big [Bud05, Lev00, Sam07, See08, Ste04, Wil07, Wil06, Iri03, Mis02, Sad06]. Bilbao [PP09]. Bilder [Heß04]. Bilderhandschriften [Wal04]. Billboards [Fos04]. Billion [Hou03]. Bilsamhället [Jør09]. Bind [Elk02]. Biogas [RV04]. Biographical [Con07]. Biographie [Men07]. Biography [Dis05, MN06, Neu06b, Wes00]. Biological [Bes08, Rol05, Sch02b]. Biology [Poo03, Sch02b, Ser05, Tho00]. Bionic [O’G05]. Biotech [Cha07]. Biotechnologie [Joh02a]. Biotechnology [Sha05b, Bud02, Cha02, Pow04]. Bioterrorism [Sha05b]. Bird [Lev05]. Birth [Bed03, Ben01a, Bri01, Cer06, Dal06b, Gam05, Hin09, Hor06, Mag05b, McB01, Mol00a, Mor08a, Mor05b, Sch05e, Sin06, Van00a, Wer05b]. bis [Ber06, Die01, Fox02, Low00a, Mom02, Neu02, Pet00, Sch05b, Zel04]. Bits [Rus09]. Black [All02, Bow03, Car09, Eri08a, Lak01, Sat07, War05b, Min00, Ham03, Pre04c]. Blackett [Rau05]. Blacks [Hor06]. Blacktop [Hol03]. Blame [Kra04]. Bleaching [dlP06]. Bletchley [Kid07]. Blind [Con01b, Jeo08]. Blogging [Tur09a]. Blue [Her01a, Tom03]. Blues [Dis06]. Boat [Thi06]. Boats [Bro07, Sve02, McG02]. Bodies [Bow00c, Bro09b, Bud02, Isr03, Pic00, Pre04a, Van00a, Wei01, Fit00, Kou02, Vol09a]. Body [Bak05, Bin06, Goo07, Ker06, Kno07b, Lin05, Mai04, Ten05b, Tho01, Ver08, Bra08c, Bul01]. Bold [Sto09a]. Bolshevik [Sid04]. Bolt [Sch04b]. Bomb [Jos03, KM04, Kin09a, Mea05, Olw01, Wal01a, Sil00, Fri09, Kru09]. Bombards [Buc05a]. Bombay [Sat04]. Bombe [All06b]. Bomber [Hou03]. Bombers [Moy06]. Bombing [Gui00, Moy02]. Bombs [Rob08a, Rol05, O’H01a]. Bonanzas [Con06b]. bonifica [Squ00]. Bonn [All03]. Bonneville [LB08]. Book [AA01, How08b, Kie09, Lon09, Bru01, Joh08]. Booklet [OS05]. Books [Alc07, Alt00, vF05]. Boom [Hor04a, Coa02]. Boomers [Wei03]. Boorstin [Skr04]. Bootstrapping [Bar02]. Border [Ell05]. Borderless [Tur08]. Borders [Mie09]. Born [Col05]. Bosch [Hor02]. Boston [Elk06, Sha01a, Gam07, Wer04]. Bound [Luc04]. Boundaries [Cho07, Och02, PB03]. Bounds [Tho06]. Bounty [KO07]. Bourbon [Reu06]. Bourgeois [Tod02]. bout [Mul01a]. Bowl [O’H07, Sti07]. Box [Bow03, Fro05, Min00, Wil07]. Boy [Alc07, Coo03a, Alc07]. Boys [Hoc03, Smi02]. Bracing [Wer07a]. Brand [Bar08]. Brandy [Bow00a]. Braun [Fos09, Cro06c]. Brazil [Bre00]. Brazilian [Sha04]. Breach [Kol07]. Breaches [PB03]. Break [Lai07]. Breaking [Big08, Mod08]. Breakthroughs [Mil04a]. Breast [Gar09, Her01c]. Breath [Pre04d]. Bred [Ten05a]. Bretagne [Bor06a]. Brickmaking [Hol09]. Bridge [Doi06, Kra04, Mil03, Mil04c, Pos08a, Win02, Wos02a, Mil03]. Bridges [Mid03, Och02, Rey01b, Mil01]. Bridging [Mid06]. Brief [Div03b, Fri09, Gre06a, Kha08]. Bright [Tra05]. Brilliant [Bow08]. Bringing [Bar09b]. Britain [Ber02, Bud09, Cro08a, Cro08b, Div03a, Far00, Gre04b, Hac07, How08b, Jaf03, Mal01, Pow02, Sch06c, Smi03, Bro00c, Bow00b, Goo07, Ric08a, War09c]. British [AKP09, Doe08b, Dor01, Dou09, Epp03, Gif07, Gle03, Gui00, Hac05a, Kno06b, Mac00, McG00c, Moh08, Wit08, Ash01, WW04, Edw03b, Mac01]. Broadcast [Whi02]. Broadcasters [New06]. Broadcasting [Kry04, Mul07, Slo08, Smu07, Soc08, Sto09a, Mor07a, Wan02]. Broken [Egl07, Ens07]. Broker [Tea08]. Bronze [McM00]. Brook [Her08b]. Brooke [McG04]. Brookhaven [Wes00]. Brooklyn [Doi06]. Brothers [Dou04, Tri06b, War06a, Mil04b]. Brought [Lai03b]. Brown [Hey08]. Brownie [Oli07]. Brunel [Coo04a, Gor08]. Brute [Jud03]. Bryant [Les07]. Buda [Fri09]. Buffalo [O’H06]. Build [Kid06]. Builder [Her01b, Mil03]. Builders [Roc02]. Building [Aik07, Big08, Edw03a, Gui08, Hof02, Hol01, Iri01, JI04, Kai02, Les08b, Low05a, Mom05, Mou08, Och02, Pet09, Sad06, See03, Sim06, SS07, Sti03, Tea09, Tom03, Vog04, vD02, vD09, Ber03b, Gar09, Jac05a, MS08, Nak07, Pet03, Ven00, Wan07, Wei04]. Built [Hof02, Mai04, Mil08, Nye05]. Bulldozer [Sch02d]. Bullet [Bug04]. Bungalow [Sim03, Ste08a]. Burden [Hug02]. Bureaucracy [Tod02]. Burgundy [Bac06a]. Burlingame [Seg09]. Burnham [Iri08a]. Burning [Flo00, Rey01b]. Bus [Sch00c]. Business [Bor09a, Buc03a, Bud05, Bug07, Col07c, Gle03, Mac02, Mie04, Moh08, Ros05a, Sha09b, Str01, Tei05, Tod02, War08, Wes03, Wil01, vR07, Dow07]. Buyways [Fos04].

C [Bro03, Lev03b, Mil03]. Cable [Bek05, Mor05a]. Cage [Sha05a]. Calculating [Kid08]. California [Bre01, Har07b, Sch05e, Rot02, Wak04b, Kir00, Wil03a]. Cambridge [Low05b]. Came [CK01, Hol03]. Camembert [Gou04b]. Camera [Bow07]. Camouflage [Mei00b]. Campaigning [Kno01a]. Campaigns [McC09c]. Camps [Van02, Mie04]. Campus [Sti08]. Can [Ham05, Wil02b]. Canada [Dur06, Epp02a, For03, JI04, Boc09, Pav09, Guc08, Mos02, Shu08, dO01a]. Canadian [Bro00a, Hov04, Jon00, Cro07, Sla02a]. Canal [vD09, Jon04c]. Canal-Building [vD09]. Canals [Gei00, Sha00b]. Cancer [Gar09]. Candace [Spa02]. Cane [War09a]. Canvas [Cer09, Low05b]. Canyon [Ell01, Sha07]. Capital [Ble07, Car08, Les01, Wer06, Kno06a]. Capitalism [Mok09, Rei05, Roc02, Sch02a, Chu04a, Sti08]. Capitalist [Dow05, Gud07]. Capitol [Sch08b]. Captioning [Bar09a]. Capturing [Kra05a]. Car [Bor08, Fri09, Sch02e, Sie09, Vol08b, Bor02, Hok07]. Cardiac [Law03]. Cardwell [Mar00c]. Care [Gar09, Law03]. Careers [Buc01a]. Caribbean [Cha01, Sol03]. Carl [Hor02, Iri08b]. Carnegie [Hin07b]. Caro [Tea08]. Carolina [Con03, Mar09]. Carriage [Bro05a]. Carriages [Sha06e]. Carrier [Gui00]. Carriers [Ham08b]. Carroll [Cer09]. Cars [Bar07, Ben05b, Bla09c, Bor09b, Cor09, Lug03, Pin08, War06a, Lun09, Mid05b]. Carson [Kar08, Kar09b]. Carter [Sch05e]. Cartographers [Har02]. Cartographic [Kar09a]. Cartographies [Clo06a]. Cartography [Clo03]. Case [Bij02, Fre07, Jac04a, Kri00b, LW05, Las06, O’H09]. Cash [Ros07]. Catalogue [SS02]. Catapult [Bac07]. Cathedrals [San00, Her01b]. Cattle [Ten05a]. Causality [Rol07]. Cause [Win08]. Caverns [Lon01b]. Cell [Mod08, Eis09, Aga09]. Cellphone [Tri06a]. Cellular [Coo03b]. Cement [Muk06]. Centennial [Con07, De 04, Ryd03, Tho09c]. Center [Hun00, Sat05, All04, Day08, Flo01]. Central [Dav00, Lin02b, Mil02, Pre00]. centrales [Bla03]. Centuries [Gor02b, Gre02, Lin02b, Mie09]. Century [Ade08, Arn01, Bra08a, Cor08, Dav00, De 03a, Ens06, Fra08, Fro05, Gam07, Gla01, Hac04, How08b, Hur04, Igl01, Iri00, Iri01, Jac02, Jac07, Jag00, Jon04a, Kha01, Kle06b, Kno06b, Kuw00, Lar03, Las06, Mar02, MS08, Mis02, O’H09, Per06b, Pet07b, Pyn09, Ras01, Rob07, Rob05b, Rol06b, Sch07a, Shu02, Sil00, Sim06, Sti09, Tym00, Wak02, Wal05, War07b, Wes08, DvdE03, Gag02, Lau08, Lon09, Muk06, Por03, Rey02b, Sch05a, Sto04b, Bas05, Bro07, Cla01, Dir07, Ede09, Fal06b, Far00, Fel07, Flo07, Fra02, Gor04b, Gre08b, HM01, Isr03, Jac00, Jon00, Kel09, Lig09, Lip06, Lon07, McE06, McE07, McG00a, Pye01, Ran01, Rau05, Rot08, Sca08, Seg08, Sha00a, Sto03a, Tea05, Tho01]. CERN [Mod08]. CERN-ATLAS [Mod08]. Certainties [Con06a]. Châtelier [Le 07]. Chadwick [Gol00a]. Chain [For03]. Chair [Coo07a, Mar05b]. Challenge [Bri08, Cro08b, Scr09, Bai01, Kar09b]. Challenged [O’H09]. Challenges [Bij09, vD02]. Chalons [Pan02a]. Chalons-sur-Marne [Pan02a]. Champagne [Gou04b]. Champion [Tom03]. Chance [Pos09a]. Chancery [Coo01a]. Chandler [Hug07]. Change [Car08, Con08, Daw07b, Epp03, Fit02, Gor06, Her00, How08a, Kai02, Lan07a, Lug03, Mor08b, Pra03, Rol03, Sch01a, Sve02, Vol04b, Zac01, Fis09, Gan03, Gor02b, Kon06, Kru02, MA09, Ver00]. Changed [Ake05, Buc05a, Hug05, Kno05a, Kra05a, McB00, McE04, Wil04]. Changing [Bud09, Hor07, Mac04, Sch06a, Sha00a, Sla02a, Yar09b, Bow00c, Kri03, Lan03a]. Chapter [Sta05, Whi05]. Chapters [Kar02]. Charcot [Pan05b]. Chariots [Men09]. Charles [Fer05, Fin03, Reu05, Sch07c, Cav08b, Poh07]. Charlotte [Cro05]. Chart [MA09]. Chasing [Bug04, Dic05]. Chaussées [Gra08]. Cheap [Ben00]. Cheaper [Joh02c]. Chemical [Col06, Gou04a, Joh06, Pow09, Rol05, Sho08]. Chemicals [For04, Ser02c]. Chemie [Fox02]. Chemins [Haf07]. chemische [Har08b]. chemischen [Joh02a]. Chemist [Joh00b, Str05]. Chemistry [For09, Hor07, Joh00b, Low04b, Mul01b, Sha00a, Low02a, McC00b, Spe04]. Chess [Tom03]. chez [Lau02]. Chicago [Iri08a, MS08, Sim03, Iri08b, Les08b, Bri06, Iri06a, Sch00b, Yar09a]. Child [Dir05, Hun02, Pre04a, Daa07]. Childhood [Aik03, Dir05]. Childlessness [Bul03b]. Children [Dir05, Dir07, Tym00, Mor00, Pas06]. Chilean [Bri07]. Chill [For03]. Chimney [Qui03a]. China [Biv03, Dan01, Ham09, Hig01, Maz03, Pet03, Sch02b, De 06b, Leo01, Per07, Vol03a, Wan08]. Chinese [Baa01, Rap01, Rei05, Spi00b, vF05, Kry04, Biv03]. Chipley [Cut02]. Chlorinated [Jac04a]. Choice [Bug01, Das08, Sch00c, Tho00, Tho09a]. Choices [Fri07b, Neu01]. choix [Pan09]. Choreographies [Kat05]. Christ [Ale03]. Christine [Bix04]. Christophe [Bas07c]. Chroniques [Lev03a]. Chrysler [Jon02a, War04]. Church [San00, Jen00]. CIA [Day04a, Joh02b]. ciencia [dSL04]. Cinema [Bra07b, Bra08c, Daw07a, Gom05, Jos03, Rom08, RP09]. Circa [Yea04, Gra08]. Circolazione [Dol09]. Circulation [Bel04a]. Citations [RK08]. cite [Mul01a, Gwy09]. Cities [Bri06, Kon06, Lig09, Luc02, MK01, Sha01a, vD03, Sim09]. Citizens [Jos01, dO01a]. Citizenship [Jac07, Bla09c]. City [Cep08, Dob08a, Flo01, Gre08b, Iri02, Iri06c, Jac05a, Jac08a, Jud05, Jud08, KM04, Kol06, Kru04, Lan07b, Les04b, Lin06a, Luc04, Luc06, Mid05b, Mol06, Ren09, Rot02, Sew06, Sha06e, Sta01, Tha04, Vee05, Wer05b, Wer06, Epp09, Fin02, Gro09, Iri01, Iri08a, Kno05b, Kol07, Mil02, Pla06, Rom03, Sch03, Sim06, Vog02, Win08]. Civic [Jac05a]. Civil [Bar03c, Col07c, Eri07, Gei00, McB00, McB04, Smi06]. Civilisation [Sch02b]. Civilization [Jen00, Cer09, Wil02a]. Civilized [Nei05]. Civilizing [Mor07b]. Cladding [Les08b]. Claim [Bra07a]. Clark [Sca05]. Class [Jac06, Lak05, Les04b, Sim03, Smi02, Zem02]. Classical [Hac06, KO07, Muk06]. Classification [Low01b]. Classroom [Cru03]. Claude [War05b]. Clay [Sch03]. Clayton [BP02]. Cleansing [Dob08a]. Clears [Les04a]. Cleavage [Her01c]. Clemens [Fis07]. Cleveland [Mil04b]. Clifford [Pet05]. Climate [Pin08]. Clippers [Fra06]. Clock [Bra03a]. Clocks [Sha06e, Ste05b]. Close [Cro06a, Kan07]. Closed [Bar09a]. Cloth [Moh04, Ako07]. Clothing [Dir05, Zem02, Ako06]. Cluss [Wer06]. Co [Dob07, Els09, Scr05]. Co-Construction [Scr05, Els09]. Co-Production [Dob07]. Coal [Kno05a, Kno06b, Lon01b, Qui05]. Coast [Tag05, Kol08, MA09]. Coastal [Bur06, DvdE03]. Coaster [War04]. Cod [Hub09]. Code [Coo04c, AA01, Tur06]. Codebreaker [Kid06]. Codebreaking [Kid07]. Coevolution [Bar02, Sto06a]. Coffee [Low04b, Owe04]. Cognitive [Sch08b, Sla04]. Cognitively [Kel04]. Coincidence [Fin09]. Coke [Kno03]. Cold [Bar03c, Ber03b, Cru03, Day01, Hec08, Hec09, Her01a, Kid08, Nad04b, Rau01, Rus00, Sch04c, Sto09a, Wei06, Bas09b, Boc09, Fre07, Lai03a, McC09b, Slo02b, Wei03, Low04a, Mul04, Smi03]. Collaboration [Ake01, Bud09]. Collaborative [Tur06]. Collapse [AST08, Hou03, Kra04]. Collapsed [Epp06]. Collecting [Gan03]. Collection [Goo09, Sha06d, Gre04b, SS03]. Collections [May00]. Colleges [Cas07]. Collier [Lev00]. Collies [Ten05a]. Collins [Bek02]. Colombia [Och01, VCM^{+}09]. Colonial [Big08, Bur00a, Col05, Gan03, Hea01, Kin09b, Mil06, Sat04, Sta01, Sto01c, Tow02, Van08, Biv07]. Colonialism [Gan03]. Colony [Bar03a]. Color [Bro05b, Rob07, Eri08a, Tra05]. Colorado [Ell05, Pri03]. Colossus [Epp00, Kid07]. Colouring [Tra02]. Colt [Smi07]. Columbia [AKP09]. Columns [Les08b]. Combat [Nei02]. combinées [Dav09]. Come [Ben05b]. Comes [Ben05b, Sch06a]. Comfort [Ber02, Ber04]. Coming [Dug09, Mor06b, Mul06, O’G04b, Smi09]. Command [Day06, Fel07, Mol02, Pow02, Fel07]. Comment [Epp02b, Scr00, Wak04b]. Comments [Tur09a]. Commerce [Bla05, Lig06, Moh04, Rey02a, Ser05]. Commercial [Bed03, Bro06c, Doe08b, Mar09, Spi04]. Commercialism [Jac02]. Commercialization [Mar00b, McC09b]. Commercializing [Mod06b]. Commission [Pop00]. Commissioners [Coo01a]. Committee [AKP09, Hec09, HW05, Jon01, SD01]. Commodification [Dir05]. Commodities [Bow00c]. Commodity [For03, Mor06a]. Common [Ble07, Min03]. communes [Ben03]. Communicating [Isr03, Mul01b]. Communication [Car02, DK00, Dow08a, Hoc03, Sch09b, Bla00b, Vai00, Yal08]. Communications [Ade08, Bil05, Bre06, Bri07, Dic05, Div00, Dow05, HA00, Hug09b, IEDB04, Joh00a, KM04, KFB00, LL04, Mul09, Pav09, Slo02b, STW01, TCOA07, dO01a, Rou05, Wha06]. Communism [Goo06]. Communist [Fre07]. Communities [Abb03, Con06b, Joh05, Mod06b, Sha00b, Bar09b, Nie07b]. Community [Mil06, Ryd01, Tur05, Rob00, Shu08]. como [dSL04]. Compact [Bas07b]. Companies [Bek05, Bra08b, Cer04, Mil04a]. Companion [Ham06b, Joh08]. Company [Bro07, Jon04a, Lan09, Lig06, Epp00, Las03, Mer05, Sch06d]. Comparative [For08, Gar09, Har02]. Comparing [Och02]. Compatible [Dow06]. Competence [JI04]. Competition [Doe08b, Mul05, Ste01c]. Competitive [Sto06a]. Completed [Haf05]. Complex [Cow06, Kim04, Sch02c, Mea05, Sch09b]. Complexity [Dow03b]. Component [Sch08b]. Component-Stimulated [Sch08b]. Comprehensive [Coo09]. Computer [Bij05, Bor00, Dow06, Eel09, Ens06, Has02, JI04, Kid01, Lig06, McG00b, Tom03, Wil00b, Wil04, Wil00b, Coo05c, Dow05, End00, Fri02b, Kid06]. Computer-Compatible [Dow06]. Computers [Ada06, Ake05, Ben05b, GS02, Kid08, Mos02, Pfa06, Spi01, Gui03, Has02, Kid07, Lig06]. Computing [Bas09b, Kid03, Kid06, Owe03, Sei03, Sha00c, Uss04, Wil04, Cro05, Bar02, Fit01, Hai03]. Comrades [Lun09]. Concentration [Mie04]. Conception [Sma05]. Concepts [Ots03, Wen01]. Conceptual [McG06]. Concerns [Pra03]. Conclusive [Rob08b]. Concorde [De 03b]. Concrete [See03, Sch03]. Condit [Iri08b]. Condition [Lig02]. Conditioning [Ber04, Bla00a]. Conditions [Goo07, Wel07]. Condon [Las03]. Conductor [Cav08b]. Confederacies [Gor01a]. Confederate [Col07b, Kno01b, Kno01a]. Conference [Cor06, Esb05, Sch06b]. Confidence [SS07]. Confidential [War08]. Confined [For05]. Conflict [O’H01b, Sat05, Sid03, Sho02]. Conflicts [Sha04]. Confusion [Bri07]. Congenial [Kel04]. Congress [Tei01]. Connecticut [Kel05, Rab04, Rob02, Ryd01, Sha06e]. Connection [Sad06]. Connellsville [Kno03]. Conquering [Poh08]. Conquest [Cox04, Ros03, Tod09]. Consciousness [Bar04a, Ree05b, Bla05]. Consequences [Low01b]. Conservation [Ald06, Hey08, Mah02]. Conservatives [McG00c]. conserveurs [Bor06a]. Conserving [Rey05]. Considering [Sho08]. Conspicuous [Str02]. Conspiracy [Tho09a]. Constant [Scr00]. Constraints [Neu01]. Construct [Sch04a]. Constructing [Day06, Dow06, Lai03a, Luc04, Zac01]. Construction [Bek04a, Edw03a, Gag02, Nak07, Poh08, Scr05, Els09, Don07]. Constructions [Bek00, Kar09a]. Constructivism [LS00]. Consumer [Bix04, Blu04, Ens06, How09b, Jac05b, Jac00, Kul03, Lan03b, Pre04b, Rii03, Sca05, Smu05, Str01, Tho09a, Coo07a, Dir05, Tak01]. Consumerism [Bix02, Mah02]. Consumers [Bor06b, Bor09b, Bud05, Dir07, Spa01, Wel08, Fit02]. Consuming [De 03b, Jac02]. Consumption [Ben01a, Jaf05, Lan07a, Str02, Zac02, Pre04b]. Containers [Rol07]. Contemporary [Biv03, Coo02a, Cru07, Der04b, HV06, O’H05, Sch02e]. Content [Rob07]. Contentious [Elk06]. Context [Bud05, DO03, Kra07b, McM06, Pos09a]. Contexts [Dow07, Sto04a]. Contingency [Pos09a]. Continuity [Sve02]. Contraception [Neu01]. Contractor [dO01b]. Contractual [Dob08c]. Contributions [Osl07].

Control [Bea02, Cro08b, Day06, Fel07, Kai02, Lin06b, Owe03, Pfa04, Sla01, Sta06b, Con01a, McB07, Yat01]. Controlled [Bla00a]. Controlling [Ste01a]. Controls [Tur08]. Controversial [Sin06]. Controversy [Her01c, Sch09b, Sha07]. Convergence [Bar09a]. Convicts [Sto02]. Cookstove [Bra01a]. Cool [Mul04, Ber04]. Cooperative [Jør08]. Coordinating [Bar00c]. Copenhagen [AKP09]. Copper [Bur00a, Bur00a]. Copyrights [Hin08]. Corner [Bas09c, Pos01b]. Corning [Mat02, McE03, Spa01]. Corona [Spi00a]. Corporate [Bro06c, Kim04, Mor08b, Ste04, Bla09b, Lic02, Phi04]. Corporation [Har05, Las06, Mis02, Ree05a, Wel08, Mat02, War04]. Corporations [Mod06b]. Corridor [Col06]. Corsets [Sha06e]. Cost [Haf03, Joh02c, Pri00, Hur04]. Cottage [Sim03]. Cotton [Coc04, Gle03, Lak01, Moh08, Bro00b]. Countercultural [Cha07]. Counterculture [Bar08, Tur05]. Counterpoints [Pre03]. Countries [Bel08, Bra08b, Mok01]. Country [Fit02, Pla06, Uss09, Ell01, Mor06a, Web08]. Countryside [Her00, Kir00, Sch02d]. Course [McB00, McE04]. Court [O’H09]. Courtroom [Dow06]. Cowan [Par05]. Cradle [All06a]. Craft [Coo07a, HV07, Iri00, McE03, Moh04, Pre06, War01b, Bel08, HV01]. Crafting [Roc02]. Craftsmanship [Mil06]. Create [Sha06a]. Created [How00]. Creating [Sch02a, Sha06d, Con01a, Pet07b]. Creation [Bla01, Bla04b, Gor03, Gre04c, Jar00, Jon04c, Mar06a, Rob07, Rou05, Sch07b, Sil00, Slo04, Vol01]. Creative [Hig01, Ste05a, Swe08]. Creativity [Jac08a, San06]. Creator [Ens07]. Credit [Dea07, Lau08]. Crescent [Kol06]. Criminal [Sla02b]. Crises [Sol04]. Crisis [Bri07, Cas03, Gor04b, Jac08b, Mor00, Bar04b, Geo06, Vol04a]. Critical [Nag08, Sch08a, Tur09a, Bre01, Jos01]. Critics [Daw07b, Poh00]. Crop [Cro06b]. Cross [Pet07a]. Crossbow [Bac06b, Bac04]. Crossbows [Bac04, Bac06b]. Crossing [Kri00b, Pos03, Doe08b, Och02]. Crow [Ham05]. Crowded [Nie02]. Crucibles [BS09]. Crusader [Bur06]. Crystal [Mol00a]. CT [How00]. Cuba [Por03]. Cuban [Bar04b, Geo06, Wat00]. Cuisine [Spi04]. Culmann [Pet01a, Pet01a]. Culmann-Texten [Pet01a]. Cult [dlP07]. Cultivation [Biv07, Ham03]. Cultural [Bek00, Bij04, Bla01, Bow00b, Cas03, Daw07b, De 03a, Dem03, Din08, Doi06, Hey06, Iri06b, Koi02, Mor07a, Nad04a, Poh00, Pyn09, Rot04a, Rus09, Sat05, Ste01a, Uek07, Wit07, Zac06]. Culture [Aga09, Arn03, Bek04b, Bla05, Bor08, Bra07b, Bud02, Bug01, Car02, Cul09, De 06b, Dob08b, Epp03, Eri08b, Fra08, Gan03, HP00, How08a, How09b, Ink09, Jac05a, KO09, Kra07a, Kra04, Kul03, Lin06b, Lon02b, Luc05b, Lym03, Mag00, Mar09, May09, Maz04a, McC08a, McG00b, McS02, MB06, Mok09, Mor01, Mul07, O’G08, OS05, Pan06, Poo03, Rii03, Sca08, Sch06c, Seg09, Sid08, Sie09, Slo02a, Sto07b, Tho00, Tod02, Vas01, Vol04b, Wak04a, Wea02, Wel07, Pet07b, Ako06, Bal00, Bar03b, Blu04, Bra03c, Bro05b, Bro06c, Cra07, Dal06a, Dow03b, Fre01, Glo02, Goo04, Jer00, Mul04, Nye04, Nye05, Pfa06, Rus08, Sch00a, Tur09a, dlP07]. Cultures [Lan07a, Neu01, Pow09, Sch09a, Ser05, Bor02, Clo06b, Moñ07, Tri09, Yat01]. Culturing [Pic00]. Culver [Rot02]. Cunard [SS07]. Cuneiform [Dal05]. Cuomo [KO09]. Cup [Owe04]. Curiosity [vL09, Jan09]. Curious [Whi05]. Curitiba [Cep08]. Currency [Rob05a]. Current [Low00b, Low02b]. Curtin [Rob00]. Curtiss [Kin00]. Customs [Jaf05]. Cutting [Hor04b]. Cyber [Vin06]. Cyberculture [Abb04, Bar08]. Cybering [O’G03]. Cybernetics [Kli06, Pic00, Bow04, Owe03]. Cyberspace [Kid03, Bow05, Gre03, Wu03]. Cyberspeak [Bow04]. Cyborgs [Ala09]. Cycling [Nix09]. Czechoslovakia [Fre07].

D [Dal04, Hug07, Rom08]. D.C [Ano08d, Cer07]. D.C. [All04, Dou04, Sti03, Vog04]. Daguerreotype [Goo09]. dal [Lon02a]. Dam [Ell01, Sha07, Sin06, LB08]. Dammed [Ell01]. Dampfmaschine [Lac03]. Dams [Jos07, Sha07]. Dancing [Kat05]. Danger [Ree05b]. Dangerous [Car03, For07, DeV03]. Dani [Vas01]. Daniel [Cox04, Iri08a, Skr04, Smi09]. dans [Bla03, Fro01b, Lyn09]. d’archives [Gou00]. Dark [Dug09]. d’armes [Lan01b]. DARPA [Sei03]. d’Arts [Pan02a]. Data [O’G08, Yar09b]. Date [Clo07a]. David [Rol06a, See08]. Davidson [Pre04c]. Dawn [Epp09, Jeo09, Olw09, Poo03, Wel07, Wes01]. Day [Bra02, Cro06a, Sel09, Whi01]. Days [Bil02, Bra03b, Dor08]. DDR [Cie01, Sto05, Fre06, Zel06]. DDT [Sta05, Kin09b]. Dead [Bar03c, Hog03]. Deadly [Gor04a, Hul03]. Deak [Sil00]. Deal [Wan07, How09b]. Dearborn [Jon04b, Hyd01]. Death [Kol07, You01, Hog03, Luc07]. debat [De 00]. Debate [Bro06a, Fan01, Ham07]. Debates [HÖ08]. Debating [Zac02]. Deceit [Gor04a]. December [Isr02]. Decision [Rol01]. Decline [Kno03, Lon07, Sch01b, Kin00]. Deconstructing [Lym03]. Decorative [Ben01b]. Deep [Hin07a, Jer07, Mid06, Sad06, Tom03, Dur06, Jos07, Wer07b]. Defeat [Nis07, Nei07]. Defeated [Tom03]. Defence [Smi03]. Defense [Bar03c, Fri02b, Sch04c, Van00b, Boy03]. Defibrillator [Law03]. Defining [Bro06a, War06b]. Deforestation [Her00]. d’égouts [Poi08]. Degradation [Moo02b]. Deindustrialization [Bla09b, Pay05]. del [Gag07, PP09]. d’Electricite [Fox00]. délétère [Fro01b]. Delight [Sma05]. Dell [Dow05]. dell’ingegneria [Vas08]. dell’invenzione [Lap06]. Delta [Dis06, Ell05, Lan02b]. Delusions [Sti08]. Demise [Wer03]. Democracy [BB00, Clo07b, Lic02, O’G03, Coo04b, Seg09, Smu07]. Democratic [Kru02, Lon02b]. democratie [Lev01]. Democratization [Hin08]. Democratizing [Nag08]. demonstration [War03]. Demythologizing [Lar04]. Denial [Gor04a]. Denkmale [Lac03]. Denmark [AKP09]. d’enseigner [War03]. Dentiste [Ben01d]. Dependency [Bur00b]. Depicting [Kon05]. Depression [Rus08]. Deregulating [Wil01]. Deregulation [Pra01]. d’Ernest [Lau02]. Desegregating [dlP06]. Desert [Bug04, Jon04c]. Design [Ben05a, Ber02, Bro00c, Col07a, Div09, Don07, Kāt06, Kat09, Kle04, McG00b, Mei05, Mor07b, Nic02, Pet04b, Spa01, Spa02, Fri07a, Iri03, Kat00, Kāt07, Ste05a, Zac06]. Designing [Wos06, McC01a, Mel06, Ryd03]. Designs [Blu04, Kul03, Les08a]. Desire [McS00, Mom05]. Desires [Mel01]. Desolation [Wos02b]. Destined [Gui00]. Destiny [Biv07, Dor07, Epp04]. Destruction [For08, Jos01, All05, Swe08]. Detectors [Col08]. Determined [Wak07]. Determinism [Cer05]. Detroit [Ano00h, Esb05, Jar00, Fis07, Hyd01]. Deutsche [Wal04, Zac03]. deutschen [All06b, Mar03, Neu02, Pet00]. Deutsches [FF00]. Deutschland [Joh02a, Mar03, Mom02, MW07, Sch05b, Wen02]. Devastation [Col04, Her05]. Development [Awt07, Bar03c, Bry07, Bug04, Con02, Das08, Dow05, Gui00, Hig01, Hin08, HÖ08, Jos08, Kel04, LW05, Lan00, Lon00, McG05, Mei00a, Och01, Pre05, Reu02, Sat08, Sha09b, Sta05, Sto09b, Tho05, Vol03b, Lak04]. Devices [Bak05, Les08a, Mel01]. Dialogues [Bra04]. Diary [dO01b]. Dibner [Rol06a]. Dickens [Pop01b]. Dictatorship [Sid08]. Dictionary [MN06, Ben01b]. Did [Coo06, Ree05b]. Diesel [Gra01, Kle08]. Differences [Fit00]. Diffusion [Ald06, Ang05b]. Dig [Wil06]. Digital [Ake05, Bar08, Bar09a, Bow05, Cer04, Eri08b, Gui03, Iri06c, Low08a, Mul05, O’G04b, Pic00, Sch07a, Tur06, Epp06, Her09a, O’G05]. Digitizing [Bar05]. Dilemma [Cas07, Skr07, Sto06b]. Dilemmas [Elk02, Kru02, Mor03, Skr00]. Dimension [Rob06a, Cro02b]. Dinge [Bak04]. Diplay [Wea02]. Diplomacy [Hug05]. Directorate [Joh02b]. Directory [Law00]. Direktor [Gis07]. Dirty [Elk06]. Disability [Kee03]. Disadvantage [Les01]. Disarmament [Olw04]. Disaster [Kel06b, Luc03, Sha06d, Wer07a, Yar08]. Disasters [Coo01b, Sha09a]. Discerning [Clo07b]. Disciples [Wei06]. Discipline [Ale06, Low05a]. Disciplined [Boc09]. Discontents [O’H04]. Discourse [Pav09]. Discourses [Lew00]. Discovery [Bla04a, Cus05, Jer07, McC09b]. Discussing [Low04b]. Discussion [Bri07]. Disease [Clo06a, Hum09]. Dish [Sad06]. Disney [Mol08, Wil09]. Dispersal [KM04]. Displaced [Jos07]. Display [Pop01a]. Disrupting [Hac09]. Dissected [Ver08]. Dissemination [Awt07, HPV06]. Distance [Har03]. Distant [Sel09, Bek04a]. Distracted [Dug09]. Distributed [Dow03a]. District [Pos06, Rob02]. Disziplin [Fox02]. Dive [Gui00]. Divergence [Maz03]. Divided [Ric08b, San06]. Divination [Dal06a]. divulgazione [Gag03]. DNA [AA01, Cow04b]. Do [Jag00]. Doctor [How09a]. Documents [Bug07]. Dodge [War06a]. Doers [War01a]. Does [Cla02c, Epp02b, Wil03b]. Dog [Coo06, Jos08]. Doing [Lar06]. Dolls [Mar00b]. Dolly [Lan08]. Domain [Epp07]. Dome [Sch08b]. Domestic [Bra01a, Pan02b, Pav09, Wal01b, Ber01]. Domesticity [Dav06]. Dominance [Pal00, Mor07b]. Donald [Mar00c, Sti07]. Door [OS05]. Doors [Ben05a]. Dorado [Dor01]. Dose [Fra02]. Double [Orr02]. Douglas [Bar02, Mor06a, Cro05]. Dover [Ash00]. Down [And09, Bar09b, Fra06, Mol04, RP09, Vog04]. Dowsing [Dym08]. Dr. [Cro06c]. drahtlosen [Pet01c]. Drainage [Dob08c]. Drais [Mos06]. Dramatize [Kri00a]. drappo [Lon02a]. Draw [Mol04]. Drawing [Ber07]. Drawings [Bro00c]. Drawn [Bro05a]. Dreadnought [McB07]. Dream [Bij03, Con07, Daa07, Gou04a, McC01b, Nei05, Hea02, Lai03b]. Dreamer [Fos09]. Dreamers [War01a]. Dreams [Alc07, Bel01, Bro00b, Daw06, Nad08, Pfa06, Bye04, Cul09, Mou08, Och01, Seg08]. Drift [And06b]. Drilling [Hin07a]. Drive [Bar04a]. Drive-By [Bar04a]. Drivin [Nie07a]. Driving [Cer03, Fra08]. Dropping [Sch01b]. Drug [Bud05, Bud06, dlP07, Bud07]. Drugs [Sli09]. Dry [Goo09]. Dry-Plate [Goo09]. DSM’s [vR07]. DuBois [Bro05b]. Dukes [Bac06a]. Duluth [Lan09]. Dung [RV04]. DuPont [Rob08a]. Duracell [Hin09]. durch [Sto05]. Durchsicht [Bey07]. Durée [Mol05, GC07]. During [Bar03c, TeB02, Bac04, Bac06b, Bel04a, Bra08b, Bro00a, Coo06, Fre07, Jon00, Kid08, Kry04, Mor03, Slo08, Str09b]. Dust [O’H07, Sti07]. Dutch [Bij02, DvdE03, Kai02, Reu02, Rob04]. Dwelling [Par02a]. Dyck [Bey06]. Dyestuffs [HP00, Tra02]. Dynamics [Din08, Fie01, Hei04, Koi02].

E.H [Col01]. Eadweard [Bro05c]. Eagle [Rob00]. Early [Bar03c, Bek04b, Bel08, BM07, Ber02, Bor06b, Bra08c, Cas08, Con00a, Coo00, De 06b, Dir07, Fri08, Goo07, Hal01, Hun02, Jaf08, Jon04b, Kle03b, Lak05, Lan00, Las06, Lin06a, Lon02b, Moh04, Nak07, Osl04, Per07, Ree05b, Rey02a, Ryd01, Sew06, Sid07, Sta05, Ste07b, Tim09, Tym00, dlP06, Bel04a, Hec08, HPV06, Sto09a, Buc05b]. Early-Nineteenth-Century [Tym00]. Earth [Bar08, Chu07, Hoh03, Hor02, Ser02a, Tra05, Cio06]. Earthmoving [See01]. Earthquake [Wer07a, Wit07]. Earthquake-Resistant [Wer07a]. Earthquakes [Wil03a]. East [Sto09b, Vol01, Zac01, Zac02]. Eastern [Sim06]. Eastman [Oli07]. Easy [Ben00, Kra07a]. Eating [Win05]. Echoes [Mol07]. Eckert [Lig06]. Eckert-Mauchly [Lig06]. Eco [Dis05]. Eco-Biography [Dis05]. Ecole [Pan05a]. Ecological [Hub09, Moo02b, vD02]. Ecologies [Bow00c]. Ecology [Abb05, Ell05, Sto03a, Wak05]. Economic [Baa01, Ber01, Buc03a, Car08, Elk02, Hin08, Igl01, Jac07, Kha08, Lak04, Mod06a, Per02, Rav01, See07]. Economics [Bix02, Mok05, Pop07, Dal06b]. Economy [Big02, Bre00, Cha01, Epp06, SHW03, Ste08b, Tur05, Ung05, Wan07, Kie03, Lan04a, Maz03, Sha04, Tur06]. ed. [Mil04c]. Eddie [McE07]. Eden [Nye04, Har07b]. Edens [Lig09]. Edge [Bar00a, Luc03, McC01b]. Edison [Cow04a, Pre01, Pre02, Pre05, Pre08a, Pre08b, Rot01, Sch07b, Pre04d, Kli00, Mar05b, Wos04]. Editor [Par02b, Sta02]. Edoardo [Pet05]. eds [Cer09]. Educación [PP09]. Education [BDD09, Pet04d, Rud09, Slo08, VCM^{+}09, Wis09b, Cru03]. Edward [Coo02c, Bac06b, Las03, Scr00, Kle03a]. Edwardian [Qui03a]. Edwin [All02, Kli08]. eeuw [Wac05]. Effect [Gom05]. Effects [Bas05, Hal01]. Efficiency [Ale06, McG00c, Pra03, Bix04]. Efficient [Ale06]. Effluent [Luc02]. efterkrigstidens [Jør09]. Egypt [Bis05]. Eighteenth [Far00, How08b, Por03, Ran01, Rob07, Wak02, Wes08, Rey02b]. Eighteenth-Century [How08b, Rob07, Wak02, Wes08, Rey02b]. Einsicht [Bey07]. Einstein [Bey09, Ste05b]. Eisenhower [Day04a, Ham07, Wan04]. Eisenstein [How08a]. Electric [Hug02, Las03, Mai04, Mar05b, McS00, MK01, Pra01, Pre01, Sch04b, Sch05a, Sch08b, Tho01, Vol05c, Mid05a, Mor08b, Pfa06]. Electrical [Bri01, Mol04, Tak01, Tod08, Wil00a]. Electricity [Tym00, Wil01, Ser02a]. Electrification [Jon09]. Electrify [Cow04a]. Electrifying [Pre08a]. Electromagnet [Sch08b]. Electronic [Cor02a, Ens06, Ens07, Flo00, Bij05]. Electronics [Ens06, Vol03b, Neb01, Tom05]. Electropolis [Cav08a]. Elefanten [Gal00]. Elegant [Mil03]. Elektrifizierung [Osi00]. Elektrizitat [Osi00]. Elektrizitatswirtschaft [Sto01b]. Elevators [Vog04, vDS05]. Elite [Rav01]. Elites [Bas09b, Pri02, Dol09]. Elizabeth [How08a]. Elizabethan [Ash06, Dea08, War02]. Eloge [Kah01]. Else [Shu09b]. Elusive [Gou04a]. Emancipation [Lak01]. Embedded [Flo03]. Embodiment [Jac08a]. Emerald [Jud08]. Emergence [Ade08, Arn00, Joh05, Kli06, McB06, Swa09]. Emerging [Dow03b, Jac04a]. Eminence [Epp07]. Eminent [Nel04]. Emotion [Mal09a]. Empire [Ale03, Bek04a, Day04a, Dic05, Gre04a, Har07b, Ink06, Mou08, Mul09, Pla06, Ste07b, Ven00, Yal08, Hof02, Ayd05, Bek04c, Bek06, De 06a, Slo09, vD03]. Empires [Mil04b, Ste05b, Cow04a]. Empirical [Hil02, Kou03]. Empiricism [BM07]. Employed [Edw03a]. Empowering [Wil00a]. Empowerment [Wan08]. Empty [For05]. Enchantments [Nad04a, Nad06]. Enclosure [Ste04]. Encounter [Cro06b, Sil00, Win08]. Encyclopedia [Sch01c, Vol09b, Cli00, Kou07, Ott08, Sch06f, Wal02]. End [Bas09a, Fre07, Her01a, Mor00, Sch08a]. Endeavor [Tri03]. Endless [Flo05, Lip00a]. Enduring [Cla07]. Enemy [Ang04, Ang05a, Nis08]. Energie [Die01]. Energy [Bow00b, DeV03, Kra09, Lai03a, Low04c, Luc02, Nie02, Pri03, Qui03b, Ser02b, Shu09b, Sla02a, The08, Kie03, Kol08, Pas06, Shu09b, Sti09, Tes09]. Engelbart [Bar02]. Engine [Bow07, Coo05a, Kid06, Rob04]. Engineer [Bri01, Bro09a, Fos09, Fre01, Kno01a, Reu08, Wis09b, Nel04, Pet07a]. Engineering [Alt00, Ang01, Big08, Bro00c, BDD09, Cas03, Dav01, DvdE03, Gei00, Gie02, Har01, Kle04, Kli08, KO08, Lev00, LB08, Lig06, Mah02, Mil02, Muk03, NV07, Owe04, Rey05, Sat05, Sha00b, Ste05a, TeB02, Ten05b, VCM^{+}09, Wer07a, Yea04, Bri01, Flo05, Goo06, Gre02, Iri05, McG00b, McM06, Pri00, Vol07a]. Engineers [Ald00, Bla00a, BDD09, Buc01a, Flo07, Gor01c, Kid08, Muk06, Och01, Och02, VCM^{+}09, Wes07, Bar03a, Bor00, Kli08]. Engineman [Gam02]. Engines [Fre01, Gre04b, Seg09, Con08, May09, Per02]. England [Bro07, Bur02, Cor06, Dem03, GC07, Jaf05, Sch05c, Sha09b, Ung05, Zai08, Bac04, Bac06b, Jør08, LW05, Arn01, Ash06, Chu04a, Jaf08, Lon02b, Mil06, Per02, vD09]. English [Big02, Hal01, Mac02, Ran01]. Enhancement [Bes08]. ENIAC [End00]. Enlightened [Maz04b]. Enlightenment [Ale03, For09, How08b, Rey05, Ver08, Ald00, Jan09, Maz04a, Mol04, Wal06b]. Enola [Pos04]. enquêtes [Pan03b]. Enriching [Hor02]. Enrico [Lap06]. Enrollment [Wei04]. enseñanza [PP03]. Enterprise [Jon09, Lan07b, Mer05, Per02, Phi04, Pri02, Ryd01, Sal05, Sha09b, Spa02, War06b, War09c].

Enterprises [Dow05, Pan09]. Entertainment [HV06]. Enthusiasts [Pos07]. Entomologie [Har08b]. entre [Bel04b, Haf09]. Entrepreneur [And06a]. Entrepreneurs [Con00a, Flo00, Gro08, Mar09]. Entrepreneurship [Mac02, Per07, Slo04]. Entwicklung [Lac03, Wen07]. Envelope [WF03]. Environment [Bla00a, Dob04a, Doe08c, Kle06a, Ols07, Par02a, Sol03, Vol03a, Ell05, Bor09b, Don06, Jos07, Luc02, Mar06b, Mil08, Min03, Sti09]. Environmental [Aik07, All07b, All07a, Ang05a, Bri06, Cio06, Col04, Elk06, Gor04b, Guc08, Hal07, Jac04a, Jos01, Jud05, Jud08, Kar08, Kli04, Kol08, Kri03, Lan07a, MA09, Osl04, Pla06, Rey02b, Sol04, Uek07]. Environmentalism [Sch02d, Sin06]. Environmentalists [Gor01c]. Environs [Gor02b]. Envy [Jac05b]. Epic [Ens06, Jos07, Poh08]. Epoch [Kie01]. Equally [SHW03]. Equipment [See01]. Equipped [Bix02]. Era [Eel09, Whi01, Bas09b, HPV06, Kry04, McC09b, Pre02]. Erfinderbiographie [Gis07]. Erie [Ste01a]. Erosion [Dug09, Rau03]. Error [Epp00, Wak03]. Errors [Wri06]. Erudition [BM07]. Escalators [Vog04]. Escape [Wer03]. Escuela [PP09]. espace [Fro01b]. Especial [PP09]. Espionage [Far00, Bra01b, Day04a]. espionnage [Gil06]. Essay [dO01a]. Essays [Clo03, Con00a, Con07, Cut02, KO02, McC01b, Par02a, Bug07, Pet05]. Essential [Kra05b]. Establishing [Las06]. Establishment [Wal01c]. Estey [Rii03]. eta [Squ00]. Ether [Bra01b, Doe08b]. Ethernet [Abb03]. Ethical [Rau03, Sto09b, Pow04]. Ethics [Pri00, Web08, Bud02, Kou07]. Ethiopians [dlP06]. Ethnicity [Zem02]. ethnographie [Dav09]. Ethnography [Kle04]. Eugene [Hou04b]. Europa [Die01, Dol09]. Europe [Big01, Hår08a, HP00, Joh00b, Maz03, Phi03, Poo03, Sto01a, Tim09, Wen01, Luc05a, Maz04b, Pre00, Sch06b, BM07, Ham08a, Kno06b, Kon05, Mae09, Nor09, Rey02a, Rob07, Tra02, Vee05]. europea [Squ00]. European [Ash05, Bek05, Bel06, Jac00, Kra01, Mac05a, Mar00a, Mir08, Nak07, Kra07b, Kri00b, Mom05, Gor06]. Eutrophication [Ste01a]. Eve [Wos04]. Eventually [Elk06]. Ever [Rob08b, Rol06b]. Every [Soc08, Dan03]. Everyday [Aga09, Ben01a, Kat00, Mor01, Zac06, Jer00, Kle04, Rot04a]. Everything [Shu09b, Tri06a]. Evidence [Luc05a, Cor02b]. Evolution [Bru01, Gui00, Mor05a, Tom05, Vol09a, WW00, Kle08, Sch09a]. Evolutionary [Bas02, Woo05]. Excavations [Gof09]. Excellence [Mui03]. Excise [Jaf05]. Excreta [Luc04]. Excursionists [Men09]. Exercises [Wol08]. Exhibiting [Skr07, Skr00]. Exhibition [Bal00, Pop01a, Ryd03, Tym00]. Exhibitions [Jon00, Wea02]. Exhibits [Wos02b]. Exit [Wer03]. Exorcising [Ham06a]. Expand [AST08]. Expanding [WF03]. Expansion [Wel02]. Expectations [Vol05c]. Expedition [Bek02]. Expensive [Elk06]. Experience [Fra09, Gre05, Lin05, Mac00, McB00, Vol09a, McS00]. Experiences [Nei02]. Experiencing [Ste07b]. Experiment [Cav08b, Gro09, SW07, Tym00]. Experimentation [Low02a]. Experiments [Eze02, Yea04, dlP06]. Expert [Kri03, Sch05c]. expertis [Jør09]. Expertise [Ash00, Ash06, Cor08, Par09].

Experts [Bis05, GS02, Ham06a, Ste01b]. Exploration [Dor01, Jer07, Osl07, And00, Gli00, Pyn09]. Exploring [Cer03, Joh04, Lan04a]. Explosive [Buc05a]. Explosives [Bac08, DeV03]. Expo [Nad07]. Exposed [Rob06b]. Exposing [Neb01]. Express [Hof02]. Extension [Slo08]. exterieurs [Bel04b]. Extravagance [SHW03]. Eye [Ben08, Bro06c, Lon04a, Mei00b, War09b, Spi00a]. Eyes [Hum09]. Eyewitness [Rob06b]. Eyewitnessing [Cor02b]. Ezekiel [Mor00].

F [All04, Dow05]. F.B [Isr05]. Fabbriche [Vas08]. Fabricating [Ben02]. Face [Eri08a, Lon07, Zai08]. Faced [Bar04b]. Faces [Sho02]. Fachleute [Zac03]. Fact [Gol00c, Yea04]. Factories [Por03]. Factory [Dan03, Jos05, Lew01, Yar09a]. Facts [Sch01c]. Faculty [Gie02, War06b]. Fail [Nei07]. Failed [Epp00, Wis09b]. Failure [Big08, Fri07a, Har05, Lip03, McC01a, McI00, Pet01b, Bea02]. Fair [Bro02, MS08]. Fairs [Bro02, Wea02, Sta00]. Faith [Cas05, Kol07]. Fakes [Orv09]. Fall [Ens07, Joh06, Kie03, Lic01, Sch01b, Sha06e]. Falling [Mai09]. Fallout [Glo02, Sil00]. False [Sau05]. Falun [For09]. Family [Bij03, Hul09, Kno05a, Mac04, Oli07, Rab04, Ryd01, You09]. Famines [Jon02b]. Farben [Str09b]. Farbfernsehfrage [Sto05]. Farewell [Kha08]. Farm [And09, Dan03, Bur02, Jos05]. Farmer [And06a]. Farmers [And05]. Farming [Moo02a]. Fashion [Cla02a]. Fast [Bel01, Wei03, Zei03, Moy06]. Fastenings [Thi06]. Faster [Joh02c]. Fatalities [Ten00]. Fate [Nye04, Sch04b, Sid08, Spi04, Spi00b]. Fates [Moo00]. Fathoming [Jer07]. Faux [Orv09]. Fazlur [Iri05]. FDR [Wil00a]. Fears [Ros05b]. Features [Sla01]. February [Sch07a]. Federal [Igl01]. Feedback [Min00, Owe03]. Feenberg [Nag08]. Feet [Jac08a]. Female [Sha09b, Moh08]. Feminism [Kat00, Shu02]. Feminisms [Shu09a]. Fer [Bla03]. Ferguson [Hou04b]. Fermentations [Sch02b]. Fernlenkwaffentechnik [Neu02]. Fernsehempfanger [Pet00]. ferreux [Pan00]. Fertility [How09a]. Fertilizer [vR07]. Fessenden [Fro01a]. Fetal [Cow03]. Fiber [Fin02]. Fiction [All02, Bas07a, Dow08b, Fra08, Nad08, Hen09, O’H05]. Fictions [Sla04]. Field [Rey07, Wal06a]. Fifteenth [Lon09]. Fifteenth-Century [Lon09]. Fiftieth [Bij09, Pos08a]. Fifty [Sta09]. Fighting [NE09, Ser02c, Epp09]. Figurations [Pre04a]. fil [Mul01a]. File [Sch01c]. Film [Bro09b, Gom00, Mag01, Rom08, Jos03, Lon01b]. Filmmaking [Mul07]. Final [Jos04]. Finance [Jon09]. Financial [Bow07, Lau08, Nis08]. Financing [Mie09]. Finding [Kon09]. Findlay [Bro00b]. Fineness [Gag02]. Fingerprinting [Col05]. Fingerprints [Sla02b]. Finland [Sch06b]. Fir [Mor06a]. Fire [Bro00b, Hac03, Hal01, Hea02, Her05, Kno07b, McB07, Men09, Mol00a, Per06b, Rol00a, Rol05, Sch01b, Wer03, Win05, Pow09]. Firearms [Hac05b, Smi05a]. Fireplace [Bra01a]. Fireproof [Iri01]. Fires [Wak04b]. Fireside [Slo02a]. Firms [Bre00, Sto06a, Gle03]. First [Ang06, Bek05, Bil02, Cla01, Clo05, Coa02, Cow03, End00, Gor07, Hof02, Lev08, McG00a, Mis02, O’H01a, Pav09, Sch02e, Sid06b, Spi01, Sti09, Tri04, dO01b, Con03, Neu06a]. Fischer [Cro05]. Fish [Jos08, Sti04b]. Fisheries [Sve02, Hub09]. Fishery [Sve03]. Fitch [Mil08]. Fits [McC09c]. Five [Cut09b, Dor08]. Five-Million-Pound [Cut09b]. Fix [Sha06a]. Fixing [Kra04]. Flag [Sch06d]. Fleeming [Bri01]. Fleet [McB05, Tap02]. Flesh [Jen05]. fleuve [Pri04]. Flexibility [Big00]. Flight [Bye04, Con07, De 04, Goo04, Gui03, Jon04b, Sat05, Tri04, Tri06b, War05a, WF03, Alc07, Bil02, Her07b, Her09a, Mei00a, Tom05]. Flipping [Coo05b]. Floating [vDS05]. Flooding [Abb05]. Floor [Hoh03]. Florida [Sta06b, War09a]. Flourished [Hur04]. Flower [War08]. Flowers [War08]. Flowing [Hof00]. Fluid [Bla08]. Flushing [Dob08b]. Flux [Her08b]. Fly [Con03, Gui03]. Fly-By-Wire [Gui03]. Flying [Bed05b, Nei05, Fra06, Gel04, Van00a]. Fold [Orr02]. Folgen [Zel04]. Follow [Sch06d]. Folly [Cas05, Iri03]. fonds [Gou00]. Food [App03, Bel01, Fri02a, Kle09, KO07, Moo02a, Sch02b, Sha05b, Spi04, Van02, App03, Hor02, Lan03b]. Fool [Mei00b]. Foragers [Dob04a]. Forbidden [Kry04]. Force [Bug04, Jud03, Low04a]. Forces [Nei07, Pow02, Sid02]. Ford [Fer05, Jon04b, Jon04a, McE06, Mom02, Pie06]. Fordism [McI00]. Forecast [War07a]. Foreign [Ang04, Bri08, Vie06]. Foresaw [Mul06]. Foresights [Bak00]. Forest [Ald06, Bar04a, Qui03b]. Forge [Wal06a, Le 05]. Forging [Flo00, Lan09, Wer05a]. Forgotten [Mag05b]. Forlanini [Lap06]. Form [Epp04, Flo07, Kie02, Rob07]. formation [Gui06, Sto07b]. Formative [Wha06]. formazione [Gag03]. Forms [Pet09]. Forschen [Zac03]. Forschung [Sto07a]. Fortress [HV07]. Forty [Mar01]. Fossil [Pri03]. Fouls [Kru02]. Foundation [Rav01, Sta05]. Foundations [Har01, McC07a]. Founded [Goo06]. Founder [Hug09a]. Founding [Sti04b]. Fragile [HV01, vL09]. Frame [MA09]. Framing [Sla02b, Epp03]. francaise [Gou04c, Pan03b]. France [Ben01a, Ben01c, Ben02, Bud09, Far00, Fox00, Gou00, Lyn09, Pan03a, Pic01, Wak02, Wak05, Gui08, Hei04, Muk06, De 03a, Pan03b, WH09]. Francis [McC07a, McC09a, dSL04]. Francisco [Wer07a]. Franco [Vee03]. Franco-German [Vee03]. Frank [Buc01a]. Frankenstein [Tym00]. Frankfurt [Cep08]. Franklin [Gor05, Lev08, Mol04, Sch04b, Ste07a]. Frankreich [Fox02]. Franz [Kel09]. Frauen [Fre06]. Fraunhofer [War01b]. Fred [Low05a]. Frederick [Bix04]. Free [Rus09]. Freedom [Seg08]. Freer [Har03]. Freiberg [Dym08]. French [Dic05, Gou04b, Pan07, Pri02, Rey05, Sla09, Sto02, Zei03, Rey02b, AA01, Ald00]. Frenzied [Coo03b]. Frequencies [Kry04]. frères [Pan09]. Friends [Lon04a, Sch07e]. Fritz [Cut09b, Hor02, Joh06, Str05]. Froese [Cro05]. Front [Nei02]. Frontier [Bug05, De 06a, Jos04, Osl04, Sim06, Ang06, Flo05]. Frontiers [And00, Sti04a]. Frontinu [KO02, KO02]. FRS [Fin03]. fruhen [Low00a, Pet01c]. Fruits [Ake01, Har07b, Kir00, Fro09]. FSA [Pan06]. Fuel [Wor01, Ell05]. Fuels [Pri03]. Fulfilled [Coo04b]. Fulton [Hea02]. Fund [Mal09a]. Fundable [Van00b]. fundamentacao [dSL04]. Funding [Tei01]. Funken [Pet01c]. Fusion [Nad04b]. Future [Bak05, Bar00a, Day04b, Dow09, Lai03a, McM06, Mul06, Rus07, Wel02, Wil03b, Zim04, Kat02, McC08b, vL09].

G [Bac02, Dow05, For09, Kan07]. Gahn [For09]. Gaining [Tya04, Wer04]. Galactic [Hen09]. Galileo [Lon04a, Dea07]. Gallery [Bro09c]. Game [Sch07a]. Games [Lig08]. Gaming [Sch07a]. Ganz [Heß04]. Garde [Lar08, Gan05]. Garden [Bed05b, Don06, Kou01b, Lip00b, Lon03b, O’H05, Pos06, Mei03]. Gas [Sch08b, Wor01]. Gas-Lighter [Sch08b]. Gatekeepers [Wit08]. Gauge [Men09, Sch05e]. Gay [Pos04]. Gegna [Bur06]. geheime [All06b]. Geigerzahler [Hes04]. Geistesgeschichte [Osi00]. geistigen [Has02]. Gender [BB00, Bij03, Bul03b, Coo02b, Gor02a, Mai07, Mal01, Mel01, New00, Sin00, Zem02, Rot04c]. Gendered [Bix02]. Genealogy [Gou04a, Lan08]. Geneaology [Sch08a]. General [Bye04, Dic04, Ben03]. générale [Pan03b]. Generations [Mat02, Pye01]. Genesis [Mac09, Tur09b]. Genetic [AA01, Gar09]. Genghis [Sta06a]. Genius [Ens07, Fro05, Hea02, Sch07c, Bey09, Jon02a, Lev08]. Genocide [Mie04]. Genomes [Bow00c]. Gentle [Kar08]. Gentlemen [Buc01a]. Genuine [Orv09]. Geoffrey [Mis08]. Geographic [Jac08a]. Geographical [Gag02, Yar07]. Geographies [Dob08a]. Geography [Boc09, Dor01, Kar02]. Geometry [Her01b, San02]. Geopolitics [Joh00a]. Georeferencing [Yar07]. Georg [Gis07]. George [Tap02, Oli07, See07]. Georgia [Bro00b, Fal06a, Col07b]. German [Bek07, Bey08, Cav08a, Nei02, Str05, Vee03, Zac06, Zel01, Zel02, Mer05]. Germany [Bek04a, Cas07, Mor03, Pal02, Ric08b, Sch06c, Tod02, Zac01, Zet08, Ale06, Eck05, Wer06, Zac02, ZH00, Bra07b, Bud06, Hey08, Tod09, Uek07, War07b]. Germs [Moo00]. gescheiterten [Gis01]. Geschichte [All06b, Ber06, Bey07, Gis01, Hes04, Lac03, Low00a, Pet01c, Sto01b, Vee00, Mom04, Orl02, Pet04a]. Geschiedenis [Mok01]. Geschlecht [Fre06]. Get [For07, Pav09, Les04a]. Getting [Jac00]. Ghosts [Hac06, Ham06a, Mei05, Nad08]. Giant [Bla04a, Tat05]. Giants [Mil04b]. Giedion [Mol02]. Gifts [Wil03b]. Gilded [Gre08a, Poh08]. Gimpel [Bac05]. Gin [Coc04]. Ginger [Coo04c]. Girlhood [Mar00b]. GIS [Yar09b]. Giving [Vog02]. Giza [Edw03a]. ’Glades [War09a]. Glaha [Ell01]. Glare [Sid03]. Glass [Les08b, Mei00b, War05b, War07a, Ste08b, Whi04]. Glassforming [McC03]. Glassmaking [HV01]. Glen [Ell01]. Global [Cla07, Cus05, Doe08c, Dow08a, Gre08c, Jac08b, Mat02, McC09c, Mer05, Pur08, Sha04, Smi05a, Ste04, Ste01c, Sto09a, Ver00, Wal01a, War09a, Joh00a, Jon09, Pfa04]. Globalization [Bij09, Kra07b, Pfa04, Slo02b, Sol04, Yal08]. Globe [Bul03b, Clo07a]. Globes [SS02, SS02]. gloire [Lan04b]. Glory [Gui00]. Go [Str02, Tho08]. God [Smi09]. Gods [Gel04]. Gold [Car08, Gag02]. Golden [You09]. Goldsmith [Fis03]. Gono [Rav01]. Good [Ash01, Coo01b, Kou01a, Mah02, Par02b, Sha01b, War08]. Goods [Ber01, Tak01, Mil09]. Google [Low08b]. Got [Fei09]. Gotham [Jac05a, Poh08]. Gotthard [Els09]. Governance [Ash01, Wu03, Kri03]. Government [Bud05, Coo05c, Kle03b, Low01a, Cro08a]. Governments [Mos02]. GPS [Zim04]. Grace [Cro06b, Vin06]. Grafton [Cav08b]. Grain [Cha02, vDS05]. Gramophone [Mag01]. Grand [Mar05a, Mil02]. grandeur [Zet08]. Grant [McM06]. Granville [Pre04c]. Graphic [Kat09]. Graphics [McG00b, Ham09]. graphische [Pet01a]. Graveyard [Dal02]. Gravity [Bas05]. Gray [Fin09]. Great [Gre04b, Hoa01, Kol07, Lak01, Mal01, Maz03, Mol00b, Mor05b, O’H07, Pop01a, Rey01a, Ric08a, Rus08, Sho05, Sid06a, Tho07, Vee03, Vol08b, Wel02, Yar08, Bro00c, Edw03a, Nei02]. Greatest [Gam07]. Greatness [Kin00]. Greece [Lan02a, KO08]. Greek [Rol00a, KO09, Rol05]. Green [Hey08, Ols07, Pos09b, Sto03a, Wak05]. Greenfield [Hyd01]. Greenwich [SS02]. Greenwood [Cut07]. Grooves [Pre02]. Grossprojekte [Gal00]. Ground [Flo03, Wer04]. Groundwater [Jac04a]. Group [Ake01]. Growth [And05, Ell05, Kno03, Mod06a]. Growth-Regulator [And05]. Guiana [Sto02]. Guide [Hov04, Mac01, Rey07, McS02]. Guided [Das08]. Guilds [Bel04a, Bel08]. Guitar [McS00]. Gulf [Kol08]. Gunnery [McB07]. Gunpowder [DeV03, Gor03, Bac08, Buc05a, Hac05b, Rol00a]. Guns [Sto04b, Bek06, Moo00, Phi03]. Gusher [Feh04]. Gutenberg [How06, Rei05]. Gutters [Jør08].

H [Pre04c]. Habana [Por05]. Haber [Hor02, Joh06, Str05]. Hacker [Bar03b]. Hagley [Asn03, Chu04b, Kul03]. Half [Fis07, Sil00, Sch07a]. Half-Century [Sil00]. Hallmark [Goo09]. Ham [Har03, Cra07]. Hamburgers [Rus07]. Hand [Ake05, Cer06, Fri08]. Hand-Loom [Cer06]. Handbook [Cut09a, KO07, KO09, Coo09]. Handloom [Gle08]. Hands [Gor06]. Happen [Mil04a]. Happy [Bar03a]. Harbor [Ash00, Elk06, McB05, Nis08]. Hard [Low00a, Mei05, O’H07]. Hardware [Seg09]. Hardy [Pet07a]. Harmonious [War07b]. Harness [McG05]. Harold [Mul09, dO05]. Harriman [Col01]. Hartree [Cro05]. Harvest [Lan01a, Rou02]. Hate [Tho09b]. Hath [Smi09]. Haushaltstechnisierung [Ber03a]. Haven [Sha06e]. Hazard [Wil03a]. Hazardous [Abb05]. Hazards [For04, Hei04]. Hazy [All04]. Health [All07a, And06b, Biv07, Bro09b, Gar09, Gol00a, Lan07b, Law03, Sat04, App03, Sli01]. Healthis [WW04]. Heard [Gom00]. Hearing [Wer07b]. Heartland [Gol00b]. Hearts [Law03]. Heat [Kra09]. Heaven [Arn00, Mou08]. Heavenly [Dal06a]. Heavy [Bro00b, Moy06]. Hedgehog [Tom04]. Hegemony [Ham08a]. Heidegger [Dob04b]. Heirs [Wal06b]. Helicopters [Ter00]. Hell [Bor08]. Hells [Sha07]. Help [How00]. Hennebique [Wer02]. Henri [Le 07]. Henry [Fer05, Isr02, Jon04a, McE06, Nie07a, Pet03, Pie06, Bac04, Jon04b, Pos05]. Her [Ash09, KO07, Uss09]. Herbert [Gif07, Tho06]. Herbicides [And05]. Herders [Rus07]. Herman [Kru08]. Hermann [Gis01]. Hero [Coo03a, Hec08, McE07]. Heroes [Jon04b]. Herrera [Mis08]. Herstellung [Heß04]. Hetch [Sin06]. Hetchy [Sin06]. Hidden [Day04a, Dow01, Jer00, Slo08, Wak03]. Hierarchy [Sti09]. High [Baa01, Bas09c, Bra08b, Bug05, Gui08, Hul03, Kno07a, Mok01, Sha07, Vol08a, Daw06, Pri03, Sha04]. High-Altitude [Pri03]. High-Industrial [Bra08b]. High-Lead [Hul03]. High-Speed [Gui08, Daw06]. High-Tech [Sha04]. High-Technology [Kno07a]. Highland [Gre04c]. Highway [Fei09, Fra09, Mom05, Vol07b]. Highway-Network [Mom05]. Highways [Tea05]. Hill [Bla04a, Gor05]. Hilton [Ber03b]. Hindle [McG04]. Hiroshima [Sil00, Kru09]. Histoire [Dav09, Fox00]. Historia [BM07]. Historian [Bes08, Flo03, Kat02, Lip00b, Sil00, Uss09]. Historians [Wil00c]. Historical [Bel06, Big01, Bro06a, Cli00, Cor02b, Cru07, Don06, For08, Les08a, Mor07a, OS05, Sha08, Tho05, Wes08, Yar09b, Clo05]. Histories [Gof09, Iri06a, Lon04b, Wya04, Iri06b]. Historiography [Sto04a]. Historische [Heß04, Lac03, Rie05]. History [AA01, Aik03, AKP09, Ang05a, Ash09, Asn03, Bar00a, Bek03, Bek04c, Ben01a, Bla04a, Blu01, Bow00b, Bow04, Bra02, Bra04, Bro01, BDD09, Buc05a, Bug05, Bug07, Bul03a, CK04, Cer05, Cio06, Clo03, Clo07a, Col08, Col04, Coo03a, Coo09, Coo05c, Cor06, Day08, Div03b, Dob08b, Elk06, Esb05, Fin00, Fit01, Fra02, Fri09, Gel04, Gom00, Gor05, Gor07, Gro09, Gui03, Hac06, Hai03, Har04, Hey06, Hey08, Hof00, HW05, Hov04, Hub09, Igl01, Joh08, Joh00b, Joh04, Jon01, Jon09, Jud08, Kee03, Kha08, Kid03, Kid04, Kie09, Kle06c, Kol08, Kra01, Kra05b, Law00, Lin02a, Lon05a, Low02a, Low05b, Low00b, Low02b, Luc04, McC03, McC07b, McC08b, McG03, McG05, McS02, Mer05, Mid05b, Mil01, Mil03, Mol06, NV07, Och02, Old06, Olw04, Osl04, Osl07]. History [Pac02, Pas06, Per02, Pet05, Poh00, Pos01a, Pri03, Pur08, Pye01, Pyn09, Qui03a, Ree05a, Rit07, Rol00a, Roz09, RK08, Ryd03, Sch05c, Sch06d, Sch02e, Sch06e, See07, Sha00c, Sha01b, Sin05b, SD01, Smi05a, Spe04, Spi01, Sta05, Sta06b, Sto09b, Ten00, Tho07, Tom04, Tow02, Tra02, Vai00, Wal02, Wan02, War04, War07a, War09b, Wat00, Wer04, Wha06, Whi05, Wos04, Wri00, Yar09b, Zac06, Zel01, Zel02, dO01a, vdV08, Cas04a, Fra03, Wei03, McB00, Abb04, Ako07, All07b, Bac07, Bac08, Bas07b, Buc03a, Div09, Doi06, Edg08, Guc08, Hac03, Hug00b, Hug02, Kat09, Kie02, Lon04a, Luc06, Lue05, McG00a, Pla06, Qui05, Ras01, Sch00b, Sch07d, Sha06c, Sho05, Sli09, Tho09c, Uek07, Vol05a, Wal00, Whi04, Wil04, Woo05]. Hitler [All06b]. Hoax [Sch04b]. Hobby [Har03]. Hockey [Mok05]. Hoechst [Str09b]. Hold [Ham05]. Holding [Sha06b, Kou01a]. Holland [Nak07, vD02]. Hollywood [Fra06, Gre04a, Per00, Bas09a]. Holocaust [All02]. Holocausts [Jon02b]. Holodeck [O’H05]. Holographic [McC07b]. Holography [Joh05]. holprige [Mom04]. Holyoke [War06b]. Holznot [Die01]. Home [Bix02, Daw07b, End01, Gof09, Jag00, Par02b, Rit07, Dav06, Daw07a, End01, Har06b]. Homespun [Bla04b]. hommes [Bel04b]. Homogeneity [Gor03]. Honduras [Lan07a]. Honest [Sch00c]. Honor [Con07, Cut02]. Hoover [Ell01, Mah02]. Hope [Ede09]. Hopes [Ros05b]. Hopper [Vin06]. Horace [Mid06]. Horoscopy [Dal06a]. Horse [Bro05a, Gre08b, Cor09]. Horse-Drawn [Bro05a]. Horses [Men09, MK01, Ten05a]. Horticultural [Fro09]. Hot [Bor09a, Jos08, Smi03]. Hotel [Ben05a]. Hotels [Ber03b]. House [Dea08, Low04b, Mar00b, Sew06]. Household [Bix04]. Houses [Roc02, Sim03, Ste08a]. Housework [Zac02]. Housing [Bij03]. Houston [Kol08]. Howard [Coo02c, Wil00b, Wil00b]. Howe [Smi09]. Hub [Fin00]. Hubris [Hey06]. Huggies [Sca05]. Hugh [Kan07]. Hughes [Cut02, GS02, Cut02]. Human [Ala09, Bes08, Bla00b, Cro02b, Gli00, Her09a, Moo00, O’H01b, Owe03, Qui05, Rus07, Sch09a, Sli01, Ada06, Nye05]. Human-Built [Nye05]. Humanism [Kel06a]. Humanities [O’G08]. Humanity [Pas06]. Humans [O’G05, Ser02c]. Hundred [Clo05, Sch02e]. Hunsaker [Lev03b]. Hunter [Bro03]. Hunters [Rus07]. Hurricane [Kol07]. Hybrids [Hey06]. Hydraulic [Hof00, Muk06, TeB02]. Hygiene [Biv07]. Hype [See08]. Hysplex [Lie01]. Hysteria [Pan05b, Wos00].

IBM [Ake01, All02]. Icarus [Bes08]. Ice [Car03, Lon00, Ree05b]. Icelandic [Sve02]. ICOHTEC [GC07]. Iconography [Pan05b, Vol07b]. Icons [Hec08]. Idea [FF00, Sin09, Vai00, Doe08a, McM06]. Ideal [Bra01a, Dan03]. Idealism [Sto09b]. Ideals [Rau01]. Ideas [Sla02a, Ang05b, Pre03]. Identification [Sla02b]. Identities [VCM^{+}09]. Identity [BB00, Bra08b, De 03a, Els09, Lau08, Pic01, Smi06, Gou04b, Els09, Lin06a]. Ideologi [Jør09]. Ideological [Sch08a]. Ideologie [Fox02]. Ideologies [Sca08]. ideologische [Sto05]. Ideology [Bla01, Kuw00, Pal00, Ots03]. Idol [Bac02]. IG [Str09b]. II [And06b, Blu01, Cow06, Dor08, Sam07, Bac06b, McE04, Mei00b, Pic01, Vin03]. III [Bac04, Tap02]. Illuminating [Rom03]. Illusion [BW04, Gol00c]. Illusions [Tur08]. Illustrated [Ben01b, Joh04, Mid05b, Rit07, Sch00b, Sch02e]. im [Boc03, Gal00, Osi00, Sto07a, Wal04, Wen09, Joh02a]. Image [Dob08c, Mic08, Sin05b]. Imagery [Day01]. Images [Cor02b, Dea07, Hal07, How00, McG03, Sto04a]. Imagination [Bra08a, Coo02c, Cul09, Jos03, Kli04, May09, Sid04, Gre01, Mol08, Nei05, Seg00]. Imaging [New00]. Imagining [Goo04, Spa01]. Imitating [O’G08]. Imitation [Bur00b]. Imitative [Ben00]. Immediate [Kat02]. Immersion [BW04]. Immersive [RP09]. Impact [Bar09b, Bra03b, Bul03a, How06, Kno06b, Sch09a, Wal01a]. Imperative [Vie08]. Imperatives [Mor07b, Blu01]. Imperial [Ash04, Dou09, McB06, Sho05, Tag05]. Imperialism [Pan02b]. Implantable [Law03]. Implications [Kle06c]. Impossible [Gor08]. Imprint [Col05]. Imprisonment [Car09]. Improbable [Vin03]. Improvement [Jag00, Kle03b]. Improvements [Sha09a]. Impure [Ser05]. Incarnations [Bra08c]. Inclusive [Rob08b]. Incommensurable [Reu08]. Incorporare [Gag07]. Incredibly [Elk06]. Index [Ano00a, Ano00i, Ano02a, Ano02h, Ano02j, Ano03d, Ano04f, Ano05g, Ano06d]. India [Twe05, Bas09b, Sha08, Tya04, Hac05b, Alv04, Col05, Con08, Sto01c, Sto04a, Wal01a]. Indian [Bas09b]. Indies [Sto09b]. Indigenous [Har02, Hex07]. Individuals [Ang01]. Indonesia [Van08]. Industrial [All06a, Ash08, Bal00, Big02, Bra08b, Cho07, Cut06, Dan03, Dem03, Ede09, Fal06a, Fin00, Goo07, Gor04a, HP00, Hor07, Hyd00, Jos05, Kie09, Kir00, Kno06b, Lak05, Lan00, Las03, Luc05a, Mac09, Mac05b, Mae09, Mar02, McB04, McG00c, McN02, Mil04c, Mil09, Per07, Qui03b, Rey07, Rob05a, Sch02a, Sch02c, Sch00d, Sha00b, Sha06e, Sla01, Tra02, Tri03, Vol09b, Wea02, Wen01, Yar09a, Zac06, vR07, Wal06a, Far00, Jon06, Sch07c, Tod02]. industriale [Vas08]. Industriales [PP09]. industrialisation [Bel04b, Fri08]. Industrialisierungsgeschichte [Mar03]. Industrialization [Bre00, Coc03, Eel09, Lip00a, Lon03b, Pan07, Sve03, Big00, Ble04, Cox04, Pre06, Wit08]. Industrialized [Jud03]. Industrializing [Don07, Mac02, Woo05]. Industrie [Har08b, Haf09, Sto07a]. Industriegeschichte [Lac03]. industriel [Gil06]. Industrielle [Joh02a, Gou00, Pan09, Lyn09]. Industries [Con00a, Dav01, Div00, Gle03, Kno07a, Moh08, Ree05b, Ake05, Ens06, Pie06]. Industrin [Har06a]. Industrious [Wer05a]. Industry [App03, Bek06, Bel06, Bor09a, Bro00b, Bud05, Bud09, Bur00a, Bur00b, Cro07, Dir05, Gif07, Hor06, Ink06, Kle06c, Kno01b, Luc02, Mad06, Maz01, McI00, Mer05, Moh04, Pan02a, Ran01, Spi00b, Ste01c, Ste08b, Tak00, Tei05, Vol01, Vol03b, Zem02, Bed03, Bla01, Ble07, Bra08b, Cha07, Epp03, Kno03, Kra00, Lic02, Lin06b, Lug03, Mac04, Pra03, See01, Tom04]. Inequality [Sla09]. Infertility [Bul03b]. Infinity [McC01b]. Inflame [Bra07a]. Influence [Moh08]. Influences [App03]. Informatics [Pic00]. Information [Bar00a, CK01, CK04, Cer06, Coo09, Dow01, Dow08b, Epp06, Hea01, Hoc05, Iri06c, Jac08a, Kou01a, Lip06, Low04c, Mar00a, Mol00a, Mol03, Pfa04, Pop01b, Wil00c, Cas03, Kli06, Lip03, Eri08b, Kid01, Pop07, Sha06c, Yar07]. Infrastructure [Sta01, Sti03, Rey07]. Infrastructures [Hår08a, Lig02, Poi08]. ingegnere [Lap06]. Ingenieros [PP09]. Ingenieur [Gis07]. Ingenieure [Har00a]. ingénieurs [Fro01b, Gui06]. Ingenuity [Fre01, Fou06]. Inglorious [Jon00]. Injuries [Ten00]. Injury [Col07a]. Innis [Mul09, dO05]. Innocence [Cla07]. Innovate [Mil04a]. Innovation [Ang04, Bak00, Bar05, Bas02, Bur00b, Cla01, Eel09, Fre07, Joh02c, Lar04, LB08, Lug03, Moy06, Ryd01, Sli01, Spa01, Swe08, Vol04a, Wel07, vDS05, vL09, vR07, Lev03a, Kle04, McE03, Mei00a, Moñ07]. Innovations [Mod08, Pet07b, vD02]. Innovationsschwache [Cie01]. Innovative [Dow05, dWvdESvO02]. innovativen [Wen07]. Innover [Haf07]. Inquiry [Fri08]. Insatiable [Moo02b, vL09]. Inscriptions [vF05]. Insects [Kin09b, Ser02c]. Inspired [Orv09]. Instantaneous [Bri07]. Institute [Bas09b, Hof00]. Instituten [Sto07a]. Institutional [Div00, Kai02, Ser02b, Wea02]. Institutions [Leo01, Mod06a, Sla02a, Div00, Sto06a]. Instrument [HM01, Law00, Tap02, War02, War07b]. Instrumental [Mod06b, Ste06]. Instrumentalisierung [Sto05]. Instruments [Ash08, Dea07, DvdE03, Dob08d, Lan02a, Law00, Ste06, Cav05, Cli00, Low02a, McS00]. Insular [Tag05]. Insull [Sch07b]. Insurance [Lip06]. Integration [Edw03b, Kra07b].

Intellect [Pop02]. Intellectual [Abb04, Bra05, Lew00, Mac00, Mac05b]. Intellectuals [Sch04c]. Intelligence [Ang04, Kel04, Ric08b, Sei03, Sla09, Ang04]. Intelligent [Kāt06, Muk03]. intérêt [Ben03]. Interactions [Dob04a]. Interests [Sla02a]. Interface [Kri00b]. Interferon [Bud07]. Interior [Gud07, Ben09]. Internal [Kle03b]. International [Bar00c, Ber03b, Bri07, Clo07a, Jon01, Jon09, McG00c, Mis08, Ryd03, Sha06c, Vol03b, Wea02, Wha06, dO01b, AKP09, Cor06, Esb05, HW05, SD01]. Internationalism [Eck05]. Internet [Dow09, Tur08, Tur09a, Wan08, Wu03, RK08, Bas09c, Kou01b, O’G03, Rol00b]. Internetworks [Dow01]. Interpreting [Van02]. Interwar [Lew01, Kry04]. Intimate [Smu07]. Introducing [Woo05, vDS05]. Introduction [Bac02, Par02b, Pre00, Pre03, Rob04]. Intuitive [Kru08]. Invented [Pre08b, Wil04, Lig09]. inventer [Haf07]. Inventing [All03, Ell01, Fro01a, Tri04, Coc04, Ens06, Fan01, Mar06b, Reu05, Rol00b, Tri06b]. Invention [Ald06, Ber02, Bra03b, DO03, Dal05, Dou04, Fri08, Fro05, Hin08, Hol06, Lau08, Lon01a, Mac01, McB00, Mol07, Nor07, Pan03a, Sch08b, Tra05, Coo05a, Fab06, Kli00, Pan05b, Smi07]. Inventions [Bra05, Buc03b]. Inventive [Fou06]. Inventiveness [Haf03]. Inventor [Fro05, Hin09, San03]. Inventore [Twa00]. Inventors [MN06, Pre04c, Kha01]. Investigations [Bud05]. Invincible [DvdE03]. Invisible [Mil04b, Wan02, Wor01]. Inviting [Luc03]. Iowa [Gol00b, Hof00, And05]. iPod [McC09b]. Ipswich [Mil06]. Ireland [How08b]. Iron [Bro00b, Coh07, Gre02, Kno01b, Kno06b, Men09, Mil04c, Rob00, Ryd01, Seg09, Spi00b, Buc03a, Gor01a, Lan00]. Iron-Making [Lan00]. Ironclad [McB00]. Ironclads [McB04]. Ironies [Her01c]. Ironmakers [Twe05]. Ironmaking [Rob02]. Ironworkers [Wer05a]. Ironworking [Awt07, Dav00]. Ironworks [Rob00]. Irony [Coh07, Tod08]. Irradiation [Spi04]. Irrigation [Sha08, Van08]. Isambard [Coo04a]. Islam [Sal05]. Islamic [Bul03a, Sal01, Mir08]. Island [Bro07]. Isolation [Tho08]. Issue [Ano00d, Ano00e, Ano00f, Ano00g, Ano01b, Ano01c, Ano01d, Ano02d, Ano02e, Ano02f, Ano02g, Ano03a, Ano03b, Ano03c, Ano04c, Ano04d, Ano04e, Ano05c, Ano05d, Ano05e, Ano05f, Ano06b, Ano06c, Ano06f, Ano06g, Ano07d, Ano07b, Ano08b, Ano08c, Ano09d, Ano09b, Ano09c, Wal01b, Ano09d]. Issues [HPV06, Pow04, Skr00]. Istanbul [BE04, Din08, Ber06]. Italia [Lon02a]. Italian [Car08, Lon01a]. Italy [Bel04a, Gag02]. Ivy [Tei05].

J [Dal06b, Pre04c, Rot04a, Skr04]. J. [For09, Her07a]. Jack [Coo02c]. Jackson [Rot04a]. Jacksonian [Chu04a]. Jacquard [Cer06]. Jaffa [Men09]. Jahre [Die01]. Jahren [Sto01b]. Jahrhundert [Boc03, Gal00, Low00a, Wal04, Zel04]. Jahrhunderts [Ber06]. James [Mor08a, Mil08]. January [Isr02, Pre05, Ano00j, Ano01f, Ano02k, Ano03e, Ano04g, Ano05h, Ano06h, Ano07e, Ano08f, Ano09g]. Janus [Zai08]. Janus-face [Zai08]. Japan [Bek04b, Cla01, Jer00, Nis08, Sam07, Tak01, Koi02, Tak00, Coo02a, Ink09, Kuw00, Rav01, Sto01a]. Japanese [Flo00, Hig03, McB06, Nad08, Ots03, Tak01, Wit07, Wit08, Nis07]. Java [Van08]. Jaywalking [Nor07]. Jazz [New06]. Jean [Bac05]. Jean-Antoine [War03]. Jenkin [Bri01]. Jerome [Lev03b]. Jeronimo [San03]. Jersey [Gor05]. Jerusalem [Men09]. Jew [Str05]. Jewel [Dea08]. Jewelry [Car08]. Jobs [Fan01]. Joel [Wil03b]. John [How09a, Kno01a, Nie07a, Bac04, Luc09, Ros03, Sti04b]. Joined [Smi05b]. Joneses [Jac05b]. Jose [Ano02i]. Joseph [Bra04, Isr02, Swe08, War01b, Per06a]. Journalism [Mul07]. Journals [RK08]. Journey [Hac04, Whi01]. JPL [Sat07, Wes07]. Jr [Cer09]. Jr. [Hug07, Rol03]. JSTOR [Lue05]. Juanelo [Alt00]. July [Ano00l, Ano01h, Ano02m, Ano03g, Ano04i, Ano05j, Ano06j, Ano07g, Ano08h, Ano09d, Ano09i]. Junction [Zac02]. Junctions [dWvdESvO02]. Jurassic [Rot02]. Just [Fin09, Ros05a]. Justice [Gol00a, All07a]. Jutland [McB07].

Kahn [Kru08]. Kaiser [Mor03, Sto07a, Pos05]. Kaiser-Wilhelm-Instituten [Sto07a]. Kaiserreich [Fox02]. Kalter [Fre06]. Kamil [HV07]. Kampfstoff [Sto07a]. Kampfstoff-Forschung [Sto07a]. Kanawha [Rey01a]. Kanpur [Bas09b]. Karl [Mos06, Pet01a]. kartografische [Boc03]. Katrina [Kol07, Sha06d]. Kavala [Ren09]. Keep [Jeo09]. Keeper [Coh07]. Keeping [Jac05b]. Kemper [Coo04c]. Kennedy [Day08]. Kept [HV07]. Kernwaffenversuche [All06b]. Keyword [Kli06]. Khan [Sta06a, Iri05]. Kidder [Moo04a]. Kill [Pop00, For03]. Kimberly [Sca05]. Kimberly-Clark [Sca05]. Kinematics [Kel09]. King [Tap02, Bac04, Bac06b, Mid06]. Kingdom [Ale03, Cha01, Coo04a, Kno05a]. Kitchen [Par02b, Van02]. Kite [Sch04b]. Klan [Gor01a]. Kleenex [Sca05]. Knock [Mor06a]. Knocking [OS05]. Know [Ree05b, Tya04, Wal09]. Know-How [Tya04]. Knowing [Pac02, Sin00]. Knowledge [Ang04, Ash06, Bel04a, CK01, Cav05, Con99, Eck05, Gre08c, Ham09, Her05, HPV06, Kri03, Moo04b, Muk06, Pop02, Scr00, Ste06, Tim09, Van00b, Wak04a, Wei04, Har02, Low08b, Sto06a]. Knowledgeable [Con06b]. Koc [BE04]. Kommunikationssystems [Pet01c]. Konsum [Rie05]. Konsumgesellschaft [Orl02]. Kontroverse [Sto05, Zet08]. Kooperation [Sto07a]. Korea [Vol03b]. Kotex [Sca05]. Kraftfahrkorps [Has06]. Kranzberg [Cer09, Pos09a]. Krieg [Fre06]. kriegstechnische [Wal04]. Krise [Har00a]. Kultur [Sud08]. kulturelles [Pet00]. Kulturgeschichte [Ber03a, Zel06, Zet08]. Kyoto [Mac04].

L [How08a, Mis08]. Lab [McC09b]. Labor [Ble07, Coo00, Coo02b, Cox04, Dow01, Hoc03, Kno01b, Mie04, Mol06, Ste08b, Tod02, Wak02, Haf03, Gle08]. laboratoire [Le 05]. Laboratories [Smi03, Wen01]. Laboratory [Gol00b, Low04a, Sha00c, Wes00, Cut09b, Sta05]. Laborers [Gle08]. Labour [Moh08]. Labrador [Fal06b]. Labs [Cer04, Wes04]. Labyrinth [Kou06]. L’age [Lev03a]. Lake [Ell01, Hig01, Ste01a]. Lakes [Dal02]. Lamps [Bow00a]. LAN [Abb03]. Land [Bar03a, Fri07b, McM06, Pos09b, Sto03a].

Land-Grant [McM06]. Landfill [Gam07, RV04]. Landing [Con01b]. Landings [Jeo08]. Landmaking [Wer04]. Landmarks [Gre02]. Landscape [Bla09b, Coa02, Igl01, McN02, Nak07, Pas02, Rob04, Rob02, Rot04a, Tha04, Tod09, Dem03, For05, Fos04, Lan02b, Mno00, Rey07, Tea05]. Landscapes [Nie02, Nor09, Pla06, Uek07, Mel06, Twe05]. Landschaft [Boc03]. Landschaftsveranderung [Sch05b]. landskap [Hyl00]. Landwirtschaftspolitik [Har08b]. Lane [Coo01a]. Langewiesche [Flo03]. Langley [Hun00, Joh02b]. Languages [Lak05, Pre01]. Laplace [Wak07]. Lappeenranta [Sch06b]. Large [Cer09, Les08b, Sve02]. L’Art [War03]. Last [Cro08a, Kel06b, Pos03, Pos09b, Pre04d]. Late [Bal00, Bek04c, Ede09, Fri08, Tod08, Jør08, Por03, Jon02b]. Late-Eighteenth-Century [Por03]. Latimer [Pre04c]. Latin [Sol04]. Latitude [Rot04b]. Launch [Bug05, Coo03b, Sat08]. Launched [Joh06]. Launching [Sid06b]. Laundries [Mal01]. Laureate [Joh06, Str05]. Law [Col07a, Gor02a, Mac00, Mac02, Cer05, Uss03]. Lawrence [Coh07]. Laws [Sch05c, Sch05c]. Layton [Kli08]. Lead [Aik03, Hul03, Yar08]. Lead-Paint [Aik03]. leaf [WW00]. Learned [Bra03a, Kri00a, Pri00]. Learning [Ayd05, Reu02]. Leather [Orv09]. L’École [Pan02a]. Lécuyer’s [Bas07c]. Led [Cer06]. Legacies [Sch09b]. Legacy [Bac05, Jer00, Kar09b, Kha01, KO02, Per06a, Sha00c, Pet04b, War06a]. Legal [Dob08b, Wil04]. Legend [Col01, Mid06, Nie07a]. Legendary [Hok07, McB00]. Legitimacy [Muk03]. Lehigh [BS09, Cut09b]. Leicester [GC07]. Lengthening [Bra02]. L’Enquête [Lyn09]. Leo [Mei03]. Leonardo [Kel09, Lon06]. lernen [Mom02]. Lessons [Hug00a, Luc03, Pri00, Ben05a]. L’Etat [Pan03a, Lan04b]. Letter [dV02]. Levant [McM00]. Leviathan [O’H09]. Lewis [Pre04c, Wil02a]. l’histoire [Haf07]. Libavius [Pow09]. Liberal [Wis09b]. Liberalism [Wan07]. Liberty [Gor02a]. Libraries [Orr02]. Library [Ker06, Kid03, Asn03, Chu04b, Kul03]. Licensed [Pop00]. Lie [Col08]. Life [AA01, Ber08, Bix02, Bla00b, Bow00c, Col01, Coo04a, Cro05, Cro06c, Cut07, Epp07, Flo01, Gam02, Git01, Gor04b, Isr05, Jon02a, Kli00, Kou01a, Lip06, Lun09, Mat02, McC06, Mid06, Neu06a, O’H03, Pic00, Poo03, Sch07c, Sew06, You01, Mor01, Aga09, Bij03, Buc03b, For04, Her07a, Kle03a, Mah02, Wos04]. Lifting [Van00a]. Light [Bra07b, Cow04a, Fin02, Jon09, Kra09, Wak05, Pre01]. Light-Green [Wak05]. Lighter [Sch08b]. Lighthouse [Buc01b, Sch05a]. Lighting [Bow08, Bra02, Tag05]. Lightning [Coh07, Mol04, Isr05]. Lights [Rom03]. Like [Hof02, Reu08, Wis09b, Gel04]. Limited [Bry07]. Limits [Her07b, Sch08a, Sto01a, Vol07a]. Lincoln [Eri07]. Lindbergh [Fer05]. Linde [Ree05a]. Lindqvist [Hyl00]. l’industrialisation [Pan00]. L’industrie [Pan03b, Le 07]. Line [Bro05b, Cas05, Coo02b, McG00b, SS07, Clo07a]. Linen [Bel06]. Lines [Clo07a, MA09, Sto03c]. Link [RK08]. Linking [Dyb09]. Linkoping [Luc04]. Links [Dow04]. L’Invention [Gou01, Wer02]. Lisbon [Ano09e, Dio08]. Lise [Wes01]. Listening [Arn03, Kra07a, Ras01, Coo02c, Lar03]. Lit [Tag05]. Literacy [Kid01]. Literary [Ker06, Pin08]. Literature [De 04, Egl07, Hov04, Jen05, Kie01, Lon01b, O’H01a, Pic00, Jac04b, Kel06a, Low04c]. Live [Mis07, Wil02b, Bra03a]. Livelihood [Par02a]. Liverpool [Bro09c, Hor07]. Lives [Lon04b, Sti03, How00]. Living [Ala09, Gre08b, Jac06, Jaf03, Mor01, Muk06, Kou06]. Lloyd [Rau01]. Local [Kle09, Sat05]. locales [Bel04b]. Locks [Ben05a, Sha06e]. Locomotion [Gam05]. Locomotive [Kle08, Sch01b, Sch05d]. Locomotives [Jon03]. Loewy [Blu04, Kul03]. Logging [Hul03]. Logic [All05]. London [Dea08, Sch07a, War02, Div08, Dob08c, Dob08a, Fin00, Hun02, Luc06, Tym00]. Long [Bro07, Elk06, Mom05, Stu07, Yea04, Whi01]. Long-Range [Mom05, Yea04]. Longue [Mol05, GC07]. Looking [Ano08d, Ano08d, How00]. Looks [Fre01]. Loom [Cer06, Gle08]. Loose [WW00]. Loose-leaf [WW00]. Lords [Rou02]. Losing [Kin09b]. Loss [Pra01]. Lost [Lan03a, Mor08a, Sta05, Hor06]. Lou [Ham05]. Louie [Sch06e]. Louis [Ben05a, Mil03, Pre06, Sch06e, Bro03]. Louisiana [Lan02b, All06a, Sha06b]. Louse [Sta05]. L’Outil [DeV00b]. Love [All07b, Der04a, Gor08, Tho09b]. Low [Bel08, Con06a, Jeo08, Joh02c, Mok01]. Low-Cost [Joh02c]. Low-Visibility [Jeo08]. Lowell [Gro09]. Lowengard [Rob07]. Lower [All06a]. Ltd [Gif07]. Luddites [Jaf06]. Luftschifffahrt [dS09]. Lumières [Gou01]. Lunar [Eze02, Her08a]. Lustron [Har05]. Luxus [Rie05]. Lynn [Rol03]. Lynwood [Les07]. Lynx [Lon04a]. Lyon [Lon00].

M [Bek02, BE04]. Machine [Bed05b, Bor00, Chu04b, Coc04, Coo04c, Coo05c, Cor02a, Cut09b, Don06, Fre07, Gan05, Gif07, Haf03, Her09a, Lak05, Mei03, Mei05, Moo04a, Mor03, Owe03, Pie06, Pos06, Sei03, Sin05b, Ver08, Wil09, O’H00, May09, Sin00]. Machinery [Kel06a]. Machines [Alt00, Bro06b, Con00a, Der04a, Gre08b, Isr03, Kat05, Kel09, Mag05a, Mai04, Mei02, O’G08, Osl07, Pre02, Pre03, Waj00, Eri08b, Law03, Mol07, Wei01, Wei04]. Machinists [Kno07a]. Macht [Hey01]. Mackenzie [dO01b]. Mackinac [Pos08a]. Macon [Bro00b]. Made [Con08, Fis07, Her01c, Yar09a, Zet08, Lai07, Mag09, Mar00b]. Madman [Pos05]. maestra [Por05]. Magic [Arn01, Fro05, Nye06, O’H00, Nad04a]. Magical [Wos04]. Magick [Spe04]. Magnetron [Ste01b]. Magnificent [Har05]. Mail [Bed03, Wal05]. Mails [Eri07]. Main [Dob08c, How09b]. Maine [Kli04]. Mainstream [Str02]. Maize [Cro06b]. Majesty [Ash09]. Major [Bug07]. Makers [Law00, Tap02, War02, War07b, War02]. Makes [McC01a]. Makin’ [Wil00b]. Making [Ang06, Bas07c, Ben01d, Boy01, Bro05a, Cho07, Coo07a, Cow04b, Div00, Flo07, Fro01a, Gou04b, Hec08, Hor04b, Joh00b, Jon02b, Jud05, Kir00, Mac00, Maz03, McN02, Mir08, Sta06a, Str02, Tak01, Tei05, Tod09, Wel08, Ash08, Big02, Cut06, Jaf03, Kra07a, Lan00, Lug03, McB06, Neu06b, Rob00, Roc02, Tim09, Vol08a, Waj00, Wea02, Wes00, Wya04, Zem02]. Male [Neu06b]. Man [Car09, Her01c, Hol06, Isr05, McB00, Neu06a, Nie07a, Ver08]. Man-Machine [Ver08]. Man-Made [Her01c]. Management [Big07, Bij02, Cla01, DvdE03, Hei04, Kli06, Kuw00, Lig06, Lin00, Lin02b, Mac05a, Reu02, Sto01a, TeB02, Wak02, Wes07, Haf03, dV02]. Managerial [Har01]. Managing [Jør08, Mod08]. Manchester [Bri06, Qui03a]. Manhattan [Sch00c]. Mania [Bor09b]. Manifesting [Wei01]. Mankind [Gan00]. Manned [Her07b, Mac06]. Manon [Ber08]. Manufactory [Coo00]. Manufacture [Law00, Ran01, Tya04, DeV03]. Manufactured [Wor01]. Manufacturers [Gud07]. Manufacturing [Ake05, Epp00, Koi02, Las03, McG00c, Har01, Kuw00, McN02, Rii03, Ste01c]. Many [Ben05b, Pet09]. Mao [Vol03a]. Map [Clo07a, Kri00a]. Mapmaker [Kon05]. Mapmakers [MA09]. Mapping [Clo06a, Hoh03, Bek04b]. Maps [Bra07a, Clo03, Clo06a, Day04b, Kar09a, Ste05b, Yar09b, Kon09]. Marc [Pan09]. March [Pop02, Pre05, Pre08a, Rob08a, Gre04c, Seg09]. Marconi [Bow03, Fro05]. Marine [Don06, Wei06]. Mariner [Lon09]. Maritime [Bro01, SS02, Leo01]. Market [Ald06, Dir07, Dan01, Wel08]. Marketing [Van00b, De 03a, Str01]. Marketplace [Cor09]. Markets [Fri07b, Ran01, Yal08, Bow07]. Marne [Pan02a]. Marriage [Low01a]. Marshall [Sat05]. Marston [Mil08]. Martin [Sch05e]. Marvels [Rey02a]. Marx [Mei03]. Mary [Lin00]. Masculine [Waj00]. Masculinity [Smi02]. Masons [Muk06, Har02]. Masquerades [Cla07]. Mass [Big00, Dir07, For08, Lic01, Mor08b, Mul07, Pre04b, Rap01, Rus08, Sca08, Smu07]. Mass-Mediated [Smu07]. Massachusetts [Mil09]. Masses [Bre01, Jos01, Sid04, Tho05]. Master [Egl07, Her01b, Kid06, Mil03, Joh06]. Master-Slave [Egl07]. Masters [Gud07, Dor01]. Masterwork [Haf05]. Masterworks [Ste05a]. Material [Ako06, Ber01, Bla00b, Gan03, Jer00, Kra07a, Poo03, Ros03, Sch09a, Sew06, Sch00a, Sch09a]. Materialising [Els09]. Materiality [Dob06, Jon06]. Materials [Ben00, Bug01, Ben01b, Wes08]. Materiaux [Kah01]. Mathematical [Kid04]. Mathematics [Wis09a]. Mathematik [Bey06]. Matter [Ink06, Jac04b, Scr05]. Mattern [Lau02]. Matters [Mis07, Bas05]. Mauchly [Lig06]. Mavericks [Spe04]. May [Sch06b]. Mayhem [Bla09c, Spe04]. McCarthyism [Mul04]. McCullough [Mil03]. McPhee [Sti04b]. Me [Bor00, Sti03]. Meadow [Bra07a]. Mealtime [Van02]. Meaning [Abe02, Bey09, Lip03, War05b, Bas00]. Meanings [Pay05, Sch06a, vdV08]. Means [Wer07b, Sch08a]. Measure [Ash04, Jac06, Sla09, Wak03, Ash04]. Measurement [Tod08]. Measures [Ash05]. Measuring [Coo04b]. Meat [Boy01, Lan07b]. Mechanical [Ale06, Big07, Kat05, Lak01, Wos04]. Mechanics [Cor08, Pet05, Abe02]. Mechanism [Lie01]. Mechanisms [Pra03]. Mechanization [Gle08, Hol09, Mol02, Zac02]. Mechanizing [Maz04b, Uss04]. Médecins [Fro01b]. Media [Cha02, Kim04, O’G04a, Rou05, Wer07b, Yal08, Mor08b, Mul08, Poh07, Sam05]. médias [Lev01]. Mediated [Smu07]. Mediating [Mal09a]. Mediation [Reu08, Gui06]. Medical [Aik03, How00, Les08a]. Medicating [Sli09]. Medicine [Biv03, Biv07, Bro09b, Fit00, Gar09, Her07b, Ram05, Sat04, Sto01c, Sto03b, Wei01, Bro09b, Clo06a, Ker06, Lan03c, Pac02, Shu02]. Medieval [Bac05, Hac05b, Her01b, Jør08, Luc05a, Maz07, Phi03, Rol03, Sal01, Ung05, vD09, LW05, Luc05a, Wal00]. Medioevo [Lon02a]. Mediterranean [Maz04b]. Medium [Has02, Mul04, Tri06a]. Meeting [Ano00h, Ano01e, Ano02i, Ano04a, Ano05a, Ano06e, Ano07c, Ano08d, Ano09e]. Meiji [Ink09]. Meitner [Wes01]. Mel [Pos09a]. Meltdown [Luc05b]. Melts [Wil00c]. Melvin [Cer09]. Memoir [Fis07]. Memoirs [Hac04]. Memorial [McG04]. Memories [Gof09]. Memory [Bac05, Bla09b, Git01, Gom00, Gre06b, Pet05]. Men [Cor09, Dav00, Hof02, Osl07, Phi03, Sch05c, Seg09, Waj00, War06a, Har03]. Menlo [Pre05, Pre08b, Rot01]. Menschen [Bak04]. Mental [Bey08]. Mentality [Bey08]. Mercator [Clo07a]. Merchant [Rab04, Sch07b]. Merchants [Gud07, Per07, Ran01, Rey02a]. Mercury [Bea02]. Merging [Sto03c]. Meridians [Clo07a]. Merit [Sla09]. Mesopotamian [Dal06a]. Messenger [Bek00, Hoc03]. Metal [Bug01, Jen05, Pan00]. metallica [Lon03a]. metallurgie [Le 05]. Metals [Gor06]. Metaphor [Egl07]. Metaphysics [Dob04b]. Meteorological [Kri00b]. Meteorologists [Kri00a]. Meteorology [Tur09b]. Method [Awt07, Hun02]. Methodological [HPV06]. Methods [Edw03a, vR07]. Metiers [Pan02a]. Metric [Ash04]. Metro [Jen00, Tho07]. Metropolis [Abb05, Chu04b, Iri06c, Kol08, McC08a, Sch07b, Yar09a]. Metropolitan [Fri07b, Sim03, Sim09, Hoo04]. Meucci [Twa00]. Mexican [Lar08]. Mexico [Car02, Lan07b, Tow02]. Michael [Sch07c, Lon09]. Michelin [De 03a]. Michigan [Gre04c, Jon04b]. Microchip [Hol06]. Microscope [War09b]. Microscopy [Mod06b]. Mid [Sha00a, Sca08]. Mid-Century [Sca08]. Mid-Twentieth [Sha00a]. Middle [Bla08, Hur02, Jac06, Jaf03, Les04b, Ros06, vD03, HPV06, TeB02]. Middle-Class [Les04b]. Midway [Gui00]. Midwestern [And06a]. Might [Cut06]. Mighty [Sti04a]. Migration [Lak01]. Mikrosoziologie [Wen08]. Mile [Fis07]. milieu [Pan03b]. Militär [Sto07a]. Military [Ang04, Ang05b, Bek06, Blu01, Bro00a, Col07c, DeV03, Hac09, Mei02, Muk06, Rey05, Sch02c]. Military-Industrial [Sch02c]. Mill [Che04, Wak02]. Millennium [Hut00, Kou01a]. Miller [Men07]. Millers [Tri03]. Milling [Luc05a, Maz07]. Million [Cut09b]. Mills [Maz04b, Ung05]. Milwaukee [Hyd01]. Mind [Joh06, Ver08, Mai09]. Mind-Body [Ver08]. Mindedness [Pal00]. Mindful [Fri08]. Minding [Lak05]. Minds [Dob08d, Sch09a]. Mine [Gor05, Pan00]. mineraria [Dol09]. Miners [Dym08]. Mines [Lon01b, Och01]. Mining [Bur00b, Cro07, Dym08, Gor05, Hov04, Lon03a, Gof09, Kir08]. Minister [dV02]. Ministry [WW04]. Minneapolis [Sch06e, Ano06e]. Mint [Hur02]. Minutes [Low04b]. Miracle [Bud07, Bek07]. Mirror [War05b]. mise [Mic08]. Missile [Bar04b, Geo06, Hig01]. Missing [RK08, Scr00]. Mission [Bea02, Mor07b, Bro06a]. Missions [Sto03b]. Mississippi [All06a, Mil03, Har04]. Misunderstanding [Hou03]. Mitigation [Wil03a].

MITRE [Fri02b]. Mitte [Ber06]. Mittelmeers [Gis01]. mixte [Kah01]. Mixtures [Lan08]. Mobile [Aga09, Tri06a, Dow08a, Tri09]. Mobilisierung [Fre06]. Mobilität [Zel06]. Mobilities [Lig02]. Mobility [Ble07, Cor06, Dyb09, Esb05, Mom05, Bla09c]. Mobilization [Col07c, Nei02, McB04]. Model [Tho09c, WW04, Wel07]. Modèles [Gui06]. Models [Bow07, DvdE03, Rob06a]. Modern [Ade08, Alv04, Bal00, Bek04b, Bel04a, Bel08, BM07, Ben09, Ber03a, Ber02, Ber03b, Bey00, Boy03, Bro08, Cul09, De 06b, Dou01, Eis09, Fit01, For04, Gin07, Gre08a, Hai03, HPV06, Jac02, Jac06, Jaf08, Ker06, Kid06, Lak01, Lin06a, Lon02b, Lon04a, Lon04b, Mac00, Mai04, Maz03, McC08a, Mok09, Nak07, Oli07, Osl04, Par02b, Per07, Por04, Pre08b, Rey02a, Rob08a, Rou05, Rus08, Sch07b, Sho02, Sho08, Sid08, Sin06, Sli09, Spi00b, Sta06a, Sto07b, Sve03, Tho06, Tim09, Tod02, Tod09, Tur09b, Vog02, Waj00, Wel08, Nad04a]. moderna [Squ00]. Moderne [Har00a, Mol02]. Modernism [Bra08b, Kat03]. Modernist [Big07, Der04b, Pri02, Win08]. modernite [Ben03]. Modernities [Hex07, Shu09a]. Modernity [Arn03, Epp02a, Lar08, Lew00, Mae09, Ots03, Poo03, Pri02, Sch06c, Sim09, Sto04a, Wak05, Bis05, Cav08a, Hou04a]. Modernization [Cro08a, Din08, Gor03, See03, Zei03, vD02, Bek07]. Modernizing [Col07b, How09b]. Module [Rap01]. Moky [Wil03b]. Mold [Big08]. Moment [Dow03b]. Moments [Seg08]. Monasteries [vD03]. Money [Rau03, Mag05a]. Monique [dV02]. Monitor [McB00, McB00]. Monitoring [Hei04]. Monitors [Eri08a]. Monster [Pos09b]. Mont [Mag05b]. Montgolfier [Wak02]. Montreal [McN02, AST08, Poi08]. Monument [Chu04b]. Moog [Bra03b]. Moon [Eze02]. Moonwatch [Jeo09]. Moore [Cer05]. Mooring [Bur06]. Moral [Ber01, Olw01, SHW03]. Morality [Pop02, Wal05]. Morals [Tod08]. Morgan [Lan09]. Morris [Kno01a]. Morse [Isr05]. Morton [Bac02]. MOS [Cer04]. Moses [Epp07, Tea08]. Mosquito [Sta06b]. Most [For07, Fro05, Sin06, Tri06a, Zel01, Zel02, dlP07]. Mother [Par05]. Motion [Mal09a, Sam05, Vol09a]. Motion-Picture [Mal09a]. Motor [End01, Epp09, War06a, Wel07, Nor07]. Motorcycle [Tho00]. Motorcycling [Vol09a]. Motorenkonzeptes [Wen07]. Motoring [Fra09]. Motorisierung [Mom02, Has06]. Motorization [MK01, Sch00c]. Motorized [Mom05]. Motors [Dic04]. moulines [Bla03]. Mount [Jen00, War06b]. Mountains [Bar09b]. Mountaintop [Bar09b]. Mouse [Wil09]. Move [Cas04a]. Movement [Olw04, Kar08]. Movements [Hac09]. Movie [Gof09, Cer03]. Movies [Bas09a]. Moving [Bro09b, Con00b, Vog04]. MP3 [Hor04a, Mag05a]. Mr. [Eri07]. MRI [How00]. Mrs [Ber03a, Bix04]. Mud [Fei09]. Mulholland [Spo02]. Mulhouse [Gwy09]. Multhauf [Fin05]. Multilevel [vDS05]. Multinational [Jon09]. Multiplex [Daw07a]. Multitude [Vol08b]. Mumford [Wil02a]. Munchen [Bey06]. Munich [Ano01e, May00]. Munitions [Dav01, Das08]. Muntz [Pos05]. Murphy [Pet03]. Murrow [Coo02c]. Muse [O’H01a, Rii03]. Museum [All04, Asn03, BE04, Cas04a, Chu04b, Dou04, Fra03, Goo09, Gwy09, Jon04b, Kul03, Pos01a, Sch07a, Sch04d, Sti03, Vog04, Wei03, Zei01, Div08, FF00, Hoa01, Ren09, Rot02, SS02]. Museums [Bek04c, Dio08, Hut00, Pre00, RP09, ZH00, Dou01, Wya04]. Music [Bra01b, Bra03c, Git01, Mag05a, PB03, Ras01, Tho01, Arn01, Bij05, Hor04a, Kra05a]. Musical [McS00, Ras01]. Musicians [War07b]. Muybridge [Bro05c]. My [Gor01b]. Myriad [Sho02]. Myself [Sti03]. Mysteries [Bas07c]. Mystery [Rot04b]. Mystique [Alc07]. Myth [Bow05, Coc04, Coo03a, Low08b, Min00, Pan05a, Gou04b, Bla04b]. Mythical [Sma05]. Myths [Tes09].

’n [Coo02c]. NACA [Joh04, Lev00, WF03]. Naissance [Le 05, Haf09]. Nakamura [Bow08]. Name [Bra07a, Coo04c]. Named [Coo04c]. Nano [See08]. Nano-Hype [See08]. Napoleon [Ang04, Pan02a]. Napster [Hor04a]. Narrative [Pos04]. Narratives [Mar06a, Sch04a]. Narrow [Men09, Sch05e]. NASA [Bro06a, Daw06, Gui03, Her08b, Hun00, Lev00, Rol01, Van00a, Joh04, Kra00, WF03]. NASA’s [Sat05]. Nathan [Rot05]. Nation [Big08, Boc03, Car02, Coo03b, Fox02, Glo02, Hea01, Jon03, Lak04, Mos02, Ols07, Par02b, Pop01a, Wit07, dO01a, Gol00b, Pos01b, Kli04]. Nation-Building [Big08]. National [All03, Bar07, Bro00c, Dou04, Els09, Gol00b, Gou04b, Gre04b, JI04, Kle03b, MN06, McC01a, Nis07, Och01, Pic01, SS02, Sha00c, Soc08, Sti03, Sto06a, Wes00, Wes04, Zei01, Zel01, Zel02, All04, Cas04a, Fra03, Vog04, Wei03]. Nationalism [Bar03a, Car02, Zai08]. Nationalsozialismus [Sto07a]. Nationalsozialistische [Has06]. Nations [For07, Lan03b, Ham06a, Lan01b]. Native [Gor06]. Natur [Hey01]. natura [Gag07]. Natural [Ang05a, Arn01, Bar00a, Jud03, Kar09a, Kid08, Kir00, Lon04a, Sha01b, Tra02, Wor01, Ang05a, Hal07, HP00, Kli04, Sch02a]. Nature [Bar07, Clo03, Jud03, Kel06b, Kou01a, Lan04a, Mor06a, Nad07, Nye04, Ols07, Ots03, Per06b, Pic00, Ryd01, Sch05c, Sch03, Ser02c, Shu09b, Sol03, Ste07b, Tha04, Tod09, Uek07, Vol03a, Har07a, Jud05, O’H09, Rob06b, Vol04b]. Natures [Kar09a]. Naturschutz [Vee01]. nätverken [Har06a]. Navajo [Kir08]. Naval [Fel07, Hac05a, Hig03, Mad06, Min03, Smi05b, McB02]. navarro [San03]. Navigation [Sch05a, Rey01a]. Navy [Hou03, McB01, McB04, McB06, Sch01a, Sch02c, Vin03, Wol08, Yea04, Ash09]. Nazi [Hey08]. Nazis [Ols07]. Near [Coo01a, Kol07, McC01b, Sla01]. Nearest [Mou08]. Nebraska [Mil01]. Nederland [Mok01, Wac05]. Nederlands [Gwy09]. Need [Pat03, Wal09]. Needham [Per06a, Bra04]. Needle [Ben08]. Needs [Gre05]. Nef [Ros03]. Negotiating [Gag02, Hex07]. Nehru [Tya04]. Neil [Neu06a, HV07]. Neither [Bar03c]. nel [Lon05b]. nell’Italia [Gag03]. Nelson [Goo09]. Nelson-Atkins [Goo09]. Neo [HV06]. Neo-Baroque [HV06]. Neptune [Don06]. Nernst [Bey00]. Nerves [Cav08a]. Nervous [Goo07]. Net [KM04]. Netherlands [Div03b, Sto09b, Lin02b, RV04, TeB02, dWvdESvO02]. Network [Bar08, Bri07, Dow03b, Mag05b, Mom05, Sch06b, Smu07, Soc08, Tak00, Wei04, Dow04]. Networked [Lig02, Kno07a]. Networking [Kid03, Ste07a, Hår08a, Isr03]. Networks [Elk02, Gle03, Yar09a]. Netzwerke [Wen09]. Neuerer [Pet04c]. Never [Ros05a]. Newcastle [Gwy04]. Newfoundland [Fal06b, Sve03]. News [Bar05, Bri08, Kie02]. Newspapers [Bar05]. Newspeak [Bow04]. Newton [Ink06]. Next [Sch02a]. Nexus [Kle06c, Hug09b]. Niagara [Sti04a]. Nick [BP02]. Niedergang [Wen07, dS09]. Night [Bro07, Coo06, Cro06a, Lon01b, Rom03]. Nightmares [Bel01]. Nineteenth [Ade08, Bra08a, Dav00, Gam07, Gla01, Gre08b, HM01, Iri00, Iri01, Isr03, Jon00, Kha01, Kle06b, Kno06b, Lar03, Mie09, O’H09, Rob05b, Sim06, Tym00, Wal05, War07b, Lau08, Sch05a, Sto04b]. Nineteenth-Century [Ade08, Bra08a, Dav00, Gam07, Gla01, Iri00, Iri01, Kha01, Kle06b, Kno06b, Lar03, O’H09, Rob05b, Sim06, Wal05, War07b, Lau08, Sto04b]. Nino [Jon02b]. Nitrogen [vR07]. No [Ano09d, Ben05a, KO07, Wer03]. Nobel [Joh06, Str05]. Nodes [Dow04]. Nollet [War03]. nom [Ben03]. non [Pan00]. non-ferreux [Pan00]. Noncommercial [Slo08]. Nonferrous [Bur00b]. Nonweapons [Nis07]. nor [Bar03c]. Norbert [KM04]. nord [Bel04b]. Nordic [Bra08b]. Normalcy [Coo06]. normale [Heß04]. Normativities [BDD09]. Normativity [Dob07]. Norrkoping [Luc04]. North [Bek02, Con03, Gli00, Gre02, Guc08, Hub09, Ott08, Qui01, Lew01]. Northern [Cro07, Moo02a, Sha09b, Whi01, Bel04a, Boc09]. Note [Sta02]. Notes [Ano00h, Ano01e, Wil01, Cro02a]. Nothin [Gom00]. Nothing [Hof02]. nouveauté [Haf07]. nouveaux [Kah01]. novecento [Gag07]. Novel [Bra05]. Novelty [Lip00a]. November [Ano06e, Esb05]. Noxious [All07a]. Noyce [Hol06]. NSKK [Has06]. Nuclear [Bre01, Car03, Cer07, Dor08, Dur06, Ham07, Her05, Her08b, Jen01, Jos01, Lin06b, Luc05b, McB01, Mea05, O’H01a, Olw04, Olw09, Pic01, Pop00, Pow02, Sch09b, Sid02, Sil00, Sla02a, Wal01a, Wal01b, Wes01, Sch09b, Wal09]. Number [Ano00k, Ano00j, Ano00l, Ano00m, Ano01g, Ano01f, Ano01h, Ano01i, Ano02l, Ano02k, Ano02m, Ano02n, Ano03f, Ano03e, Ano03g, Ano03h, Ano04h, Ano04g, Ano04i, Ano04j, Ano05i, Ano05h, Ano05j, Ano05k, Ano06i, Ano06h, Ano06j, Ano06k, Ano07f, Ano07e, Ano07g, Ano07h, Ano08g, Ano08f, Ano08h, Ano08i, Ano09h, Ano09g, Ano09i, Ano09j]. Numbers [Tur09b, Wil00b]. Numeracy [Kid01]. Nursing [Mel01]. Nutrition [App03]. Nützliche [Har08b]. Nylon [Cla02a, Rob08a].

Oak [Mol06]. Oakland [Sha01a]. Oasis [Ell05]. Oaxaca [Moo02a]. Obduracy [Kon06]. Object [Ben05a, Jac04b]. Objectives [FF00]. Objects [Bla04b, Mai09, Tim09, Zac06, Dob04b]. Obra [Por05]. Observatories [San00]. Observatory [Smi05b, Jar00]. Observer [Dav09]. Obsession [Col08]. Obsolescence [Lai07, O’G06]. Ocean [Ash09, Bro01, Hoh03, Jer07, Jos08, Min03, Pyn09, Rit07, SHW03, Smi05b]. Oceanographers [Wei06]. Oceans [Olw09]. October [Ano00h, Ano02i, Ano04a, Ano05a, Ano07c, Ano08d, Ano09e, Cor06, Sch07a, Ano00m, Ano00n, Ano01i, Ano02n, Ano03h, Ano04j, Ano05k, Ano06k, Ano07h, Ano08i, Ano09j]. Odds [Bro08]. Odyssey [Fab06, Wak03]. Off [Coo05b, Mag00]. Office [Bed03, Ber07, Coo01a, Der04b, Gor01b, dWvdESvO02]. Officers [Ang01, Min03]. Offs [Kno07b]. Offshore [Vie08]. Ogdensburg [Gor05]. Oil [Bla04a, Coa02, Hin07a, Maz04b, Shu08, The08, Vie06, Vie08, Bla01, Feh04]. ökologische [Zel04]. Olanda [Squ00]. Old [Ben02, Mol07, Pur08, Aik03]. Olive [Maz04b]. Olkrise [Die01]. Once [Rol03]. One [Alt00, Car09, PB03, Rob08b, Sti04a, Glo02, McC09c, Sha01b]. Online [Bar05, CK04]. Only [Cer07]. Ontology [Wes08]. Oosterschelde [Bij02]. Open [Ens05, Fre07, Par08]. Open-End [Fre07]. Opening [Min00, Wil07]. Openness [Wak04a]. Operation [Jeo09]. Operational [Kri00b, Mor03]. Operations [Jeo08]. Oppenheimer [Bey09, Her07a, Olw01, Hec08]. Opportunities [Bij09, Gre05]. Opposition [Bre01]. Optic [Tag05]. Optics [Fin02, War01b]. Optimismus [Har00a]. Option [Bea02]. Orality [McG03]. Orange [Har07b]. Orbit [Clo06b, Log04]. Order [O’H09, Rey02b]. Ordnung [Boc03]. Oregon [Kli04, Mil03]. Organ [Hog03]. Organise [Lau02]. Organisms [Woo05]. Organismus [MW07]. Organization [Mac02]. Organizational [Ano00h, Ano01e, Kim04, Kra04]. Organizations [Ang01, Her05]. Organized [Las06]. organizzazioni [Vas08]. Organs [Rii03]. Orgasm [Wos00]. Orígenes [PP03]. Origin [Gre08a, Min00]. Original [Vol04b]. Originals [Pos05]. origines [Gou00]. Origins [Bar02, Bij04, Cha07, Goo09, Ham03, Her07b, Mac05b, Nad08, Pra01, Rom08, Rou05, Sch02c, Tur05, War02, Zei03, Kid03, Lon02b]. Orleans [Gor02b, Tha04]. Ornament [Por04]. Orville [Bed02]. Oscillator [Fro01a]. Oskar [Men07]. ostdeutschen [Sto05]. Other [Ash04, Ben05a, Ben05b, BP02, Dow08b, How00]. Otherness [Pat03]. Otherwise [Sha08]. Otis [Vog02]. Ottawa [Dur06]. Ottoman [Ayd05, Bek04a, Bek04c, Bek06, Bek00]. Our [Cow06, Fre01, Kon09, Law03, Pos04, Web08, Bak05]. Ourselves [Fan01]. Outer [Mar05a]. Oversights [Bak00]. Overture [KO09]. Own [Bak05, Jar00]. Ownership [Tur06]. Oxford [KO09].

P [Dic04, Fin05, Jon02a]. Pacemaker [Law03]. Pacific [Bek02, Bry07, Sch04d]. Page [Cav08b]. Paint [Aik03]. PAL [Sto05, Zet08]. PAL-SECAM-Kontroverse [Sto05, Zet08]. Palace [McE04]. Pallet [Ham05]. Panorama [Sch05b]. Paper [Bul03a, McG00b, Bul03a, Orr02]. Papermaking [Wak02]. Papers [Isr02, Pre05, Pre08a, Rot01]. Paradigm [Gwy04]. Paradise [Sol04, Bla09a]. Paradox [Fri07a]. Parallels [Clo07a]. Paris [Haf09, Pri04]. Park [Gre04c, Kid07, Lan09, McE04, Pre05, Pre08b, Rot01]. Parke [Cut02]. Parks [Bed05b, Bar07]. Parsifal [Cas08]. Parsimony [SHW03]. Parsons [Sil00]. Part [Bac06b]. Parting [Jon04c]. Parts [Bre00, Lon04b]. Passe [Rey01b]. Passenger [Sch02e]. Passengers [Sch02e]. Passes [Sha05a]. Passion [Sti04b, Vol05b]. Passione [Lap06]. Past [Bak05, Bij04, McC00b, McM06, Mor07a, Rus07, Cro06a]. Pastures [For05]. Patent [Coo01a, Kra07b, Swa09, Bra05]. Patent-System [Kra07b]. Patented [Fou06]. Patents [Bel04a, Hin08, Kra07b, Las06, Mac01, Coo07b]. Path [Mod08, Pan07]. Path-Breaking [Mod08]. Pathologies [Biv07]. Pathology [Ale06]. Paths [Arn00]. Patina [Dem03]. Patricia [Wes01]. patrie [Gou04c]. Patriot [Sha05a]. Patrons [Hor07]. Patterns [Div00]. Paul [Fis07]. Paving [Hol03, Tea09]. Payne [Mal09a]. Payson [Mol05]. Pazifist [Gis07]. Peace [Jon00, Nis07, Seg08, Wit08, Ham06a, Wan04]. Pearl [McB05, Nis08]. Peasants [Dan01, Pal00]. Pecheurs [Bor06a]. Pedagogy [Cru07]. Peerless [Gor03]. Penicillin [Tya04, Qui08]. Penn [Poh08]. Pennsylvania [Chu04b, Gam02, Kno06a, Lon07, Moh04]. People [Hei04, Jaf03, Jos07, Kir08, McE06, O’G05, Pat03, Sha06a, Ste08a, Wei04, Mea05]. Perception [Kat03, Par02a]. Perceptions [Jaf08]. Perfect [Coo05a, Daa07, Har07a, Owe04, Ash00]. Perfecting [Sch05d]. Perfection [Ten05a]. Performance [Con00b, LS00]. Performativity [LS00]. Performing [LS00, Nye06]. Peril [Bor08]. Perils [Sti08]. Period [Bek04b, Lew01, Pos04, Bra08b]. Periodicals [Rob05b]. Permanent [Sto03a]. Permissible [Fra02]. Perón [Cas08]. Perry [Bek02]. Persian [SS03]. Persistence [Gre06b]. Personal [Bar02, Epp04, Has02, Kan07]. Perspective [Bes08, McG00c, O’H03, Gui06, Ash05, Bel06, Dow08a, Les08a, Por04, Tho05]. Perspectives [Ada02, Don06, Gre08a, Joh05, Kin09a, KO08, Lak04, Mor07a, OS05, Big01, Cru07, Kee03, Mar02, Sal05]. Pesticides [And06b]. Pests [Kin09b]. Peter [Phi04]. Petitions [Lon02b]. Petroleum [Bla01, Pra03, The08, Vie08]. Petrolia [Shu08, Coa02]. Peugeot [Lau02]. Pflanzenzüchtung [MW07]. pflanzlichen [MW07]. Phanomen [Pet00]. Pharmaceutical [Bud09, Gre08c, Sli01]. Pharmacology [dlP07]. Pharmacopolitics [Bud06]. Phase [Dow04]. Philadelphia [Les04b, McB01, Roc02, Ale03, Ryd03]. Philanthropy [Las06]. Philippines [Biv07]. Phillips [Joh01]. Philosopher [Eis09]. Philosophical [Low04b, MB06, Dob08d]. philosophie [Kah01]. Philosophies [Wes07]. Philosophy [Cav05, Hil02, Kou03, McG06]. Phoenix [Epp06]. Phone [Aga09, Tri09]. Phonograph [Git01]. Photographic [Goo09, Gre04b, Pan05b]. Photographing [Wos02b]. Photographs [Ell01, Pan06]. Photography [Bro06c, Goo09, Lin00, Oli07, Rob06b, Wea02, Wri06, Bro05b, Poh00, Wri00]. Photoreconnaissance [Mei00b]. Physical [Bey00, Dow01, Sha00c, Spe04]. Physicians [How00]. Physicists [Bey08, War07b, Sid06a]. Physics [Bow00b, O’H01a, Pye01, Rau05, Wes00]. physique [Le 05, War03]. Picker [Lak01]. Pickfords [Lin00]. Picture [Cow03, Mal09a]. Picturephone [Lip03]. Pictures [Bro09b, Ham05]. Picturing [Bro06b, Pan06]. Pierre [Wak07]. Pigeons [Bri08]. Pill [Neu06b]. Pilots [Pal00]. Pines [Mad06]. Pioneer [Seg09, Wil00b]. Pioneering [Gui03]. Pioneers [Hor04a, Sel09]. Pipe [Luc04, Yar08]. Pipe-Bound [Luc04]. Piracy [Mac05b]. Pitfalls [MN06, vdV08]. Pittsburgh [Col04]. Pkw [Cie01]. Pkw-Bau [Cie01]. Place [Bek04b, Dem03, Dow07, Dyb09, Ede09, Kon09, Sti07]. Place-Based [Dyb09]. Places [Kat00, Str09a]. Placing [Yar09b]. Plain [Big08]. Plains [Fro09]. Plan [Iri08a, Sto09a]. Planet [For04]. Planets [Mar05a, Sad06]. Planning [Jac08a, Wer05b, Pow02]. Plans [BB00, Bro00c, Her01b, Iri03]. Plant [Pop00]. Plantation [Cha01, Lan02b, Por03]. Planter [Col07b]. Planting [Gud07]. Plastic [Fri01]. Plastik [Sud08]. Plate [Goo09, Les08b]. Platonic [Her01b]. Platz [Bak04]. Play [Mar00b, Ste01b]. Playback [Mag05a]. Playing [Bro08, Vol05b]. Please [Eri08b]. Plenary [Sch06b]. plip [Lev03a]. Plotting [Clo07a]. Plum [Her08b]. Plurality [Biv03]. Pocketbook [Jac07]. Poems [All03]. Poetics [Ben01c, Hor04b]. Poincare [Ste05b]. Point [Cut06, Nie07b]. Poised [Hut00]. Poison [Rol05, Olw09]. Poisoning [Aik03]. Polar [Osl07]. Pole [Gli00]. Polemical [Pow09]. Polemics [Iri08b]. Policy [Che04, Cla01, Jac05a, Jud05, Kle06c, Kli06, Par08, Pav09, Per06b, Rot08, Ser02b, Sha09a, Sha06c, Sla02a, Sti04b, Vie06, Bro01]. Politic [Goo07]. Political [Bar03c, Bre00, Cla07, Elk02, Ell05, Gag02, Kie03, Ots03, Rau03, Rou05, Sch05a, Slo02a, Ste08b, Wan07, Elk06]. Politics [Aik07, All03, All07a, App03, Bij02, Bis05, Ble07, Chu04a, Col07a, Con01b, Cow03, Dob07, Eri08b, Fie01, For04, Gar09, Gor04a, Hac09, Jac07, Kle06b, Log04, McB06, Mul05, Pre04b, Qui01, Rau03, Rau05, Sch00c, Sha01a, Slo02a, Tei01, The08, Vol03a, Wil00a, Wil03a, Wit07, Doe08c, Hac04]. Politik [Boc03, Sto01b]. politique [Lyn09, Zet08]. politisch [Sto05]. politisch-ideologische [Sto05]. Politische [Zet08, Sud08]. Pollution [Pra03, Qui03a, Gor04a, Qui01]. Polytechnique [Pan05a]. Ponts [Gra08]. Pope [Epp00]. Popular [Car02, Git01, Goo04, Kle03b, Sid04, Sca08]. Popularizing [Dow04]. Population [Wal06a]. Pork [Tei01]. Port [vDS05]. Portable [Hin09]. Portals [Bri08]. Portrait [Hul09, Wil00b]. Portraits [Mea05]. Possessed [Bek04c]. Possessors [Bek04c]. Possibilities [Big00, Kra05b]. Possible [You01]. Post [And06b, Bed03, Coo06, Kol06, McC09b]. Post-Apocalypse [Kol06]. Post-Cold [McC09b]. Post-Soviet [Coo06]. Post-World [And06b]. Postal [Ade08, Kie01]. Postcolonial [Van08]. Postcolonialities [Shu09a]. Posthuman [Bas07a, Pic00]. Postindustrial [Gro09]. Postmodern [Kra05b, O’G04b, Sch08a]. Postwar [Ber01, Bud05, Edg08, Ham08a, Kin09a, Koi02, Kra07a, Lai03b, Sto01a, Ten05b, Vie08]. Potato [Cro02a]. Poultry [Boy01]. Pound [Cut09b]. pour [Gou04c]. Poverty [Pan06]. Powell [Ell01]. Power [Arn00, Bek06, Bel08, Bow05, Bre01, Cut02, De 06b, Hac05a, Hal07, Hig03, Hin09, HÖ08, Iri08b, Jen01, Jon09, Kra07b, Kra09, Leo01, Lic02, Lin06b, Moh08, Mor05b, Muk03, Nad04a, Nie02, Pic01, Pop00, Sch07b, Sha07, Sti09, Tea08, Wal06a, Wel08, Wol08, Ash06, Car03, Flo07, Mac05b, Pra01]. Practicable [Sla01]. Practical [Con99, Lon04b, Ink06]. Practice [Bro00c, Cas07, Cru07, Fie01, Iri00, Muk06, Der04b, Tod08]. Practices [Fit00, Moh08, Tim09, Wen01]. Practitioner [Swa09]. Prague [Hut00, Jon01]. Prairie [And06a]. pratique [Lau02]. Pratt [Cox04]. Pre [Wal06a, Lyn09]. Pre-industrial [Wal06a]. pré-industrielle [Lyn09]. Precision [Das08, Sla01, War01b, War02, Zim04]. Predecessors [Tri06b]. Predict [Kri00a]. Prefabricated [Har06b]. Prefiguring [Abb04]. Preindustrial [WH09]. Premodern [Jer00, Leo01]. Prepared [Cer07]. Prescott [Owe04]. Prescription [Sli09]. Present [Gre04c, McM06, Hea01, Olw04, Sta01]. Preservation [Ald06, Bar04a, Iri08b, Mil08, Chu07]. Preserving [Ald06, Nic02]. President [Hec09]. Press [Pat03]. Pressey [Pet04d]. Pretty [Coo01b]. Preventive [For08]. Pride [Dav00]. Primer [PP09, Pet07b]. Print [Ben01c, Bul03a, How08a, Lin06a, Pan06, Rei05, Sch01d]. Printed [McG03, Rob05a]. Printing [Lon02b, Lon03a, McG03, How06]. Printmaker [Lon01a]. Privacy [Hum09, Mno00, Day04b]. Private [Div00, Lan07b, Sha07, Sti03]. Privatisation [Edw03b]. Privilege [Rey02b]. Probability [Con06a]. Probable [Edw03a]. Probe [Mod06b]. Problem [Aik07, Daw07b, Kle09]. Problematic [Per06a]. Problems [Bug07, Hug09a, Jac04a, Pop01b, Sch04c, Sha06a]. Process [Nic02, vDS05, Bas02]. Processing [KO07]. Product [Col07a]. Production [Big00, Boy01, Bur02, Dob07, Epp03, Ham09, Jaf05, Jos01, Lip00a, Maz04b, Rap01, Vol03b, Wal06a, Fox00, Lan01b, Lic01, San06]. Productions [Sim09]. Productivity [McG00c]. Products [Mil09, Nis07]. Profane [Vas01]. Profession [Fox02, Iri00]. Professional [Rau01, Swa09]. Profit [Dow05]. Profits [Vie06]. Program [Her08a, Her08b, Jen01, Sat07, Eze02, Hou03, Joh02c]. Programming [Mor05a]. Programs [Mac05a]. Progress [Col07b, Ink09, Lev05, MS08, Str01, Wil03b, Gol00c, Jon04a]. Progressives [Gor01c]. Progressivism [Kli08]. Project [Lev00, Gam07, Gui03]. Projectile [Hac03]. Projection [Clo07a]. Projects [Reu08]. Projekt [Gis01]. Proliferation [Wal01a]. Promethean [Har07a]. Prometheus [McC00a]. Promise [Bor08, Coo04b, Fri01, Kle03b, Tat05]. Promises [See08, vdV08]. Proof [Ash01, Uss04]. Propaganda [McB06]. Property [Bra05, Mac00]. Prophet [Swe08]. Propulsion [Sat08]. Prosopography [MN06]. Prosthetic [Bin06]. Prosthetics [Lon04b]. Protection [Fra02]. Protoindustrial [Rav01]. Provided [Har03]. Providence [SS07]. Provoked [Bri07]. Prozesse [Wen09]. Prüfstand [Wen07]. Psychiatry [Gre08c]. Pt [Sch02b]. Ptolemy [Kar02]. pubblico [Gag03]. Public [Ash01, Cha02, Dav06, Doe08b, Epp07, Gro09, Iri01, Jac05a, Kle09, Kle03b, Lan07b, Lon02b, O’G03, Per06b, Rey02b, Rot08, Sat04, Sha09a, Sid04, Smu07, Sti03, Tho05, Wan07, dV02, Gol00a, Sha07]. publique [Pan03a, Poi08]. Published [Bed02]. Publishers [How08b]. Publishing [De 06b]. puissance [Lan01b]. Pull [Ste07a]. Pump [Ham05]. Purgatory [AA01]. Purity [Gra08]. Pursell [Cer09]. Pursuit [Lev08, WW04]. Pushing [Vol07a]. Put [O’H09]. Putting [Pos09a]. Pyramid [Edw03a]. Pyrénées [Bla03]. Pyrotechnics [Buc05a].

Qaeda [Ang04]. Quality [Gor01c, Sla01]. Quantum [Pye01]. Quebec [Kra04, Win02]. Quest [Har07a, Hin07a, Kon05, Log04, Mid03, Moñ07, Sei03, Wos04]. Question [McB07, San06]. Questioning [Har00b]. Questions [BP02, Mis07, Poi08]. Quick [Ben00].

R [Coo02c, Iri05]. R&D [Fri02b, Kri00b]. Race [Bai01, Biv07, Bro05b, Cow04a, Gor02a, Kin09a, Kno01b, McG00c, Mor00, Rud09, Sid06b, Sto04b, Vee03, dlP06, Coo02b]. Rachel [Kar08, Kar09b]. Racial [All07a, Biv07]. Radar [Blu01]. Radiance [Pic01]. Radiant [Spi04]. Radiation [Fra02]. Radical [vDS05]. Radio [Cas08, Coo02c, Cra07, Doe08b, Doe08a, Har03, New06, Pat03, Pet00, Slo02a, Slo08, Soc08, Sto09a, Tak00, Car02, Rus08, Smu05, Smu07, Whi02]. Radio- [Pet00]. Radioactive [Olw09]. Rae [Luc09]. RAF [Hol01]. Rahmi [BE04]. Rail [Edw03b, Sma05, Gam05]. Railroad [Bry07, Dav06, Gam02, Gor02a, Gre02, Hof02, Kle06b, Lar04, Luc07, Men09, Pos03, Sch06d, Sch02e, Hof02]. Railroads [Ang01, Chu04a, Chu07, Con08, Her00, Mag09, Mil02, Mol00b, Rot08, Sto03c, Wat00, Jon03, Ott08, Uss03]. Rails [Dav06, Luc07, Mom05, Mil03]. Railway [Els09, Gwy09, Hoa01, Vee05, Whi05, Zei01, Zei03, Zel01, Zel02, dO01b, Pop01b, Bek04a, Sch06e]. Railways [Buc05b, Cro08a, Gei00, Gor01a, Hoo04, Mid05a, Mie09, Nis07, Vee03, Alv04, Div03a, Div03b, Gre01]. Rain [All07b]. Raising [Dir07, War09a]. Raj [Col05]. Raketenindustrie [Neu02]. RAND [Low04a, Lig06]. Range [Mom05, Yea04]. Rapid [Hoo04]. Rasmussen [Dal04]. Rational [Con99, Lip02]. Rationality [Har01]. Rationalization [Bro06c]. Rationalizing [Ber07]. Ray [How00, dlP06]. Raymond [Kul03, Blu04]. Rayner [Cro05]. RCA [Cho07]. Re [Kat00, Lon03a]. Re-Visioning [Kat00]. Reach [Bug05]. Reader [Rot04c]. Reading [Wis09b]. Reagan [Her01a]. Real [Dow06, Orv09, Web08]. Real-time [Dow06]. Realism [Dow08b]. Realities [Ber07]. Reality [Kou01a]. Realization [FF00]. Realizing [Con07]. Reaping [Lan01a, Pas02]. Reason [CK01, Gin07, Gre08c, Tho06, Ale03]. Reassessing [Epp07]. Rebellion [Lan07b]. Rebirth [You01]. Rebuilding [Ash00]. Recasting [Gor02a, Pie06, Pre06]. Receiver [Tak00]. Reception [PB03]. recherche [Pan03a]. Recognizing [Jac04a]. Reconfiguring [Ots03]. Reconsidering [Mar01]. Reconstructing [Dyb09]. Reconstruction [Cru03, Ham08a, Gor01a]. Record [Hei04, Lau08, Mag00, Pos09b]. Recorded [Ras01, Git01]. Recording [Hor06, Mag00, Kra07a]. Records [Hor04b]. Recovering [Hen09]. Recycling [Ros07, Pas02]. Red [Bar03c, Sid03, Jen01]. Redefining [Pra03]. Redesigning [Lew01, Bij03]. Reeds [Big08]. Reengineering [Wes07]. Reference [McS02, Gui06]. Reflections [Cer05, War09b]. Reflects [Sil00]. Reform [Bri06, McI00, Hal07]. Refrigeration [Ree05b]. Refrigerator [Nic02]. regard [Haf07]. regelskapande [Jør09]. Régent [Lyn09]. Regime [Ben02].

Regimes [Kle06a]. Region [Les01, Lon07, Col04]. Regional [Les01, Sch05e]. Regnault [Wri06]. Regulating [Lar04, Wal01b]. Regulation [Bud06, Fri07b, Pav09, Rey02b, Whi02]. Regulator [And05]. Regulatory [Pop00, Pra03]. Rehabilitated [Ste01a]. Reich [Zel01, Zel02, Fer05, Ols07, Str09b]. Reign [Bac06b]. Reigns [Bac04]. Reinforced [See03]. Reingold [Rot05]. Reinterpretation [Jos05]. Reinterpreting [Ede09]. Reinvent [Bor06b]. Reinvented [Gwy09]. Reinventing [Nye04]. Reinvention [Hog03]. Reisman [Bek07]. Rejoinder [Cla02b]. Relation [Clo07b]. Relations [Kry04, Ten00]. Relationship [Pos01a]. Relationships [Bud09, Rus07]. Relays [Kie01]. Reliability [RK08]. Reliable [Con99, Scr00]. Reliably [Con00b]. Religion [Nye06, Wal02]. Religious [McC07a]. Religiously [HV07]. Reluctant [Din08]. Remains [Low08a]. Remaking [Bla05, Hac05a, Iri08a, Lan08]. Remarkable [Fro05]. Remington [Lig06]. Removal [Bar09b, Luc04]. Renaissance [Fri08, HV01, Jan09, Kel06a, Kel09, Kon05, Lon01a, May09, Maz01, Bla08, Coo07b, Mir08, Ner08, Ros06, Wak04a]. Renewal [Col04]. Repair [Cio06, McI00]. Repeating [Gor07]. Replaceable [Ten05b]. Replicating [Les01]. Reply [BP02]. Reporting [Bri08, Dow06, Lau08]. Representation [Dob08c, Low04c, Mae09, San06, Skr00, Bel08]. Representations [Cor02a, McG00b, Rob05b, Sim09]. Representing [Pre02]. Reproduction [Bij04, Bul01, Kat05]. Reproductive [Bul03b, How09a, Ram05]. Republic [Bra08a, Kel05, Mul07, Pre04b, Wer06, Coo00, Ryd01]. Republics [Sla09]. Rescuing [McC00a]. Research [Cer04, Cho07, Hun00, Las03, Las06, Lev00, Lig06, O’G08, Pre05, Smi03, Ste06, Ste01b, WF03, Wen01, vR07, Hof00, Kid08, Mod08, Moo04b]. Researchers [Mac01]. Researching [Asn03]. réseaus [Gui06]. Reservations [Tom04]. Reserve [The08]. Reshaping [Tea05]. Residual [Mul08]. Resistant [Wer07a]. Resource [Shu08]. Resources [Reu02, Sha01a, Kle06a]. Respect [Vol04b]. Respond [Pre06]. Responsibility [Olw01, Web08]. ressources [Bel04b]. Restoration [Pan02a, Cio06]. Restructuring [For03, Pra01]. Resurrection [McB05]. Retail [Ake05]. Retailing [Jaf08]. Rethinking [Gor06, Mal09b, Min00, Fis09, Par09]. Retooling [Cas03]. Retrieval [Kha01]. Reuleaux [Kel09]. Reusable [Log04, Rol01]. review [Abb03, Abb04, Abb05, Abe02, AA01, Ada06, Ada02, Ade08, Aga09, Aik03, Aik07, Ake05, Ako06, Ako07, Ala09, Alc07, Ald00, Ale03, All03, All05, All06b, All07b, All07a, Alt00, Alv04, And00, And06a, And06b, And09, Ang04, Ang05b, Ang05a, Ang06, App03, Arn00, Arn01, Arn03, Ash04, Ash05, Ash06, Ash08, Ash09, Ayd05, Baa01, Bac05, Bac06a, Bac07, Bac08, Bac02, Bai01, Bak00, Bak04, Bak05, Bal00, Bar03a, Bar00a, Bar02, Bar03b, Bar05, Bar08, Bar09a, Bar07, Bar03c, Bar04b, Bar09b, Bar00c, Bas00, Bas02, Bas05, Bas07a, Bas09a, Bas07b, Bas09c, Bea02, Bed02, Bed03, Bed05a, Bek02, Bek03, Bek04a, Bek04b, Bek04c, Bek05, Bek06, Bel06, Bel08, Bel04b, BM07, Ben00, Ben01a, Ben01d, Ben01b, Ben01c, Ben02, Ben03, Ben05a, Ben05b, Ben08]. review [Ben09, Ber03a, Ber01, Ber02, Ber03b, Ber04, Ber06, Bey00, Bey06, Bey07, Bey08, Bey09, BS09, Big00, Big01, Big02, Big07, Bij03, Bij04, Bij05, Bil02, Bis05, Biv03, Biv07, Bix04, Bla01, Bla04a, Bla09a, Bla09b, Bla08, Bla03, Bla09c, Bla00a, Bla00b, Bla04b, Bla05, Ble04, Ble07, Blu04, Blu01, Boc03, Bor06a, Bor00, Bor02, Bor06b, Bor08, Bor09b, Bor09a, Bow00a, Bow00b, Bow03, Bow08, Bow00c, Bow04, Bow05, Bow07, Boy03, Bra01a, Bra02, Bra03a, Bra08a, Bra08b, Bra05, Bra01b, Bra03b, Bra03c, Bra07b, Bra07a, Bra08c, Bra04, Bre00, Bre01, Bri06, Bri01, Bri08, Bro01, Bro00a, Bro08, Bro00b, Bro02, Bro05a, Bro05b, Bro05c, Bro06c, Bro06a, Bro06b, Bro07, Bro09a, Bro09b, Bru01, BW04, Bry07, Buc01a, Buc01b, Buc03a, Buc03b, Buc05b, Buc05a, Bud02]. review [Bud06, Bud07, Bud09, Bug01, Bug04, Bug05, Bug07, Bul03a, Bul01, Bul03b, Bur02, Bur00a, Bye04, CK01, CK04, Car08, Car02, Car03, Car09, Cas08, Cas07, Cav05, Cav08a, Cep08, Cer03, Cer04, Cer06, Cer07, Cha01, Cha02, Cha07, Che04, Chu04a, Chu07, Cie01, Cio06, Cla01, Cla07, Cla02a, Cli00, Clo03, Clo05, Clo06a, Clo06b, Clo07a, Coa02, Coc03, Coc04, Col05, Col07a, Col08, Col03, Col01, Col07c, Col07b, Col04, Con08, Con00a, Con01a, Con02, Con03, Con07, Coo07a, Coo04a, Coo05a, Coo00, Coo01a, Coo02b, Coo02a, Coo04b, Coo07b, Coo02c, Coo03b, Coo05b, Coo09, Coo05c, Cor02a, Cor08, Cor09, Cor02b, Cow03, Cow04a, Cow04b, Cow06, Cox04, Cra07, Cro05, Cro08a, Cro02a, Cro06a, Cro06b, Cro08b, Cro02b, Cro06c, Cru03, Cru07, Cul09, CH05]. review [Cus05, Cut02, Cut06, Cut09a, Daa07, Dal02, Dal05, Dal06a, Dal06b, Dan03, Dan01, Das08, Dav00, Dav01, Dav06, Dav09, Daw06, Daw07a, Day01, Day04a, Day04b, Day06, Day08, De 00, De 03a, De 06a, DeV00b, DeV03, De 06b, Dea07, Dea08, Dem03, Der04a, Der04b, Dev00a, Dic04, Dic05, Die01, Dir05, Dir07, Dis05, Div00, Div03a, Div03b, Div09, Dob08a, Dob08b, Dob04a, Dob04b, Dob06, Dob07, Dob08d, Doe08b, Doe08a, Doe08c, Doi06, Dol09, Don06, Don07, Dor00, Dor01, Dor08, Dou01, Dou09, Dow03a, Dow03b, Dow05, Dow07, Dow08a, Dow08b, Dow09, Dug08, Dug09, Dur06, Ede09, Edg08, Edw03b, Eel09, Elk02, Elk06, Ell01, Ell05, Els09, End01, End00, Ens05, Ens06, Ens07, Epp02a, Epp03, Epp06, Epp09, Eri07, Eri08a, Eri08b, Eze02, Fab06, Fal06a, Fal06b]. review [Fan01, Far00, Feh04, Fei09, Fel07, Fer05, Fie01, Fin00, Fin02, Fin03, Fis09, Fis03, Fit00, Fit02, Fit01, Flo00, Flo01, Flo05, Flo07, For03, For05, For08, For04, Fos04, Fos09, Fou06, Fox00, Fox02, Fra02, Fra06, Fra08, Fra09, Fre01, Fre06, Fri09, Fri01, Fri07a, Fri08, Fri02b, Fri07b, Fro09, Fro01b, Fro05, Gag03, Gag07, Gal00, Gam02, Gam05, Gam07, Gan03, Gan05, Gar09, Gei00, Gel04, Geo06, Gie02, Gif07, Gil06, Gin07, Gis00, Gis01, Gis07, Git01, Gla01, Gle03, Gle08, Gli00, Glo02, Gof09, Gol00c, Gol00a, Gom00, Gom05, Goo07, Goo06, Goo04, Goo09, Gor01a, Gor01b, Gor02a, Gor05, Gor06, Gor07, Gor08, Gor01c, Gor02b, Gor04b, Gor04a, Gou00, Gou01, Gou04a, Gou04b, Gou04c, Gra01, Gre05, Gre04a, Gre08a, Gre08b, Gre08c, Gre03]. review [Gre01, Gre02, Gre04b, Gro08, Gro09, Guc08, Gud07, Gui06, Gui00, Gui03, Hår08a, Hac05a, Hac03, Hac04, Hac05b, Hac06, Hac07, Hac09, Haf07, Haf09, Hai03, Hal07, Hal01, Ham07, Ham08a, Ham06b, Ham08b, Ham09, Ham03, Har00a, Har06a, Har07a, Har02, Har01, Har05, Har06b, Har07b, Har00b, Har04, Har08b, Has06, Has02, HM01, Hea01, Hea02, Hec09, HV01, HV06, Hen09, Her01a, Her05, Her07a, Her00, Her07b, Her08b, Her09a, Her01b, Her01c, Hes04, Heß04, Hex07, Hey01, Hey06, Hey08, Hig01, Hig03, HP00, Hil02, Hin07a, Hin07b, Hin08, Hoc03, Hoc05, Hof00, Hof02, Hog03, Hoh03, Hok07, Hol06, Hol01, Hoo04, Hor04b, Hor04a, Hor06, Hor02, Hou03, Hou04a, Hov04, How06, How08a, How08b, How09a, How09b, How00, Hub09, Hug00b]. review [Hug02, Hug05, Hug09b, Hum09, Hun00, Hun02, Hur02, Hur04, Hyd00, Hyl00, Igl01, Ink06, Ink09, Iri00, Iri01, Iri03, Iri05, Iri06b, Iri06a, Iri06c, Iri08a, Isr02, Isr03, Isr05, Jac05a, Jac05b, Jac00, Jac02, Jac04b, Jac06, Jac07, Jaf03, Jaf05, Jaf06, Jaf08, Jag00, Jan09, Jar00, Jen01, Jen05, Jen08, Jeo08, Jeo09, Jer00, Jer07, Joh00a, Joh08, Joh00b, Joh02c, Joh02a, Joh02b, Joh04, Joh06, JI04, Jon00, Jon02b, Jon02a, Jon03, Jon04c, Jon04a, Jon06, Jon09, Jør09, Jos01, Jos03, Jos04, Jos05, Jos07, Jud03, Jud05, Jud08, Kah01, Kar02, Kar09a, Kar08, Kar09b, Kat00, Kat02, Kat03, Kat05, Kāt07, Kat09, Kee03, Kel05, Kel04, Kel06a, Kel09, Ker06, Kha01, Kha08, Kid01, Kid03, Kid04, Kid06, Kid07, Kid08, Kie01, Kie02, Kie09, Kie03]. review [Kim04, Kin09b, Kin00, Kir00, Kir08, Kle03a, Kle06a, Kle03b, Kle06b, Kle08, Kle04, Kle06c, Kli00, Kli04, Kno01a, Kno03, Kno05b, Kno05a, Kno06a, Kno06b, Kno07b, Kno07a, Kol07, Kol08, KO02, KO07, Kon05, Kon06, Kon09, Kou01b, Kou01a, Kou02, Kou03, Kou06, Kou07, Kra00, Kra01, Kra05a, Kra07a, Kra09, Kri00a, Kri03, Kru02, Kru04, Kru08, Kru09, Kuw00, Lac03, Lai03b, Lai07, Lak01, Lak04, Lak05, Lan08, Lan02a, Lan02b, Lan03a, Lan03b, Lan04a, Lan07a, Lan07b, Lan01b, Lan04b, Lan03c, Lan00, Lan09, Lap06, Lap09, Lar08, Lar03, Las06, Lau02, Law00, Law03, Le 05, Le 07, dSL04, Leo01, Les08a, Les04a, Les04b, Lev01, Lev03a, Lev08, Lev00, Lev03b, Lev05, Lew00, Lic01, Lic02, Lie01, Lig02, Lig06, Lig09, Lin05, Lin06a, Lin06b, Lin02a, Lip00a]. review [Lip02, Lip06, Lit03, Log04, Lon01b, Lon01a, Lon02b, Lon02a, Lon03b, Lon04a, Lon05b, Lon06, Lon07, Lon04b, Lon00, Low01a, Low04a, Low05a, Low00a, Low02a, Low04b, Low04c, Low01b, Low08b, Luc02, Luc03, Luc04, Luc06, Luc07, Lue05, Lug03, Lun09, Lyn09, Mac09, Mac00, Mac01, Mac02, Mac05b, Mac05a, Mac04, Mad06, Mae09, Mag09, Mag00, Mag01, Mag05b, Mag05a, Mah02, Mai04, Mai07, Mai09, Mal01, Mal09b, Mar01, Mar05a, Mar00a, Mar00b, Mar09, Mar03, Mar06a, Mar06b, Mar02, Mar05b, Mat02, May09, Maz03, Maz01, Maz04a, Maz07, McB00, McB01, McB02, McB04, McB05, McB06, McB07, McC08a, McC07a, McC09a, McC00a, McC00b, McC03, McC06, McC07b, McC08b, McC09c, McC01b, McE03, McE04, McE06, McE07, McG00a, McG00b, McG02, McG03, McG05, McG06, McG00c]. review [McM00, McM06, McN02, McS02, McS00, Mea05, Mei07, Mei00b, Mei00a, Mei02, Mel01, Mel06, Men07, Men09, Mer05, MS08, MA09, Mid03, Mid05a, Mid05b, Mid06, Mie04, Mie09, Mil04a, Mil01, Mil02, Mil03, Mil04c, Mil04b, Mil06, Mil08, Mil09, Min03, Mir08, Mis02, Mis07, Mno00, Mod06a, Mod08, Moh04, Moh08, Mol00a, Mol03, Mol04, Mol06, Mol07, Mol08, Mol00b, Mom02, Mom04, Moñ07, Moo00, Moo02b, Moo02a, Moo04b, Mor01, Mor05a, Mor07a, Mor08a, Mor03, Mor05b, Mor07b, Mor06a, Mor00, Mor06b, Mor08b, Mos02, Mos06, Mou08, Moy06, Mui03, Mul01a, Mul04, Mul05, Mul07, Mul08, MW07, Mul01b, Nad04a, Nad04b, Nad06, Nad08, Nag08, Nak07, Neb01, Nei02, Nei05, Nei07, Nel04, Ner08, Neu02, Neu06a, Neu01, Neu06b, New00, New06, Nie02, Nie07a]. review [Nie07b, Nis08, Nix09, Nor09, Nye04, Nye05, Nye06, O’D03, O’G03, O’G04b, O’G04a, O’G05, O’H01a, O’H01b, O’H03, O’H05, O’H07, O’H09, Och01, Och02, Ols07, Olw01, Olw04, Olw09, Orl02, Orr02, Orv09, Osi00, Osl04, Ots03, Ott08, Owe03, Pac02, Pal02, Pan00, Pan03a, Pan03b, Pan07, Pan09, Pan02b, Pan05b, Pan06, Par02a, Par09, Par08, Pas02, Pas06, Pay05, Per07, Per00, Per02, Per06b, Pet01a, Pet03, Pet04b, Pet04a, Pet04c, Pet05, Pet07a, Pet07b, Pet00, Pet01c, Pfa04, Pfa06, Phi03, Pic00, Pic06, Pic01, Pie06, Pla06, Poh08, Poh00, Poh07, Poi08, Poo03, Pop00, Pop01a, Pop02, Pop07, Por04, Por05, Pow02, Pow04, Pow09, Pra01, Pra03, Pre04a, Pre01, Pre02, Pre03, Pre04c, Pre04b, Pre05, Pre06, Pre08a, Pre08b, Pri03, Pri00]. review [Pri02, Pri04, PP03, PP09, Pur08, Pye01, Pyn09, Qui08, Qui01, Qui03a, Qui03b, Qui05, Rab04, Ram05, Ran01, Rap01, Ras01, Rau01, Rau03, Rau05, Rav01, Ree05a, Rei05, Reu05, Rey02a, Rey05, Rey01a, Rey07, Ric08a, Ric08b, Rie05, Rii03, Rit07, Rob00, Rob06a, Rob08a, Rob02, Rob06b, Rob06c, Rob05b, Roc02, Rol00a, Rol00b, Rol01, Rol05, Rom08, Rom03, Ros07, Ros06, Ros05b, Rot04a, Rot08, Rot01, Rot04b, Rot04c, Rou02, Rou05, Roz09, Rud09, Rus00, Rus07, Rus09, Rus08, Ryd03, Ryd01, RP09, Sad06, Sal01, Sal05, Sam05, Sam07, San00, San02, San06, San03, Sat07, Sat08, Sat04, Sau05, SS00, SS02, SS03, Sca05, Sca08, Sch02a, Sch08a, Sch00a, Sch04b, Sch09a, Sch02b, Sch09b, Sch05b, Sch06c, Sch05c, Sch01a, Sch02c]. review [Sch00b, Sch02d, Sch03, Sch04c, Sch07b, Sch01b, Sch04d, Sch05d, Sch06d, Sch08c, Sch02e, Sch05e, Sch06e, Sch07c, Sch01c, Sch01d, Sch06f, Sch07d, Sch00d, Scr05, See01, See03, Seg00, Seg08, Sei03, Sel09, Ser02a, Ser02b, Ser02c, Ser05, Sew06, Sha00a, Sha06a, Sha00b, Sha01a, Sha07, Sha09a, Sha00c, Sha04, Sha05a, Sha06c, Sha09b, Sha05b, Sha01b, Sha06e, Sho02, Sho05, Shu02, Shu08, Shu09a, Sid02, Sid04, Sid06b, Sid06a, Sid07, Sid08, Sil00, Sim03, Sim06, Sim09, Sin06, Sin09, Sin00, Sin05a, Sin05b, Skr00, Sla04, Sla02a, Sla09, Sli01, Sli09, Slo02a, Slo04, Slo09, Sma05, Smi02, Smi03, Smi05a, Smi05b, Smi06, Smi07, Smu05, Smu07, Sol03, Sol04, Spa01, Spa02, Spe04, Spi00a, Spi01, Spi00b, Spo02, Squ00, Sta00, Sta01, Sta06a, Sta06b]. review [Ste06, Ste07a, Ste01a, Ste05a, Ste01c, Ste05b, Ste07b, Ste08a, Ste08b, Sti04a, Sti08, Sti09, Sto01a, Sto01b, Sto05, Sto06a, Sto07a, Sto03a, Sto01c, Sto02, Sto03b, Sto04a, Sto07b, Sto09b, SW07, Sto03c, Str09a, Str05, Str09b, Str01, Sud08, Sve03, Swe08, Tak01, Tap02, Tat05, Tea05, Tea09, Tei01, Tei05, Ten00, Ten05a, Ten05b, Ter00, Tes09, Tha04, The08, Thi06, Tho06, Tho01, Tho05, Tho07, Tho08, Tho09b, Tho09c, Tim09, Tod02, Tod08, Tod09, Tom03, Tom04, Tom05, Tow02, Tra02, Tra05, Tri06a, Tri09, Tri04, Tri06b, Tri03, Tur06, Tur08, Tur09b, Tur09a, Twa00, Twe05, Tym00, Uek07, Ung05, Uss03, Uss04, Vai00, Van00a, Van00b, Vas08, Vas01, Vee00, Vee01, Vee03, Vee05, Ven00, Ver00, Vie06, Vie08, Vin03, Vin06]. review [Vog02, Vol01, Vol03a, Vol03b, Vol04a, Vol04b, Vol05a, Vol05c, Vol05b, Vol07a, Vol07b, Vol08a, Vol09a, Vol09b, Waj00, Wak02, Wak03, Wak04a, Wak07, Wak05, Wal02, Wal05, Wal01a, Wal09, Wal01c, Wal00, Wal04, Wal06a, Wal06b, Wan02, Wan04, Wan07, Wan08, War04, War06a, War05a, War01a, War01b, War02, War03, War05b, War06b, War07a, War07b, War08, War09b, War09a, War09c, WF03, Wat00, Wat02, Wea02, Web08, Wei01, Wei04, Wei06, Wel02, Wel08, Wen01, Wen02, Wen07, Wen08, Wen09, Wer02, Wer04, Wer05b, Wer05a, Wer07a, Wer07b, Wes08, Wes00, Wes01, Wes03, Wes04, Wha06, Whe00, Whi01, Whi02, Whi04, WH09, Wil00a, Wil00b, Wil03a, Wil04, Wil09, Win05, Win08, Win02, Wis09a, Wit07, Wit08, Wol08, Woo05, Wor01, Wos00, Wos02a, Wos04, Wri00]. review [Wri04, Wri06, Wu03, Wya04, Yal08, Yar07, Yar08, Yar09a, Yar09b, Yat01, You01, You09, Zac01, Zac03, Zac06, Zei03, Zel01, Zel02, Zel04, Zel06, Zem02, Zet08, Zim04, dS09, dlP07, dO01a, dO01b, dO05, vD03, vD09, vL09, vF05]. Revisionary [Fin00]. Revisions [Iri06a, Wri00]. Revival [Van08]. Revolt [Kli08]. Revolution [All06a, And09, Big02, Bow08, CK01, Coo03b, Fei09, Fin00, Gor02a, Gor07, Hun00, JI04, Kno06b, Luc05a, Mor05a, Mor08a, O’G04b, Pan07, Per00, See07, Vol09b, Wak02, Wil00c, Wit08, Zim04, Bor06a, Gou00, Pan09, Mil04c, Ash08, Bar00a, Ben08, Cla02a, Cow04b, Dea08, Fro05, Gou04a, Gra01, Har01, Hoc05, Hoh03, How09a, Jen05, Jos05, Kid01, Lar08, Lin05, Qui03b, Rey05, Sca05, Sch02a, Sch00d, Ste07b, Wer05a]. Revolutionaries [Mol03]. Revolutionary [Coo05c, Vol03a]. Revolutionizing [Sli01]. Revolutions [Big01, Kra09]. Rewards [Sti08]. Reworking [Sch03, Sto01a]. Reyner [Kat02]. Rhetoric [Bul01]. Rhine [Dis05]. Rhodes [Lon09]. Rhumb [Clo07a]. Rice [Ham03]. Richard [Buc03b]. Rickenbacker [McE07]. Rickover [Mui03]. Ride [Epp02a]. Ridge [Mol06]. Riding [War04]. Rife [Wes01]. Rifles [Gor07]. Rinascimento [Lon05b]. Ring [Mor05b, Sto09a]. Rio [Fra06]. Rise [Bar08, Big07, Bix04, Bla05, Cer04, Cut06, Dav06, Dir05, Ens07, Epp06, Fer05, Hig03, Joh06, Kar08, Kid08, Kie03, Lev03b, Lic01, May09, Mol07, Mor05a, Mor08b, Rab04, Rus08, Sch02d, Sha06e, Sla02b, Spo02, Str01, Vog02, Vol08a, War06b, Wer03]. Risk [Luc05b, Pri00, Uss04, Wil03a, Kno07b, Lip02]. Risks [Lin06b]. Rituals [Van02]. Rivalry [Vee03, Bro09a]. Rivals [Nor07]. River [Dur06, Ell05, Har04, HÖ08, Luc06, Sha09a, Sti04a, Tha04, You01, All06a, Bro05c, Reu05]. Rivers [Mid06, Bro03, Mal09b]. Rivertown [Mal09b]. Road [Baa01, Cer03, End01, Gre02, Tea09, Wel07]. Roads [Bar07, Mar00a, Mom05, Nor09]. Roadside [Bel01]. Roam [O’H06]. Roanoke [Pos03]. Robert [Bro00b, Hea02, Her07a, Hol06, Epp07, Fin05, Hec08, Tea08, Nel04]. Roberts [Buc03b]. RoboCop [Ver08]. Robot [Kou01b, Nad08]. Robots [Ala09].

Rock [Hor04b, How09a, Kno05a]. Rockefeller [Sta05]. Rocket [Gam05, Sel09, Sid03]. Rocketry [Log04]. Rockets [Sto02]. Rod [Bor09a]. Rode [Luc07]. Roger [Seg09]. Rogers [See07, Sin09]. Rogue [New06]. roi [Lan04b]. Roland [Rob00]. Role [Con99, vR07]. Roles [Sha00a]. Roller [War04]. Romae [KO02]. Roman [vD03, KO09]. Romance [Ber04]. Romanticism [Van08]. Rome [KO07, KO08, Lan02a]. Röntgenstrahlen [Bey07]. Rooms [Coo01a]. Root [Wu03]. Roots [Ble04, Rey02b]. Rosalind [Cas03]. Rose [Lip00b]. Rotterdam [vDS05]. Round [Bil02, Tea08]. Round-the-World [Bil02]. Roundhouse [Hyd01]. Route [Vol07b]. Routes [Cep08]. Royal [Hun02]. Ruben [Hin09]. Rufiji [HÖ08]. Ruin [Cut06]. Ruins [Jon06, Kru04, Pay05]. Rule [Bis05]. Rules [Dic04]. Ruling [Wu03]. Rum [Rab04]. Rumor [Coo03a, Lau08]. Rundfunkindustrie [Pet00]. Running [Gam02]. Rural [Fit02, For05]. rurales [Ben03]. Russia [Jen01, Lan00, Sid04, Jos01, Sid08]. Russian [Coo06, Gor03, Sid02]. russischen [Pet04c]. Rust [Ede09]. Rustbelt [Les04a]. Rusting [Col06]. Rustungsindustrie [Zac03]. Ruth [Par05]. Ruttman [Pos05]. RWE [Sto01b]. RWE-Geschichte [Sto01b].

S [Hou04b]. Sacred [Vas01]. Sacrifice [Gin07]. Safe [Sha05b]. Safeguards [Wal01b]. Safety [Bla09c, Col07a, Con01a, Iri01, Luc07, Ree05b, Wos06]. SAGE [Fri02b]. Sail [Lan03c]. Sails [McG05]. Sale [Mil09, Sti08]. Salisbury [Rob02]. Salpetriere [Pan05b]. Salvaging [McB05]. Sam [Jon03, Sch07b]. Samuel [Isr05, Owe04, Hin09, Joh01, Smi07]. San [Wer07a, Ano02i]. Sands [Che04]. Sanitary [Dob08a, Owe04, Sta01]. Sanitation [Jør08]. Santé [Poi08]. Sarah [Rob07]. sardinière [Bor06a]. Satellite [Day06, Kri00b, Pav09, Slo02b, Mar01]. Satellites [Clo06b, Sid06b, Spi00a]. Satisfaction [Wos00]. Sausage [Lan07b]. Save [How00]. Saved [Wel02]. Savines [Ben03]. Sawmill [Gre04d]. Saxon [Fis03]. Say [Pav09]. Scale [Fab06, Vol05b]. Scandinavia [Jør08]. Schädlinge [Har08b]. Schatten [Joh02a]. Schedule [Pri00]. Scheitern [Gal00]. Schemes [Fro01a]. Scholars [Dym08]. Scholarship [Yar09b]. School [Iri08b, Och01, Pan02a]. Schoolhouse [OS05]. Schumpeter [Swe08]. Schwab [Sch07c]. Schwartz [Par05]. Schweiz [Har08b]. Schweizer [Vee01]. Science [All02, Arn01, Ash09, Bac05, Bas07a, Bas09a, Biv07, Boc09, Bow00b, Boy01, Bra04, Bro00a, Bud07, Car09, Cas07, Cli00, Clo07b, Coo06, Cro05, Cru03, Cru07, Cut09a, Dio08, Dow07, Eze02, Fis09, Gan00, Gie02, Gin07, Gol00b, Goo07, Hac04, Hac09, Ham07, Ham08a, Hec09, Hen09, Hoh03, Hug00a, Ink06, Jen08, Joh02b, Jos04, Kle06c, Kli06, Kou07, Kru08, Lev00, Low01a, Low04c, Mai09, Mal09a, Maz04a, McC09a, McE04, Mir08, Mol02, Moo02a, Moo04b, Mor08a, Nad04b, Nad08, Nye06, O’H05, Pac02, Per06b, Por03, Pre03, Rey02a, Ric08b, Rus00, Sal05, Sch07a, Sch01c, Shu02, Sid04, Sid06a, Smi03, Sol03, Ste06, Sto03b, Sto04a, Tei01, Wal02, Wal01c, Wal06b, Wes08, Wes04, Wil03a, Gou04c, Le 07, Pan03a, Dow04, Ayd05, Bey00, Coo02a, Dor00, Dou01].

Science [Her08b, Hey06, Hug00b, McC07b, New00, Par08, Rau01, Rau03, Rob06a, Rob06b, Rob05b, Sam07, Sch02b, Sch06f, Sid04, Sti08, Sto01c, Sto07b, Wal01c, War06b, Wei06]. Science-Industry [Kle06c]. Sciences [Lyn09, Rob05b, De 00, Gou04c, Shu09a]. Scientific [Ash08, Big07, Cav05, Cow04b, Dea08, Har02, Hec08, Kuw00, Lev08, Lin05, Lin00, Sch05c, Ste07b, War06b, Yea04]. scientifica [Gag03]. Scientist [Bri01, Olw01, Wak07]. Scientists [Bro00a, Hac09, Kid08, Min03, Rus00, Vin03, Cru03, Wes03]. scienza [Gag03, Lon05b]. Scope [CH05]. Scorpion [Rol05]. SCOT [BP02, Cla02c, Epp02b]. Scotland [Ash05]. Scottish [Hal01, How08b]. Scrap [Ros07]. Screw [DO03, SD09, Sha01b]. Screwdriver [Sha01b]. Script [Ben01c]. Scripts [Pre02]. Sea [Dic05, Fel07, Leo01, Wol08, Jer07, Roz09, Vin06]. Seagoing [McM00]. Seamanship [McM00]. Seamstresses [Ben02]. Search [Koi02, Rol01, Vie08]. Searching [Reu06, Hum09]. Seating [Ale06]. Seattle [Ste08a, Jud08]. SECAM [Sto05, Zet08]. secoli [Dol09]. Second [Ber08, Bro00a, Coc03, Esb05, Hoc05, Lak01, Mar06a, Pic00, Sch06b]. Secrecy [Wak04a, Dea07]. Secret [Geo06, Mac05a, McE04, Day04a, Sam07]. Secrets [HV07, Kid07, Lap09, Mac05b, McC00b]. Sector [Leo01]. Secular [Nad04a]. Security [Coo09, Epp04, Hug09b, The08, Vie06]. Seduced [Lap09]. Seeing [Reu08, Wer07b, Mor01]. Segregation [Pre04c]. Seguin [Gou00, Pan09]. Seicento [Lon02a]. Seine [Pri04]. Seismicity [Wit07]. Selbstverstandlichkeit [Heß04]. Selected [Mil08]. Self [Cro08b]. Self-Control [Cro08b]. Selling [Bud07, Fro01a, Lan03b, Lug03, Mei07, Bix04, Gla01]. Semantics [Old06]. Semiconductor [Vol01]. Sennacherib [DO03]. Sense [Big02, Jac04b, Sti07]. Senses [Kat03]. Sensibilities [Ber02]. Sentenced [Car09]. Separating [Pow09]. September [Wos02b]. Serafina [KO09]. Serialized [Rob05b]. Series [OS05]. Seriously [Lig08]. Service [Doe08b, Fox00, Ink06, Lan04b]. Services [CK04, Kra09]. ses [Pan09]. Set [Gam02]. seta [Lon02a]. Settlement [BS09]. Seven [Wak03]. Seven-Year [Wak03]. Seventeenth [Arn01, Muk06]. Seventeenth-Century [Arn01, Muk06]. Seventh [Jon01]. Sewage [Dob08b]. Sewerage [Luc04]. Sewn [Thi06]. Sex [Gel04]. Sexual [Wos00]. Shad [Sti04b]. Shadow [Hec09, Olw01, Sam07, San06]. Shadows [Bro05c]. Shaffers [Pos03]. Shanghai [Rei05]. Shanly [Buc01a]. Shape [Bow07, Ros05b]. Shaped [Coo04b, Hea01, Lak01, Luc06]. Shaping [Bal00, Hår08a, Mar00a, McS00, Wer06]. Share [Ake01]. Sheeler [Poh07]. Shelby [Pre04c]. Shell [Vie08]. Shelter [Glo02]. Shifting [Joh05]. Ship [Don07, Wit08, Thi06]. Shipbuilding [Buc03a, Don07, McG06]. Ships [McG02, McM00, Sve02, McG06]. Shivers [RP09]. Shock [Cav08a, Pur08, Bri06]. Shockley [Ens07]. Shop [Coo07a]. Shore [Fal06a, Fal06b]. Shore-Station [Fal06b]. Shoreham [Pop00]. Short [McG05]. Shorthorn [Ten05a]. SHOT [Bij09, Hug09a, OS05, Pos01a, Sta09]. SHOT/AHA [OS05]. Shots [Bra08c]. Should [Moy02, Sch07e, PB03].

Shouldering [Lin06b]. Shows [Nye06]. Shu [Ham05]. Shuji [Bow08]. Shuttle [Rol01, Pri00, Van00a]. Sibling [Bro09a]. Sick [Aik07]. Side [Sie09]. Sidewalks [Vog04]. Sidewinder [Hig01]. Sidney [Pet04d]. siècle [Gou01, Pan03b, Bor06a, Fro01b, Lev01, Poi08]. siècles [Bla03, Gui06]. Siege [Nis08, Whe00]. Siegeszug [Mom04]. Siemens [Gis00]. Sierra [Moo02a]. Sigfried [Mol02]. Siglo [PP09]. Signalman [Pop01b]. Significance [Rus09]. Signifying [Old06]. Signor [Fro05]. Silent [Kar08, Ser02c]. Silicon [Bas07c, Col03, Goo06, Hol06, Low05a, Les01, Pic00]. Silk [Gro08, Mac04, Maz01, vF05]. Silver [Bug04]. Silvio [Pos08b]. Simon [Tho06, Wak07]. since [And09, Coo07b, Cro08b, Fel07, Gle03, Joh00a, Mar01, Mul07, Pfa04, Pur08, Ste07a, Ten05a, Vie06, War05a]. Sing [Bin06, Tho01]. Singer [Phi04]. Single [Log04]. Sino [Kry04]. Sir [Hac05a, Fin03]. sistemi [Vas08]. Sisters [Gor01b]. Sites [Wen01]. Sits [Bor00]. Situating [Bra04]. Situation [Les08b]. Sixteenth [Fel07]. Sixth [SD01]. Skies [Jeo09]. Skill [Sto04b, Par02a]. Skule [Gie02]. Sky [Jon04b, Spi00a, De 06a, Kno05b, Smi05b]. Skyline [Iri02]. Skyscraper [Iri06b, Les08b]. Slave [Egl07, Mie04, Sha05a]. Sleepwalking [Bla09a]. Sloan [Dic04, Dic04]. Sludge [RV04]. Sluice [vD02]. Small [Sve02]. Smallholder [Sto06b]. Smelter [Qui01]. Smelting [For09, Lon03a]. Smil [Shu09b]. Smithson [Mor08a]. Smithsonian [Pos01a, Sha06d, Mor08a, Skr00]. Smoke [Les04a, Qui01, Qui03a, Win05]. Smokeless [Gor03]. Smokestacks [Gor01c]. Snapshot [Oli07]. Soaring [Mil03]. Social [Bra08a, Car08, Clo07a, Dob08b, Fie01, Fis09, Fit02, Gol00a, Hac09, Har03, Her00, Joh00b, LS00, Lin02a, Lip03, Mal09a, Mar00a, Mol06, Nic02, Rol03, Sla01, Sve02, Ten00, Wan02, Wri04, Sch01d, Vol04b]. Socialism [All03, Zac01]. Socialist [Zac02]. société [Bla03]. Societies [Jac00, Lan00, Lan03b, Moo00]. Society [Big02, Coo02a, Dan01, Dow08a, Gui08, Hun02, Jac05b, Jon00, Lan04a, Low04b, OS05, Pre03, Sch09a, Sch06f, Sti09, Tes09, Tho07, Tod02, Wak05, Wan08, Dob07, Mar00a, Nix09, Sch01c, Shu09b]. Sociology [Har02]. Sociotechnical [Kon06]. Södertälje [SW07]. Soft [Bar00a, Sha05a]. Software [Tom04, Rus09]. Solar [San00, Ser02a, Ser02b]. Soldiers [Hac06]. Solid [Wil00c]. Solomon [Hun02]. Solve [Sha06a]. Solved [Rot04b]. Solvents [Jac04a]. Some [Kel06b]. Sonic [Tom04, Hor04a]. Sophocles [Reu06]. Sorgel [Gis01]. Sorting [Sim06, Low01b]. Soul [Chu07, Moo04a, HV07]. Sound [Bij04, Bro07, Kra05a, Mag00, Per00, Mor06b]. Sounds [Bra03c, Hor06, Mor08b]. Soundscape [Arn03]. Source [Ens05]. South [Bek05, Fal06a, Fis07, Gud07, Hac05b, Hin07a, Lak04, Mar09, Mid06, Sha08, And06b, Lak01, Mad06]. Southeast [Tag05]. Southern [Bar09b, Bry07, Coc03, Cox04, Eel09, Gor01a, Gud07, Lak04, Mar09, Sto04b]. Soviet [Bai01, Bow04, Goo06, Jen05, Jen08, Jos05, Lun09, Mar01, Sid04, Sid06a, Coo06, Hug00a, Jen00, Pal00, Sid03]. Sowjetischen [Zac03, Neu02]. Sozial [Ber03a]. Sozial- [Ber03a]. Soziale [Wen09]. Space [All04, And00, Arn00, Bai01, Bek04b, Cro06c, Day04a, Day06, Day08, Dou04, Fos09, Her07b, Her09a, Jeo09, Joh02c, Joh04, Kra00, Kra01, Les04b, Log04, Mac05a, Mol02, Mul09, O’G03, Pri00, Rol01, Sad06, Sat05, Sat07, Sat08, Ser02a, Sid06b, Sim09, Van00a, War01a, Kim04, Seg00, Sto02]. Space-Launch [Sat08]. Spacecraft [Sid07]. Spacefaring [Cro02b]. Spaceflight [Mac06, McC01b, Hun00]. Spaces [Dow01, Jon06, Kat00]. Spain [MB06]. Spanish [Cha01, Sol03, Ste07b]. Spans [Mil03, Mil01]. Sparked [Fro05]. Sparrows [Cut06, Nie07b]. Spartanburg [Mar09]. spaten [Low00a]. Spatial [Dow07, Har03, Yar09b]. Speaking [Vai00]. Special [Bas05, Hal01, Pos01a]. Specialty [Lip00a]. Specification [Yea04]. Spectacles [Ner08]. Spectrum [War01b]. Speed [Bor09a, Flo07, Gam05, Gui08, Per00, Pos09b, Daw06]. Spending [Jac00]. Sphere [Lon02b]. Spheres [SS02]. Spied [Goo06]. Spillman [Dal06b]. Spindletop [Bla04a]. Spine [RP09]. Spinning [Cor02a, Fre07]. Spiral [Cav08b]. Spirit [Fou06, O’D03]. Spirited [Spe04]. Splintering [Lig02]. Spoorwegmuseum [Gwy09]. Sprawl [Sch02d, Tea05, Bas07b]. Sprawling [Str09a]. Spreadsheets [Kid04]. Spring [Kar08, Ser02c]. Sputnik [Hec09, Hun00, Mar01]. Spy [Lap09, Spi00a]. Spy-Tech [Lap09]. Spying [Day04b, Day06, Ric08b]. Square [Ham05]. Squaw [Bra07a]. SS [Mie04, Ric08a]. St [Mil03, Pre06, Sch06e]. Staat [Sto05]. Stability [Mid03]. Stadia [Lie01]. Stadt [Die01]. Stage [Log04]. Stagecoach [Whi01]. Stalin [Goo06, Jen01, Zac03, Jen08, Neu02, Sid06a]. Standard [Gag02, Abb03, Jac06]. Standardization [Bar00c, WW04]. Standards [Ash01]. Stanford [Low05a]. Stanley [Hok07]. Star [Her01a]. Stargazer [McC06]. Starling [Coh07]. Stars [Wan02, War01a]. Start [Cer04]. Start-Up [Cer04]. Starting [Lie01]. Stasi [Lap09]. State [Bac08, Col07c, Dic05, Div00, Gol00b, Hac05a, Hac07, Hum09, Kle06b, Low04a, Mou08, Muk03, Rau01, Sid04, Slo08, Sto07b, Sve03, Tea09, Wan08, Wes03, Kle06a, Ram05, Ven00]. States [Aik03, Bro01, Bud06, Bur00a, Coo04b, Cro08b, Div03a, Fei09, Hea01, Jos01, Kli04, Lit03, Mal01, Mis02, Moo02b, Mor05a, Nor09, Par08, Pow02, Sch01a, Slo02a, Vie06, Wel08, Wer05b, Whi02, Bas09b, Bri07, Bro00c, Eck05, Hol03, Soc08, Spi04, Bro02, Bye04, Col07a, Hyd00, Kle03b, Lan07a, Low01a, Sol03]. Statik [Pet01a]. Station [Fal06b, Her08b, Poh08, Chu04b]. Stationary [Gre04b]. Stations [Fal06a]. Statistique [Pan03b]. Stealth [Hou03]. Steam [Bro09c, Coo05a, Gam05, Gre04b, Hoa01, Hok07, Hul03, Low04b, Pos03, Rob04, Sch01b, Sch05d, Mal01]. Steamboat [Sha09a, Whi01]. Steamboats [Bro03]. Steamer [Hok07]. Steamers [SS07]. Steamship [Bro07, SHW03, Thi06]. Steel [All03, Cut06, Har05, Lan09, Mis02, Moo00, Nie07b, SS03, See01, BS09, Nie07a]. Steelmaking [Nie07b]. Stem [Mod08]. Stenographers [Dow06]. Stenography [Bar09a]. Stepchildren [Her08a]. Stephenson [Nel04]. Stereoscopic [Rom08]. Sterling [Gor05]. Stern [Coo02c]. Steuerung [Wen09]. Steve [Coo04c]. Steven [Lar04, All04]. Stevensons [Buc01b].

Stewart [Bar08]. Stick [Jos08, Mok05]. Still [Geo06, Uss06]. Stimulated [Sch08b]. Stirrups [Rol03]. Stock [GC07]. Stone [Eis09, Oli07, Vas01]. Stone-Age [Oli07]. Stood [Geo06]. Stop [Dow09]. Stores [Mad06]. Storia [Vas08]. Storie [Gag07]. Stories [Bla04b, Clo07a, Cut07, Her09b, LS00]. Storm [Bij02, Bug04]. Story [Ash04, Car09, Cla02a, Day04a, Der04a, Ens06, Eri07, Fin02, Fis07, Fri02b, Gam07, Gie02, Jeo09, McB00, McC09b, Nie07a, O’H07, See01, Ser02a, Spi00a, Ste05a, Tri06a, Tri03, Wel02, Bed05a, Dug08, Elk06]. Strahlenschutzes [Hes04]. Straits [Pos08a]. Strange [Mai04, Bra03c]. Stranger [All02]. Strasse [Sch05b]. Strategic [Eck05, Epp00, Hug09b, Sid02, The08, Wol08, Gre03, Sei03]. Strategies [Fro01a]. Strategy [Hol01]. Street [How09b, McE04, Nor07, Tap02, Whi05, Reu06]. Streets [Jør08, Mol00b]. Streit [Wen02]. Strength [Zai08]. Strictly [Wil01]. Strike [Mol00b]. Strong [Par02b, DvdE03]. Structural [Pet04b]. Structure [Moy05]. Structuring [Lip06]. Struggle [Haf03, Jos07, Kid06, Mui03]. Struggling [Les04a]. STS [Pre03, MB06]. Studies [Bac05, Bar00c, Coo06, Har02, How08a, Kli08, Mal09a, Moo04b, Nad04b, Rot04a, Pre03, Cut09a, Sal01]. Study [Baa01, Bro09a, Fal06a, Las06, Lig06, McG06, Sch05e, Wei04, Gre05]. Stuff [Ben05b, Con99, Scr00]. Sturdy [Low05b]. Style [Bla05, Pop07, Win08]. Styles [Bro00c]. Styx [McC08a]. Subject [Ano00i, Ano02j]. Subjects [Cow03]. Sublime [Bow05]. Submarine [Lon00]. Submarines [Mei07, Wei03]. submersion [Ben03]. Substance [Bla05, Pop07]. Subterranean [Qui03b]. Subtitling [Bar09a]. Suburban [Sch02d, Har05]. Suburbanizing [Tho05]. Suburbia [Hen09]. Suburbs [Mei07]. Subversive [Kar08]. Subway [Tho07, Wel02]. Subways [Mid05b]. Success [Elk06, Ens05, Pet01b, Ste07a, Tri03, Fri07a]. Succor [Stu07]. sud [Bor06a]. sueurs [Haf09]. Suez [Jon04c]. Suffering [Gin07]. Sugar [Cha01, Lan02b, War09a, Dan01, Wat00]. Suitable [Sla02b]. Sultan [Bek06, Bek00]. sum [Low00a]. Sumer [Kid04, Dal05]. Summer [Jon01, Van02]. Sun [Pas06, San00]. Sunburst [Hig03]. Sunset [Bry07]. suo [Twa00]. Superhighways [Mar00a]. Superman [Bas05]. Supersonic [Daw06]. Supple [Low05b]. Supplement [Ano00n, Ano02o]. Supply [Din08, Luc04, Ste04]. Surface [Sch00c]. Surge [Bij02]. Surgeons [CH05]. Surgical [Bac02, Law00]. Surprising [Mil04a]. Surveillance [Day04b, Hum09, Sha05a]. Survey [Luc05a, Sch08a]. Surveyed [Dor01]. Surveying [Cro07, Kra05b, Lan02a]. Survival [Kru04]. Survived [O’H07]. Suspension [Mid03]. Sustainable [Cep08]. Svante [Hyl00]. Sverige [Jør09]. Sweden [Lan00, Luc04, Ber07, SW07]. Swedish [Twe05]. Sweringen [Mil04b]. Swift [Dow05]. Swimming [O’G06]. Swindon [Hoa01]. Swiss [Els09, Pet04b]. Switch [Coo05b]. Switching [Baa01]. Swivel [Gwy04]. Symbol [Mil04c]. Symbolgeschichte [Osi00]. Symposia [HW05]. Symposium [Jon01, SD01]. Synchronous [Soc08]. Syndrome [Aik07]. Synthese [Joh02a]. Synthesis [Biv03]. Synthesizer [Bra03b].

Synthetic [For04]. System [Coo02b, Dob08c, Kai02, Kra07b, Mis08, Pav09, Rot08, Wes04, WW00, Kie01, Pra01]. Systematic [Ber07]. Systematically [GS02]. Systems [GS02, Jac08a, Kai02, Mac05a, Par08, Qui03b, Sat05, Day06, Dow08b].

T [Bac02, Eri07, Kli08, Pre04c, Tho09c, Wel07]. T&C [Ano09d]. T-Mails [Eri07]. Tables [Kid04]. Tacit [Muk06]. Tacoma [Mid03]. Tactical [Mor03]. Tail [Ham05]. Tailings [Che04]. Take [Gui03]. Taken [Pan07]. Takes [Mol02]. Taking [GC07, Lig08, Web08, Bre00, Dow07, Eze02, Tri04]. taktische [Wal04]. Tale [Con06b, Fab06]. Talents [Sla09]. Talk [Bul01]. Talkie [Per00]. Talking [Doe08a, Gom00]. Taming [Joh01, TeB02, Wu03]. Tanavoli [SS03]. Tank [Sha08]. Tanks [Moy06, Van08]. Tanzania [HÖ08]. Tappan [Jar00]. Tapping [Mad06]. Target [Con00b, Scr00, Sil00]. Tarnished [Bac02]. Taste [Lan03b]. Taylor [See07]. Taylorism [Kan07, Kan07]. taylorisme [Lau02]. Taylorized [Big07]. Teaching [Ber08, Wis09a]. Tech [Lap09, Sha04, Vol08a]. Techne [Sch08a]. Technical [Bas09b, Bel04a, Ber07, Bix02, Blu01, Bug01, Cra07, Cul09, Eck05, Egl07, Ham09, Har03, HPV06, Hut00, Joh05, Pet07b, Sch00c, Sve02, Tya04, Wak04a, Wer07b, Kru02]. Technics [Wil02a]. Techniek [Mok01, Wac05]. Technik [Bey06, Fre06, Har00a, Osi00, Sch06a, Sto01b, Zet08, Wen08]. Technikgeschichte [Wen02]. Techniktheorien [Bak04]. Technique [Muk06, Bla03, Gou01, Ben01c, Lev03a]. Techniques [Ben01b, Fit00, De 00, Lyn09]. Technische [Lac03]. technischen [Has02]. Technischer [Gis07, Gal00, Wen09]. technisches [Pet00]. Techno [Bis05, Eis09, Lig09, Zai08]. Techno-Analogy [Eis09]. Techno-Cities [Lig09]. Techno-nationalism [Zai08]. Techno-Politics [Bis05]. Technoculture [Sch08a]. Technografie [Wen08]. Technographies [Cut07]. Technological [Abb03, Bac08, Bro05c, Cer05, Daw07b, Fan01, Gor06, JI04, Kno07b, Kru02, Lar08, Lig02, Mor03, Mor07b, Mor08b, Pra03, Scr09, See07, Sha06a, Tho00, Tri03, Van08, Wak05, Wil02b, vD02, Bur00b, Bak00, Bas02, Big01, Ros05b, Sch01a, Wan08]. Technologies [Bla08, CK01, Coo01b, Cut07, Daw07a, Day04b, Dow01, Lai03a, Les08a, Mei02, MK01, Neu01, New00, Sta05, Van02, dWvdESvO02, Epp04, Bul03b, Cut02, Pas02, Ros05b, Sin05b]. Technologietransfer [Neu02]. technologique [Gil06]. Technology [Aga09, Ald06, AKP09, Ang05b, Asn03, Ayd05, Baa01, Bac05, Bar03a, Bar00b, Bar00c, Bas09b, Bas09c, Bek03, Ben01c, BB00, Bij02, Boy01, Boy03, Bra01a, Bra03c, Bra07b, Bro00a, BDD09, Bul01, Cer05, Cer09, Con06b, Coo02b, Coo02a, Coo06, Cor08, Cut09a, Dan01, Dio08, Dob04a, Dob07, Don06, Dor07, Dow07, Epp03, Eri08a, Far00, Fie01, Fis09, Fit02, Fre01, Fre07, Gar09, Gwy04, Hac03, Her01c, Hey06, Hoc03, HW05, Hug00a, Hug00b, Iri01, Jud03, Kat03, Kel06b, Kle06a, Kle06b, Kli06, Kno01b, Kno06b, Kno07a, Koi02, KO08, KO09, Kou03, Kou07, Kra05a, Kra05b, Lai07, Lew00, Lig06, Lip06, Log04, Lon05a, Low04c, Low00b, Low02b, Lym03, Mag00, Mal01, Mal09a, Mar06a, May00, McB00, McC08b, Mel01, MB06, Mok01, Mol04, Moñ07, Moo04b, Nag08].

Technology [Neu06b, Nis07, Nye05, O’H04, O’H05, Old06, OS05, Pac02, Par08, Pat03, Pet07b, Pfa04, PB03, Por03, Pre02, Pre03, Pur08, RV04, Ree05a, Reu02, Rob08b, Rol03, Rot02, Rot04c, Sat08, Sau05, Sch06a, Sch08a, Sch05a, Sch02b, Sch05e, Sch01c, Sch06f, Ser02b, Sha06a, Sha08, Sha06c, Sho02, Shu02, Sid03, Sid07, Smi02, Ste05a, Sti04b, Sto01a, Sto06a, Sto01c, Sto03b, Tho01, Twe05, Uss09, Van08, Vol01, Vol05c, Waj00, Whi02, Wil00c, Wos00, Zac01, ZH00, dlP06, vD03, vdV08, Jon01, LS00, Pos01a, SD01, Ada02, Bak05, Bas00, Bas07a, Bow08, Bra02, Bri08, Cas07, Cer04, Cla01, Coo05b, Edg08, Eel09, Gre05, Ham06b, Har00b, Hil02, Hou04a, Ink09, Iri06c, Joh00a, Joh02b, Kou01a, Kou02, Kou06, Luc03, Mai07, Maz07, Mis07, Mis08]. Technology [Mno00, Mod06a, Mul05, Nad06, Nag08, O’D03, Rud09, Sal01, Sch06c, Sch07d, Scr05, Sto09b, Tat05, Van00b, Ver00, Vol03b, Vol04a, Wil09, Wit08, Wri04]. Technomanifestos [Mol03]. Technophobia [Bas07a]. Technopolitics [Daw06]. Technoscience [Aik07]. técnica [PP03]. Tecnología [PP09, dSL04]. tecnologico [Dol09]. Teknikens [Hyl00]. teknikhistorisk [Hyl00]. tekniska [Har06a]. Telecom [Mos02]. Telecommunications [Baa01, Mos02, Bar00c, Kid03, Slo09]. Telefunken [Gis07]. Telegraph [Bek02, Eri07, Gor01b, Hug05, Sch08b, Hoc03]. Telegraphic [Gor01b, Dow08b]. Telegraphy [Bek00, Bek03, Yea04]. Telepistemology [Kou01b]. Telerobotics [Kou01b]. Telescope [McC01a, McC08b, McC06, War07a]. Telescopes [Dea07, Tat05, Ner08]. Television [Lai03b, Mor01, Mul04, Mul05, Tak00, Whi02, Bar09a, Bro09b, Mag05b]. Televisual [Clo06b]. Telling [Sha08]. tempo [Twa00]. Temporality [Bra08a, Smi06]. temps [Bla03]. Ten [Rap01]. Tender [Pan02b]. Tennessee [Mol06]. Tension [MK01]. Tensions [Kra05b, Sch06b]. Tercio [PP09]. Terman [Low05a]. Terminal [Mil02]. Terms [Mul06]. Territories [Dow05]. Terror [Sid03, Jos04]. Terrorism [Wal01b]. Tesla [Cow04a]. Test [Bij02, Cer07, Ham07]. Test-Ban [Ham07]. Testimony [Sch05c]. Testing [Cut09b, Her07b, Joh04, Wol08]. Texas [Bla04a, Feh04]. Text [Bar09a, Ham09]. Textbooks [Mul01b]. Texten [Pet01a]. Textildruck [Low00a]. Textildruck- [Low00a]. Textile [Ble07, Con00a, Ran01]. Textiles [Tra02, Low05b]. Textilfarbeindustrie [Low00a]. Texts [Tim09]. TGV [Zei03]. TH [Bey06]. Thames [Buc03a, Luc06]. Theatre [Hal01]. Their [How08b, Law00, Mul01b, Sha00a, Sha00b, Smi02, Tri06b]. Themes [Hug09a]. Theology [McC09a]. Theoretical [Gra08, Kar02]. Theory [Der04b, Gra08, Nag08, Pin08, McG03]. Therapy [AST08]. There [Ben05a, Her01a, You09]. Theremin [Bra01b]. thermique [Fox00]. Thermonuclear [Kru08]. These [Pat03]. Thing [Her09b, Mou08, Cav05]. Things [Ben01a, Ben05b, Coo00, Jac04b, Jer00, Low01b, Mor01, Rap01, Sch09a, Wak03, Kat00]. Think [Con06b, Nye05]. Thinking [Bul03b, GS02, Sin00]. Third [Cor06, Fer05, Jon02b, Ols07, Rob06a, Str09b]. Thirteen [McG00c, Tes09]. Thirty [Jac08a]. Thomas [Cut02, Pre05, Pre08a, Pre08b, Rot01, Sch07b, Sch05e, GS02]. Those [O’H07]. Thought [Fri02a, McC09a, Sch09a, Arn00, McC07a, O’H03]. Thousand [Jac08a, Rap01, Rob08b]. Thread [Ble07]. Threat [Pri02]. Three [Clo05, Pos05, Wos02b]. Throwing [Hac03]. Tidal [DvdE03]. Tides [Ash09]. Ties [Bek04a, Elk02]. Tiger [Vol01]. Tignes [Ben03]. tillagnad [Hyl00]. Tilling [Ham05]. Time [Bra03a, Bra08a, Fra03, Gla01, Hof00, Les04b, McC09c, Mil01, Mol02, Mul09, O’H07, Pos04, Smi06, Ste01c, Wer07b, Dow06, Cow06, Mor00, O’H01b, Ste05b, Tow02]. Timekeeping [Gla01]. Times [Bur00a, Coo04a, Cro06a, Hea01, Jon02a, Kin09b, Mat02, McC06, McG00a, Sha00a, Sid06a, Sta01, Sve03, Nak07]. TIN [Wac05]. TIN-20 [Wac05]. Tinkerers [Tak00]. Tinkering [Bor06b]. Tit [Bra07a]. Tits [SW07]. Toasters [Ben05b]. Tobacco [Ren09]. Today [Jen01]. Toilets [Ben05b]. Tokugawa [Bek04b]. Toledo [Hyd01]. Tolls [Dyb09]. Tom [SW07]. Tomato [Lan04a]. Tomorrow [Sch08c]. Tonenden [Pet01c]. Tool [Gif07, Dob04b]. Tool-Being [Dob04b]. Tools [Dob08d, Gra08, Jac08a, Wis09a]. Tootin [Sch06e]. Topics [Str02]. Topografien [Boc03]. Toronto [Bal00, Gie02, Jud05]. Total [Nei02]. Tour [Mar05a]. Tourismus [Vee01]. Tourists [Fra06]. Tower [LW05, Der04b]. Towers [Flo01]. Town [Shu08, Lan09, Sew06]. Towns [Kel05, Ste08b, Car08]. Toxic [And06b]. Toys [Smi02]. trace [Pri04]. Traces [McC00b]. Track [Zei03]. Tracks [Hor04b]. Tracy [Moo04a]. Trade [Bas07c, Bro05a, Flo01, Jaf05, Kno07b, Sha04, War02, War09a, Mac05b]. Trade-Offs [Kno07b]. Trader [Ash01]. Trading [Cor09]. Tradition [Wal02]. Traditional [Mac04, Spi00b]. Traffic [Con01a, Cor06, Epp09, Esb05, NE09, Tho08]. Tragedies [End00]. Tragedy [Qui08]. Train [Gwy09, Sch08c, Vee03]. Training [Bix02, Kno01a]. Trains [Vol05b, Cro08a]. Traitteur [Pet04c]. Traktate [Wal04]. Transatlantic [Cio06]. Transborder [Qui01]. Transcontinental [Hof02]. Transcription [Dow06]. Transfer [Eck05, Far00, Fre07, Gwy04, Twe05, Vol03b, Moo04b]. Transformation [All06a, And06a, And09, Bri06, Don07, For05, Fro09, Hor02, Jud03, Pow09, Shu08, War09a, Mis08, Sha00b]. Transformations [Lan04a]. Transformed [Las03, Rob02, Tri06a, Wak03, Hea01]. Transforming [Les04b, Mei02, Mis08, Gor02b, Her00]. Transgressive [Str02]. Transistor [Pet00]. Transistorisierung [Pet00]. Transistors [Cho07]. Transit [Dyb09, Gre06a, Hoo04, Mid05b, Sch00c, Sch00b, Sma05, War09c]. Transition [Bey00, Dow06, Baa01]. Transitions [Dow04]. Translation [Kar02]. Transmitting [Mor07a]. Transnational [Hår08a, vdV08]. Transplants [Hog03]. Transport [Cor06, Div00, Div08, Esb05, Tho05, Wya04, dV02, Div09, Sma05]. Transportation [Ake05, Daw06, Dyb09, Fri07b, Rot08, Sch00c, See07, Tho08]. trasferimento [Dol09]. Trash [Ros07, Lin02a]. travail [Dav09]. Travel [Div09]. Traveler [Pre00]. Tredegar [Gwy04]. Triads [War07b]. Trial [O’H09]. Tricksters [Har02]. Trident [Wei04]. Tripp [Gre04d]. Tritium [Car03]. Triumph [Abb03, Dic04, Har00a, Mol08, Qui08, Rau03]. Triumphant [Elk06]. Triumphs [End00]. Trophy [Lev00]. Tropical [Biv07, Moo02b]. Tropics [Sto02]. Trouble [AA01, Sol04]. Troubled [dlP07, Sha09a]. Troy [Pos05]. Truck [Gwy04]. Trucks [MK01, Rot08]. True [Gla01, Uss09, Sau05]. Truesdell [Pet05]. Trust [Car08, Tod08, SS07, Day01, Uss04, Wel08]. Truth [Cha02, Mil04a]. Trying [O’H09]. Tuberculosis [AST08]. Tuchlebarone [Low00a]. Tunneling [Wel02]. Tunnels [Poh08]. Tupperware [Fri01]. Turing [Kid06, Sha00c]. Turk [Bek07]. Turkey [Bek07]. Turn [Hil02, Kou01a, Sha01b, Kou03]. Turning [HÖ08]. Turriano [Alt00]. Tuxedo [McE04]. TV [Daw07b, Mor01, Mul05]. Twentieth [Bro07, Cor08, De 03a, Dir07, Flo07, Fra02, Fra08, Hac04, Hur04, Igl01, Jac00, Jac07, Jag00, Kuw00, Las06, Lig09, Lip06, Lon07, McE07, Mie09, Per06b, Pet07b, Pye01, Pyn09, Rau05, Rot08, Seg08, Sha00a, Shu02, Sto03a, Tho01, DvdE03, Gag02, Mar02, Fal06b, Tea05]. Twentieth-Century [Cor08, De 03a, Fra08, Hac04, Igl01, Jac07, Jag00, Kuw00, Las06, Per06b, Pyn09, Shu02, DvdE03, Gag02, Fal06b, Tea05]. Twenty [Alt00, Cla01, Jon01, McG00a, SD01, Sti09]. Twenty-First [Cla01, McG00a, Sti09]. Twenty-One [Alt00]. Twenty-Seventh [Jon01]. Twenty-Sixth [SD01]. Twice [Hog03]. Twin [Flo01]. twintigste [Wac05]. Two [Con06b, Goo06, Gre02, Lin02b, Pos07, Sti04a, Tea08, Rus09]. Tycoon [McE04, McE06]. Typefounders [Pre06]. Typewriter [Mag01]. Typhus [Sta05]. Typography [McG03]. Tysons [Bas09c].

U [Ros03]. U.S [Ang04, Bug04, Bug05, Joh02c, Kid08, Lan09, Lug03, McB01, McB04, Mei02, Moy06, Nis08, Pan02b, Pra03, Sat08, Sch09b, Sch02c, Sha04, Smi05b, Sto01a, Sto03b, The08, Vin03, Wea02, Wol08]. U.S. [Ang01, Yea04]. U.S.-Brazilian [Sha04]. U.S.A. [Kāt07]. Ubaye [Ben03]. Udvar [All04]. Udvar-Hazy [All04]. UFOs [Mor00]. Ugly [Ste04]. Ultrasound [Cow03, How00]. umfangreichen [Pet01a]. Umwelt [Sto01b, Hey01]. Unappeasable [Pas06]. Unbuilding [Kon06]. Uncertainty [Aik07, Lip02, Scr09, Mor01]. Uncle [Jon03]. Uncommon [Ham08b]. Uncovering [Ang04]. Undead [Nad04b]. Underdeveloped [For07]. Underground [Glo02, Jen00]. Underworlds [McC08a]. undiscovered [Lin00]. Unemployment [Fan01]. Unending [Osl04]. Unequal [Tya04]. unglaublichen [Mos06]. Uniformity [McC09c]. Union [Bai01, Sch04d, Sid03]. Unique [Elk06]. Unit [Hul03]. United [Aik03, Bro01, Bro02, Bud06, Bur00a, Bye04, Col07a, Coo04b, Cro08b, Div03a, Fei09, Hea01, Hyd00, Jos01, Kle03b, Lan07a, Lit03, Low01a, Mal01, Mis02, Moo02b, Mor05a, Nor09, Par08, Pow02, Sch01a, Slo02a, Wel08, Wer05b, Whi02, Bas09b, Bri07, Bro00c, Eck05, Ham06a, Hol03, Soc08, Spi04, Vie06]. Units [Ash04]. Universal [Cut09b, Low08b]. Universities [Mod06b, Slo08, Tei01]. University [Cut09b, Gie02, Low05a, Ser05, Tei05, Gol00b, Hor07, O’G04b]. Unnatural [Roz09]. Unorthodox [Gel04]. Unprepossessing [Cas04b]. Unraveling [Fit00, McC00b]. Unreliable [Con99, Scr00]. Untamed [Coo04b]. Unthinkable [Sho08]. Untold [Eri07, Nie07a, O’H07]. Upcountry [Mar09]. Upheaval [Hoh03]. Upper [Har04]. Uranium [Che04, Kir08]. urbain [Fro01b]. Urban [Abb05, All07a, Elk02, Ell05, Hol03, Iri03, Jac08a, KM04, Kel06b, Kon06, Lig02, Mal09b, McC08a, Sch04c, Sha00b, Sha09b, Sta01, Tho05, Tho08, Win05]. urbanisation [Poi08]. Urbanism [Lig02, Hex07]. urbis [KO02]. Ursprung [Wen07]. Use [Jac08a, Kel04, Kri00b, NV07, Sha06a, War05b, Fri07b]. Used [Eri07]. Useful [Pop02]. User [Ake01]. Users [Scr05]. Uses [Cor02b, Muk03, Vas01]. Usher [Mol05]. usines [Bel04b]. USRA [Jon03]. USS [McB00]. Usselman [Lar04]. USSR [Neu02, Sie09]. Utility [Elk02, Pra01]. Utopia [Kno06a]. Utopian [Seg08]. Utopianism [Bar08]. Utrecht [Gwy09].

V [Neu02, Rau01]. V-2 [Neu02]. Vacations [You09]. Vaclav [Shu09b]. Valencia [Bur06]. Valley [Bas07c, BS09, Les01, Col03, Goo06, Hol06, Low05a]. Valuable [Zel01, Zel02]. Value [Bla05, Gre08c]. Values [Gle03, Pre03, O’H01b, Ser02b]. Valuing [Rey01b]. Vandergrift [Kno06a]. Vanishing [Lan02b]. vapeurs [Haf09]. Variation [Rot04b]. Varying [vR07]. Vauban [Rey05, Lan04b]. VCR [Gre04a]. Veblen [O’H03]. Vegas [Ano07c, Bro08]. Vehicle [Hug02, Vol05c, Rol01]. Vehicles [Bro05a, Sat08, Bug05]. veille [Gil06]. Venetian [McG02]. Venezia [Squ00]. Veni [Gre04a]. Venice [HV01, Hur02, Lin06a, Nak07, Maz01]. Venus [War09c]. Verify [Day01]. Verkehrstechnik [Vee01]. Verkehrswege [Sch05b]. Vernetzte [Wen09]. Versus [Mei00b, Tea08, Zet08]. Very [Con06a, Pos01a]. vetenskapen [Har06a]. Vibrator [Wos00]. Vice [dV02]. Victor [Wri06]. Victoria [AKP09]. Victorian [Bal00, Bow00b, Bra05, Bri01, Dob08a, Dow08b, Gre01, Jon02b, Qui03a, Rii03, Rob06b, Shu08, Tod08, War09c]. Victorians [Pop02]. Victrola [Mag05a]. Video [Daw07b, Gre04a]. Vidi [Gre04a]. Vienna [Hut00]. Vietnam [Bug04, Big08]. View [Nie02, RP09]. Viewing [Bar04a]. Views [Fie01]. Village [Hyd01, Pie06, Ben03]. Villain [Coo03a]. Villard [Bac05]. ville [Fro01b]. Vinci [Kel09]. Violence [Gor01a, Pan02b]. Virginia [Bar09b, Her00, Pos03]. Virtual [Ber08, Dow01, Pic00, Tur05, Bla09a, BW04]. Visibility [Jeo08]. Visible [Coo01b, New00, Uss06, Wea02]. Vision [Gis01, Iri05, Ner08, Osi00, Owe04, Sto09a]. Visioning [Kat00]. Visions [Bas07a, Ber07, Cer03, Cro07, Hal07, Lan03a, McC07b, Mol03, Pan02b, Ros05b, Pre03, Mae09]. Visual [Bar04a, Bro05b, McG00b, Sin05b]. Visualization [Bek04c]. visuellen [Heß04]. VOA [Sto09a]. Voice [Slo08]. Voices [Mol07, Nie07b, Kel05]. Voile [Low05b]. Vol [Fox00, Isr02, Pre05, Sch02b, Kie09, Pre08a, Rot01, Zel02]. Vol. [Isr02]. Volksgemeinschaft [Has06]. Volta [Maz04a]. Volume [Ano00k, Ano00j, Ano00l, Ano00m, Ano00n, Ano01g, Ano01f, Ano01h, Ano01i, Ano02h, Ano02l, Ano02k, Ano02m, Ano02n, Ano02o, Ano03d, Ano03f, Ano03e, Ano03g, Ano03h, Ano04f, Ano04h, Ano04g, Ano04i, Ano04j, Ano05g, Ano05i, Ano05h, Ano05j, Ano05k, Ano06d, Ano06i, Ano06h, Ano06j, Ano06k, Ano07f, Ano07e, Ano07g, Ano07h, Ano08e, Ano08g, Ano08f, Ano08h, Ano08i, Ano09d, Ano09f, Ano09h, Ano09g, Ano09i, Ano09j, Rob08b]. Voluntarism [Ake01]. vom [Fox02, Low00a, Zel04]. Vorarlberg [Vee00]. vorindustriellen [Die01]. Vorsicht [Bey07]. Voyager [Mar05a]. Vulnerability [Bij09].

W [Bro05b, Cer09, Iri08b, Phi04]. Wabash [Sch06d]. Wachhund [Hes04]. Wagon [Fri09]. Wake [Mid03]. Wakon [Koi02]. Waldo [Lon00]. Walker [Bro09c, Smi09]. Walking [BB00]. Wall [McE04]. Walloon [Awt07]. Walt [Mol08]. Walter [Buc01a, Jon02a, Rob06c]. Walther [Bey00, Bey06]. Wampler [Kno01a]. Wankel [Wen07]. Want [Lin02a]. War [And05, And06b, Ang04, Ang06, Bas09b, Ber03b, Blu01, Bro00a, Col07c, Cow06, Cru03, Dor08, Eck05, Hec09, Hug09b, Hul03, Kid08, Kin09b, Kle06b, Low04a, McB00, McB04, McC09b, McE04, Mei00b, Mul04, Nei02, Nei05, Nis07, Pic01, Rau01, Rau05, Sam07, Sch04c, Ser02c, Sho05, Smi03, Smi06, Ste06, Vie06, Vin03, Vol03a, Wal06b, Wei06, Whi05, Wit08, Boc09, Fre07, Hec08, Lai03a, McB00, Slo02b, Sto09a, Wei03, Rus00, Ang05a, Bar03c, Bek05, Bro00a, Cer07, Day01, Dor08, Eri07, Fos09, Her01a, Kru08, Mei02, Mor05b, Ser02c, Sho02, Tri04]. Warfare [Gre03, Hac05b, McB02, Nei05, Phi04, Rol05, Sch04c, Sho08, Wal00, Hac05b, Hac07, Joh06, Whe00, Zim04]. Warm [Che04]. Warming [Cus05]. Warnings [Sau05]. Warp [Ham09]. Warriors [Vin03, Phi04]. Wars [Clo07a, Her01a, Per06b, Sha04, Sta06b, Jen08]. Warships [McG00a]. Was [Ree05b, Col05]. Washington [All04, Ano08d, Bar07, Cer07, Dou04, Sti03, Tho07, Vog04]. Wasser [Ber06]. Wasserversorgung [Ber06]. Waste [Kno03, Mel06, Olw09, Lin02a]. Wasted [McG00c]. Wasteland [Bla09b]. Watch [Ste01c]. Watcher [Cro02a]. Watching [Jeo09, Tho08]. Water [Big08, Bij02, Bla08, DO03, Din08, Ell05, Ham05, HÖ08, Jac08b, Jos07, Kai02, Lin02b, Luc04, Maz07, Nak07, Reu02, Reu08, Sha01a, Ste04, TeB02, Tod09, Van08, Yar08, dV02, vD03]. Waters [Dur06, Elk06, Sha09a]. Watertown [Kan07]. Waterways [vD09]. Waterwolf [TeB02]. Waterworks [vD03]. Watson [Cow04b]. Watt [Coo05a]. Wave [Coc03]. Wavelength [Soc08]. Waves [McB06]. Way [Ang06, Gor04b, Tea09, Her01a]. Wayne [Dal04]. Ways [Ang04, Pac02]. Wealth [Kno03]. Weaponry [Hac05b]. Weapons [Bek06, Bug04, Car03, For07, For08, Her05, Jos01, Sam07, Sch09b, Sho08, Wal09, Sil00, Das08]. Weather [Kri00a, Tur09b]. Weathering [Dem03]. Weaver [Moh04]. Weavers [Gle08, Mac04]. Weaving [Ham05]. Web [Cer06]. Webs [Dow01]. Wedgwood [Spa01]. Weeds [And05]. Week [Geo06]. Weft [Ham09]. Weights [Ash05]. Weimar [All03]. Weisse [Gal00]. Welfare [Sch04c]. Well [Cro08b, Olw09, Tur05]. Well-Being [Cro08b]. Welland [Sha00b]. Welt [Bak04]. Weltgeschichte [Hey01]. Were [Ada06, Flo00, Ols07]. Werkzeug [Has02]. Wernher [Cro06c]. West [Ako07, Bar09b, Bry07, Dav00, Her00, Hov04, Mag09, Wil00a, Zac06, Bro08, Bro05c, Che04, Dev00a, Hov04, Igl01, Wal06a, Whi01]. Westdeutschland [Sud08]. Western [Big00, Bro03, Cer09, Hoa01, Low05b, Nei02, Nye04, Pow02, Ric08b, Wal02, War05b, Rob08b, Sat04, SS00]. Westfalen [Zel04]. Westinghouse [Cow04a, Las03, Ste01b]. WHA [Doe08a]. Whale [O’H09]. Whaling [Fal06a, Fal06b]. Wheatstone [Fin03]. Wheel [Bre00, Cor02a]. Wheelchair [WW04]. Wheeler [Spa02]. Wheels [Bor08, Jon04a]. Where [O’H06, Tur05, Ben05b]. Whirlwind [Fri02b]. White [Rol03]. Whiteness [Biv07]. Whittling [Coo03a]. Who [Hof02, Joh06, O’H07, AA01, Tur08, Wil04]. Whole [Bar08, Her05]. Whorehouse [Bra07a]. Whose [Fro05, McB00]. Wiener [KM04]. Wilbur [Bed02]. Wild [Bro05c]. Wilderness [Bar07, Coo04b, Pyn09]. Wilhelm [Sto07a, Pet04c]. Will [Epp06]. William [Bac02, Ens07, Hac05a, dO01b, Flo03, Dal06b, Sin09, Spo02]. Williams [Cas03]. Willis [Zai08]. Win [Eri07]. Wind [All07b, McG05, Bla08, Maz07, Nie02]. Windmills [LW05]. Windshield [Nor09, Bar07]. Windsor [Coo07a]. Windsor-Chair [Coo07a]. Wine [Gou04b]. Wing [Lev05]. Wings [Bug01, Van00a, Bug01]. Winners [Lev00]. Winter [KO07]. Wir [MW07]. Wire [Bek02, Gui03, Hug05]. Wired [Nad08]. Wireless [Slo04, Yea04, Bow03, Coo03b]. Wires [Whi05]. Wirtschaft [Mar03]. wirtschaftshistorische [dS09]. Wisconsin [Doe08a]. Wise [Her01b]. Wissenschaft [Mar03]. Wissenschaftliche [MW07]. Wissenschaftsorganisation [Bey06]. Within [How00]. Without [Van00a, Bla09c, Mom05, Pic06, Sch08a]. Wives [Nie07b]. Wizard [Pre08b, Rot01]. Wizards [Joh02b]. Wolfman [Coo02c]. Woman [Ber03a, New00]. Women [Aik07, Ben02, Ben08, BB00, Dav01, Dav06, Fra08, Gor01b, Hen09, Kha01, Mac06, Moh08, Nic02, Nie07b, Och02, Sha09b, Vin03, Waj00, Wan02, War05a, War06b, Wos00, Cor02a, Dou09, Mai07, Sch00d, Sha00a, WH09]. Won [Jac02]. Wonder [Jan09, Nye06]. Wood [Ald06, Bug01, Mor06a]. Wood-Preserving [Ald06]. Wooden [McG06]. Woods [Pre04c]. Wool [Ran01]. Work [Ake05, Bel08, Ben08, Ber07, Dav01, Mac04, Mal01, Maz07, Mol06, Par05, WH09, Ash00, Dob08d, Dow03a]. Worker [Ale06, Moh08, Rob00]. Workers [Dow01, Lin00, Aik07]. Working [Bro00c, Sch07c, Sim03, Lon03b]. Workplace [Lew01]. Works [Bro00b, Epp07, Her08a, Hyd00, Kel06b, Kle03b, Lan03a, Ryd01, Sti03, Wan07, dV02, Zei01]. Workshops [Wen01]. World [And06b, Bek02, Bek05, Ber08, Bil02, Blu01, Bro00a, Bro02, Coo00, Cow06, Cro06b, DO03, Dan01, Doe08c, Dor08, End00, Fie01, Flo01, Geo06, Her05, Hof02, Hor02, Hug00b, Hug09b, Jon04a, Kid08, Kru09, Lak04, Lin06a, Maz03, McE04, Mei00b, MS08, MA09, Mie09, Mor08a, Nad07, Nei05, Nor09, Nye05, Olw04, Pic01, Qui03a, Ros03, Rot08, Sam07, Sch07d, Ser02c, Ser05, Sid06b, Tom03, Tri06a, Tri04, Tya04, Vie06, Vin03, Whi04, Wil02b, dlP07, Eck05, For07, Pic00, Bas09a, Big00, Buc05a, Bul03a, Cow04a, Jon02b, Jud03, Kar09a, Kha08, Kno05a, Kon09, Lap09, Moo02b, Nie02, Osl04, Pre08b, Rol05, Slo04, Sta00, Sta06a, Ste06, Sto03b, Tur08, Wak03]. Worlds [Kon05, Kru08, Luc05a, Pre04a]. Worse [Low01a]. Worst [O’H07, Rol06b]. Worster [Sti07]. Worthy [Wer06]. Would [Les08b]. Wright [Dou04, Kin00, Tri06b, Bed02]. Wrights [De 04]. Write [Mei05]. Writers [Lon03b]. Writing [Dal05, Dal06a, Jac08b, McG03, Pre02, Sch00d]. Writings [Bed02, Mil08, Jaf06]. Written [Lau08, vF05]. Wrong [Lar06]. Wrote [AA01]. Wrought [De 04, Har04, Smi09]. Wurzeln [Has02].

X [How00, dlP06]. X-Ray [How00, dlP06]. XIIIe [Bla03]. XIV [Ben05a]. XIX [Dol09]. XIXe [Bor06a, Fro01b, Poi08]. XVIe [Bla03]. XVIII [Dol09]. XVIIIe [Fro01b, Gui06]. XX [PP09]. XXe [Gui06].

Yankee [Fra06, Coo03a]. Yard [McB01]. Year [Col03, Poo03, Wak03]. Years [Clo05, Mag05a, Mar01, McG00c, Orv09, Rob08b, Sch02e, Wal01c, Ber01, Sid07, Wha06]. Yellow [See01]. York [All07a, Chu04b, Cor06, Flo01, Jac05a, Mid05b, Mil02, O’H09, Pre08a, Sch03, Tea09, Wos02b, Iri02, Les01, Zei01, Wel02]. Young [Sch00c]. Yourself [Jag00].

Zapotec [Moo02a]. Zecca [Hur02]. Zeitalter [Wen09]. Zentralverwaltungswirtschaften [Cie01]. Zeppelin [Pal02]. Zero [Tur09a]. Zigong [Per07]. Zoned [Fri07b]. Zoomscape [Sam05]. zum [Zel04]. Zur [Ber03a, Ber06, Cie01, Low00a, Osi00, Sto07a, Wen08, Die01, Heß04, Mar03]. Zweiten [Fox02]. zwischen [Sto01b, Sto05]. Zwischenkriegszeit [Fox02].

References

Abir-Am:2001:FDT

[AA01] Pnina G. Abir-Am. French DNA: Trouble in Purgatory, and: Who Wrote the Book of Life? A History of the Genetic Code (review). Technology and Culture, 42(2):390–392, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

              tronic). URL https://muse.jhu.edu/pub/1/article/33702.

Abbate:2003:TET

[Abb03] Janet Abbate. The Triumph of Ethernet: Technological Communities and the Battle for the LAN Standard (review). Technology and Culture, 44(3):635–637, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

              URL https://muse.jhu.edu/pub/1/article/46080.

Abbate:2004:PCI

[Abb04] Janet Abbate. Prefiguring Cyberculture: An Intellectual History (review). Technology and Culture, 45(4):901–902, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175731.

Abbott:2005:HMF

[Abb05] Carl Abbott. Hazardous Metropolis: Flooding and Urban Ecology in Los Angeles (review). Technology and Culture, 46(2): 460–461, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184300.

Abell:2002:MMA

[Abe02] Anne B. Abell. Mechanics and Meaning in Architecture (review). Technology and Culture, 43(1):159–160, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33851.

Adams:2002:APT

[Ada02] Robert McCormick Adams. Anthropological Perspectives on Technology (review). Technology and Culture, 43(1):155–157, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33852.

Adam:2006:WCW

[Ada06] Alison Adam. When Computers Were Human (review). Technology and Culture, 47(4):875–877, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204965.

Adelman:2008:PAE

[Ade08] Joseph M. Adelman. The Postal Age: The Emergence of Modern Communications in Nineteenth-Century America (review). Technology and Culture, 49(1):272–273, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230892.

Agar:2009:CPC

[Aga09] Jon Agar. Cell Phone Culture: Mobile Technology in Everyday Life (review). Technology and Culture, 50(3):726–728, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270554.

Aiken:2003:OPM

[Aik03] Katherine G. Aiken. Old Paint: A Medical History of Childhood Lead-Paint Poisoning in the United States to 1980 (review). Technology and Culture, 44(2):405–407, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/42003.

Aiken:2007:SBS

[Aik07] Katherine G. Aiken. Sick Building Syndrome and the Problem of Uncertainty: Environmental Politics, Technoscience, and Women Workers (review). Technology and Culture, 48 (2):444–446, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215375.

Akera:2001:VFC

[Ake01] Atsushi Akera. Voluntarism and the fruits of collaboration: The IBM user group, share. Technology and Culture, 42(4): 710–736, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33799.

Akera:2005:DHH

[Ake05] Atsushi Akera. The Digital Hand: How Computers Changed the Work of American Manufacturing, Transportation, and Retail Industries (review). Technology and Culture, 46(2):439– 441, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184301.

Akou:2006:CMC

[Ako06] Heather Marie Akou. Clothing as Material Culture (review). Technology and Culture, 47(4):824–826, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204966.

Akou:2007:CWA

[Ako07] Heather Marie Akou. Cloth in West African History (review). Technology and Culture, 48(2):431–433, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215376.

Ampuja:2009:ICH

[AKP09] Outi Ampuja, Jeremy Kinney, and Stefan Poser. International Committee for the History of Technology: Copenhagen, Denmark, 2007; Victoria, British Columbia, 2008. Technology and Culture, 50(4):883–893, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/363419.

Alac:2009:BHL

[Ala09] Morana Alac. Beyond Human: Living with Robots and Cyborgs (review). Technology and Culture, 50(2):468–469, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263338.

Alcorn:2007:BBB

[Alc07] Aaron Alcorn. Boys’ Books, Boys’ Dreams, and the Mystique of Flight (review). Technology and Culture, 48(4):899–900, October 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/223339.

Alder:2000:FAE

[Ald00] Ken Alder. French Architects and Engineers in the Age of Enlightenment (review). Technology and Culture, 41(3):588– 589, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33545.

Aldrich:2006:FCM

[Ald06] Mark Aldrich. From forest conservation to market preservation: Invention and diffusion of wood-preserving technology, 1880–1939. Technology and Culture, 47(2):311–340, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200832.

Alexander:2003:PEK

[Ale03] Jennifer Karns Alexander. Philadelphia’s Enlightenment, 1740–1800: Kingdom of Christ, Empire of Reason (review).

Technology and Culture, 44(4):825–827, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49677.

Alexander:2006:EPM

[Ale06] Jennifer Karns Alexander. Efficiency and pathology: Mechanical discipline and efficient worker seating in Germany, 1929–1932. Technology and Culture, 47(2):286–310, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200833.

Allen:2002:STS

[All02] Michael Thad Allen. Stranger than science fiction: Edwin black, IBM, and the Holocaust. Technology and Culture, 43(1): 150–154, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33853.

Allen:2003:PSN

[All03] Michael Thad Allen. Poems in Steel: National Socialism and the Politics of Inventing from Weimar to Bonn (review). Technology and Culture, 44(2):414–415, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/42004.

Allen:2004:SFU

[All04] Catherine W. Allen. The Steven F. Udvar-Hazy Center: National Air and Space Museum, Washington, D.C. Technology and Culture, 45(2):358–362, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55653.

Allen:2005:AAA

[All05] Michael Thad Allen. Architects of Annihilation: Auschwitz and the Logic of Destruction (review). Technology and Culture, 46 (2):449–451, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184302.

Allen:2006:CRI

[All06a] Barbara L. Allen. Cradle of a revolution? The industrial transformation of Louisiana’s Lower Mississippi River. Technology and Culture, 47(1):112–119, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196024.

Allen:2006:HBG

[All06b] Michael Thad Allen. Hitler’s Bombe: Die geheime Geschichte der deutschen Kernwaffenversuche (review). Technology and Culture, 47(3):671–672, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/201438.

Allen:2007:NNY

[All07a] Barbara L. Allen. Noxious New York: The Racial Politics of Urban Health and Environmental Justice (review). Technology and Culture, 48(4):868–870, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223340.

Allen:2007:LWR

[All07b] Barbara L. Allen. “To Love the Wind and the Rain”: African Americans and Environmental History (review). Technology and Culture, 48(1):200–201, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209819.

Altshuler:2000:TOB

[Alt00] Jose Altshuler. The Twenty-One Books of Engineering and Machines of Juanelo Turriano (review). Technology and Culture, 41(3):576–578, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33546.

Alvarez:2004:RMI

[Alv04] Samantha Alvarez. Railways in Modern India (review). Technology and Culture, 45(1):180–181, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51733.

Anderson:2000:FSE

[And00] John David Anderson. Frontiers of Space Exploration (review). Technology and Culture, 41(2):380, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33509.

Anderson:2005:WWI

[And05] J. L. Anderson. War on weeds: Iowa farmers and growthregulator herbicides. Technology and Culture, 46(4):719–744, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192282.

Anderson:2006:PFE

[And06a] J. L. Anderson. From Prairie Farmer to Entrepreneur: The Transformation of Midwestern Agriculture (review). Technology and Culture, 47(1):205–206, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196025.

Anderson:2006:TDP

[And06b] J. L. Anderson. Toxic Drift: Pesticides and Health in the Post-World War II South (review). Technology and Culture, 47(2): 452–454, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200834.

Anderson:2009:RFT

[And09] J. L. Anderson. A Revolution Down on the Farm: The Transformation of American Agriculture since 1929 (review). Technology and Culture, 50(4):927–929, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363427.

Angevine:2001:IOE

[Ang01] Robert G. Angevine. Individuals, organizations, and engineering: U.S. Army officers and the American railroads, 1827–1838. Technology and Culture, 42(2):292–320, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33703.

Angevine:2004:IWK

[Ang04] Robert G. Angevine. Intelligence in War: Knowledge of the Enemy from Napoleon to Al-Qaeda, and: Uncovering Ways of War: U.S. Intelligence and Foreign Military Innovation, 1918–1941 (review). Technology and Culture, 45(4):881–883, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175732.

Angevine:2005:NEN

[Ang05a] Robert G. Angevine. Natural Enemy, Natural Ally: Toward an Environmental History of War (review). Technology and Culture, 46(3):657–659, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186600.

Angevine:2005:DMT

[Ang05b] Robert G. Angevine. The Diffusion of Military Technology and Ideas (review). Technology and Culture, 46(1):220–222, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179745.

Angevine:2006:FWW

[Ang06] Robert G. Angevine. The First Way of War: American War Making on the Frontier (review). Technology and Culture, 47 (2):423–425, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200835.

Anonymous:2000:AI

[Ano00a] Anonymous. Author index. Technology and Culture, 41 (Supplement):154–171, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33650.

Anonymous:2000:A

[Ano00b] Anonymous. Awards. Technology and Culture, 41(3):549– 557, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33547.

Anonymous:2000:B

[Ano00c] Anonymous. Bibliography. Technology and Culture, 41 (Supplement):5–153, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33651.

Anonymous:2000:Ia

[Ano00d] Anonymous. In this issue. Technology and Culture, 41(1):ix, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33479.

Anonymous:2000:Ib

[Ano00e] Anonymous. In this issue. Technology and Culture, 41(2):viii– ix, April 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33527.

Anonymous:2000:Ic

[Ano00f] Anonymous. In this issue. Technology and Culture, 41(3):ix– x, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33566.

Anonymous:2000:Id

[Ano00g] Anonymous. In this issue. Technology and Culture, 41(4):ix–x, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33618.

Anonymous:2000:OND

[Ano00h] Anonymous. Organizational notes: The Detroit Meeting, 7– 10 October 1999. Technology and Culture, 41(3):558–565, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33584.

Anonymous:2000:SI

[Ano00i] Anonymous. Subject index. Technology and Culture, 41 (Supplement):172–202, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33652.

Anonymous:2000:VNJa

[Ano00j] Anonymous. Volume 41, number 1, January 2000. Technology and Culture, 41(1):??, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2000:VNA

[Ano00k] Anonymous. Volume 41, number 2, April 2000. Technology and Culture, 41(2):??, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2000:VNJb

[Ano00l] Anonymous. Volume 41, number 3, July 2000. Technology and Culture, 41(3):??, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2000:VNO

[Ano00m] Anonymous. Volume 41, number 4, October 2000. Technology and Culture, 41(4):??, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2000:VSO

[Ano00n] Anonymous. Volume 41, supplement, October 2000. Technology and Culture, 41(Supplement):??, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL /journal/194.

Anonymous:2001:A

[Ano01a] Anonymous. Awards. Technology and Culture, 42(3):531– 544, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33748.

Anonymous:2001:Ia

[Ano01b] Anonymous. In this issue. Technology and Culture, 42(2): ix, April 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33717.

Anonymous:2001:Ib

[Ano01c] Anonymous. In this issue. Technology and Culture, 42(3): ix, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33770.

Anonymous:2001:Ic

[Ano01d] Anonymous. In this issue. Technology and Culture, 42(4):ix, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33819.

Anonymous:2001:ONM

[Ano01e] Anonymous. Organizational notes: The Munich Meeting, 17– 20 August 2000. Technology and Culture, 42(3):545–558, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33780.

Anonymous:2001:VNJa

[Ano01f] Anonymous. Volume 42, number 1, January 2001. Technology and Culture, 42(1):??, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2001:VNA

[Ano01g] Anonymous. Volume 42, number 2, April 2001. Technology and Culture, 42(2):??, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2001:VNJb

[Ano01h] Anonymous. Volume 42, number 3, July 2001. Technology and Culture, 42(3):??, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2001:VNO

[Ano01i] Anonymous. Volume 42, number 4, October 2001. Technology and Culture, 42(4):??, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2002:AI

[Ano02a] Anonymous. Author index. Technology and Culture, 43 (Supplement):122–135, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33999.

Anonymous:2002:A

[Ano02b] Anonymous. Awards. Technology and Culture, 43(1):110–127, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33854.

Anonymous:2002:B

[Ano02c] Anonymous. Bibliography. Technology and Culture, 43 (Supplement):5–121, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/34000.

Anonymous:2002:Ia

[Ano02d] Anonymous. In this issue. Technology and Culture, 43(1):ix–x, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33870.

Anonymous:2002:Ib

[Ano02e] Anonymous. In this issue. Technology and Culture, 43(2): ix, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33949.

Anonymous:2002:Ic

[Ano02f] Anonymous. In this issue. Technology and Culture, 43(3): ix, July 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33966.

Anonymous:2002:Id

[Ano02g] Anonymous. In this issue. Technology and Culture, 43(4):ix, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34016.

Anonymous:2002:IV

[Ano02h] Anonymous. Index volume 43 (2002). Technology and Culture, 43(4):846–870, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34021.

Anonymous:2002:SJM

[Ano02i] Anonymous. The San Jose Meeting, 4–7 October 2001. Technology and Culture, 43(1):128–138, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33881.

Anonymous:2002:SI

[Ano02j] Anonymous. Subject index. Technology and Culture, 43 (Supplement):136–156, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/34003.

Anonymous:2002:VNJa

[Ano02k] Anonymous. Volume 43, number 1, January 2002. Technology and Culture, 43(1):??, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2002:VNA

[Ano02l] Anonymous. Volume 43, number 2, April 2002. Technology and Culture, 43(2):??, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2002:VNJb

[Ano02m] Anonymous. Volume 43, number 3, July 2002. Technology and Culture, 43(3):??, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2002:VNO

[Ano02n] Anonymous. Volume 43, number 4, October 2002. Technology and Culture, 43(4):??, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2002:VSA

[Ano02o] Anonymous. Volume 43, supplement, April 2002. Technology and Culture, 43(Supplement):??, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL /journal/194.

Anonymous:2003:Ia

[Ano03a] Anonymous. In this issue. Technology and Culture, 44(1):ix, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179439.

Anonymous:2003:Ib

[Ano03b] Anonymous. In this issue. Technology and Culture, 44(3):ix– x, July 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/46095.

Anonymous:2003:Ic

[Ano03c] Anonymous. In this issue. Technology and Culture, 44(4):ix–x, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49699.

Anonymous:2003:IV

[Ano03d] Anonymous. Index: Volume 44 (2003). Technology and Culture, 44(4):859–884, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49700.

Anonymous:2003:VNJa

[Ano03e] Anonymous. Volume 44, number 1, January 2003. Technology and Culture, 44(1):??, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2003:VNA

[Ano03f] Anonymous. Volume 44, number 2, April 2003. Technology and Culture, 44(2):??, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2003:VNJb

[Ano03g] Anonymous. Volume 44, number 3, July 2003. Technology and Culture, 44(3):??, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2003:VNO

[Ano03h] Anonymous. Volume 44, number 4, October 2003. Technology and Culture, 44(4):??, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2004:AMO

[Ano04a] Anonymous. The Atlanta Meeting, 16–19 October 2003. Technology and Culture, 45(1):142–153, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51764.

Anonymous:2004:A

[Ano04b] Anonymous. Awards. Technology and Culture, 45(1):123–141, January 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/51734.

Anonymous:2004:Ia

[Ano04c] Anonymous. In this issue. Technology and Culture, 45(2):v– vi, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55703.

Anonymous:2004:Ib

[Ano04d] Anonymous. In this issue. Technology and Culture, 45(3):ix– x, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172977.

Anonymous:2004:Ic

[Ano04e] Anonymous. In this issue. Technology and Culture, 45(4):i–ii, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175754.

Anonymous:2004:IV

[Ano04f] Anonymous. Index: Volume 45 (2004). Technology and Culture, 45(4):931–959, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175753.

Anonymous:2004:VNJa

[Ano04g] Anonymous. Volume 45, number 1, January 2004. Technology and Culture, 45(1):??, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2004:VNA

[Ano04h] Anonymous. Volume 45, number 2, April 2004. Technology and Culture, 45(2):??, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2004:VNJb

[Ano04i] Anonymous. Volume 45, number 3, July 2004. Technology and Culture, 45(3):??, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2004:VNO

[Ano04j] Anonymous. Volume 45, number 4, October 2004. Technology and Culture, 45(4):??, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2005:AMO

[Ano05a] Anonymous. The Amsterdam Meeting, 8–10 October 2004. Technology and Culture, 46(1):149–163, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179773.

Anonymous:2005:A

[Ano05b] Anonymous. Awards. Technology and Culture, 46(1):137–148, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179746.

Anonymous:2005:Ia

[Ano05c] Anonymous. In this issue. Technology and Culture, 46(1):i–ii, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179761.

Anonymous:2005:Ib

[Ano05d] Anonymous. In this issue. Technology and Culture, 46(2):ix– x, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184321.

Anonymous:2005:Ic

[Ano05e] Anonymous. In this issue. Technology and Culture, 46(3):ix– x, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186620.

Anonymous:2005:Id

[Ano05f] Anonymous. In this issue. Technology and Culture, 46(4):ix–x, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192327.

Anonymous:2005:IV

[Ano05g] Anonymous. Index: Volume 46 (2005). Technology and Culture, 46(4):882–909, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192298.

Anonymous:2005:VNJa

[Ano05h] Anonymous. Volume 46, number 1, January 2005. Technology and Culture, 46(1):??, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2005:VNA

[Ano05i] Anonymous. Volume 46, number 2, April 2005. Technology and Culture, 46(2):??, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2005:VNJb

[Ano05j] Anonymous. Volume 46, number 3, July 2005. Technology and Culture, 46(3):??, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2005:VNO

[Ano05k] Anonymous. Volume 46, number 4, October 2005. Technology and Culture, 46(4):??, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2006:A

[Ano06a] Anonymous. Awards. Technology and Culture, 47(1):120–140, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196027.

Anonymous:2006:Ia

[Ano06b] Anonymous. In this issue. Technology and Culture, 47(1):ix–x, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196075.

Anonymous:2006:Ib

[Ano06c] Anonymous. In this issue. Technology and Culture, 47(2):ix– x, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200852.

Anonymous:2006:IV

[Ano06d] Anonymous. Index to volume 47 (2006). Technology and Culture, 47(4):882–912, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/204982.

Anonymous:2006:MMN

[Ano06e] Anonymous. The Minneapolis Meeting, 3–6 November 2005. Technology and Culture, 47(1):141–155, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196056.

Anonymous:2006:Ic

[Ano06f] Anonymous. This issue. Technology and Culture, 47(3):ix– x, July 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/201481.

Anonymous:2006:Id

[Ano06g] Anonymous. This issue. Technology and Culture, 47(4):ix–x, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/205011.

Anonymous:2006:VNJa

[Ano06h] Anonymous. Volume 47, number 1, January 2006. Technology and Culture, 47(1):??, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2006:VNA

[Ano06i] Anonymous. Volume 47, number 2, April 2006. Technology and Culture, 47(2):??, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2006:VNJb

[Ano06j] Anonymous. Volume 47, number 3, July 2006. Technology and Culture, 47(3):??, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2006:VNO

[Ano06k] Anonymous. Volume 47, number 4, October 2006. Technology and Culture, 47(4):??, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2007:A

[Ano07a] Anonymous. Awards. Technology and Culture, 48(1):117–136, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209820.

Anonymous:2007:Ib

[Ano07b] Anonymous. In this issue. Technology and Culture, 48(2):ix– x, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215417.

Anonymous:2007:VMO

[Ano07c] Anonymous. The Las Vegas Meeting, 12–15 October 2006. Technology and Culture, 48(1):137–152, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209847.

Anonymous:2007:Ia

[Ano07d] Anonymous. This issue. Technology and Culture, 48(1):ix–x, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209862.

Anonymous:2007:VNJa

[Ano07e] Anonymous. Volume 48, number 1, January 2007. Technology and Culture, 48(1):??, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2007:VNA

[Ano07f] Anonymous. Volume 48, number 2, April 2007. Technology and Culture, 48(2):??, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2007:VNJb

[Ano07g] Anonymous. Volume 48, number 3, July 2007. Technology and Culture, 48(3):??, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2007:VNO

[Ano07h] Anonymous. Volume 48, number 4, October 2007. Technology and Culture, 48(4):??, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2008:A

[Ano08a] Anonymous. Awards. Technology and Culture, 49(1):133–178, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230894.

Anonymous:2008:Ia

[Ano08b] Anonymous. In this issue. Technology and Culture, 49(1):a–b, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230934.

Anonymous:2008:Ib

[Ano08c] Anonymous. In this issue. Technology and Culture, 49(3):a– b, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243096.

Anonymous:2008:LBL

[Ano08d] Anonymous. “Looking Back, Looking Beyond”: The Washington, D.C., Meeting, 17—21 October 2007. Technology and Culture, 49(1):179–201, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230919.

Anonymous:2008:V

[Ano08e] Anonymous. Volume 49 (2008). Technology and Culture, 49(4): 1106–1137, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255453/pdf.

Anonymous:2008:VNJa

[Ano08f] Anonymous. Volume 49, number 1, January 2008. Technology and Culture, 49(1):??, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2008:VNA

[Ano08g] Anonymous. Volume 49, number 2, April 2008. Technology and Culture, 49(2):??, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2008:VNJb

[Ano08h] Anonymous. Volume 49, number 3, July 2008. Technology and Culture, 49(3):??, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2008:VNO

[Ano08i] Anonymous. Volume 49, number 4, October 2008. Technology and Culture, 49(4):??, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2009:A

[Ano09a] Anonymous. Awards. Technology and Culture, 50(1):127–154, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258551.

Anonymous:2009:Ia

[Ano09b] Anonymous. In this issue. Technology and Culture, 50(1):a–b, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258561.

Anonymous:2009:Ib

[Ano09c] Anonymous. In this issue. Technology and Culture, 50(3):i– ii, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270550.

Anonymous:2009:IJV

[Ano09d] Anonymous. In this issue, July 2009: Volume 50, No. 4: T&C’s 200th Issue. Technology and Culture, 50(4):a–c, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363411.

Anonymous:2009:LMO

[Ano09e] Anonymous. The Lisbon Meeting, 11–14 October 2008. Technology and Culture, 50(1):155–174, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258550.

Anonymous:2009:V

[Ano09f] Anonymous. Volume 50 (2009). Technology and Culture, 50(4): 987–1015, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363460/pdf.

Anonymous:2009:VNJa

[Ano09g] Anonymous. Volume 50, number 1, January 2009. Technology and Culture, 50(1):??, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2009:VNA

[Ano09h] Anonymous. Volume 50, number 2, April 2009. Technology and Culture, 50(2):??, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2009:VNJb

[Ano09i] Anonymous. Volume 50, number 3, July 2009. Technology and Culture, 50(3):??, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2009:VNO

[Ano09j] Anonymous. Volume 50, number 4, October 2009. Technology and Culture, 50(4):??, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Apple:2003:FPH

[App03] Rima D. Apple. Food Politics: How the Food Industry Influences Nutrition and Health (review). Technology and Culture, 44(4):839–841, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49679.

Arnold:2000:BPH

[Arn00] David C. Arnold. Beyond the Paths of Heaven: The Emergence of Space Power Thought (review). Technology and Culture, 41 (4):851–852, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33599.

Arns:2001:MSN

[Arn01] Robert Arns. Music, Science, and Natural Magic in Seventeenth-Century England (review). Technology and Culture, 42(1):190–192, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33654.

Arns:2003:SMA

[Arn03] Robert Arns. The Soundscape of Modernity: Architectural Acoustics and the Culture of Listening in America, 1900–1933 (review). Technology and Culture, 44(2):399–401, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42005.

Ash:2000:PAW

[Ash00] Eric H. Ash. “A perfect and an absolute work”: Expertise, authority, and the rebuilding of Dover harbor, 1579–1583. Technology and Culture, 41(2):239–268, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33510.

Ashworth:2001:BTP

[Ash01] William J. Ashworth. “Between the trader and the public”: British alcohol standards and the proof of good governance. Technology and Culture, 42(1):27–50, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33655.

Ashworth:2004:MMS

[Ash04] William J. Ashworth. Measure for Measure: The Story of Imperial, Metric, and Other Units (review). Technology and Culture, 45(2):438–439, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55654.

Ashworth:2005:WMS

[Ash05] William J. Ashworth. Weights and Measures in Scotland: A European Perspective (review). Technology and Culture, 46 (3):640–642, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186601.

Ashworth:2006:PKE

[Ash06] William J. Ashworth. Power, Knowledge, and Expertise in Elizabethan England (review). Technology and Culture, 47(1): 188–190, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196026.

Ashworth:2008:MSI

[Ash08] William J. Ashworth. Making Scientific Instruments in the Industrial Revolution (review). Technology and Culture, 49(1): 267–268, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230893.

Ashworth:2009:THO

[Ash09] William J. Ashworth. Tides of History: Ocean Science and Her Majesty’s Navy (review). Technology and Culture, 50(3):705– 706, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270543.

Asner:2003:RHT

[Asn03] Glen Asner. Researching the history of technology at the Hagley Museum and Library. Technology and Culture, 44(4): 762–777, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/49680.

Adams:2008:CEA

[AST08] Annmarie Adams, Kevin Schwartzman, and David Theodore. Collapse and expand: Architecture and tuberculosis therapy in Montreal, 1909, 1933, 1954. Technology and Culture, 49(4): 908–942, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255441.

Awty:2007:DDW

[Awt07] Brian G. Awty. The development and dissemination of the walloon method of ironworking. Technology and Culture, 48(4): 783–803, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223341.

Aydin:2005:STL

[Ayd05] Cemil Aydin. Science, Technology, and Learning in the Ottoman Empire (review). Technology and Culture, 46(1):249– 250, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179747.

Baark:2001:CRH

[Baa01] Erik Baark. The Chinese Road to High Technology: A Study of Telecommunications Switching Technology in the Economic Transition (review). Technology and Culture, 42(1):178–180, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33656.

Bacon:2002:TIW

[Bac02] Douglas R. Bacon. Tarnished Idol: William T. G. Morton and the Introduction of Surgical Anesthesia (review). Technology and Culture, 43(4):793–794, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34005.

Bachrach:2004:CKC

[Bac04] David Steward Bachrach. Crossbows for the King: The Crossbow during the reigns of John and Henry III of England. Technology and Culture, 45(1):102–119, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51735.

Bachrach:2005:VLS

[Bac05] David Steward Bachrach. Villard’s Legacy: Studies in Medieval Technology, Science, and Art in Memory of Jean Gimpel (review). Technology and Culture, 46(3):635–636, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186602.

Bachrach:2006:ADB

[Bac06a] David Steward Bachrach. The Artillery of the Dukes of Burgundy 1363–1477 (review). Technology and Culture, 47(4): 832–833, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204967.

Bachrach:2006:CKP

[Bac06b] David Steward Bachrach. Crossbows for the King, part II: The crossbow during the reign of Edward I of England (1272–1307). Technology and Culture, 47(1):81–90, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196028.

Bachrach:2007:CHR

[Bac07] David Steward Bachrach. The Catapult: A History (review). Technology and Culture, 48(4):846–848, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223342.

Bachrach:2008:GES

[Bac08] David Stewart Bachrach. Gunpowder, Explosives, and the State: A Technological History (review). Technology and Culture, 49(3):785–786, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/243087.

Baird:2001:CAS

[Bai01] Andrew Baird. Challenge to Apollo: The Soviet Union and the Space Race, 1945–1974 (review). Technology and Culture, 42(4):817–818, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33800.

Bakker:2000:TIO

[Bak00] Martijn Bakker. Technological Innovation: Oversights and Foresights (review). Technology and Culture, 41(1):124–125, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33464.

Bakker:2004:TPD

[Bak04] Lara Tauritz Bakker. Techniktheorien: Der Platz der Dinge in der Welt des Menschen (review). Technology and Culture, 45(4):903–904, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175733.

Bakker:2005:OOD

[Bak05] Lara Tauritz Bakker. Our Own Devices: The Past and Future of Body Technology (review). Technology and Culture, 46 (2):470–471, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184303.

Ball:2000:BMT

[Bal00] Norman R. Ball. Becoming Modern in Toronto: The Industrial Exhibition and the Shaping of a Late Victorian Culture (review). Technology and Culture, 41(3):627–628, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33548.

Barnes:2000:SEN

[Bar00a] Susan B. Barnes. The Soft Edge: A Natural History and Future of the Information Revolution (review). Technology and Culture, 41(2):390–392, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33511.

Barnstone:2000:TA

[Bar00b] Tony Barnstone. Technology as addiction. Technology and Culture, 41(1):190–193, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33465.

Barry:2000:CTS

[Bar00c] Andrew Barry. Coordinating Technology: Studies in the International Standardization of Telecommunications (review). Technology and Culture, 41(1):153–154, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33466.

Barnes:2002:BDE

[Bar02] Susan B. Barnes. Bootstrapping: Douglas Engelbart, Coevolution, and the Origins of Personal Computing (review). Technology and Culture, 43(1):200–201, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33855.

Barker:2003:EHL

[Bar03a] Joshua Barker. Engineers of Happy Land: Technology and Nationalism in a Colony (review). Technology and Culture, 44 (2):412–413, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42006.

Barnes:2003:HCR

[Bar03b] Susan B. Barnes. Hacker Culture (review). Technology and Culture, 44(1):194–196, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40142.

Barnhart:2003:NDN

[Bar03c] Megan Barnhart. Neither Dead nor Red: Civil Defense and American Political Development During the Early Cold War (review). Technology and Culture, 44(1):204–205, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40141.

Barnett:2004:DVV

[Bar04a] Gabrielle R. Barnett. Drive-by viewing: Visual consciousness and forest preservation in the Automobile Age. Technology and Culture, 45(1):30–54, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51736.

Barnhart:2004:AAH

[Bar04b] Megan Barnhart. Awaiting Armageddon: How Americans Faced the Cuban Missile Crisis (review). Technology and Culture, 45(3):635–636, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/172962.

Barnes:2005:DNI

[Bar05] Susan B. Barnes. Digitizing the News: Innovation in Online Newspapers (review). Technology and Culture, 46(2):434– 435, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184304.

Barnett:2007:WWC

[Bar07] Gabrielle R. Barnett. Windshield Wilderness: Cars, Roads, and Nature in Washington’s National Parks (review). Technology and Culture, 48(4):875–876, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223343.

Barnes:2008:CCS

[Bar08] Susan B. Barnes. From Counterculture to Cyberculture: Stewart Brand, the Whole Earth Network, and the Rise of Digital Utopianism (review). Technology and Culture, 49(3):824– 826, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243102.

Barnes:2009:CCS

[Bar09a] Susan B. Barnes. Closed Captioning: Subtitling, Stenography, and the Digital Convergence of Text with Television (review). Technology and Culture, 50(2):469–471, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263343.

Barnhill:2009:BMI

[Bar09b] John H. Barnhill. Bringing Down the Mountains: The Impact of Mountaintop Removal on Southern West Virginia Communities (review). Technology and Culture, 50(3):701–703, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270556.

Basalla:2000:MTR

[Bas00] George Basalla. Meaning in Technology (review). Technology and Culture, 41(3):575–576, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33549.

Basalla:2002:TIE

[Bas02] George Basalla. Technological Innovation as an Evolutionary Process (review). Technology and Culture, 43(2):403–404, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33899.

Basalla:2005:MGS

[Bas05] George Basalla. Matters of Gravity: Special Effects and Superman in the 20th Century (review). Technology and Culture, 46 (3):684–685, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186603.

Basalla:2007:TSF

[Bas07a] George Basalla. Technophobia! Science Fiction Visions of Posthuman Technology (review). Technology and Culture, 48 (4):904–906, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/223344.

Basmajian:2007:SCH

[Bas07b] Carlton Basmajian. Sprawl: A Compact History (review). Technology and Culture, 48(2):422–424, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215377.

Bassett:2007:MTC

[Bas07c] Ross Knox Bassett. The mysteries of the trade: Christophe Lécuyer’s making silicon valley. Technology and Culture, 48 (2):401–403, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215378.

Basalla:2009:HSM

[Bas09a] George Basalla. Hollywood Science: Movies, Science, and the End of the World (review). Technology and Culture, 50(2):473– 474, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263344.

Bassett:2009:AIC

[Bas09b] Ross Bassett. Aligning India in the Cold War Era: Indian technical elites, the Indian Institute of Technology at Kanpur, and computing in India and the United States. Technology and Culture, 50(4):783–810, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/363414.

Bassett:2009:IAH

[Bas09c] Ross Bassett. Internet Alley: High Technology in Tysons Corner, 1945–2005 (review). Technology and Culture, 50(2):483– 484, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263340.

Bijker:2000:WWT

[BB00] Wiebe E. Bijker and Karin Bijsterveld. Women walking through plans: Technology, democracy, and gender identity. Technology and Culture, 41(3):485–515, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33550.

Brown:2009:NEE

[BDD09] John K. Brown, Gary Lee Downey, and Maria Paula Diogo. The normativities of engineers: Engineering education and history of technology. Technology and Culture, 50(4):737–752, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363412.

Bix:2004:RMK

[BE04] Amy Sue Bix and Taner Edis. The Rahmi M. Koç Museum, Istanbul. Technology and Culture, 45(3):590–596, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172965.

Beamish:2002:FOM

[Bea02] Thomas D. Beamish. Failure Is not an Option: Mission Control from Mercury to Apollo 13 and Beyond (review). Technology and Culture, 43(3):645–646, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33950.

Bednarek:2002:PWW

[Bed02] Janet R. Daly Bednarek. The Published Writings of Wilbur and Orville Wright (review). Technology and Culture, 43(3):639– 640, July 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33951.

Bednarek:2003:AAM

[Bed03] Janet R. Daly Bednarek. Airlines and Air Mail: The Post Office and the Birth of the Commercial Aviation Industry (review). Technology and Culture, 44(1):182–183, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/40143.

Bednarek:2005:AAA

[Bed05a] Janet R. Daly Bednarek. America from the Air: An Aviator’s Story (review). Technology and Culture, 46(1):227–228, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179748.

Bednarek:2005:FMG

[Bed05b] Janet R. Daly Bednarek. The flying machine in the garden: Parks and airports, 1918–1938. Technology and Culture, 46 (2):350–373, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184305.

Bektas:2000:SMC

[Bek00] Yakup Bektas. The Sultan’s messenger: Cultural constructions of Ottoman telegraphy, 1847–1880. Technology and Culture, 41(4):669–696, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33600.

Bektas:2002:WWP

[Bek02] Yakup Bektas. To Wire the World: Perry M. Collins and the North Pacific Telegraph Expedition (review). Technology and Culture, 43(1):189–191, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33856.

Bektas:2003:HTT

[Bek03] Yakup Bektas. A History of Telegraphy: Its Technology and Application (review). Technology and Culture, 44(2):396–397, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42007.

Bektas:2004:DTG

[Bek04a] Yakup Bektas. Distant Ties: Germany, the Ottoman Empire, and the Construction of the Baghdad Railway (review). Technology and Culture, 45(4):872–874, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175734.

Bektas:2004:MEM

[Bek04b] Yakup Bektas. Mapping Early Modern Japan: Space, Place, and Culture in the Tokugawa Period, 1603–1868 (review). Technology and Culture, 45(1):175–176, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51737.

Bektas:2004:PPM

[Bek04c] Yakup Bektas. Possessors and Possessed: Museums, Archaeology, and the Visualization of History in the Late Ottoman Empire (review). Technology and Culture, 45(3):666–668, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172963.

Bektas:2005:ECC

[Bek05] Yakup Bektas. The European Cable Companies in South America before the First World War (review). Technology and Culture, 46(4):828–830, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/192283.

Bektas:2006:GSM

[Bek06] Yakup Bektas. Guns for the Sultan: Military Power and the Weapons Industry in the Ottoman Empire (review). Technology and Culture, 47(2):405–407, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200836.

Bektas:2007:GTM

[Bek07] Yakup Bektas. “The German–Turk Miracle”: Arnold Reisman’s Turkey’s Modernization. Technology and Culture, 48(4): 840–843, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223345.

Belasco:2001:RDF

[Bel01] Warren James Belasco. Roadside dreams, fast food nightmares. Technology and Culture, 42(4):767–770, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33801.

Belfanti:2004:GPC

[Bel04a] Carlo Belfanti. Guilds, patents, and the circulation of technical knowledge: Northern Italy during the Early Modern Age. Technology and Culture, 45(3):569–589, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/172964.

Beltran:2004:BUH

[Bel04b] Alain Beltran. Des barrages, des usines, et des hommes : L’industrialisation des Alpes du nord entre ressources locales et apports exterieurs (review). Technology and Culture, 45(4): 879–880, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175735.

Belfanti:2006:ELI

[Bel06] Marco Belfanti. The European Linen Industry in Historical Perspective (review). Technology and Culture, 47(1):193–195, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196029.

Belfanti:2008:CGE

[Bel08] Carlo Marco Belfanti. Craft Guilds in the Early Modern Low Countries: Work, Power, and Representation (review). Technology and Culture, 49(4):1049–1050, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255454.

Benhamou:2000:CQE

[Ben00] Reed Benhamou. Cheap, Quick, and Easy: Imitative Architectural Materials, 1870–1930 (review). Technology and Culture, 41(2):396–398, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33512.

Benhamou:2001:HET

[Ben01a] Reed Benhamou. A History of Everyday Things: The Birth of Consumption in France, 1600–1800 (review). Technology and Culture, 42(2):343–345, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33704.

Benhamou:2001:MTD

[Ben01b] Reed Benhamou. Materials and Techniques in the Decorative Arts: An Illustrated Dictionary (review). Technology and Culture, 42(3):571–573, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33749.

Benhamou:2001:TTS

[Ben01c] Reed Benhamou. Technique and Technology: Script, Print, and Poetics in France, 1470–1550 (review). Technology and Culture, 42(4):780–782, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33802.

Benhamou:2001:MDR

[Ben01d] Reed Benhamou. The Making of the Dentiste, 1650–1760 (review). Technology and Culture, 42(1):200–202, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33657.

Benhamou:2002:FWS

[Ben02] Reed Benhamou. Fabricating Women: The Seamstresses of Old Regime France, 1675–1791 (review). Technology and Culture, 43(3):593–595, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33952.

Benhamou:2003:MAV

[Ben03] Reed Benhamou. La modernite au village: Tignes, Savines, Ubaye. … La submersion de communes rurales au nom de l’intérêt general, 1920–1970 (review). Technology and Culture, 44(4):812–814, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49681.

Benhamou:2005:DWT

[Ben05a] Reed Benhamou. By Design: Why There Are No Locks on the Bathroom Doors in the Hotel Louis XIV and Other Object Lessons (review). Technology and Culture, 46(3):630–631, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186604.

Benhamou:2005:WSC

[Ben05b] Reed Benhamou. Where Stuff Comes From: How Toasters, Toilets, Cars, Computers and Many Other Things Come to Be as They Are (review). Technology and Culture, 46(2):471– 473, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184306.

Benhamou:2008:NEW

[Ben08] Reed Benhamou. The Needle’s Eye: Women and Work in the Age of Revolution (review). Technology and Culture, 49(1): 268–270, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230895.

Benhamou:2009:MIR

[Ben09] Reed Benhamou. The Modern Interior (review). Technology and Culture, 50(4):954–955, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/363443.

Berger:2001:DGM

[Ber01] Molly W. Berger. Domestic Goods: The Material, the Moral, and the Economic in the Postwar Years (review). Technology and Culture, 42(3):585–587, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33750.

Berger:2002:ICS

[Ber02] Molly W. Berger. The Invention of Comfort: Sensibilities and Design in Early Modern Britain and Early America (review). Technology and Culture, 43(4):795–796, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/34006.

Berendsen:2003:MMW

[Ber03a] Marja Berendsen. “Mrs. Modern Woman”: Zur Sozial- und Kulturgeschichte der Haushaltstechnisierung (review). Technology and Culture, 44(4):835–836, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49682.

Berger:2003:BCW

[Ber03b] Molly W. Berger. Building the Cold War: Hilton International Hotels and Modern Architecture (review). Technology and Culture, 44(1):205–207, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40144.

Berger:2004:CCA

[Ber04] Molly W. Berger. Cool Comfort: America’s Romance with Air-Conditioning (review). Technology and Culture, 45(2):452– 454, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55655.

Bernhardt:2006:IWG

[Ber06] Christoph Bernhardt. Istanbul und das Wasser: Zur Geschichte der Wasserversorgung und Abwasserentsorgung von der Mitte des 19. Jahrhunderts bis 1966 (review). Technology and Culture, 47(2):407–408, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200837.

Berner:2007:RTW

[Ber07] Boel Berner. Rationalizing technical work: Visions and realities of the “Systematic drawing office” in Sweden, 1890–1940. Technology and Culture, 48(1):20–42, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209821.

Berger:2008:MSL

[Ber08] Molly W. Berger. Manon of second life: Teaching in the virtual world. Technology and Culture, 49(2):430–441, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236967.

Bess:2008:IHP

[Bes08] Michael Bess. Icarus 2.0: A historian’s perspective on human biological enhancement. Technology and Culture, 49(1): 114–126, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230896.

Beyler:2000:WNT

[Bey00] Richard H. Beyler. Walther Nernst and the Transition to Modern Physical Science (review). Technology and Culture, 41(2): 395–396, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33513.

Beyler:2006:WDM

[Bey06] Richard H. Beyler. Walther von Dyck (1856–1934): Mathematik, Technik und Wissenschaftsorganisation an der TH Munchen (review). Technology and Culture, 47(1):233–234, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196030.

Beyler:2007:DEV

[Bey07] Richard H. Beyler. Durchsicht, Einsicht, Vorsicht: Eine Geschichte der Röntgenstrahlen, 1896–1963 (review). Technology and Culture, 48(1):198–200, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209822.

Beyler:2008:MAM

[Bey08] Richard H. Beyler. The Mental Aftermath: The Mentality of German Physicists, 1945–1949 (review). Technology and Culture, 49(2):484–486, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236972.

Beyler:2009:EOM

[Bey09] Richard H. Beyler. Einstein and Oppenheimer: The Meaning of Genius (review). Technology and Culture, 50(3):722– 723, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270547.

Biggs:2000:WPF

[Big00] Lindy Biggs. World of Possibilities: Flexibility and Mass Production in Western Industrialization (review). Technology and Culture, 41(2):366–368, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33514.

Biggs:2001:TRE

[Big01] Lindy Biggs. Technological Revolutions in Europe: Historical Perspectives (review). Technology and Culture, 42(2):335– 336, April 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33705.

Biggs:2002:MSI

[Big02] Lindy Biggs. Making Sense of the Industrial Revolution: English Economy and Society, 1700–1850 (review). Technology and Culture, 43(3):595–596, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33953.

Biggs:2007:TBM

[Big07] Lindy Biggs. The Taylorized Beauty of the Mechanical: Scientific Management and the Rise of Modernist Architecture (review). Technology and Culture, 48(3):643–644, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218694.

Biggs:2008:BCM

[Big08] David Biggs. Breaking from the colonial mold: Water engineering and the failure of nation-building in the Plain of Reeds, Vietnam. Technology and Culture, 49(3):599–623, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243118.

Bijker:2002:OSS

[Bij02] Wiebe E. Bijker. The Oosterschelde Storm Surge Barrier: a test case for Dutch water technology, management, and politics. Technology and Culture, 43(3):569–584, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33954.

Bijsterveld:2003:RAD

[Bij03] Karin Bijsterveld. Redesigning the American Dream: Gender, Housing, and Family Life (review). Technology and Culture, 44(4):833–835, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49678.

Bijsterveld:2004:APC

[Bij04] Karin Bijsterveld. The Audible Past: Cultural Origins of Sound Reproduction (review). Technology and Culture, 45(2):441– 442, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55656.

Bijsterveld:2005:ECM

[Bij05] Karin Bijsterveld. Electronic and Computer Music (review). Technology and Culture, 46(4):867–868, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192284.

Bijker:2009:GVC

[Bij09] Wiebe E. Bijker. Globalization and vulnerability: Challenges and opportunities for SHOT around its fiftieth anniversary. Technology and Culture, 50(3):600–612, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270536.

Bilstein:2002:AWD

[Bil02] Roger E. Bilstein. Around the World in 175 Days: The First Round-the-World Flight (review). Technology and Culture, 43 (4):817–819, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34004.

Bilstein:2005:C

[Bil05] Roger E. Bilstein. Communications. Technology and Culture, 46(1):263–264, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/179749.

Bingle:2006:SBP

[Bin06] Gwen Bingle. I sing the body prosthetic. Technology and Culture, 47(3):578–583, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/201439.

Bishop:2005:REE

[Bis05] Elizabeth Bishop. Rule of Experts: Egypt, Techno-Politics, Modernity (review). Technology and Culture, 46(4):876–878, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192285.

Bivins:2003:CMC

[Biv03] Roberta E. Bivins. Chinese Medicine in Contemporary China: Plurality and Synthesis (review). Technology and Culture, 44 (3):606–607, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46081.

Bivins:2007:CPA

[Biv07] Roberta E. Bivins. Colonial Pathologies: American Tropical Medicine, Race, and Hygiene in the Philippines, and: The Cultivation of Whiteness: Science, Health, and Racial Destiny in Australia (review). Technology and Culture, 48(4):866–868, October 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/223346.

Bix:2002:ELG

[Bix02] Amy Sue Bix. Equipped for life: Gendered technical training and consumerism in home economics, 1920–1980. Technology and Culture, 43(4):728–754, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34007.

Bix:2004:SMC

[Bix04] Amy Sue Bix. Selling Mrs. Consumer: Christine Frederick and the Rise of Household Efficiency (review). Technology and Culture, 45(4):906–907, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175736.

Blaszczyk:2000:ACA

[Bla00a] Regina Lee Blaszczyk. Air-Conditioning America: Engineers and the Controlled Environment, 1900–1960 (review). Technology and Culture, 41(3):593–595, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33551.

Blaszczyk:2000:MLH

[Bla00b] Regina Lee Blaszczyk. The Material Life of Human Beings: Artifacts, Behavior, and Communication (review). Technology and Culture, 41(4):793–795, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33602.

Black:2001:OIC

[Bla01] Brian Black. Oil and Ideology: The Cultural Creation of the American Petroleum Industry (review). Technology and Culture, 42(1):143–144, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33658.

Blanchard:2003:TMF

[Bla03] Ian Blanchard. Le temps des moulines: Fer, technique et société dans les Pyrénées centrales, XIIIe–XVIe siècles (review). Technology and Culture, 44(4):807–808, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49683.

Black:2004:GUH

[Bla04a] Brian Black. Giant Under the Hill: A History of the Spindletop Oil Discovery at Beaumont, Texas, in 1901 (review). Technology and Culture, 45(1):191–192, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51738.

Blaszczyk:2004:AHO

[Bla04b] Regina Lee Blaszczyk. The Age of Homespun: Objects and Stories in the Creation of an American Myth (review). Technology and Culture, 45(1):186–188, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51739.

Blaszczyk:2005:SSH

[Bla05] Regina Lee Blaszczyk. The Substance of Style: How the Rise of Aesthetic Value is Remaking Commerce, Culture, and Consciousness (review). Technology and Culture, 46(1):261–262, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179750.

Blaine:2008:WWM

[Bla08] Bradford Blaine. Wind and Water in the Middle Ages: Fluid Technologies from Antiquity to the Renaissance (review). Technology and Culture, 49(1):261–263, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230897.

Black:2009:VAS

[Bla09a] Brian Black. Virtual America: Sleepwalking through Paradise (review). Technology and Culture, 50(4):961–962, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363447.

Blackmar:2009:CWL

[Bla09b] Elizabeth Blackmar. Corporate Wasteland: The Landscape and Memory of Deindustrialization (review). Technology and Culture, 50(1):228–230, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/258589.

Blanke:2009:MMS

[Bla09c] David Blanke. Mobility without Mayhem: Safety, Cars, and Citizenship (review). Technology and Culture, 50(1):240–242, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258586.

Blewett:2004:RAI

[Ble04] Mary H. Blewett. The Roots of American Industrialization (review). Technology and Culture, 45(4):837–838, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175730.

Blewett:2007:CTL

[Ble07] Mary H. Blewett. A Common Thread: Labor, Politics, and Capital Mobility in the Textile Industry (review). Technology and Culture, 48(4):857–859, October 2007. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223347.

Blumtritt:2001:RHW

[Blu01] Oskar Blumtritt. A Radar History of World War II: Technical and Military Imperatives (review). Technology and Culture, 42 (2):372–373, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33706.

Blum:2004:RLD

[Blu04] Martina Blum. Raymond Loewy: Designs for a Consumer Culture (review). Technology and Culture, 45(4):854–855, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175737.

Bengoetxea:2007:HEE

[BM07] Juan Bautista Bengoetxea and Carl Mitcham. Historia: Empiricism and Erudition in Early Modern Europe (review). Technology and Culture, 48(3):617–618, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218693.

Boch:2003:TNP

[Boc03] Ralph Boch. Topografien der Nation: Politik, kartografische Ordnung, und Landschaft im 19. Jahrhundert (review). Technology and Culture, 44(4):817–819, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49684.

Bocking:2009:DGA

[Boc09] Stephen Bocking. A disciplined geography: Aviation, science, and the Cold War in Northern Canada, 1945–1960. Technology and Culture, 50(2):265–290, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263326.

Borg:2000:MMA

[Bor00] Kevin L. Borg. The Machine in Me: An Anthropologist Sits among Computer Engineers (review). Technology and Culture, 41(1):120–121, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33467.

Borg:2002:CCR

[Bor02] Kevin L. Borg. Car Cultures (review). Technology and Culture, 43(3):636–637, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33955.

Borde:2006:RSP

[Bor06a] Christian Borde. La Révolution sardinière: Pecheurs et conserveurs en Bretagne sud au XIXe siècle (review). Technology and Culture, 47(1):200–201, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196031.

Borg:2006:TCR

[Bor06b] Kevin L. Borg. Tinkering: Consumers Reinvent the Early Automobile (review). Technology and Culture, 47(2):436–438, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200838.

Borg:2008:HWP

[Bor08] Kevin L. Borg. Hell on Wheels: The Promise and Peril of America’s Car Culture, 1900–1940 (review). Technology and Culture, 49(2):466–467, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/236973.

Borg:2009:BSH

[Bor09a] Kevin Borg. The Business of Speed: The Hot Rod Industry in America, 1915–1990 (review). Technology and Culture, 50 (3):715–717, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270551.

Borg:2009:AMC

[Bor09b] Kevin L. Borg. Auto Mania: Cars, Consumers, and the Environment (review). Technology and Culture, 50(1):238–240, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258563.

Bowers:2000:BBL

[Bow00a] Brian Bowers. Brandy, Balloons and Lamps: Ami Argand, 1750–1803 (review). Technology and Culture, 41(4):812–814, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33603.

Bowers:2000:SEC

[Bow00b] Brian Bowers. The Science of Energy: A Cultural History of Energy Physics in Victorian Britain (review). Technology and Culture, 41(2):362–363, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33515.

Bowker:2000:CLG

[Bow00c] Geoffrey C. Bowker. Changing Life: Genomes, Ecologies, Bodies, Commodities (review). Technology and Culture, 41(1): 179–181, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33468.

Bowers:2003:WMB

[Bow03] Brian Bowers. Wireless: From Marconi’s Black Box to the Audion (review). Technology and Culture, 44(2):398–399, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
42008.

Bowker:2004:NCH

[Bow04] Geoffrey C. Bowker. From Newspeak to Cyberspeak: A History of Soviet Cybernetics (review). Technology and Culture, 45(1): 208–209, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51740.

Bowker:2005:DSM

[Bow05] Geoffrey C. Bowker. The Digital Sublime: Myth, Power, and Cyberspace (review). Technology and Culture, 46(2):436–437, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184307.

Bowker:2007:ECH

[Bow07] Geoffrey C. Bowker. An Engine, Not a Camera: How Financial Models Shape Markets (review). Technology and Culture, 48(4): 906–907, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223348.

Bowers:2008:BSN

[Bow08] Brian Bowers. Brilliant! Shuji Nakamura and the Revolution in Lighting Technology (review). Technology and Culture, 49 (2):479–481, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236974.

Boyd:2001:MMS

[Boy01] William Boyd. Making meat: Science, technology, and American poultry production. Technology and Culture, 42(4):631– 664, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33803.

Boyd:2003:AMT

[Boy03] William Boyd. Agriculture and Modern Technology: A Defense (review). Technology and Culture, 44(2):423–424, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42009.

Bijker:2002:SAO

[BP02] Wiebe E. Bijker and T. J. Pinch. SCOT answers, other questions: A reply to Nick Clayton. Technology and Culture, 43 (2):361–369, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33900. See [Cla02c].

Braden:2001:FCT

[Bra01a] Donna R. Braden. From Fireplace to Cookstove: Technology and the Domestic Ideal in America (review). Technology and Culture, 42(4):809–810, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33804.

Braun:2001:TEM

[Bra01b] Hans-Joachim Braun. Theremin: Ether Music and Espionage (review). Technology and Culture, 42(4):802–803, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33805.

Braden:2002:LDH

[Bra02] Donna R. Braden. Lengthening the Day: A History of Lighting Technology (review). Technology and Culture, 43(1):193–194, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33857.

Braden:2003:THA

[Bra03a] Donna R. Braden. On Time: How America Has Learned to Live by the Clock (review). Technology and Culture, 44(3):640– 642, July 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/46082.

Braun:2003:ADI

[Bra03b] Hans-Joachim Braun. Analog Days: The Invention and Impact of the Moog Synthesizer (review). Technology and Culture, 44 (3):632–634, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46083.

Braun:2003:SSM

[Bra03c] Hans-Joachim Braun. Strange Sounds: Music, Technology, and Culture (review). Technology and Culture, 44(2):401–402, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
42010.

Bray:2004:SHS

[Bra04] Francesca Bray. Situating the History of Science: Dialogues with Joseph Needham (review). Technology and Culture, 45(1): 173–175, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51741.

Brattin:2005:PII

[Bra05] Joel J. Brattin. Patent Inventions: Intellectual Property and the Victorian Novel (review). Technology and Culture, 46(3): 649–650, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186605.

Braun:2007:STW

[Bra07a] Lindsay Frederick Braun. From Squaw Tit to Whorehouse Meadow: How Maps Name, Claim, and Inflame (review). Technology and Culture, 48(2):418–419, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215379.

Braun:2007:CLC

[Bra07b] Marta Braun. A Culture of Light: Cinema and Technology in 1920s Germany (review). Technology and Culture, 48(2):435– 436, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215380.

Braden:2008:RTT

[Bra08a] Donna R. Braden. A Republic in Time: Temporality and Social Imagination in Nineteenth-Century America (review). Technology and Culture, 49(4):1076–1077, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255499.

Brandt:2008:IMC

[Bra08b] Thomas Brandt. Industry and Modernism: Companies, Architecture, and Identity in the Nordic and Baltic Countries during the High-Industrial Period (review). Technology and Culture, 49 (4):1092–1094, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255535.

Braun:2008:BSE

[Bra08c] Marta Braun. Body Shots: Early Cinema’s Incarnations (review). Technology and Culture, 49(4):1082–1084, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
255500.

Bresciani:2000:TWA

[Bre00] Luis Paulo Bresciani. Taking the Wheel: Auto Parts Firms and the Political Economy of Industrialization in Brazil (review). Technology and Culture, 41(4):806–807, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33604.

Breyman:2001:CMO

[Bre01] Steve Breyman. Critical Masses: Opposition to Nuclear Power in California, 1958–1978 (review). Technology and Culture, 42(1):141–143, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33659.

Brende:2006:C

[Bre06] Eric Brende. Communications. Technology and Culture, 47 (3):704–706, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201440.

Brittain:2001:VSE

[Bri01] James E. Brittain. A Victorian Scientist and Engineer: Fleeming Jenkin and the Birth of Electrical Engineering (review). Technology and Culture, 42(1):170–171, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33660.

Brimblecombe:2006:SCE

[Bri06] Peter Brimblecombe. Shock Cities: The Environmental Transformation and Reform of Manchester and Chicago (review). Technology and Culture, 47(3):665–666, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/201441.

Britton:2007:CPI

[Bri07] John A. Britton. “The Confusion Provoked by Instantaneous Discussion”: The new international communications network and the Chilean crisis of 1891–1892 in the United States. Technology and Culture, 48(4):729–757, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223349.

Britton:2008:PNP

[Bri08] John A. Britton. From Pigeons to News Portals: Foreign Reporting and the Challenge of New Technology (review). Technology and Culture, 49(4):1074–1076, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255501.

Brooks:2000:SWC

[Bro00a] Randall C. Brooks. The Science of War: Canadian Scientists and Allied Military Technology during the Second World War (review). Technology and Culture, 41(3):630–631, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33552.

Brown:2000:CFD

[Bro00b] John K. Brown. Cotton, Fire, and Dreams: The Robert Findlay Iron Works and Heavy Industry in Macon, Georgia, 1839– 1912 (review). Technology and Culture, 41(1):143–145, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33469.

Brown:2000:DPW

[Bro00c] John K. Brown. Design plans, working drawings, national styles: Engineering practice in Great Britain and the United States, 1775–1945. Technology and Culture, 41(2):195–238, April 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33516.

Broeze:2001:AOH

[Bro01] Frank Broeze. Abandoned Ocean: A History of United States Maritime Policy (review). Technology and Culture, 42(3):611– 612, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33751.

Brown:2002:FAW

[Bro02] Julie K. Brown. Fair America: World’s Fairs in the United States (review). Technology and Culture, 43(2):429–431, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33901.

Brown:2003:LCH

[Bro03] John K. Brown. Louis C. Hunter, Steamboats on the Western Rivers. Technology and Culture, 44(4):786–793, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49685.

Brown:2005:CTM

[Bro05a] John K. Brown. The Carriage Trade: Making Horse-Drawn Vehicles in America (review). Technology and Culture, 46(3):677– 679, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186606.

Brown:2005:PCL

[Bro05b] Julie K. Brown. Photography on the Color Line: W. E. B. DuBois, Race, and Visual Culture (review). Technology and Culture, 46(3):687–688, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/186607.

Brown:2005:RSE

[Bro05c] Julie K. Brown. River of Shadows: Eadweard Muybridge and the Technological Wild West (review). Technology and Culture, 46(2):426–427, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184308.

Brown:2006:DNH

[Bro06a] Alexander Brown. Defining NASA: The Historical Debate over the Agency’s Mission (review). Technology and Culture, 47(4): 867–869, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204968.

Brown:2006:PMR

[Bro06b] John K. Brown. Picturing Machines, 1400–1700 (review). Technology and Culture, 47(1):186–188, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196032.

Brown:2006:CEP

[Bro06c] Julie K. Brown. The Corporate Eye: Photography and the Rationalization of American Commercial Culture (review).

Technology and Culture, 47(3):648–649, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/201442.

Brown:2007:NES

[Bro07] John K. Brown. The New England Steamship Company: Long Island Sound Night Boats in the Twentieth Century (review). Technology and Culture, 48(2):466–468, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215381.

Brosnan:2008:POV

[Bro08] Kathleen A. Brosnan. Playing the Odds: Las Vegas and the Modern West (review). Technology and Culture, 49(4):1059– 1060, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255455.

Brown:2009:AES

[Bro09a] John K. Brown. Architect and Engineer: A Study in Sibling Rivalry (review). Technology and Culture, 50(2):456–457, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263341.

Brown:2009:MMP

[Bro09b] Julie K. Brown. Medicine’s Moving Pictures: Medicine, Health, and Bodies in American Film and Television (review). Technology and Culture, 50(2):474–476, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263339.

Brown:2009:AAS

[Bro09c] Rowan Julie Brown. Art in the Age of Steam: Walker Art Gallery, Liverpool, 18 April–10 August 2008. Technology and Culture, 50(2):418–425, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/263334.

Brugger:2001:EBR

[Bru01] Robert J. Brugger. The Evolution of the Book (review). Technology and Culture, 42(4):778–780, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33806.

Bryant:2007:SLS

[Bry07] Keith L. Bryant. Sunset Limited: The Southern Pacific Railroad and the Development of the American West, 1850–1930 (review). Technology and Culture, 48(1):213–214, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209823.

Bezis-Selfa:2009:BCL

[BS09] John Bezis-Selfa. Backcountry Crucibles: The Lehigh Valley from Settlement to Steel (review). Technology and Culture, 50 (1):223–225, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/258567.

Buchanan:2001:GEC

[Buc01a] R. A. Buchanan. Gentlemen Engineers: The Careers of Frank and Walter Shanly (review). Technology and Culture, 42(1): 165–166, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33661.

Buchanan:2001:LSR

[Buc01b] R. A. Buchanan. The Lighthouse Stevensons (review). Technology and Culture, 42(3):603–604, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33752.

Buchanan:2003:IST

[Buc03a] R. A. Buchanan. Iron Shipbuilding on the Thames, 1832– 1915: An Economic and Business History (review). Technology and Culture, 44(3):610–612, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46084.

Buchanan:2003:LIR

[Buc03b] R. A. Buchanan. Life and Inventions of Richard Roberts, 1789– 1864 (review). Technology and Culture, 44(4):819–820, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
49686.

Buchanan:2005:GAB

[Buc05a] Brenda J. Buchanan. Gunpowder: Alchemy, Bombards, and Pyrotechnics — The History of the Explosive that Changed the World (review). Technology and Culture, 46(2):405–406, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184309.

Buchanan:2005:ERR

[Buc05b] R. A. Buchanan. Early Railways 2 (review). Technology and Culture, 46(1):207–208, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179751.

Bud:2002:BCB

[Bud02] Robert Bud. Biotechnology and Culture: Bodies, Anxieties, Ethics (review). Technology and Culture, 43(4):839–840, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34009.

Bud:2005:ABB

[Bud05] Robert Bud. Antibiotics, big business, and consumers: The context of government investigations into the postwar American drug industry. Technology and Culture, 46(2):329–349, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184310.

Bud:2006:PDR

[Bud06] Robert Bud. Pharmacopolitics: Drug Regulation in the United States and Germany (review). Technology and Culture, 47(1): 231–232, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196033.

Bud:2007:ISS

[Bud07] Robert Bud. Interferon: The Science and Selling of a Miracle Drug (review). Technology and Culture, 48(1):197–198, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209824.

Bud:2009:CPI

[Bud09] Robert Bud. Collaboration in the Pharmaceutical Industry: Changing Relationships in Britain and France, 1935– 1965 (review). Technology and Culture, 50(4):975–976, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363455.

Bugos:2001:WWW

[Bug01] Glenn E. Bugos. Wings of Wood, Wings of Metal: Culture and Technical Choice in American Airplane Materials, 1914–1945 (review). Technology and Culture, 42(2):370–372, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33707.

Bugos:2004:CSB

[Bug04] Glenn E. Bugos. Chasing the Silver Bullet: U.S. Air Force Weapons Development from Vietnam to Desert Storm (review). Technology and Culture, 45(3):630–631, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/172966.

Bugos:2005:RHF

[Bug05] Glenn E. Bugos. To Reach the High Frontier: A History of U.S. Launch Vehicles (review). Technology and Culture, 46(1): 233–234, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179752.

Bugos:2007:MPA

[Bug07] Glenn E. Bugos. Major Problems in American Business History: Documents and Essays (review). Technology and Culture, 48(4):907–909, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/223350.

Bullough:2001:BTR

[Bul01] Vern L. Bullough. Body Talk: Rhetoric, Technology, Reproduction (review). Technology and Culture, 42(1):199–200, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33662.

Bulliet:2003:PBP

[Bul03a] Richard W. Bulliet. Paper before Print: The History and Impact of Paper in the Islamic World (review). Technology and Culture, 44(1):160–161, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40145.

Bullough:2003:IAG

[Bul03b] Vern L. Bullough. Infertility around the Globe: New Thinking on Childlessness, Gender, and Reproductive Technologies (review). Technology and Culture, 44(4):852–853, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49687.

Burt:2000:CAU

[Bur00a] Roger Burt. Copper for America: The United States Copper Industry from Colonial Times to the 1990s (review). Technology and Culture, 41(1):138–139, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33470.

Burt:2000:IIT

[Bur00b] Roger Burt. Innovation or imitation? Technological dependency in the American nonferrous mining industry. Technology and Culture, 41(2):321–347, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33517.

Burnette:2002:FPE

[Bur02] Joyce Burnette. Farm Production in England, 1700–1914 (review). Technology and Culture, 43(2):416–418, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33902.

Burns:2006:GCM

[Bur06] Robert Ignatius Burns. Gegna: Coastal mooring in crusader Valencia. Technology and Culture, 47(4):777–786, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
204969.

Bryan-Wilson:2004:VAI

[BW04] Julia Bryan-Wilson. Virtual Art: From Illusion to Immersion (review). Technology and Culture, 45(3):670–671, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172961.

Byers:2004:DFG

[Bye04] Richard Byers. Dreams of Flight: General Aviation in the United States (review). Technology and Culture, 45(3):629– 630, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172967.

Carey:2002:RNC

[Car02] Elaine Carey. Radio Nation: Communication, Popular Culture, and Nationalism in Mexico, 1920–1950 (review). Technology and Culture, 43(3):625–626, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33956.

Carlisle:2003:TID

[Car03] Rodney P. Carlisle. Tritium on Ice: The Dangerous New Alliance of Nuclear Weapons and Nuclear Power (review). Technology and Culture, 44(4):846–848, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49688.

Cardoza:2008:GWT

[Car08] Anthony L. Cardoza. In Gold We Trust: Social Capital and Economic Change in the Italian Jewelry Towns (review). Technology and Culture, 49(4):1091–1092, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255502.

Caruso:2009:SSO

[Car09] David Caruso. Sentenced to Science: One Black Man’s Story of Imprisonment in America (review). Technology and Culture, 50(4):972–973, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363453.

Castells:2003:CCE

[Cas03] Manuel Castells. The cultural crisis of engineering in the Information Age: Rosalind Williams’s Retooling. Technology and Culture, 44(3):586–590, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46085.

Casey:2004:AMN

[Cas04a] Robert Casey. America on the move at the National Museum of American History. Technology and Culture, 45(4):812–816, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175738.

Casey:2004:UA

[Cas04b] Robert Casey. An unprepossessing aircraft. Technology and Culture, 45(2):476–478, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/55657.

Casey:2005:LBF

[Cas05] Robert Casey. The line between faith and folly. Technology and Culture, 46(2):475–478, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184311.

Castonguay:2007:TDA

[Cas07] Stéphane Castonguay. Technology’s Dilemma: Agricultural Colleges between Science and Practice in Germany, 1860– 1934 (review). Technology and Culture, 48(2):433–435, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215382.

Castagnaro:2008:PPE

[Cas08] Mario Joseph Castagnaro. From Parsifal to Perón: Early Radio in Argentina, 1920–1944 (review). Technology and Culture, 49(2):501–503, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236975.

Cavicchi:2005:TKP

[Cav05] Elizabeth Cavicchi. Thing Knowledge: A Philosophy of Scientific Instruments (review). Technology and Culture, 46(1): 243–245, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179744.

Cavicchi:2008:BES

[Cav08a] Elizabeth Cavicchi. Berlin Electropolis: Shock, Nerves, and German Modernity (review). Technology and Culture, 49(1): 256–257, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230898.

Cavicchi:2008:CGP

[Cav08b] Elizabeth Cavicchi. Charles Grafton Page’s experiment with a spiral conductor. Technology and Culture, 49(4):884–907, October 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/255440.

Cephas:2008:ARS

[Cep08] Jana Cephas. Alternative Routes to the Sustainable City: Austin, Curitiba, and Frankfurt (review). Technology and Culture, 49(2):518–519, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236976.

Ceruzzi:2003:DVE

[Cer03] Paul E. Ceruzzi. Driving Visions: Exploring the Road Movie (review). Technology and Culture, 44(1):172–173, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40146.

Ceruzzi:2004:DAR

[Cer04] Paul E. Ceruzzi. To the Digital Age: Research Labs, Start-Up Companies, and the Rise of MOS Technology (review). Technology and Culture, 45(4):892–893, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175739.

Ceruzzi:2005:MLT

[Cer05] Paul E. Ceruzzi. Moore’s Law and technological determinism: Reflections on the history of technology. Technology and Culture, 46(3):584–593, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186608.

Ceruzzi:2006:JWH

[Cer06] Paul E. Ceruzzi. Jacquard’s Web: How a Hand-Loom Led to the Birth of the Information Age (review). Technology and Culture, 47(1):197–198, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196034.

Ceruzzi:2007:OTH

[Cer07] Paul E. Ceruzzi. This Is Only a Test: How Washington, D.C. Prepared for Nuclear War (review). Technology and Culture, 48(3):664–666, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218695.

Ceruzzi:2009:LCM

[Cer09] Paul E. Ceruzzi. “A Large Canvas”: Melvin Kranzberg and Carroll W. Pursell Jr., eds., Technology in Western Civilization. Technology and Culture, 50(3):658–668, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270532.

Curtin:2005:SSR

[CH05] Catherine Curtin and Joel D. Howell. Surgeons and the Scope (review). Technology and Culture, 46(3):690–691, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186612.

Chandra:2001:ASK

[Cha01] Siddharth Chandra. American Sugar Kingdom: The Plantation Economy of the Spanish Caribbean, 1898–1934 (review). Technology and Culture, 42(2):352–354, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33708.

Charles:2002:GTM

[Cha02] Daniel Charles. A Grain of Truth: The Media, the Public, and Biotechnology (review). Technology and Culture, 43(1): 215–217, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33858.

Charles:2007:BCO

[Cha07] Daniel Charles. Biotech: The Countercultural Origins of an Industry (review). Technology and Culture, 48(3):666–668, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218696.

Chenoweth:2004:WSU

[Che04] William L. Chenoweth. Warm Sands: Uranium Mill Tailings Policy in the Atomic West (review). Technology and Culture, 45(4):866–868, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175740.

Choi:2007:BIR

[Cho07] Hyungsub Choi. The boundaries of industrial research: Making transistors at RCA, 1948–1960. Technology and Culture, 48(4): 758–782, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223351.

Churella:2004:CPR

[Chu04a] Albert J. Churella. Capitalism, Politics, and Railroads in Jacksonian New England (review). Technology and Culture, 45(4): 846–847, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175742.

Churella:2004:MMM

[Chu04b] Albert J. Churella. Machine, monument, and metropolis: New York’s Pennsylvania Station at the Hagley Museum and Library. Technology and Culture, 45(4):817–823, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175741.

Churella:2007:AER

[Chu07] Albert J. Churella. Allies of the Earth: Railroads and the Soul of Preservation (review). Technology and Culture, 48(1): 211–213, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209825.

Ciesla:2001:PBD

[Cie01] Burghard Ciesla. Pkw-Bau in der DDR: Zur Innovationsschwache von Zentralverwaltungswirtschaften (review). Technology and Culture, 42(4):832–833, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33808.

Cioc:2006:ERT

[Cio06] Mark Cioc. Earth Repair: A Transatlantic History of Environmental Restoration (review). Technology and Culture, 47(4): 860–862, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204970.

Campbell-Kelly:2001:WIC

[CK01] Martin Campbell-Kelly. When Information Came of Age: Technologies of Knowledge in the Age of Reason and Revolution, 1700–1850 (review). Technology and Culture, 42(4): 773–774, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33807.

Campbell-Kelly:2004:HOI

[CK04] Martin Campbell-Kelly. A History of Online Information Services, 1963–1976 (review). Technology and Culture, 45(3):661– 662, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172969.

Clancey:2001:TIJ

[Cla01] Gregory K. Clancey. Technology and Innovation in Japan: Policy Management for the Twenty-First Century (review). Technology and Culture, 42(2):396–397, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33709.

Clarke:2002:NSF

[Cla02a] Alison J. Clarke. Nylon: The Story of a Fashion Revolution (review). Technology and Culture, 43(2):446–448, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33903.

Clayton:2002:R

[Cla02b] Nick Clayton. Rejoinder. Technology and Culture, 43(2):369– 370, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33905.

Clayton:2002:SDI

[Cla02c] Nick Clayton. SCOT: Does it answer? Technology and Culture, 43(2):351–360, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu. edu/pub/1/article/33904. See comment [Epp02b] and reply [BP02].

Clancey:2007:EIG

[Cla07] Gregory K. Clancey. Enduring Innocence: Global Architecture and its Political Masquerades (review). Technology and Culture, 48(1):177–179, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209826.

Clifton:2000:ISH

[Cli00] Gloria Clifton. Instruments of Science: An Historical Encyclopedia (review). Technology and Culture, 41(2):399–401, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33518.

Cloud:2003:NNM

[Clo03] John Cloud. The New Nature of Maps: Essays in the History of Cartography (review). Technology and Culture, 44(3):647– 649, July 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/46086.

Cloud:2005:HAF

[Clo05] John Cloud. Historical Atlases: The First Three Hundred Years, 1570–1870 (review). Technology and Culture, 46(3):646– 647, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186609.

Cloud:2006:CDM

[Clo06a] John Cloud. Cartographies of Disease: Maps, Mapping, and Medicine (review). Technology and Culture, 47(4):840–842, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204971.

Cloud:2006:COS

[Clo06b] John Cloud. Cultures in Orbit: Satellites and the Televisual (review). Technology and Culture, 47(3):646–648, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201443.

Cloud:2007:PGS

[Clo07a] John Cloud. Plotting the Globe: Stories of Meridians, Parallels, and the International Date Line, and: Rhumb Lines and Map Wars: A Social History of the Mercator Projection (review). Technology and Culture, 48(2):420–422, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215383.

Cloud:2007:DRB

[Clo07b] John Cloud. Discerning the relation between American science and American democracy. Technology and Culture, 48(3):589– 593, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218697.

Coates:2002:PLA

[Coa02] Peter A. Coates. Petrolia: The Landscape of America’s First Oil Boom (review). Technology and Culture, 43(1):187–189, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33859.

Coclanis:2003:SWS

[Coc03] Peter A. Coclanis. The Second Wave: Southern Industrialization from the 1940s to the 1970s (review). Technology and Culture, 44(3):625–626, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/46087.

Coclanis:2004:ICG

[Coc04] Peter A. Coclanis. Inventing the Cotton Gin: Machine and Myth in Antebellum America (review). Technology and Culture, 45(4):834–835, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175743.

Cohen:2007:IIS

[Coh07] Emily Jane Cohen. Of iron and irony: Starling Lawrence’s The Lightning Keeper. Technology and Culture, 48(2):393– 400, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215384.

Collins:2001:LLH

[Col01] Theresa M. Collins. The Life and Legend of E.H. Harriman (review). Technology and Culture, 42(3):601–603, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33753.

Coleman:2003:AAY

[Col03] Gabriella Coleman. Artifacts: An Archaeologist’s Year in Silicon Valley (review). Technology and Culture, 44(1):196–197, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40147.

Colten:2004:DRE

[Col04] Craig E. Colten. Devastation and Renewal: An Environmental History of Pittsburgh and Its Region (review). Technology and Culture, 45(3):640–641, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172968.

Cole:2005:IRH

[Col05] Simon A. Cole. Imprint of the Raj: How Fingerprinting was Born in Colonial India (review). Technology and Culture, 46(1): 252–253, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179753.

Colten:2006:RCC

[Col06] Craig E. Colten. Rusting of the chemical corridor. Technology and Culture, 47(1):95–101, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196035.

Cole:2007:IPP

[Col07a] Simon A. Cole. Injury: The Politics of Product Design and Safety Law in the United States (review). Technology and Culture, 48(2):450–451, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/215385.

Collins:2007:PPM

[Col07b] Steven G. Collins. Planters’ Progress: Modernizing Confederate Georgia (review). Technology and Culture, 48(1):201–203, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209827.

Collins:2007:BCW

[Col07c] Steven G. Collins. The Business of Civil War: Military Mobilization and the State, 1861–1865 (review). Technology and Culture, 48(2):453–455, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/215386.

Cole:2008:LDH

[Col08] Simon A. Cole. The Lie Detectors: The History of an American Obsession (review). Technology and Culture, 49(1):254–255, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230899.

Constant:1999:RKU

[Con99] Edward W. Constant. Reliable knowledge and unreliable stuff: On the practical role of rational beliefs. Technology and Culture, 40(2):324–357, April 1999. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33365. See comment [Scr00].

Conrad:2000:AEM

[Con00a] James Lawson Conrad. Artisans, Entrepreneurs, and Machines: Essays on the Early Anglo-American Textile Industries, 1770–1840s (review). Technology and Culture, 41(4):799–801, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33605.

Constant:2000:PMT

[Con00b] Edward W. Constant. Performance is a moving target, reliably. Technology and Culture, 41(4):776–782, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33606.

Conway:2001:CSA

[Con01a] Erik M. Conway. Creating Safety in Air Traffic Control (review). Technology and Culture, 42(3):613–615, July 2001. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33754.

Conway:2001:PBL

[Con01b] Erik M. Conway. The politics of blind landing. Technology and Culture, 42(1):81–106, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33663.

Conway:2002:AAA

[Con02] Erik M. Conway. America’s Airports: Airfield Development, 1918–1947 (review). Technology and Culture, 43(4):819–820, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34010.

Conway:2003:FFN

[Con03] Erik M. Conway. First to Fly: North Carolina and the Beginnings of Aviation (review). Technology and Culture, 44(1): 180–181, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40148.

Constant:2006:CVL

[Con06a] Edward W. Constant. Certainties of very low probability. Technology and Culture, 47(1):249–252, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196036.

Constant:2006:TTB

[Con06b] Edward W. Constant. A tale of two bonanzas: How knowledgeable communities think about technology. Technology and Culture, 47(2):253–285, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/200839.

Conway:2007:RDF

[Con07] Erik M. Conway. Realizing the Dream of Flight: Biographical Essays in Honor of the Centennial of Flight, 1903–2003 (review). Technology and Culture, 48(1):232–234, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209828.

Conlon:2008:ECR

[Con08] Frank F. Conlon. Engines of Change: The Railroads That Made India (review). Technology and Culture, 49(2):471–472, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236977.

Cooper:2000:AMA

[Coo00] Carolyn C. Cooper. The American Manufactory: Art, Labor, and the World of Things in the Early Republic (review). Technology and Culture, 41(1):140–141, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33471.

Cooper:2001:RNC

[Coo01a] Carolyn C. Cooper. Rooms Near Chancery Lane: The Patent Office under the Commissioners, 1852–1883 (review). Technology and Culture, 42(3):578–580, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33755.

Coopersmith:2001:PGT

[Coo01b] Jonathan Coopersmith. Pretty good technologies and visible disasters. Technology and Culture, 42(1):204–207, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33664.

Cooper:2002:STS

[Coo02a] Gail Cooper. Science, Technology, and Society in Contemporary Japan (review). Technology and Culture, 43(1):217–218, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33860.

Cooper:2002:RLG

[Coo02b] Patricia Cooper. Race on the Line: Gender, Labor, and Technology in the Bell System, 1880–1980 (review). Technology and Culture, 43(3):632–634, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33957.

Coopersmith:2002:LRA

[Coo02c] Jonathan Coopersmith. Listening In: Radio and the American Imagination, from Amos ’n’ Andy and Edward R. Murrow to Wolfman Jack and Howard Stern (review). Technology and Culture, 43(2):436–438, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33906.

Cooper:2003:MRH

[Coo03a] Carolyn C. Cooper. Myth, rumor, and history: The Yankee whittling boy as hero and villain. Technology and Culture, 44 (1):82–96, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40149.

Coopersmith:2003:WNF

[Coo03b] Jonathan Coopersmith. Wireless Nation: The Frenzied Launch of the Cellular Revolution in America (review). Technology and Culture, 44(3):631–632, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46088.

Cookson:2004:BLT

[Coo04a] Gillian Cookson. Brunel: The Life and Times of Isambard Kingdom Brunel (review). Technology and Culture, 45(1): 183–184, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51742.

Cooper:2004:MAH

[Coo04b] Carolyn C. Cooper. Measuring America: How an Untamed Wilderness Shaped the United States and Fulfilled the Promise of Democracy (review). Technology and Culture, 45(2):436– 438, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55658.

Coopersmith:2004:MNG

[Coo04c] Jonathan Coopersmith. A machine named ginger: Steve kemper, code name ginger. Technology and Culture, 45(3):603– 607, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172970.

Cookson:2005:WPE

[Coo05a] Gillian Cookson. Watt’s Perfect Engine: Steam and the Age of Invention (review). Technology and Culture, 46(3):651– 652, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186610.

Coopersmith:2005:BFS

[Coo05b] Jonathan Coopersmith. Better Off: Flipping the Switch on Technology (review). Technology and Culture, 46(3):623–624, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186611.

Coopey:2005:GMR

[Coo05c] R. Coopey. The Government Machine: A Revolutionary History of the Computer (review). Technology and Culture, 46(4): 857–859, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192286.

Coopersmith:2006:DDB

[Coo06] Jonathan Coopersmith. The dog that did not bark during the night: The “Normalcy” of Russian, Soviet, and post-Soviet science and technology studies. Technology and Culture, 47 (3):623–637, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201444.

Cooke:2007:WCM

[Coo07a] Edward S. Cooke. Windsor-Chair Making in America: From Craft Shop to Consumer (review). Technology and Culture, 48 (3):653–655, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218698.

Cooper:2007:PSR

[Coo07b] Carolyn C. Cooper. Patents since the Renaissance (review). Technology and Culture, 48(3):677–679, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218699.

Coopersmith:2009:HIS

[Coo09] Jonathan C. Coopersmith. The History of Information Security: A Comprehensive Handbook (review). Technology and Culture, 50(1):262–263, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258545.

Corn:2002:WMR

[Cor02a] Wanda M. Corn. Women and the Machine: Representations from the Spinning Wheel to the Electronic Age (review). Technology and Culture, 43(4):832–833, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34011.

Corwin:2002:EUI

[Cor02b] Sharon Corwin. Eyewitnessing: The Uses of Images as Historical Evidence (review). Technology and Culture, 43(4):830–831, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34012.

Cornillie:2006:IAH

[Cor06] Thomas C. Cornillie. The International Association for the History of Transport, Traffic, and Mobility: Third Annual Conference, York, England, 6-9 October 2005. Technology and Culture, 47(2):369–372, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/200840.

Corn:2008:AMT

[Cor08] Joseph J. Corn. Auto Mechanics: Technology and Expertise in Twentieth-Century America (review). Technology and Culture, 49(2):464–465, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236978.

Corn:2009:HTA

[Cor09] Joseph Corn. Horse Trading in the Age of Cars: Men in the Marketplace (review). Technology and Culture, 50(2):493– 495, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263342.

Cowan:2003:BFP

[Cow03] Ruth Schwartz Cowan. Baby’s First Picture: Ultrasound and the Politics of Fetal Subjects (review). Technology and Culture, 44(4):850–851, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49689.

Cowan:2004:ELE

[Cow04a] Neil M. Cowan. Empires of Light: Edison, Tesla, Westinghouse, and the Race to Electrify the World (review). Technology and Culture, 45(3):652–653, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172971.

Cowan:2004:WDM

[Cow04b] Neil M. Cowan. Watson and DNA: Making a Scientific Revolution (review). Technology and Culture, 45(2):457–458, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55659.

Cowan:2006:WCW

[Cow06] Neil M. Cowan. The War Complex: World War II in Our Time (review). Technology and Culture, 47(4):834–835, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204972.

Cox:2004:CLD

[Cox04] Dwayne Cox. The Conquest of Labor: Daniel Pratt and Southern Industrialization (review). Technology and Culture, 45(1): 190–191, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51743.

Craig:2007:HRT

[Cra07] Douglas B. Craig. Ham Radio’s Technical Culture (review). Technology and Culture, 48(4):901–902, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223352.

Crosby:2002:NPW

[Cro02a] Alfred W. Crosby. Notes of a Potato Watcher (review). Technology and Culture, 43(3):620–622, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33958.

Crouch:2002:SHD

[Cro02b] Tom D. Crouch. Spacefaring: The Human Dimension (review). Technology and Culture, 43(4):828–830, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/34013.

Croarken:2005:DRH

[Cro05] Mary Croarken. Douglas Rayner Hartree: His Life in Science and Computing by Charlotte Froese Fischer (review). Technology and Culture, 46(4):859–860, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192288.

Crosby:2006:DCN

[Cro06a] Alfred W. Crosby. At Day’s Close: Night in Times Past (review). Technology and Culture, 47(3):640–641, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201446.

Crosby:2006:MGA

[Cro06b] Alfred W. Crosby. Maize and Grace: Africa’s Encounter with a New World Crop, 1500–2000 (review). Technology and Culture, 47(1):190–191, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196037.

Crouch:2006:DSL

[Cro06c] Tom D. Crouch. Dr. Space: The Life of Wernher von Braun (review). Technology and Culture, 47(2):455–456, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200841.

Cronin:2007:NVA

[Cro07] Marionne Cronin. Northern visions: Aerial surveying and the Canadian mining industry, 1919–1928. Technology and Culture, 48(2):303–330, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215387.

Crompton:2008:GRM

[Cro08a] Gerald Crompton. Government, the Railways, and the Modernization of Britain: Beeching’s Last Trains (review). Technology and Culture, 49(2):469–470, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/236979.

Cross:2008:CAS

[Cro08b] Gary Cross. The Challenge of Affluence: Self-Control and Well-Being in the United States and Britain since 1950 (review). Technology and Culture, 49(3):815–817, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243092.

Crumpton:2003:SCC

[Cru03] Amy C. Crumpton. Scientists in the Classroom: The Cold War Reconstruction of American Science Education (review). Technology and Culture, 44(1):202–203, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40150.

Crumpton:2007:PPS

[Cru07] Amy C. Crumpton. Pedagogy and the Practice of Science: Historical and Contemporary Perspectives (review). Technology and Culture, 48(1):190–191, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209829.

Culver:2009:TIA

[Cul09] Melissa M. Culver. The Technical Imagination: Argentine Culture’s Modern Dreams (review). Technology and Culture, 50(1): 208–209, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258556.

Cushman:2005:DGW

[Cus05] Gregory T. Cushman. The Discovery of Global Warming (review). Technology and Culture, 46(2):456–458, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184312.

Cutcliffe:2002:TPE

[Cut02] Stephen H. Cutcliffe. Technologies of Power: Essays in Honor of Thomas Parke Hughes and Agatha Chipley Hughes (review). Technology and Culture, 43(3):649–651, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33959.

Cutcliffe:2006:MSS

[Cut06] Stephen H. Cutcliffe. Making Steel: Sparrows Point and the Rise and Ruin of American Industrial Might (review). Technology and Culture, 47(1):215–217, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196038.

Cutcliffe:2007:GTL

[Cut07] Stephen H. Cutcliffe. The Greenwood technographies: Life stories of technologies. Technology and Culture, 48(1):165–168, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209830.

Cutcliffe:2009:HST

[Cut09a] Stephen H. Cutcliffe. The Handbook of Science and Technology Studies (review). Technology and Culture, 50(2):453–454, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263350.

Cutcliffe:2009:LUF

[Cut09b] Stephen H. Cutcliffe. Lehigh University’s Fritz Laboratory and the five-million-pound universal testing machine. Technology and Culture, 50(2):391–398, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263330.

Daar:2007:DPC

[Daa07] Judith Daar. The Dream of the Perfect Child (review). Technology and Culture, 48(1):193–195, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209831.

Daley:2002:GLR

[Dal02] Matthew L. Daley. Graveyard of the Lakes (review). Technology and Culture, 43(1):172–173, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33861.

Dalrymple:2004:WDR

[Dal04] Dana G. Dalrymple. Wayne D. Rasmussen, 1915–2004. Technology and Culture, 45(3):683–686, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/172972.

Dalley:2005:ICW

[Dal05] Stephanie Dalley. The Invention of Cuneiform: Writing in Sumer (review). Technology and Culture, 46(2):408–409, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184313.

Dalley:2006:HWD

[Dal06a] Stephanie Dalley. The Heavenly Writing: Divination, Horoscopy, and Astronomy in Mesopotamian Culture (review).

Technology and Culture, 47(2):409–410, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/200842.

Dalrymple:2006:WJS

[Dal06b] Dana G. Dalrymple. William J. Spillman and the Birth of Agricultural Economics (review). Technology and Culture, 47 (1):201–203, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196039.

Daniels:2001:SSC

[Dan01] Christian Daniels. Sugar and Society in China: Peasants, Technology, and the World Market (review). Technology and Culture, 42(2):354–355, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33710.

Danbom:2003:EFF

[Dan03] David B. Danbom. Every Farm a Factory: The Industrial Ideal in American Agriculture (review). Technology and Culture, 44 (4):836–838, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49690.

Daso:2008:WCD

[Das08] Dik A. Daso. Weapons of Choice: The Development of Precision Guided Munitions (review). Technology and Culture, 49 (1):252–254, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/230900.

David:2000:PMI

[Dav00] Nicholas David. Pride of Men: Ironworking in Nineteenth-Century West Central Africa (review). Technology and Culture, 41(4):797–798, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33607.

Davidson:2001:MTM

[Dav01] Janet F. Davidson. More than Munitions: Women, Work and the Engineering Industries, 1900–1950 (review). Technology and Culture, 42(3):580–581, July 2001. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33756.

Davidson:2006:HRW

[Dav06] Janet F. Davidson. Home on the Rails: Women, the Railroad, and the Rise of Public Domesticity (review). Technology and Culture, 47(1):218–219, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196040.

Davis:2009:OTH

[Dav09] Jennifer J. Davis. Observer le travail: Histoire, ethnographie, approches combinées (review). Technology and Culture, 50(4): 980–981, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363458.

Dawson:2006:HSD

[Daw06] Virginia P. Dawson. High-Speed Dreams: NASA and the Technopolitics of Supersonic Transportation, 1945–1999 (review). Technology and Culture, 47(4):866–867, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204973.

Dawson:2007:BMC

[Daw07a] Max Dawson. Beyond the Multiplex: Cinema, New Technologies, and the Home (review). Technology and Culture, 48(2): 436–438, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215388.

Dawson:2007:HVT

[Daw07b] Max Dawson. Home video and the “TV problem”: Cultural critics and technological change. Technology and Culture, 48 (3):524–549, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218700.

Day:2001:TVI

[Day01] Dwayne A. Day. Trust but Verify: Imagery Analysis in the Cold War (review). Technology and Culture, 42(4):822–823, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33809.

Day:2004:SEE

[Day04a] Dwayne A. Day. Secret Empire: Eisenhower, the CIA, and the Hidden Story of America’s Space Espionage (review). Technology and Culture, 45(1):204–206, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51744.

Day:2004:SMS

[Day04b] Dwayne A. Day. Spying with Maps: Surveillance Technologies and the Future of Privacy (review). Technology and Culture, 45(3):675–676, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172974.

Day:2006:SSC

[Day06] Dwayne A. Day. Spying from Space: Constructing America’s Satellite Command and Control Systems (review). Technology and Culture, 47(3):645–646, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201447.

Day:2008:HKS

[Day08] Dwayne A. Day. A History of the Kennedy Space Center (review). Technology and Culture, 49(3):813–815, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243119.

DeSyon:2000:STD

[De 00] Guillaume De Syon. Des sciences et des techniques: Un debat (review). Technology and Culture, 41(3):573–575, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33593.

DeSyon:2003:MMA

[De 03a] Guillaume De Syon. Marketing Michelin: Advertising and Cultural Identity in Twentieth-Century France (review). Technology and Culture, 44(2):416–417, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42011.

DeSyon:2003:CC

[De 03b] Guillaume De Syon. Consuming Concorde. Technology and Culture, 44(3):650–654, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/46120.

DeSyon:2004:WWW

[De 04] Guillaume De Syon. What the Wrights wrought: The centennial of flight in recent literature. Technology and Culture, 45 (2):350–357, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55661.

DeSyon:2006:SFA

[De 06a] Guillaume De Syon. Sky as Frontier: Adventure, Aviation, and Empire (review). Technology and Culture, 47(4):864–865, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/205009.

DeWeerdt:2006:PCP

[De 06b] Hilde Godelieve Dominique De Weerdt. Publishing, Culture, and Power in Early Modern China (review). Technology and Culture, 47(1):192–193, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196078.

Dear:2007:GIC

[Dea07] Peter Robert Dear. Galileo’s Instruments of Credit: Telescopes, Images, Secrecy (review). Technology and Culture, 48(3):620– 622, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218701.

Dear:2008:JHE

[Dea08] Peter R. Dear. The Jewel House: Elizabethan London and the Scientific Revolution (review). Technology and Culture, 49 (3):793–794, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243127.

Dempsey:2003:PPC

[Dem03] Claire W. Dempsey. The Patina of Place: The Cultural Weathering of a New England Industrial Landscape (review). Technology and Culture, 44(2):404–405, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42012.

Dermody:2004:ALS

[Der04a] Robert J. Dermody. Artificial Love: A Story of Machines and Architecture (review). Technology and Culture, 45(2):470– 471, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55660.

Dermody:2004:TOM

[Der04b] Robert J. Dermody. Tower and Office: From Modernist Theory to Contemporary Practice (review). Technology and Culture, 45(4):852–854, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175744.

Deverell:2000:AWR

[Dev00a] William Francis Deverell. The Atomic West (review). Technology and Culture, 41(3):611–612, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33553.

DeVries:2000:LR

[DeV00b] Kelly DeVries. L’Outil (review). Technology and Culture, 41(1): 132–134, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33472.

DeVries:2003:DEA

[DeV03] Kelly DeVries. Dangerous Energy: The Archaeology of Gunpowder and Military Explosives Manufacture (review). Technology and Culture, 44(2):381–382, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42013.

Dicke:2004:SRA

[Dic04] Thomas S. Dicke. Sloan Rules: Alfred P. Sloan and the Triumph of General Motors (review). Technology and Culture, 45(1):

192–194, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51745.

Dickinson:2005:CEA

[Dic05] John Alexander Dickinson. Chasing Empire across the Sea: Communications and the State in the French Atlantic, 1713– 1763 (review). Technology and Culture, 46(3):642–644, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186613.

Dierikx:2001:ESE

[Die01] M. L. J. Dierikx. Energie und Stadt in Europa: Von der vorindustriellen “Holznot” bis zur Olkrise der 1970er Jahre (review). Technology and Culture, 42(1):146–147, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33665.

Dinckal:2008:RMC

[Din08] Noyan Dinçkal. Reluctant modernization: The cultural dynamics of water supply in Istanbul, 1885–1950. Technology and Culture, 49(3):675–700, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243085.

Diogo:2008:MST

[Dio08] Maria Paula Diogo. Museums of science and technology in Lisbon. Technology and Culture, 49(3):764–772, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243115.

Dirks:2005:CCC

[Dir05] Jacqueline K. Dirks. The Commodification of Childhood: The Children’s Clothing Industry and the Rise of the Child Consumer (review). Technology and Culture, 46(4):822–823, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192289.

Dirks:2007:RCC

[Dir07] Jacqueline K. Dirks. Raising Consumers: Children and the American Mass Market in the Early Twentieth Century (review). Technology and Culture, 48(1):195–197, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209832.

Disco:2005:REB

[Dis05] Cornelis Disco. The Rhine: An Eco-Biography, 1815–2000 (review). Technology and Culture, 46(2):464–466, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184314.

Disco:2006:DB

[Dis06] Cornelis Disco. Delta blues. Technology and Culture, 47(2):341– 348, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200843.

Divall:2000:ITC

[Div00] Colin Divall. Institutions in the Transport and Communications Industries: State and Private Actors in the Making of Institutional Patterns, 1850–1990 (review). Technology and Culture, 41(3):635–636, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33554.

Divall:2003:RBU

[Div03a] Colin Divall. Railways in Britain and the United States, 1830– 1940 (review). Technology and Culture, 44(2):387–389, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
42014.

Divall:2003:RNB

[Div03b] Colin Divall. Railways in the Netherlands: A Brief History, 1834–1994 (review). Technology and Culture, 44(1):178–180, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40152.

Divall:2008:LTM

[Div08] Colin Divall. The London Transport Museum. Technology and Culture, 49(4):1010–1017, October 2008. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255447.

Divall:2009:TDT

[Div09] Colin Divall. Transport Design: A Travel History (review). Technology and Culture, 50(2):495–497, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263345.

Dorn:2000:C

[DK00] Harold Dorn and Daniel Kuehl. Communication. Technology and Culture, 41(1):187–189, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33473.

delaPena:2006:BED

[dlP06] Carolyn Thomas de la Pena. “Bleaching the ethiopians”: Desegregating race and technology through early X-ray experiments. Technology and Culture, 47(1):27–55, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
196064.

delaPena:2007:CPH

[dlP07] Carolyn Thomas de la Peña. The Cult of Pharmacology: How America Became the World’s Most Troubled Drug Culture (review). Technology and Culture, 48(3):668–670, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218722.

denOtter:2001:CNE

[dO01a] A. A. den Otter. Citizens and Nation: An Essay on History, Communications, and Canada (review). Technology and Culture, 42(3):587–588, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33784.

denOtter:2001:DWM

[dO01b] A. A. den Otter. The Diary of William Mackenzie: The First International Railway Contractor (review). Technology and Culture, 42(1):166–168, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33682.

Dalley:2003:SAW

[DO03] Stephanie Dalley and John Peter Oleson. Sennacherib, Archimedes, and the water screw: The context of invention in the ancient world. Technology and Culture, 44(1):1–26, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40151.

denOtter:2005:HIR

[dO05] A. A. den Otter. Harold Innis (review). Technology and Culture, 46(2):415–416, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/184335.

Dobres:2004:AFE

[Dob04a] Marcia-Anne Dobres. African Foragers: Environment, Technology, Interactions (review). Technology and Culture, 45(3): 641–643, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172975.

Dobres:2004:TBH

[Dob04b] Marcia-Anne Dobres. Tool-Being: Heidegger and the Metaphysics of Objects (review). Technology and Culture, 45(1): 232–234, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51746.

Dobres:2006:MR

[Dob06] Marcia-Anne Dobres. Materiality (review). Technology and Culture, 47(4):823–824, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/204974.

Dobres:2007:IPT

[Dob07] Marcia-Anne Dobres. Inside the Politics of Technology: Agency and Normativity in the Co-Production of Technology and Society (review). Technology and Culture, 48(1):179–181, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209833.

Dobraszczyk:2008:CCS

[Dob08a] Paul Dobraszczyk. Cleansing the City: Sanitary Geographies in Victorian London (review). Technology and Culture, 49(4): 1055–1056, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255456.

Dobraszczyk:2008:CFS

[Dob08b] Paul Dobraszczyk. The Culture of Flushing: A Social and Legal History of Sewage (review). Technology and Culture, 49(1): 284–285, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230901.

Dobraszczyk:2008:IAC

[Dob08c] Paul Dobraszczyk. Image and audience: Contractual representation and London’s main drainage system. Technology and Culture, 49(3):568–598, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243117.

Dobres:2008:PIM

[Dob08d] Marcia-Anne Dobres. Philosophical Instruments: Minds and Tools at Work (review). Technology and Culture, 49(3):791– 792, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243104.

Doerksen:2008:TWR

[Doe08a] Cliff Doerksen. 9XM Talking: WHA Radio and the Wisconsin Idea (review). Technology and Culture, 49(2):503–504, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236980.

Doerksen:2008:CEB

[Doe08b] Cliff Doerksen. Crossing the Ether: British Public Service Radio and Commercial Competition, 1922–1945 (review). Technology and Culture, 49(4):1081–1082, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255503.

Doern:2008:GEW

[Doe08c] Bruce Doern. The Global Environment and World Politics (review). Technology and Culture, 49(4):1069–1070, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
255504.

Doig:2006:BBC

[Doi06] Jameson W. Doig. The Brooklyn Bridge: A Cultural History (review). Technology and Culture, 47(4):847–848, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
204975.

Dolza:2009:EMC

[Dol09] Luisa Dolza. Europa mineraria: Circolazione delle élites e trasferimento tecnologico (secoli XVIII–XIX) (review). Technology and Culture, 50(4):937–939, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363433.

Donovan:2006:MNG

[Don06] Arthur Donovan. The Machine in Neptune’s Garden: Historical Perspectives on Technology and the Marine Environment (review). Technology and Culture, 47(3):692–694, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201448.

Donovan:2007:IAS

[Don07] Arthur Donovan. Industrializing American Shipbuilding: The Transformation of Ship Design and Construction (review). Technology and Culture, 48(3):646–647, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218702.

Dorn:2000:ASR

[Dor00] Harold Dorn. The Architecture of Science (review). Technology and Culture, 41(4):854–856, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33608.

Dorn:2001:MAT

[Dor01] Harold Dorn. Masters of All They Surveyed: Exploration, Geography, and a British El Dorado (review). Technology and Culture, 42(3):589–590, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33757.

Dorn:2007:TD

[Dor07] Harold Dorn. Technology as destiny. Technology and Culture, 48(3):598–602, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218703.

Dorn:2008:FDA

[Dor08] Harold Dorn. Five Days in August: How World War II Became a Nuclear War (review). Technology and Culture, 49(3):812– 813, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243120.

Douglas:2001:MMS

[Dou01] Deborah G. Douglas. Museums of Modern Science (review). Technology and Culture, 42(3):624–625, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33758.

Douglas:2004:WBI

[Dou04] Deborah G. Douglas. The Wright Brothers and the Invention of the Aerial Age: At the National Air and Space Museum, Washington, D.C. Technology and Culture, 45(2):363– 367, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55662.

Douglas:2009:WBI

[Dou09] Deborah G. Douglas. Women in British Imperial Airspace, 1922–1937 (review). Technology and Culture, 50(3):717–719, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270561.

Downey:2001:VWP

[Dow01] Greg Downey. Virtual webs, physical technologies, and hidden workers: The spaces of labor in information internetworks. Technology and Culture, 42(2):209–235, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33711.

Downey:2003:DWR

[Dow03a] Gregory John Downey. Distributed Work (review). Technology and Culture, 44(3):629–630, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46090.

Downey:2003:MCE

[Dow03b] Gregory John Downey. The Moment of Complexity: Emerging Network Culture (review). Technology and Culture, 44(1): 223–225, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40153.

Downey:2004:NLP

[Dow04] Gregory John Downey. Nodes, links, and phase transitions: Popularizing “Network Science”. Technology and Culture, 45 (1):162–167, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51747.

Downey:2005:TPC

[Dow05] Gregory John Downey. Territories of Profit: Communications, Capitalist Development, and the Innovative Enterprises of G. F. Swift and Dell Computer (review). Technology and Culture, 46(4):855–857, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192290.

Downey:2006:CCC

[Dow06] Gregory John Downey. Constructing “Computer-Compatible” stenographers: The transition to real-time transcription in courtroom reporting. Technology and Culture, 47(1):1–26, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196041.

Downey:2007:TPS

[Dow07] Gregory John Downey. Taking Place: The Spatial Contexts of Science, Technology, and Business (review). Technology and Culture, 48(2):416–418, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/215389.

Downey:2008:MCS

[Dow08a] Greg Downey. Mobile Communication and Society: A Global Perspective (review). Technology and Culture, 49(2):512–514, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236981.

Downey:2008:TRV

[Dow08b] Greg Downey. Telegraphic Realism: Victorian Fiction and Other Information Systems (review). Technology and Culture, 49(4):1054–1055, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255457.

Downey:2009:FIH

[Dow09] Greg Downey. The Future of the Internet — And How to Stop It (review). Technology and Culture, 50(4):962–964, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363448.

deSyon:2009:ANL

[dS09] Guillaume de Syon. Aufstieg und Niedergang der Luftschifffahrt: Eine wirtschaftshistorische Analyse (review). Technology and Culture, 50(2):497–498, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263363.

Leitao:2004:FBF

[dSL04] Henrique de Sousa Leitao. Francis Bacon e a fundamentaçao da ciencia como tecnologia (review). Technology and Culture, 45(1):170–171, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51758.

Dugdale:2008:ASR

[Dug08] Timothy Dugdale. Avatars of Story (review). Technology and Culture, 49(2):510–512, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/236982.

Dugdale:2009:DEA

[Dug09] Timothy Dugdale. Distracted: The Erosion of Attention and the Coming Dark Age (review). Technology and Culture, 50 (2):510–511, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/263346.

Durant:2006:DWO

[Dur06] Darrin Durant. Deep Waters: The Ottawa River and Canada’s Nuclear Adventure (review). Technology and Culture, 47(3): 673–674, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201449.

deVries:2002:LMV

[dV02] Monique de Vries. A letter from Monique de Vries, Vice Minister of Transport, Public Works, and Water Management. Technology and Culture, 43(3):473–474, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33960.

Disco:2003:SIA

[DvdE03] Cornelis Disco and Johannes Cornelis Maria van den Ende. “Strong, invincible arguments?” tidal models as management instruments in Twentieth-Century Dutch coastal engineering. Technology and Culture, 44(3):502–535, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46089.

deWit:2002:IJO

[dWvdESvO02] Onno de Wit, Johannes Cornelis Maria van den Ende, Johan Schot, and Ellen van Oost. Innovative junctions: Office technologies in the Netherlands, 1880–1980. Technology and Culture, 43(1):50–72, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33862.

Dyble:2009:RTL

[Dyb09] Louise Nelson Dyble. Reconstructing transportation: Linking tolls and transit for place-based mobility. Technology and Culture, 50(3):631–648, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/270540.

Dym:2008:SMD

[Dym08] Warren Dym. Scholars and miners: Dowsing and the Freiberg Mining Academy. Technology and Culture, 49(4):833–859, October 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/255438.

Eckert:2005:SIT

[Eck05] Michael Eckert. Strategic internationalism and the transfer of technical knowledge: The United States, Germany, and aerodynamics after World War I. Technology and Culture, 46(1): 104–131, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179756.

Edensor:2009:HRR

[Ede09] Tim Edensor. Hope and Rust: Reinterpreting the Industrial Place in the Late 20th Century (review). Technology and Culture, 50(1):230–232, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/258578.

Edgerton:2008:TPA

[Edg08] David Edgerton. Technology in Postwar America: A History (review). Technology and Culture, 49(2):491–493, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236983.

Edwards:2003:BGP

[Edw03a] James Frederick Edwards. Building the Great Pyramid: Probable construction methods employed at Giza. Technology and Culture, 44(2):340–354, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/42015.

Edwards:2003:BRI

[Edw03b] Roy Edwards. British Rail 1974–97: From Integration to Privatisation (review). Technology and Culture, 44(3):612– 613, July 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/46091.

Eelman:2009:TIS

[Eel09] Bruce Eelman. Technology, Innovation, and Southern Industrialization: From the Antebellum Era to the Computer Age (review). Technology and Culture, 50(4):930–932, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363429.

Eglash:2007:BMM

[Egl07] Ron Eglash. Broken metaphor: The master-slave analogy in technical literature. Technology and Culture, 48(2):360–369, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215390.

Eisler:2009:MLP

[Eis09] Matthew N. Eisler. “A Modern ‘Philosopher’s Stone”’: Technoanalogy and the Bacon Cell. Technology and Culture, 50(2): 345–365, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/263329.

Elkind:2002:TBE

[Elk02] Sarah S. Elkind. Ties that Bind: Economic and Political Dilemmas of Urban Utility Networks, 1800–1990 (review). Technology and Culture, 43(3):612–614, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33961.

Elkind:2006:PWL

[Elk06] Sarah S. Elkind. Political Waters: The Long, Dirty, Contentious, Incredibly Expensive but Eventually Triumphant History of Boston Harbor — a Unique Environmental Success Story (review). Technology and Culture, 47(1):223–224, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196042.

Ellison:2001:HDP

[Ell01] Karin Ellison. Hoover Dam: The Photographs of Ben Glaha, and: Glen Canyon Dammed: Inventing Lake Powell and the Canyon Country (review). Technology and Culture, 42(1): 155–158, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33666.

Ellison:2005:BOW

[Ell05] Karin Ellison. Border Oasis: Water and the Political Ecology of the Colorado River Delta, 1940–1975, and: Fuel for Growth: Water and Arizona’s Urban Environment (review). Technology and Culture, 46(2):462–464, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184315.

Elsasser:2009:MIC

[Els09] Kilian T. Elsasser. Materialising Identity: The Co-construction of the Gotthard Railway and Swiss National Identity (review). Technology and Culture, 50(4):939–940, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363434.

Endersby:2000:ETT

[End00] Linda Eikmeier Endersby. ENIAC: The Triumphs and Tragedies of the World’s First Computer (review). Technology and Culture, 41(4):825–826, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33610.

Endelman:2001:HRM

[End01] Judith E. Endelman. Home on the Road: The Motor Home in America (review). Technology and Culture, 42(3):612–613, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33759.

Ensmenger:2005:SOS

[Ens05] Nathan Ensmenger. The Success of Open Source (review). Technology and Culture, 46(4):860–862, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192291.

Ensmenger:2006:IEC

[Ens06] Nathan Ensmenger. Inventing the Electronic Century: The Epic Story of the Consumer Electronics and Computer Industries (review). Technology and Culture, 47(3):680–681, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201445.

Ensmenger:2007:BGR

[Ens07] Nathan Ensmenger. Broken Genius: The Rise and Fall of William Shockley, Creator of the Electronic Age (review). Technology and Culture, 48(4):887–889, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223353.

Epperson:2000:FCS

[Epp00] Bruce Epperson. Failed Colossus: Strategic error at the Pope Manufacturing Company, 1878–1900. Technology and Culture, 41(2):300–320, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33519.

Epperson:2002:RMB

[Epp02a] Bruce Epperson. Ride to Modernity: The Bicycle in Canada, 1869–1900 (review). Technology and Culture, 43(2):425–426, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33908.

Epperson:2002:DSA

[Epp02b] Bruce Epperson. Does SCOT answer? A comment. Technology and Culture, 43(2):371–373, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33907. See [Cla02c].

Epperson:2003:FPT

[Epp03] Bruce Epperson. Framing Production: Technology, Culture, and Change in the British Bicycle Industry (review). Technology and Culture, 44(2):389–390, April 2003. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42016.

Epperson:2004:FDT

[Epp04] Bruce Epperson. Is form destiny? Technologies of personal security. Technology and Culture, 45(3):613–617, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172973.

Epperson:2006:DPW

[Epp06] Bruce Epperson. Digital Phoenix: Why the Information Economy Collapsed and How It Will Rise Again (review). Technology and Culture, 47(3):683–685, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201450.

Epperson:2007:EDR

[Epp07] Bruce Epperson. Eminence domain: Reassessing the life and public works of Robert Moses. Technology and Culture, 48(4): 816–823, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223354.

Epperson:2009:FTD

[Epp09] Bruce Epperson. Fighting Traffic: The Dawn of the Motor Age in the American City (review). Technology and Culture, 50(1): 235–237, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258547.

Erickson:2007:MLM

[Eri07] Joe Erickson. Mr. Lincoln’s T-Mails: The Untold Story of How Abraham Lincoln Used the Telegraph to Win the Civil War (review). Technology and Culture, 48(4):861–862, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
223355.

Erickson:2008:CMB

[Eri08a] Joe Erickson. Color Monitors: The Black Face of Technology in America (review). Technology and Culture, 49(1):295–296, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230902.

Erickson:2008:IPC

[Eri08b] Joe Erickson. Information Please: Culture and Politics in the Age of Digital Machines (review). Technology and Culture, 49 (2):509–510, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236984.

Esbester:2005:IAH

[Esb05] Mike Esbester. The International Association for the History of Transport, Traffic and Mobility: Second Annual Conference, Detroit, 4–7 November 2004. Technology and Culture, 46(1): 164–167, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179757.

Ezell:2002:TSM

[Eze02] Linda Neuman Ezell. Taking Science to the Moon: Lunar Experiments and the Apollo Program (review). Technology and Culture, 43(4):827–828, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34014.

Fabel:2006:TSO

[Fab06] John Fabel. The Tale of the Scale: An Odyssey of Invention (review). Technology and Culture, 47(2):470–471, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200844.

Falnes:2006:SWS

[Fal06a] Jørund Falnes. The Shore Whaling Stations at South Georgia: A Study in Antarctic Industrial Archaeology (review). Technology and Culture, 47(3):694–695, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201451.

Falnes:2006:TCS

[Fal06b] Jørund Falnes. Twentieth-Century Shore-Station Whaling in Newfoundland and Labrador (review). Technology and Culture, 47(4):869–870, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/204976.

Fano:2001:IOJ

[Fan01] Ester Fano. Inventing Ourselves Out of Jobs?: America’s Debate over Technological Unemployment, 1929–1981 (review). Technology and Culture, 42(3):583–585, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33760.

Farnie:2000:IET

[Far00] D. A. Farnie. Industrial Espionage and Technology Transfer: Britain and France in the Eighteenth Century (review). Technology and Culture, 41(2):355–357, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33520.

Fehr:2004:OTG

[Feh04] Kregg M. Fehr. Oil in Texas: The Gusher Age, 1895– 1945 (review). Technology and Culture, 45(3):656–657, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172976.

Fein:2009:HRH

[Fei09] Michael R. Fein. The Highway Revolution, 1895–1925: How the United States Got Out of the Mud (review). Technology and Culture, 50(2):488–489, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263348.

Felker:2007:CSN

[Fel07] Craig C. Felker. Command at Sea: Naval Command and Control since the Sixteenth Century (review). Technology and Culture, 48(3):648–649, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/218704.

Ferguson:2005:AAH

[Fer05] Robert G. Ferguson. The American Axis: Henry Ford, Charles Lindbergh, and the Rise of the Third Reich (review). Technology and Culture, 46(2):451–453, April 2005. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184317.

Fehlhammer:2000:DMI

[FF00] Wolf Peter Fehlhammer and Wilhelm Fuessl. The Deutsches Museum: Idea, realization, and objectives. Technology and Culture, 41(3):517–520, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33555.

Field:2001:SDT

[Fie01] Gregory Field. The Social Dynamics of Technology: Practice, Politics, and World Views (review). Technology and Culture, 42 (2):336–338, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33712.

Fine:2000:LHI

[Fin00] Ben Fine. London, Hub of the Industrial Revolution: A Revisionary History, 1775–1825 (review). Technology and Culture, 41(2):357–358, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33521.

Finn:2002:CLS

[Fin02] Bernard S. Finn. City of Light: The Story of Fiber Optics (review). Technology and Culture, 43(1):194–196, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33863.

Finn:2003:SCW

[Fin03] Bernard S. Finn. Sir Charles Wheatstone, FRS, 1802–1875 (review). Technology and Culture, 44(4):820–822, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49691.

Finn:2005:RPM

[Fin05] Bernard S. Finn. Robert P. Multhauf, 1919–2004. Technology and Culture, 46(1):265–273, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179758.

Finn:2009:BGJ

[Fin09] Bernard S. Finn. Bell and Gray: Just a coincidence? Technology and Culture, 50(1):193–201, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258554.

Fisher:2003:AAS

[Fis03] Genevieve Fisher. The Art of the Anglo-Saxon Goldsmith (review). Technology and Culture, 44(3):600–601, July 2003. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46092.

Fishman:2007:HSP

[Fis07] Robert Fishman. Half the story: Paul Clemens, made in Detroit: a south of 8 mile memoir. Technology and Culture, 48 (2):404–406, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215391.

Fischer:2009:RST

[Fis09] Claude S. Fischer. Rethinking Science, Technology, and Social Change (review). Technology and Culture, 50(3):679–680, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270545.

Fitzgerald:2000:DMU

[Fit00] Gerard J. Fitzgerald. Differences in Medicine: Unraveling Practices, Techniques, and Bodies (review). Technology and Culture, 41(1):178–179, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33474.

Fitzpatrick:2001:HMC

[Fit01] Anne Claire Fitzpatrick. A History of Modern Computing (review). Technology and Culture, 42(3):617–618, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33761.

Fitzgerald:2002:CCT

[Fit02] Deborah Kay Fitzgerald. Consumers in the Country: Technology and Social Change in Rural America (review). Technology and Culture, 43(1):163–165, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33864.

Florida:2000:WWB

[Flo00] Richard L. Florida. We Were Burning: Japanese Entrepreneurs and the Forging of the Electronic Age (review). Technology and Culture, 41(3):620–621, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33556.

Florman:2001:TTL

[Flo01] Samuel C. Florman. Twin Towers: The Life of New York City’s World Trade Center (review). Technology and Culture, 42(1): 159–160, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33667.

Florman:2003:EHW

[Flo03] Samuel C. Florman. The embedded historian: William langewiesche, American ground. Technology and Culture, 44(4):794– 798, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/49692.

Florman:2005:EEF

[Flo05] Samuel C. Florman. Engineering: An Endless Frontier (review). Technology and Culture, 46(3):628–630, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186614.

Florman:2007:PSF

[Flo07] Samuel C. Florman. Power, Speed, and Form: Engineers and the Making of the Twentieth Century (review). Technology and Culture, 48(3):635–636, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218705.

Foran:2003:KCR

[For03] Maxwell Foran. Kill and Chill: Restructuring Canada’s Beef Commodity Chain (review). Technology and Culture, 44(2): 425–426, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42017.

Fortun:2004:SPC

[For04] Kim Fortun. Synthetic Planet: Chemicals, Politics, and the Hazards of Modern Life (review). Technology and Culture, 45 (3):643–645, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172978.

Foran:2005:EPC

[For05] Maxwell Foran. Empty Pastures: Confined Animals and the Transformation of the Rural Landscape (review). Technology and Culture, 46(2):466–468, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184318.

Forden:2007:HWM

[For07] Geoffrey E. Forden. How the World’s most underdeveloped nations get the World’s most dangerous weapons. Technology and Culture, 48(1):92–103, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209834.

Forden:2008:PAW

[For08] Geoffrey E. Forden. Preventive Attack and Weapons of Mass Destruction: A Comparative Historical Analysis (review). Technology and Culture, 49(1):251–252, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230903.

Fors:2009:FJG

[For09] Hjalmar Fors. “Away! away to Falun!”: J. G. Gahn and the application of enlightenment chemistry to smelting. Technology and Culture, 50(3):549–568, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270559.

Foster:2004:BBA

[Fos04] Mark S. Foster. Buyways: Billboards, Automobiles, and the American Landscape (review). Technology and Culture, 45(4): 857–858, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175745.

Foster:2009:BDS

[Fos09] Amy Foster. Von Braun: Dreamer of Space, Engineer of War (review). Technology and Culture, 50(2):464–466, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263351.

Fouche:2006:ISA

[Fou06] Rayvon Fouché. The Inventive Spirit of African Americans: Patented Ingenuity (review). Technology and Culture, 47(1): 221–222, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196043.

Fox:2000:HDS

[Fox00] Robert Fox. Histoire du Service de la production thermique d’Electricite de France. Vol. 1, 1946–1973 (review). Technology and Culture, 41(1):154–156, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33475.

Fox:2002:DPN

[Fox02] Robert Fox. Disziplin, Profession und Nation: Die Ideologie der Chemie in Frankreich vom Zweiten Kaiserreich bis in die Zwischenkriegszeit (review). Technology and Culture, 43(4): 811–812, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34015.

Frame:2002:PDH

[Fra02] Paul W. Frame. Permissible Dose: A History of Radiation Protection in the Twentieth Century (review). Technology and Culture, 43(1):209–210, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33865.

Franz:2003:TNM

[Fra03] Kathleen Franz. On time, the National Museum of American History. Technology and Culture, 44(1):142–146, January 2003.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/40154.

Franz:2006:FRH

[Fra06] Kathleen Franz. Flying Down to Rio: Hollywood, Tourists, and Yankee Clippers (review). Technology and Culture, 47(2):438– 440, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200845.

Franz:2008:DWF

[Fra08] Kathleen Franz. Driving Women: Fiction and Automobile Culture in Twentieth-Century America (review). Technology and Culture, 49(1):277–279, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230904.

Franz:2009:MHE

[Fra09] Kathleen Franz. Motoring: The Highway Experience in America (review). Technology and Culture, 50(1):234–235, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258549.

Fredrich:2001:EOI

[Fre01] Augustine J. Fredrich. The Engines of Our Ingenuity: An Engineer Looks at Technology and Culture (review). Technology and Culture, 42(1):160–161, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33668.

Freeze:2006:MFT

[Fre06] Karen Johnson Freeze. Mobilisierung der Frauen: Technik, Geschlecht und Kalter Krieg in der DDR (review). Technology and Culture, 47(1):237–238, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196044.

Freeze:2007:ITT

[Fre07] Karen Johnson Freeze. Innovation and technology transfer during the Cold War: The case of the open-end spinning machine from Communist Czechoslovakia. Technology and Culture, 48 (2):249–285, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215392.

Friedel:2001:TPP

[Fri01] Robert D. Friedel. Tupperware: The Promise of Plastic in 1950s America (review). Technology and Culture, 42(1):183– 184, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33669.

Friedel:2002:FT

[Fri02a] Robert D. Friedel. Food for thought. Technology and Culture, 43(4):842–845, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34008.

Friedewald:2002:WMR

[Fri02b] Michael Friedewald. From Whirlwind to MITRE: The R&D Story of the SAGE Air Defense Computer (review). Technology and Culture, 43(1):203–204, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33866.

Friedel:2007:STF

[Fri07a] Robert D. Friedel. Success through Failure: The Paradox of Design (review). Technology and Culture, 48(2):414–416, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215393.

Frisch:2007:ZRM

[Fri07b] Michael Frisch. Zoned Out: Regulation, Markets, and Choices in Transportation and Metropolitan Land Use (review). Technology and Culture, 48(4):880–882, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223356.

Friedel:2008:MHI

[Fri08] Robert Friedel. The Mindful Hand: Inquiry and Invention from the Late Renaissance to Early Industrialisation (review). Technology and Culture, 49(4):1045–1046, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255459.

Fridlund:2009:BWB

[Fri09] Mats Fridlund. Buda’s Wagon: A Brief History of the Car Bomb (review). Technology and Culture, 50(1):242–243, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258544.

Frost:2001:ISS

[Fro01a] Gary Lewis Frost. Inventing schemes and strategies: The making and selling of the Fessenden oscillator. Technology and Culture, 42(3):462–488, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33762.

Frost:2001:VDM

[Fro01b] Robert L. Frost. La ville délétère: Médecins et ingénieurs dans l’espace urbain, XVIIIe–XIXe siècle (review). Technology and Culture, 42(3):568–570, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33763.

Frost:2005:SMM

[Fro05] Gary Lewis Frost. Signor Marconi’s Magic Box: The Most Remarkable Invention of the 19th Century and the Amateur Inventor Whose Genius Sparked a Revolution (review). Technology and Culture, 46(4):827–828, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192292.

Frohlich:2009:FPH

[Fro09] Xaq Frohlich. Fruits and Plains: The Horticultural Transformation of America (review). Technology and Culture, 50(4): 921–922, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363423.

Gaggio:2002:NGS

[Gag02] Dario Gaggio. Negotiating the gold standard: The geographical and political construction of gold fineness in Twentieth-Century Italy. Technology and Culture, 43(2):291–314, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33909.

Gaggio:2003:PPS

[Gag03] Dario Gaggio. Un pubblico per la scienza: La divulgazione scientifica nell’Italia in formazione (review). Technology and Culture, 44(4):815–817, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49693.

Gaggio:2007:INS

[Gag07] Dario Gaggio. Incorporare la natura: Storie ambientali del novecento (review). Technology and Culture, 48(3):630–632, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218706.

Gall:2000:WEA

[Gal00] Alexander Gall. Weisse Elefanten: Anspruch und Scheitern technischer Grossprojekte im 20. Jahrhundert (review). Technology and Culture, 41(3):615–616, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33557.

Gamst:2002:SRL

[Gam02] Frederick C. Gamst. Set Up Running: The Life of a Pennsylvania Railroad Engineman, 1904–1949 (review). Technology and Culture, 43(2):441–442, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33910.

Gamst:2005:RSS

[Gam05] Frederick C. Gamst. Rail, Steam, and Speed: The “Rocket” and the Birth of Steam Locomotion (review). Technology and Culture, 46(3):666–668, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186615.

Gamst:2007:BBB

[Gam07] Frederick C. Gamst. Boston’s Back Bay: The Story of America’s Greatest Nineteenth-Century Landfill Project (review). Technology and Culture, 48(3):655–656, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218707.

Ganz:2000:SAM

[Gan00] Cheryl Ganz. Science advancing mankind. Technology and Culture, 41(4):783–787, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33611.

Ganz:2003:CCM

[Gan03] Cheryl Ganz. Collecting Colonialism: Material Culture and Colonial Change (review). Technology and Culture, 44(2):410– 412, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42018.

Ganz:2005:AAM

[Gan05] Cheryl Ganz. Assembling Art: The Machine and the American Avant-Garde (review). Technology and Culture, 46(3):682– 684, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186616.

Gardner:2009:BGM

[Gar09] Kirsten E. Gardner. Building Genetic Medicine: Breast Cancer, Technology, and the Comparative Politics of Health Care (review). Technology and Culture, 50(2):478–479, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263347.

Gucciardo:2007:TSL

[GC07] Dorotea Gucciardo and John Crosmun. Taking stock of la longue durée: ICOHTEC 2006, Leicester, England, 15–20 August 2006. Technology and Culture, 48(1):153–157, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209836.

Geiger:2000:CEC

[Gei00] Reed G. Geiger. The Civil Engineering of Canals and Railways before 1850 (review). Technology and Culture, 41(2):370– 371, April 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33522.

Gelzer:2004:LSG

[Gel04] Christian Gelzer. Like Sex with Gods: An Unorthodox History of Flying (review). Technology and Culture, 45(3):624– 625, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172979.

George:2006:WWS

[Geo06] Alice L. George. The Week the World Stood Still: Inside the Secret Cuban Missile Crisis (review). Technology and Culture, 47(2):454–455, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200846.

Giebelhaus:2002:SSU

[Gie02] August W. Giebelhaus. The Skule Story: The University of Toronto Faculty of Applied Science and Engineering, 1873– 2000 (review). Technology and Culture, 43(3):629–630, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33962.

Giffard:2007:AHL

[Gif07] Hermione Giffard. Alfred Herbert Ltd. and the British Machine Tool Industry, 1887–1983 (review). Technology and Culture, 48 (3):644–646, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218708.

Gillispie:2006:EIV

[Gil06] Charles Coulston Gillispie. De l’espionnage industriel à la veille technologique (review). Technology and Culture, 47(4):839–840, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204977.

Ginorio:2007:SSR

[Gin07] Angela B. Ginorio. Suffering for Science: Reason and Sacrifice in Modern America (review). Technology and Culture, 48(4): 864–866, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223357.

Gispen:2000:SR

[Gis00] Kees Gispen. Siemens, 1918–1945 (review). Technology and Culture, 41(3):606–608, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33558.

Gispen:2001:APG

[Gis01] Kees Gispen. Das Atlantropa-Projekt–Die Geschichte einer gescheiterten Vision: Hermann Sorgel und die Absenkung des Mittelmeers (review). Technology and Culture, 42(3):596–598, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33764.

Gispen:2007:GAI

[Gis07] Kees Gispen. Georg von Arco (1869–1940) — Ingenieur, Pazifist, Technischer Direktor von Telefunken: Eine Erfinderbiographie (review). Technology and Culture, 48(1):224–226, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209835.

Gitelman:2001:RMA

[Git01] Lisa Gitelman. Recorded Music in American Life: The Phonograph and Popular Memory, 1890–1945 (review). Technology and Culture, 42(1):173–174, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33670.

Glasmeier:2001:STT

[Gla01] Amy Glasmeier. Selling the True Time: Nineteenth-Century Timekeeping in America (review). Technology and Culture, 42 (4):790–791, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33810.

Glen:2003:FNB

[Gle03] Robert Glen. Firms, Networks and Business Values: The British and American Cotton Industries since 1750 (review). Technology and Culture, 44(3):623–625, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46093.

Glen:2008:LLL

[Gle08] Robert Glen. Labor and Laborers of the Loom: Mechanization and Handloom Weavers, 1780–1840 (review). Technology and Culture, 49(4):1095–1097, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255536.

Glines:2000:ANP

[Gli00] Carroll V. Glines. By Airship to the North Pole: An Archaeology of Human Exploration (review). Technology and Culture, 41(4):845–846, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33612.

Gloster:2002:ONU

[Glo02] David Gloster. One Nation Underground: The Fallout Shelter in American Culture (review). Technology and Culture, 43 (2):452–453, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33911.

Goff:2009:MHM

[Gof09] Robert Goff. Mining the Home Movie: Excavations in Histories and Memories (review). Technology and Culture, 50(4): 956–957, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363444.

Goldman:2000:PHS

[Gol00a] Joanne Abel Goldman. Public Health and Social Justice in the Age of Chadwick (review). Technology and Culture, 41(1): 176–177, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33477.

Goldman:2000:NSN

[Gol00b] Joanne Abel Goldman. National science in the Nation’s heartland: The Ames Laboratory at Iowa State University. Technology and Culture, 41(3):435–459, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33559.

Goldman:2000:PFI

[Gol00c] Steven L. Goldman. Progress: Fact or Illusion? (review). Technology and Culture, 41(1):116–117, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33476.

Gomery:2000:YAH

[Gom00] Douglas Gomery. You Ain’t Heard Nothin’ Yet: The American Talking Film: History and Memory, 1927–1949 (review). Technology and Culture, 41(4):819–821, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33613.

Gomery:2005:CER

[Gom05] Douglas Gomery. The Cinema Effect (review). Technology and Culture, 46(3):686–687, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186617.

Goodyear:2004:IFA

[Goo04] Anne Collins Goodyear. Imagining Flight: Aviation and Popular Culture (review). Technology and Culture, 45(3):627– 629, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172980.

Goodman:2006:ECH

[Goo06] Seymour E. Goodman. Engineering Communism: How Two Americans Spied for Stalin and Founded the Soviet Silicon Valley (review). Technology and Culture, 47(3):676–678, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201452.

Gooday:2007:NCS

[Goo07] Graeme Gooday. Nervous Conditions: Science and the Body Politic in Early Industrial Britain (review). Technology and Culture, 48(4):854–856, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223358.

Goodyear:2009:OAP

[Goo09] Frank H. Goodyear. The Origins of American Photography: From Daguerreotype to Dry-Plate, 1839–1885: The Hallmark Photographic Collection at the Nelson-Atkins Museum of Art (review). Technology and Culture, 50(2):471–473, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263352.

Gordon:2001:ICS

[Gor01a] Sarah Gordon. Iron Confederacies: Southern Railways, Klan Violence, and Reconstruction (review). Technology and Culture, 42(2):366–367, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33713.

Gordon:2001:MST

[Gor01b] Sarah Gordon. My Sisters Telegraphic: Women in the Telegraph Office, 1846–1950 (review). Technology and Culture, 42 (4):786–787, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33811.

Gorman:2001:SPE

[Gor01c] Hugh S. Gorman. Smokestacks and Progressives: Environmentalists, Engineers, and Air Quality in America, 1881–1951 (review). Technology and Culture, 42(1):149–150, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33671.

Gordon:2002:RAL

[Gor02a] Sarah Gordon. Recasting American Liberty: Gender, Race, Law, and the Railroad Revolution, 1865–1920 (review). Technology and Culture, 43(4):807–808, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34017.

Gorman:2002:TNO

[Gor02b] Hugh S. Gorman. Transforming New Orleans and Its Environs: Centuries of Change (review). Technology and Culture, 43(3):617–618, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33963.

Gordin:2003:MPH

[Gor03] Michael D. Gordin. A modernization of “Peerless homogeneity”: The creation of Russian smokeless gunpowder. Technology and Culture, 44(4):677–702, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49694.

Gorman:2004:DDD

[Gor04a] Hugh S. Gorman. Deceit and Denial: The Deadly Politics of Industrial Pollution (review). Technology and Culture, 45(1): 219–221, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51748.

Gorman:2004:AWL

[Gor04b] Hugh S. Gorman. From Apocalypse to Way of Life: Environmental Crisis in the American Century (review). Technology and Culture, 45(3):638–639, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172981.

Gordon:2005:MHF

[Gor05] Robert B. Gordon. Mine Hill in Franklin and Sterling Hill in Ogdensburg, New Jersey: Mining History, 1765–1900 (review). Technology and Culture, 46(4):843–844, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192293.

Gordon:2006:EMN

[Gor06] Robert B. Gordon. European Metals in Native Hands: Rethinking Technological Change, 1640–1683 (review). Technology and Culture, 47(2):415–417, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200847.

Gordon:2007:RAH

[Gor07] Robert B. Gordon. A Revolution in Arms: A History of the First Repeating Rifles (review). Technology and Culture, 48 (2):457–458, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215394.

Gordon:2008:BLI

[Gor08] Robert B. Gordon. Brunel: In Love with the Impossible (review). Technology and Culture, 49(2):467–469, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236985.

Gough:2000:FAS

[Gou00] Jerry B. Gough. Les fonds d’archives Seguin: Aux origines de la revolution industrielle en France, 1790–1860 (review). Technology and Culture, 41(3):586–587, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33560.

Gough:2001:LTA

[Gou01] Jerry B. Gough. L’Invention technique au siècle des Lumières (review). Technology and Culture, 42(4):782–783, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33812.

Gough:2004:AED

[Gou04a] Jerry B. Gough. Affinity, That Elusive Dream: A Genealogy of the Chemical Revolution (review). Technology and Culture, 45 (3):645–647, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172982.

Gough:2004:WCB

[Gou04b] Jerry B. Gough. When Champagne Became French: Wine and the Making of a National Identity, and: Camembert: A National Myth (review). Technology and Culture, 45(2):423– 426, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55663.

Gouzevitch:2004:PSP

[Gou04c] Irina Gouzevitch. “Par la science, pour la patrie”: L’association francaise pour l’avancement des sciences (1872–1914) (review). Technology and Culture, 45(4):878–879, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175746.

Grant:2001:DRR

[Gra01] H. Roger Grant. The Diesel Revolution (review). Technology and Culture, 42(2):369–370, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33714.

Graber:2008:PTT

[Gra08] Frédéric Graber. Purity and theory: Theoretical tools at Ponts et Chaussées, circa 1800. Technology and Culture, 49(4): 860–883, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255439.

Greuther:2001:RVI

[Gre01] Marc Greuther. Railways and the Victorian Imagination (review). Technology and Culture, 42(3):604–606, July 2001. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33765.

Greuther:2002:LIR

[Gre02] Marc Greuther. Landmarks on the Iron Road: Two Centuries of North American Railroad Engineering (review). Technology and Culture, 43(2):440–441, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33913.

Grenier:2003:SWC

[Gre03] John Grenier. Strategic Warfare in Cyberspace (review). Technology and Culture, 44(1):190–191, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40155.

Greenberg:2004:VVV

[Gre04a] Joshua M. Greenberg. Veni, Vidi, Video: The Hollywood Empire and the VCR (review). Technology and Culture, 45(1): 211–212, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51749.

Greuther:2004:SSE

[Gre04b] Marc Greuther. Stationary Steam Engines of Great Britain: The National Photographic Collection (review). Technology and Culture, 45(1):181–183, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51751.

Greuther:2004:PCH

[Gre04c] Marc Greuther. Present at the creation: Highland park, Michigan, 23 March 1914, 9:20 A.M. Technology and Culture, 45(1): 120–122, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/51750.

Greuther:2004:TS

[Gre04d] Marc Greuther. Tripp’s sawmill, 1926. Technology and Culture, 45(4):808–811, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175747.

Green:2005:TAA

[Gre05] Venus Green. Technology and the African-American Experience: Needs and Opportunities for Study (review). Technology and Culture, 46(1):198–200, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179759.

Greuther:2006:BT

[Gre06a] Marc Greuther. Brief transit. Technology and Culture, 47(4): 787–790, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204978.

Greuther:2006:PM

[Gre06b] Marc Greuther. Persistence of memory. Technology and Culture, 47(3):566–569, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/201453.

Greenberg:2008:GAP

[Gre08a] Brian Greenberg. The Gilded Age: Perspectives on the Origin of Modern America (review). Technology and Culture, 49(1): 274–275, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230905.

Greene:2008:HCL

[Gre08b] Ann N. Greene. The Horse in the City: Living Machines in the Nineteenth Century (review). Technology and Culture, 49 (2):462–463, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236986.

Greenslit:2008:PRK

[Gre08c] Nathan Greenslit. Pharmaceutical Reason: Knowledge and Value in Global Psychiatry (review). Technology and Culture, 49(3):828–829, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243100.

Gross:2008:ASE

[Gro08] Laurence F. Gross. American Silk, 1830–1930: Entrepreneurs and Artifacts (review). Technology and Culture, 49(3):796– 798, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243106.

Gross:2009:LEP

[Gro09] Laurence F. Gross. The Lowell Experiment: Public History in a Postindustrial City (review). Technology and Culture, 50(1): 226–228, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258590.

Gordin:2002:TST

[GS02] Michael D. Gordin and Sam Schweber. Thinking systematically: Thomas and Agatha Hughes, systems, experts, and computers. Technology and Culture, 43(2):390–397, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33912.

Gucciardo:2008:CAN

[Guc08] Dorotea Gucciardo. Canada and Arctic North America: An Environmental History (review). Technology and Culture, 49 (2):498–499, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236987.

Gudmestad:2007:PCS

[Gud07] Robert H. Gudmestad. Planting a Capitalist South: Masters, Merchants, and Manufacturers in the Southern Interior, 1790– 1860 (review). Technology and Culture, 48(2):451–453, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215395.

Guilmartin:2000:ABA

[Gui00] John Francis Guilmartin. American and British Aircraft Carrier Development, 1919–1941, and: Destined for Glory: Dive Bombing, Midway, and the Evolution of Carrier Airpower (review). Technology and Culture, 41(4):841–843, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33614.

Guilmartin:2003:CTF

[Gui03] John Francis Guilmartin. Computers Take Flight: A History of NASA’s Pioneering Digital Fly-By-Wire Project (review). Technology and Culture, 44(1):185–186, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40156.

Guigueno:2006:FIP

[Gui06] Vincent Guigueno. La formation des ingénieurs en perspective: Modèles de référence et réseaus de médiation, XVIIIe– XXe siècles (review). Technology and Culture, 47(1):198–200, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196045.

Guigueno:2008:BHS

[Gui08] Vincent Guigueno. Building a high-speed society: France and the Aérotrain, 1962–1974. Technology and Culture, 49(1):21– 40, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230906.

Gwyn:2004:TNB

[Gwy04] David Gwyn. Tredegar, Newcastle, Baltimore: The swivel truck as paradigm of technology transfer. Technology and Culture, 45(4):778–794, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/175748.

Gwynne:2009:RMR

[Gwy09] Robert Gwynne. The railway museum reinvented: The Cité du Train (Mulhouse) and the Nederlands Spoorwegmuseum (Utrecht). Technology and Culture, 50(4):876–882, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363418.

Holley:2000:C

[HA00] I. B. Holley and Robert McCormick Adams. Communications. Technology and Culture, 41(2):402–403, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33525.

Hacker:2003:TFP

[Hac03] Barton C. Hacker. Throwing Fire: Projectile Technology through History (review). Technology and Culture, 44(4):802– 803, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/49695.

Hacker:2004:MTC

[Hac04] Barton C. Hacker. Memoirs: A Twentieth-Century Journey in Science and Politics (review). Technology and Culture, 45 (2):460–461, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55664.

Hackemer:2005:ASS

[Hac05a] Kurt Hackemer. Arms and the State: Sir William Armstrong and the Remaking of British Naval Power, 1854–1914 (review). Technology and Culture, 46(3):652–654, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186618.

Hacker:2005:GFW

[Hac05b] Barton C. Hacker. Gunpowder and Firearms: Warfare in Medieval India, and: Warfare and Weaponry in South Asia 1000– 1800 (review). Technology and Culture, 46(4):880–881, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
192294.

Hacker:2006:SGH

[Hac06] Barton C. Hacker. Soldiers and Ghosts: A History of Battle in Classical Antiquity (review). Technology and Culture, 47(1): 181–182, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196046.

Hacker:2007:WSB

[Hac07] Barton C. Hacker. Warfare State: Britain, 1920–1970 (review). Technology and Culture, 48(2):459–460, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215396.

Hacker:2009:DSS

[Hac09] Barton C. Hacker. Disrupting Science: Social Movements, American Scientists, and the Politics of the Military, 1945– 1975 (review). Technology and Culture, 50(4):947–949, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363439.

Hafter:2003:CIL

[Haf03] Daryl M. Hafter. The cost of inventiveness: Labor’s struggle with Management’s machine. Technology and Culture, 44(1): 102–113, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40157.

Hafter:2005:MC

[Haf05] Daryl M. Hafter. Masterwork completed. Technology and Culture, 46(4):813–816, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192295.

Hafter:2007:CNI

[Haf07] Daryl M. Hafter. Les Chemins de la nouveauté: Innover, inventer au regard de l’histoire (review). Technology and Culture, 48(4):850–852, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/223359.

Hafter:2009:NIP

[Haf09] Daryl Hafter. La naissance de l’industrie à Paris entre sueurs et vapeurs 1780–1830 (review). Technology and Culture, 50(1): 213–214, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258571.

Haigh:2003:AHM

[Hai03] Thomas Haigh. From 0 to 1: An Authoritative History of Modern Computing (review). Technology and Culture, 44(4): 841–842, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/49696.

Hall:2001:TFS

[Hal01] Bert S. Hall. Theatre of Fire: Special Effects in the Early English and Scottish Theatre (review). Technology and Culture, 42(1):192–193, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33672.

Hales:2007:NVP

[Hal07] Peter B. Hales. Natural Visions: The Power of Images in American Environmental Reform (review). Technology and Culture, 48(2):442–444, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/215397.

Hammonds:2003:BRA

[Ham03] Evelynn Maxine Hammonds. Black Rice: The African Origins of Rice Cultivation in the Americas (review). Technology and Culture, 44(1):162–163, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40158.

Hammers:2005:HCI

[Ham05] Roslyn Lee Hammers. “How can it be that a crow’s tail can hold water?”: The square pallet pump in Lou Shu’s pictures of tilling and weaving. Technology and Culture, 46(1):132–136, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179755.

Hamblin:2006:EGA

[Ham06a] Jacob Darwin Hamblin. Exorcising ghosts in the Age of Automation: United Nations experts and Atoms for Peace. Technology and Culture, 47(4):734–756, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204979.

Hamilton:2006:CAT

[Ham06b] Shane Hamilton. A Companion to American Technology (review). Technology and Culture, 47(2):442–443, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200848.

Hamblin:2007:ESA

[Ham07] Jacob Darwin Hamblin. Eisenhower, Science Advice, and the Nuclear Test-Ban Debate, 1945–1963 (review). Technology and Culture, 48(4):892–893, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223360.

Hamblin:2008:AHP

[Ham08a] Jacob Darwin Hamblin. American Hegemony and the Postwar Reconstruction of Science in Europe (review). Technology and Culture, 49(2):489–491, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/236988.

Hamilton:2008:UCR

[Ham08b] Shane Hamilton. Uncommon Carriers (review). Technology and Culture, 49(3):801–803, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243098.

Hammers:2009:GTP

[Ham09] Roslyn Lee Hammers. Graphics and Text in the Production of Technical Knowledge in China: The Warp and the Weft (review). Technology and Culture, 50(1):202–204, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258553.

Hard:2000:OIT

[Har00a] Mikael Hard. Der Optimismus der Ingenieure: Triumph der Technik in der Krise der Moderne um 1900 (review). Technology and Culture, 41(3):597–598, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33561.

Hartley:2000:QTR

[Har00b] Peter Hartley. Questioning Technology (review). Technology and Culture, 41(4):852–854, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33615.

Harris:2001:MRE

[Har01] Howell John Harris. Manufacturing Rationality: The Engineering Foundations of the Managerial Revolution (review). Technology and Culture, 42(3):582–583, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33766.

Harpold:2002:MTC

[Har02] Terry Harpold. Masons, Tricksters, and Cartographers: Comparative Studies in the Sociology of Scientific and Indigenous Knowledge (review). Technology and Culture, 43(2):398–401, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33914.

Haring:2003:FMH

[Har03] Kristen Haring. The “Freer Men” of ham radio: How a technical hobby provided social and spatial distance. Technology and Culture, 44(4):734–761, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49697.

Harvey:2004:RWW

[Har04] Mark W. T. Harvey. The River We Have Wrought: A History of the Upper Mississippi (review). Technology and Culture, 45 (2):432–433, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55665.

Harris:2005:SSM

[Har05] Richard Harris. Suburban Steel: The Magnificent Failure of the Lustron Corporation, 1945–1951 (review). Technology and Culture, 46(4):850–852, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192296.

Hard:2006:AIV

[Har06a] Mikael Hard. Artefakter: Industrin, vetenskapen och de tekniska nätverken (review). Technology and Culture, 47(3): 702–703, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201455.

Harris:2006:PHR

[Har06b] Richard Harris. The Prefabricated Home (review). Technology and Culture, 47(4):852–853, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/204980.

Harkness:2007:PAA

[Har07a] Deborah E. Harkness. Promethean Ambitions: Alchemy and the Quest to Perfect Nature (review). Technology and Culture, 48(1):192–193, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209837.

Hartig:2007:OEC

[Har07b] Anthea M. Hartig. Orange Empire: California and the Fruits of Eden (review). Technology and Culture, 48(1):215–216, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209838.

Haard:2008:NET

[Hår08a] Mikael Hård. Networking Europe: Transnational Infrastructures and the Shaping of Europe, 1850–2000 (review). Technology and Culture, 49(4):1073–1074, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255505.

Harwood:2008:NSA

[Har08b] Jonathan Harwood. Nützliche Schädlinge: Angewandte Entomologie, chemische Industrie und Landwirtschaftspolitik in der Schweiz, 1874–1952 (review). Technology and Culture, 49(1): 257–259, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230907.

Hashagen:2002:CWM

[Has02] Ulf Hashagen. Der Computer als Werkzeug und Medium: Die geistigen und technischen Wurzeln des Personal Computers (review). Technology and Culture, 43(1):201–203, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33867.

Hascher:2006:MVN

[Has06] Michael Hascher. Motorisierung und “Volksgemeinschaft”: Das Nationalsozialistische Kraftfahrkorps (NSKK), 1931–1945 (review). Technology and Culture, 47(4):857–858, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
204981.

Headrick:2001:NTI

[Hea01] Daniel R. Headrick. A Nation Transformed by Information: How Information Has Shaped the United States from Colonial Times to the Present (review). Technology and Culture, 42(4): 774–776, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33814.

Headrick:2002:FHG

[Hea02] Daniel R. Headrick. The Fire of His Genius: Robert Fulton and the American Dream (review). Technology and Culture, 43 (2):428–429, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33915.

Hecht:2008:AHR

[Hec08] David K. Hecht. The atomic hero: Robert Oppenheimer and the making of scientific icons in the Early Cold War. Technology and Culture, 49(4):943–966, October 2008. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255442.

Hecht:2009:SSP

[Hec09] David Hecht. In Sputnik’s Shadow: The President’s Science Advisory Committee and Cold War America (review). Technology and Culture, 50(4):944–945, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363437.

Heide:2004:MPD

[Hei04] Lars Heide. Monitoring people: Dynamics and hazards of record management in France, 1935–1944. Technology and Culture, 45(1):80–101, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/51752.

Henrickson:2009:GSR

[Hen09] Heidi Allene Henrickson. Galactic Suburbia: Recovering Women’s Science Fiction (review). Technology and Culture, 50(4):959–960, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363446.

Herrin:2000:TAC

[Her00] Dean A. Herrin. Transforming the Appalachian Countryside: Railroads, Deforestation, and Social Change in West Virginia, 1880–1920 (review). Technology and Culture, 41(4):804–805, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33616.

Herken:2001:WTB

[Her01a] Gregg Herken. Way Out There in the Blue: Reagan, Star Wars, and the End of the Cold War (review). Technology and Culture, 42(2):385–386, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33715.

Herschman:2001:WMB

[Her01b] Joel Herschman. The Wise Master Builder: Platonic Geometry in Plans of Medieval Abbeys and Cathedrals (review). Technology and Culture, 42(3):565–567, July 2001. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33767.

Herzig:2001:CTC

[Her01c] Rebecca M. Herzig. Cleavage: Technology, Controversy, and the Ironies of the Man-Made Breast (review). Technology and Culture, 42(3):573–575, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33768.

Herken:2005:WWF

[Her05] Gregg Herken. Whole World on Fire: Organizations, Knowledge, and Nuclear Weapons Devastation (review). Technology and Culture, 46(2):453–455, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184319.

Herken:2007:JRO

[Her07a] Gregg Herken. J. Robert Oppenheimer: A Life (review). Technology and Culture, 48(4):889–891, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223361.

Hersch:2007:TLA

[Her07b] Matthew H. Hersch. Testing the Limits: Aviation Medicine and the Origins of Manned Space Flight (review). Technology and Culture, 48(4):897–899, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223362.

Hersch:2008:ASN

[Her08a] Matthew H. Hersch. Apollo’s stepchildren: New works on the American lunar program. Technology and Culture, 49(2):449– 455, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236969.

Hersch:2008:SFN

[Her08b] Matthew H. Hersch. Science in Flux: NASA’s Nuclear Program at Plum Brook Station, 1955–2005 (review). Technology and Culture, 49(1):246–247, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230908.

Hersch:2009:DAH

[Her09a] Matthew Hersch. Digital Apollo: Human and Machine in Space Flight (review). Technology and Culture, 50(2):466–467, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263356.

Herzig:2009:TS

[Her09b] Rebecca Herzig. A thing for stories. Technology and Culture, 50(3):613–622, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270535.

Hessenbruch:2004:WAG

[Hes04] Arne Hessenbruch. Wachhund des Atomzeitalters: Geigerzahler in der Geschichte des Strahlenschutzes (review). Technology and Culture, 45(4):869–871, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175749.

Hessler:2004:GNB

[Heß04] Martina Heßler. Ganz normale Bilder: Historische Beitrage zur visuellen Herstellung von Selbstverstandlichkeit (review). Technology and Culture, 45(4):904–905, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175750.

Heximer:2007:IMN

[Hex07] Christina Heximer. Indigenous Modernities: Negotiating Architecture and Urbanism (review). Technology and Culture, 48(4): 879–880, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223363.

Heymann:2001:NMW

[Hey01] Matthias Heymann. Natur und Macht: Eine Weltgeschichte der Umwelt (review). Technology and Culture, 42(1):148–149, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33673.

Heymann:2006:HHC

[Hey06] Matthias Heymann. Hubris and Hybrids: A Cultural History of Technology and Science (review). Technology and Culture, 47(3):700–702, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201456.

Heymann:2008:GBH

[Hey08] Matthias Heymann. The Green and the Brown: A History of Conservation in Nazi Germany (review). Technology and Culture, 49(3):806–808, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/243123.

Higham:2001:SCM

[Hig01] Robin D. S. Higham. Sidewinder: Creative Missile Development at China Lake (review). Technology and Culture, 42(2): 384–385, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33716.

Higham:2003:SRJ

[Hig03] Robin D. S. Higham. Sunburst: The Rise of Japanese Naval Air Power, 1909–1941 (review). Technology and Culture, 44(1): 186–188, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40159.

Hilde:2002:ETP

[Hil02] Thomas C. Hilde. The Empirical Turn in the Philosophy of Technology (review). Technology and Culture, 43(4):840–841, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34018.

Hinton:2007:DAQ

[Hin07a] Diana Davids Hinton. Drilling Ahead: The Quest for Oil in the Deep South, 1945–2005 (review). Technology and Culture, 48(1):203–204, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209839.

Hintz:2007:ACR

[Hin07b] Eric S. Hintz. Andrew Carnegie (review). Technology and Culture, 48(4):884–885, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223364.

Hintz:2008:DIP

[Hin08] Eric S. Hintz. The Democratization of Invention: Patents and Copyrights in American Economic Development, 1790– 1920 (review). Technology and Culture, 49(1):270–272, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230909.

Hintz:2009:PPI

[Hin09] Eric S. Hintz. Portable power: Inventor Samuel Ruben and the birth of Duracell. Technology and Culture, 50(1):24–57, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258559.

Head-Miller:2001:AIB

[HM01] Jeanine Head-Miller. America’s Instrument: The Banjo in the Nineteenth Century (review). Technology and Culture, 42(4): 805–806, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33813.

Hoag:2008:TWP

[HÖ08] Heather J. Hoag and May-Britt Öhman. Turning water into power: Debates over the development of Tanzania’s Rufiji River Basin, 1945–1985. Technology and Culture, 49(3):624–651, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243114.

Hoadley:2001:SMG

[Hoa01] Steven Hoadley. Steam — the Museum of the Great Western Railway, Swindon. Technology and Culture, 42(4):764–766, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33816.

Hochfelder:2003:TMB

[Hoc03] David Hochfelder. Telegraph Messenger Boys: Labor, Communication, and Technology, 1850–1950 (review). Technology and Culture, 44(3):628–629, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46094.

Hochfelder:2005:SIR

[Hoc05] David Hochfelder. The Second Information Revolution (review). Technology and Culture, 46(4):835–836, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
192297.

Hoffman:2000:FTT

[Hof00] Abraham Hoffman. Flowing through Time: A History of the Iowa Institute of Hydraulic Research (review). Technology and Culture, 41(2):373–374, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33524.

Hofsommer:2002:NLI

[Hof02] Donovan L. Hofsommer. Nothing Like It in the World: The Men Who Built the Transcontinental Railroad, 1863–1869, and: Empire Express: Building the First Transcontinental Railroad (review). Technology and Culture, 43(1):169–170, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33868.

Hogle:2003:TDO

[Hog03] Linda F. Hogle. Twice Dead: Organ Transplants and the Reinvention of Death (review). Technology and Culture, 44(2):433– 435, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42019.

Hohler:2003:UAO

[Hoh03] Sabine Hohler. Upheaval from the Abyss: Ocean Floor Mapping and the Earth Science Revolution (review). Technology and Culture, 44(2):435–437, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42020.

Hoke:2007:SSA

[Hok07] Donald Hoke. The Stanley Steamer: America’s Legendary Steam Car (review). Technology and Culture, 48(2):468–470, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215398.

Holley:2001:BAS

[Hol01] I. B. Holley. Bases of Air Strategy: Building Airfields for the RAF, 1914–1945 (review). Technology and Culture, 42(4): 814–815, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33817.

Holley:2003:BHA

[Hol03] I. B. Holley. Blacktop: How asphalt paving came to the urban United States. Technology and Culture, 44(4):703–733, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49698.

Holbrook:2006:MBM

[Hol06] Daniel U. Holbrook. The Man behind the Microchip: Robert Noyce and the Invention of Silicon Valley (review). Technology and Culture, 47(3):678–680, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201458.

Holley:2009:MB

[Hol09] I. B. Holley, Jr. The mechanization of brickmaking. Technology and Culture, 50(1):82–102, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258587.

Hood:2004:MRR

[Hoo04] Clifton Hood. Metropolitan Railways: Rapid Transit in America (review). Technology and Culture, 45(2):446–447, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55666.

Horrocks:2002:EEF

[Hor02] Sally M. Horrocks. Enriching the Earth: Fritz Haber, Carl Bosch, and the Transformation of World Food (review). Technology and Culture, 43(3):622–623, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33964.

Horning:2004:SBN

[Hor04a] Susan Schmidt Horning. Sonic Boom: Napster, MP3, and the New Pioneers of Music (review). Technology and Culture, 45 (4):890–891, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175751.

Horning:2004:PRC

[Hor04b] Susan Schmidt Horning. The Poetics of Rock: Cutting Tracks, Making Records (review). Technology and Culture, 45(2):442– 444, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55667.

Horning:2006:LSB

[Hor06] Susan Schmidt Horning. Lost Sounds: Blacks and the Birth of the Recording Industry, 1890–1919 (review). Technology and Culture, 47(3):651–653, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201459.

Horrocks:2007:ICC

[Hor07] Sally M. Horrocks. Industrial chemistry and its changing patrons at the University of Liverpool, 1926–1951. Technology and Culture, 48(1):43–66, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209840.

Houchin:2003:BMC

[Hou03] Roy F. Houchin. The $5 Billion Misunderstanding: The Collapse of the Navy’s A-12 Stealth Bomber Program (review). Technology and Culture, 44(1):188–189, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40160.

Houltz:2004:MTR

[Hou04a] Anders Houltz. Modernity and Technology (review). Technology and Culture, 45(2):471–473, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55668.

Hounshell:2004:EF

[Hou04b] David A. Hounshell. Eugene S. Ferguson, 1916–2004. Technology and Culture, 45(4):911–921, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175752.

Hovis:2004:MWB

[Hov04] Logan W. Hovis. The Mining West: A Bibliography and Guide to the History and Literature of Mining in the American and Canadian West (review). Technology and Culture, 45(3):658– 659, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172983.

Howell:2000:LWH

[How00] Joel D. Howell. Looking Within: How X-Ray, CT, MRI, Ultrasound, and Other Medical Images are Created, and How They Help Physicians Save Lives (review). Technology and Culture, 41(4):815–816, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33617.

Howard:2006:GIP

[How06] Nicole Howard. Gutenberg and the Impact of Printing (review). Technology and Culture, 47(2):412–414, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/200849.

Howard:2008:ACP

[How08a] Nicole Howard. Agent of Change: Print Culture Studies after Elizabeth L. Eisenstein (review). Technology and Culture, 49 (4):1047–1048, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255487.

Howard:2008:EBS

[How08b] Nicole Howard. The Enlightenment and the Book: Scottish Authors and Their Publishers in Eighteenth-Century Britain, Ireland, and America (review). Technology and Culture, 49(1): 263–265, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230910.

Howard:2009:FDJ

[How09a] Nicole Howard. The Fertility Doctor: John Rock and the Reproductive Revolution (review). Technology and Culture, 50(4): 973–975, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363454.

Howard:2009:MMS

[How09b] Vicki Howard. Modernizing Main Street: Architecture and Consumer Culture in the New Deal (review). Technology and Culture, 50(3):694–696, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/270549.

Hilaire-Perez:2000:NDI

[HP00] Liliane Hilaire-Perez. Natural Dyestuffs and Industrial Culture in Europe, 1750–1880 (review). Technology and Culture, 41 (2):360–362, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33523.

Hilaire-Perez:2006:DTK

[HPV06] Liliane Hilaire-Perez and Catherine Verna. Dissemination of technical knowledge in the Middle Ages and the Early Modern Era: New approaches and methodological issues. Technology and Culture, 47(3):536–565, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201457.

Hubbard:2009:CEH

[Hub09] Jennifer Hubbard. Cod: The Ecological History of the North Atlantic Fisheries (review). Technology and Culture, 50(1): 249–250, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258570.

Hughes:2000:LSS

[Hug00a] Thomas Parke Hughes. Lessons from Soviet science and technology. Technology and Culture, 41(2):348–352, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33526.

Hugill:2000:STW

[Hug00b] Peter J. Hugill. Science and Technology in World History (review). Technology and Culture, 41(3):566–568, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33562.

Hugill:2002:EVB

[Hug02] Peter J. Hugill. The Electric Vehicle and the Burden of History (review). Technology and Culture, 43(1):183–184, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33869.

Hugill:2005:UWH

[Hug05] Peter J. Hugill. Under the Wire: How the Telegraph Changed Diplomacy (review). Technology and Culture, 46(2):422–424, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184320.

Hughes:2007:ADC

[Hug07] Thomas Parke Hughes. Alfred D. Chandler, Jr., 1918–2007. Technology and Culture, 48(4):910–912, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223365.

Hughes:2009:SFT

[Hug09a] Thomas P. Hughes. SHOT founders’ themes and problems. Technology and Culture, 50(3):594–599, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270534.

Hugill:2009:NSC

[Hug09b] Peter J. Hugill. Nexus: Strategic Communications and American Security in World War I (review). Technology and Culture, 50(4):949–950, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363440.

Hull:2003:MDT

[Hul03] Tom Hull. “More deadly than war”: High-lead steam logging unit. Technology and Culture, 44(2):355–358, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42021.

Hull:2009:FP

[Hul09] Tom Hull. Family portrait. Technology and Culture, 50(4): 873–875, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363417.

Humphreys:2009:SEP

[Hum09] Margaret Humphreys. Searching Eyes: Privacy, the State, and Disease Surveillance in America (review). Technology and Culture, 50(2):480–481, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/263358.

Hunley:2000:SRN

[Hun00] J. D. Hunley. Spaceflight Revolution: NASA Langley Research Center from Sputnik to Apollo (review). Technology and Culture, 41(1):166–168, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33478.

Hunter:2002:SCM

[Hun02] Michael Cyril William Hunter. Solomon’s Child: Method in the Early Royal Society of London (review). Technology and Culture, 43(4):787–789, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34019.

Hurlburt:2002:ZMV

[Hur02] Holly S. Hurlburt. Zecca: The Mint of Venice in the Middle Ages (review). Technology and Culture, 43(3):587–589, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33965.

Hurt:2004:AAT

[Hur04] R. Douglas Hurt. American Agriculture in the Twentieth Century: How It Flourished and What It Cost (review). Technology and Culture, 45(3):655–656, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172984.

Hutchins:2000:PNM

[Hut00] Louis P. Hutchins. Poised for the new millennium: The technical museums of Prague, Vienna, and Berlin. Technology and Culture, 41(3):530–536, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33563.

Henninger-Voss:2001:GRV

[HV01] Mary J. Henninger-Voss. Glassmaking in Renaissance Venice: The Fragile Craft (review). Technology and Culture, 42(4): 795–797, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33815.

Henninger-Voss:2006:NBA

[HV06] Mary J. Henninger-Voss. Neo-Baroque Aesthetics and Contemporary Entertainment (review). Technology and Culture, 47 (3):657–658, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201454.

Henninger-Voss:2007:CSR

[HV07] Mary J. Henninger-Voss. Craft secrets religiously kept: Neil Kamil’s Fortress of the Soul. Technology and Culture, 48(4): 836–839, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223384.

Horning:2005:RSI

[HW05] Susan Schmidt Horning and James C. Williams. Recent symposia of the International Committee for the History of Technology, 2002–2004. Technology and Culture, 46(3):594–603, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186619.

Hyde:2000:WIA

[Hyd00] Charles K. Hyde. The Works: The Industrial Architecture of the United States (review). Technology and Culture, 41(3):591– 593, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33564.

Hyde:2001:DTM

[Hyd01] Charles K. Hyde. Detroit, Toledo, and Milwaukee Roundhouse, Greenfield Village, Dearborn. Technology and Culture, 42(4): 754–759, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33818.

Hyldtoft:2000:TLT

[Hyl00] Ole Hyldtoft. Teknikens landskap: En teknikhistorisk antologi tillagnad Svante Lindqvist (review). Technology and Culture, 41(3):580–581, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33565.

Isler:2004:C

[IEDB04] Martin Isler, James Frederick Edwards, Thomas R. DeGregori, and William Boyd. Communications. Technology and Culture, 45(1):236–240, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51753.

Igler:2001:FLE

[Igl01] David Igler. The Federal Landscape: An Economic History of the Twentieth-Century West (review). Technology and Culture, 42(3):620–622, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33769.

Inkster:2006:PMN

[Ink06] Ian Inkster. Practical Matter: Newton’s Science in the Service of Industry and Empire 1687–1851 (review). Technology and Culture, 47(2):414–416, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200850.

Inkster:2009:TCP

[Ink09] Ian Inkster. Technology and the Culture of Progress in Meiji Japan (review). Technology and Culture, 50(1):204–205, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258565.

Irish:2000:CPP

[Iri00] Sharon Irish. From Craft to Profession: The Practice of Architecture in Nineteenth-Century America (review). Technology and Culture, 41(3):590–591, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33567.

Irish:2001:FBT

[Iri01] Sharon Irish. The Fireproof Building: Technology and Public Safety in the Nineteenth-Century American City (review). Technology and Culture, 42(2):359–361, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33718.

Irish:2002:NYC

[Iri02] Sharon Irish. New York city skyline, about 1915. Technology and Culture, 43(1):100–101, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33871.

Irish:2003:BPA

[Iri03] Sharon Irish. Big Plans: The Allure and Folly of Urban Design (review). Technology and Culture, 44(1):212–214, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40162.

Irish:2005:EAV

[Iri05] Sharon Irish. Engineering Architecture: The Vision of Fazlur R. Khan (review). Technology and Culture, 46(3):633–634, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186621.

Irish:2006:CAH

[Iri06a] Sharon Irish. Chicago Architecture: Histories, Revisions, Alternatives (review). Technology and Culture, 47(3):663–665, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201460.

Irish:2006:ASC

[Iri06b] Sharon Irish. The American Skyscraper: Cultural Histories (review). Technology and Culture, 47(2):431–433, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200851.

Irish:2006:DCA

[Iri06c] Sharon Irish. The Digital City: The American Metropolis and Information Technology (review). Technology and Culture, 47 (4):850–851, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/204983.

Irish:2008:PCD

[Iri08a] Sharon Irish. The Plan of Chicago: Daniel Burnham and the Remaking of the American City (review). Technology and Culture, 49(1):279–280, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/230912.

Irish:2008:PPP

[Iri08b] Sharon Irish. Preservation, polemics, and power: Carl W. Condit, the Chicago School of Architecture. Technology and Culture, 49(1):202–214, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230911.

Israel:2002:PJH

[Isr02] Paul Israel. The Papers of Joseph Henry, Vol. 7, January 1847– December 1849, and: Vol. 8, January 1850–December 1853 (review). Technology and Culture, 43(4):791–793, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/34022.

Israel:2003:NCB

[Isr03] Paul Israel. Networking: Communicating with Bodies and Machines in the Nineteenth Century (review). Technology and Culture, 44(2):394–396, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42022.

Israel:2005:LMA

[Isr05] Paul Israel. Lightning Man: The Accursed Life of Samuel F.B. Morse (review). Technology and Culture, 46(2):420–422, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184322.

Jacobson:2000:GSE

[Jac00] Lisa Jacobson. Getting and Spending: European and American Consumer Societies in the Twentieth Century (review). Technology and Culture, 41(1):181–183, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33480.

Jacobson:2002:ACC

[Jac02] Lisa Jacobson. An All-Consuming Century: Why Commercialism Won in Modern America (review). Technology and Culture, 43(1):167–168, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33872.

Jackson:2004:REE

[Jac04a] R. E. Jackson. Recognizing emerging environmental problems: The case of chlorinated solvents in groundwater. Technology and Culture, 45(1):55–79, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51754.

Jacobson:2004:STO

[Jac04b] Lisa Jacobson. A Sense of Things: The Object Matter of American Literature (review). Technology and Culture, 45(3):672– 673, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172985.

Jackson:2005:BGC

[Jac05a] Kenneth T. Jackson. Building Gotham: Civic Culture and Public Policy in New York City, 1898–1938 (review). Technology and Culture, 46(1):205–207, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179760.

Jacobs:2005:KJE

[Jac05b] Meg Jacobs. Keeping Up with the Joneses: Envy in American Consumer Society, 1890–1930 (review). Technology and Culture, 46(4):820–822, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192299.

Jacobson:2006:SLM

[Jac06] Lisa Jacobson. Standard of Living: The Measure of the Middle Class in Modern America (review). Technology and Culture, 47 (2):440–441, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200853.

Jacobson:2007:PPE

[Jac07] Lisa Jacobson. Pocketbook Politics: Economic Citizenship in Twentieth-Century America (review). Technology and Culture, 48(2):424–426, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215399.

Jackson:2008:CTT

[Jac08a] Shannon Jackson. The city from thirty thousand feet: Embodiment, creativity, and the use of geographic information systems as urban planning tools. Technology and Culture, 49 (2):325–346, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236962.

Jackson:2008:WGW

[Jac08b] Steven J. Jackson. Writing the global water crisis. Technology and Culture, 49(3):773–778, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243110.

Jaffe:2003:MLM

[Jaf03] James Alan Jaffe. Making a Living in the Middle Ages: The People of Britain, 850-1520 (review). Technology and Culture, 44(4):805–806, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/49701.

Jaffe:2005:CET

[Jaf05] James Alan Jaffe. Customs and Excise: Trade, Production, and Consumption in England, 1640–1845 (review). Technology and Culture, 46(1):247–249, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179762.

Jaffe:2006:WLR

[Jaf06] James Alan Jaffe. Writings of the Luddites (review). Technology and Culture, 47(1):195–196, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196047.

Jaffe:2008:PRE

[Jaf08] James Jaffe. Perceptions of Retailing in Early Modern England (review). Technology and Culture, 49(4):1050–1051, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
255488.

Jager:2000:DIY

[Jag00] Ronald Jager. Do It Yourself: Home Improvement in Twentieth-Century America (review). Technology and Culture, 41(2):398–399, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33528.

Janacek:2009:CWR

[Jan09] Bruce Janacek. Curiosity and Wonder from the Renaissance to the Enlightenment (review). Technology and Culture, 50(4): 978–979, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363457.

Jarrell:2000:CHO

[Jar00] Richard A. Jarrell. A Creation of His Own: Tappan’s Detroit Observatory (review). Technology and Culture, 41(2):376– 378, April 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33529.

Jenks:2000:MMU

[Jen00] Andrew L. Jenks. A metro on the mount: The underground as a Church of Soviet civilization. Technology and Culture, 41(4): 697–724, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33619.

Jenks:2001:RAR

[Jen01] Andrew L. Jenks. Red Atom: Russia’s Nuclear Power Program from Stalin to Today (review). Technology and Culture, 42(4): 833–834, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33820.

Jenks:2005:FMS

[Jen05] Andrew L. Jenks. Flesh to Metal: Soviet Literature and the Alchemy of Revolution (review). Technology and Culture, 46(4): 838–840, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192300.

Jenks:2008:SSS

[Jen08] Andrew Jenks. Stalin and the Soviet Science Wars (review). Technology and Culture, 49(2):488–489, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/236989.

Jeon:2008:BLL

[Jeo08] Chihyung Jeon. Blind Landings: Low-Visibility Operations in American Aviation, 1918–1958 (review). Technology and Culture, 49(3):805–806, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/243124.

Jeon:2009:KWS

[Jeo09] Chihyung Jeon. Keep Watching the Skies: The Story of Operation Moonwatch and the Dawn of the Space Age (review). Technology and Culture, 50(4):946–947, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363438.

Jeremy:2000:ETP

[Jer00] David J. Jeremy. Everyday Things in Premodern Japan: The Hidden Legacy of Material Culture (review). Technology and Culture, 41(1):122–123, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33481.

Jerolimov:2007:FOD

[Jer07] Douglas Jerolimov. Fathoming the Ocean: The Discovery and Exploration of the Deep Sea (review). Technology and Culture, 48(1):184–186, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209841.

Jones-Imhotep:2004:CRC

[JI04] Edward Jones-Imhotep. The Computer Revolution in Canada: Building National Technological Competence (review). Technology and Culture, 45(3):659–661, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172987.

John:2000:GCS

[Joh00a] Richard R. John. Global Communications since 1844: Geopolitics and Technology (review). Technology and Culture, 41 (3):568–570, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33568.

Johnson:2000:MCS

[Joh00b] Ann Johnson. The Making of the Chemist: The Social History of Chemistry in Europe, 1789–1914 (review). Technology and Culture, 41(1):136–138, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33482.

Johnson:2001:SPT

[Joh01] Stephen B. Johnson. Samuel Phillips and the taming of Apollo. Technology and Culture, 42(4):685–709, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33821.

Johnson:2002:ISC

[Joh02a] Jeffrey Allan Johnson. Im Schatten der chemischen Synthese: Industrielle Biotechnologie in Deutschland (1900–1970) (review). Technology and Culture, 43(3):604–605, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33967.

Johnson:2002:WLI

[Joh02b] Loch K. Johnson. The Wizards of Langley: Inside the CIA’s Directorate of Science and Technology (review). Technology and Culture, 43(3):648–649, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33968.

Johnson:2002:FBC

[Joh02c] Stephen B. Johnson. Faster, Better, Cheaper: Low-Cost Innovation in the U.S. Space Program (review). Technology and Culture, 43(3):646–647, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33969.

Johnson:2004:TAE

[Joh04] Stephen B. Johnson. Testing Aircraft, Exploring Space: An Illustrated History of NACA and NASA (review). Technology and Culture, 45(3):632–633, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172986.

Johnston:2005:SPH

[Joh05] Sean Johnston. Shifting perspectives: Holography and the emergence of technical communities. Technology and Culture, 46(1):77–103, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/179763.

Johnson:2006:MMR

[Joh06] Jeffrey Allan Johnson. Master Mind: The Rise and Fall of Fritz Haber, the Nobel Laureate Who Launched the Age of Chemical Warfare (review). Technology and Culture, 47(4):835–837, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204984.

Johns:2008:CHB

[Joh08] Adrian Johns. A Companion to the History of the Book (review). Technology and Culture, 49(3):782–784, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243084.

Jones:2000:IAP

[Jon00] Elwood H. Jones. The Inglorious Arts of Peace: Exhibitions in Canadian Society during the Nineteenth Century (review). Technology and Culture, 41(3):629–630, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33569.

Jones:2001:PST

[Jon01] Russell Douglass Jones. Prague Summer: Twenty-Seventh Symposium of the International Committee for the History of Technology. Technology and Culture, 42(2):329–332, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33719.

Jones:2002:WPC

[Jon02a] Russell Douglass Jones. Walter P. Chrysler: The Life and Times of an Automotive Genius (review). Technology and Culture, 43(1):181–183, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33873.

Jones:2002:LVH

[Jon02b] Toby Jones. Late Victorian Holocausts: El Nino Famines and the Making of the Third World (review). Technology and Culture, 43(2):463–464, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33916.

Jones:2003:USL

[Jon03] Russell Douglass Jones. Uncle Sam’s Locomotives: The USRA and the Nation’s Railroads (review). Technology and Culture, 44(3):615–616, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46096.

Jones:2004:WWH

[Jon04a] Russell Douglass Jones. Wheels for the World: Henry Ford, His Company, and a Century of Progress (review). Technology and Culture, 45(4):856–857, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175755.

Jones:2004:HSA

[Jon04b] Russell Douglass Jones. Heroes of the sky: Adventures in early flight: At the Henry Ford Museum, Dearborn, Michigan. Technology and Culture, 45(2):368–372, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/55669.

Jones:2004:PDC

[Jon04c] Toby Jones. Parting the Desert: The Creation of the Suez Canal (review). Technology and Culture, 45(4):871–872, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175756.

Jones:2006:IRS

[Jon06] Robert A. Jones. Industrial Ruins: Spaces, Aesthetics, and Materiality (review). Technology and Culture, 47(4):821–822, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204985.

Jones:2009:GEM

[Jon09] Christopher Jones. Global Electrification: Multinational Enterprise and International Finance in the History of Light and Power, 1878–2007 (review). Technology and Culture, 50(3): 698–699, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270568.

Jorgensen:2008:CSM

[Jør08] Dolly Jørgensen. Cooperative sanitation: Managing streets and gutters in Late Medieval England and Scandinavia. Technology and Culture, 49(3):547–567, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243112.

Jorgensen:2009:BIE

[Jør09] Finn Arne Jørgensen. Bilsamhället: Ideologi, expertis och regelskapande i efterkrigstidens Sverige (review). Technology and Culture, 50(3):714–715, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270552.

Josephson:2001:CMC

[Jos01] Paul R. Josephson. Critical Masses: Citizens, Nuclear Weapons Production, and Environmental Destruction in the United States and Russia (review). Technology and Culture, 42 (2):380–382, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33720.

Josephson:2003:ABC

[Jos03] Paul R. Josephson. Atomic Bomb Cinema: The Apocalyptic Imagination on Film (review). Technology and Culture, 44(1): 170–171, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40163.

Josephson:2004:FFS

[Jos04] Paul R. Josephson. The Final Frontier: Science, America, and Terror (review). Technology and Culture, 45(1):201–202, January 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/51755.

Josephson:2005:FFR

[Jos05] Paul R. Josephson. Farm to Factory: A Reinterpretation of the Soviet Industrial Revolution (review). Technology and Culture, 46(4):837–838, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192301.

Josephson:2007:DWE

[Jos07] Paul R. Josephson. Deep Water: The Epic Struggle Over Dams, Displaced People, and the Environment (review). Technology and Culture, 48(1):210–211, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209842.

Josephson:2008:OHD

[Jos08] Paul R. Josephson. The ocean’s hot dog: The development of the fish stick. Technology and Culture, 49(1):41–61, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230913.

Judd:2003:INB

[Jud03] Richard William Judd. Industrialized Nature: Brute Force Technology and the Transformation of the Natural World (review). Technology and Culture, 44(3):591–593, July 2003. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46097.

Judd:2005:NCM

[Jud05] Richard William Judd. Nature and the City: Making Environmental Policy in Toronto and Los Angeles (review). Technology and Culture, 46(3):659–660, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186622.

Judd:2008:ECE

[Jud08] Richard W. Judd. Emerald City: An Environmental History of Seattle (review). Technology and Culture, 49(4):1062–1063, October 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/255489.

Kahlow:2001:EDM

[Kah01] Andreas Kahlow. Eloge du mixte: Materiaux nouveaux et philosophie ancienne (review). Technology and Culture, 42 (3):570–571, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33771.

Kaijser:2002:SBB

[Kai02] Arne Kaijser. System building from below: Institutional change in Dutch water control systems. Technology and Culture, 43 (3):521–548, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33970.

Kanigel:2007:TCP

[Kan07] Robert Kanigel. Taylorism up close and personal: Hugh G. Aitken’s Taylorism at Watertown Arsenal. Technology and Culture, 48(1):158–164, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209843.

Karrow:2002:PGA

[Kar02] Robert W. Karrow. Ptolemy’s “Geography”: An Annotated Translation of the Theoretical Chapters (review). Technology and Culture, 43(3):589–591, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33971.

Karvonen:2008:GSR

[Kar08] Andy Karvonen. The Gentle Subversive: Rachel Carson, Silent Spring, and the Rise of the Environmental Movement (review). Technology and Culture, 49(1):242–244, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230914.

Karrow:2009:NMC

[Kar09a] Robert W. Karrow, Jr. The Natures of Maps: Cartographic Constructions of the Natural World (review). Technology and Culture, 50(4):919–920, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/363422.

Karvonen:2009:RCL

[Kar09b] Andrew Karvonen. Rachel Carson: Legacy and Challenge (review). Technology and Culture, 50(4):924–926, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363425.

Katz:2000:DFR

[Kat00] Barry Katz. Design and Feminism: Re-Visioning Spaces, Places, and Everyday Things (review). Technology and Culture, 41(4):790–791, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33620.

Katz:2002:RBH

[Kat02] Barry Katz. Reyner Banham: Historian of the Immediate Future (review). Technology and Culture, 43(4):798–800, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34023.

Katz:2003:SMT

[Kat03] Barry Katz. The Senses of Modernism: Technology, Perception, and Aesthetics (review). Technology and Culture, 44(4): 855–856, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/49702.

Katz:2005:DMC

[Kat05] Barry Katz. Dancing Machines: Choreographies of the Age of Mechanical Reproduction (review). Technology and Culture, 46 (2):429–430, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184323.

Katz:2006:ID

[Kāt06] Barry Kātz. Intelligent design. Technology and Culture, 47 (2):381–390, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200854.

Katz:2007:DUR

[Kāt07] Barry Kātz. Design in the U.S.A. (review). Technology and Culture, 48(4):862–864, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223366.

Katz:2009:GDN

[Kat09] Barry M. Katz. Graphic Design: A New History (review). Technology and Culture, 50(1):258–260, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258546.

Keelan:2003:NDH

[Kee03] Jennifer Keelan. The New Disability History: American Perspectives (review). Technology and Culture, 44(4):853–855, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49703.

Keller:2004:AID

[Kel04] Janet Dixon Keller. Artefactual Intelligence: The Development and Use of Cognitively Congenial Artefacts (review). Technology and Culture, 45(3):668–669, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172988.

Kelleher:2005:VNR

[Kel05] Tom Kelleher. Voices of the New Republic: Connecticut Towns, 1800–1832 (review). Technology and Culture, 46(1):200–202, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179764.

Keller:2006:HMR

[Kel06a] Alex Keller. Humanism, Machinery, and Renaissance Literature (review). Technology and Culture, 47(2):410–412, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200855.

Kelman:2006:NBL

[Kel06b] Ari Kelman. Nature bats last: Some recent works on technology and urban disaster. Technology and Culture, 47(2):391– 402, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200856.

Keller:2009:MLV

[Kel09] Alexander G. Keller. The Machines of Leonardo da Vinci and Franz Reuleaux: Kinematics of Machines from the Renaissance to the 20th Century (review). Technology and Culture, 50(2): 454–456, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/263359.

Kerwin:2006:BLL

[Ker06] William Kerwin. The Body in the Library: A Literary Anthology of Modern Medicine (review). Technology and Culture, 47 (3):690–692, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201461.

Konig:2000:C

[KFB00] Wolfgang Konig, Larrie D. Ferreiro, and John K. Brown. Communications. Technology and Culture, 41(4):857–861, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33621.

Khan:2001:RLN

[Kha01] Zorina Khan. The Retrieval of a Legacy: Nineteenth-Century American Women Inventors (review). Technology and Culture, 42(3):577–578, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33772.

Khan:2008:FAB

[Kha08] Zorina Khan. A Farewell to Alms: A Brief Economic History of the World (review). Technology and Culture, 49(3):779– 781, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243095.

Kidwell:2001:IAL

[Kid01] Peggy Aldrich Kidwell. Information Ages: Literacy, Numeracy, and the Computer Revolution (review). Technology and Culture, 42(1):184–185, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33674.

Kidwell:2003:OCL

[Kid03] Peggy Aldrich Kidwell. Origins of Cyberspace: A Library on the History of Computing, Networking, and Telecommunications (review). Technology and Culture, 44(1):193–194, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40164.

Kidwell:2004:HMT

[Kid04] Peggy Aldrich Kidwell. A History of Mathematical Tables: From Sumer to Spreadsheets (review). Technology and Culture, 45(3):662–664, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172989.

Kidwell:2006:ATA

[Kid06] Peggy Aldrich Kidwell. Alan Turing’s Automatic Computing Engine: The Master Codebreaker’s Struggle to Build the Modern Computer (review). Technology and Culture, 47(2):460– 462, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200857.

Kidwell:2007:CSB

[Kid07] Peggy Aldrich Kidwell. Colossus: The Secrets of Bletchley Park’s Codebreaking Computers (review). Technology and Culture, 48(3):663–664, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/218709.

Kidwell:2008:CNW

[Kid08] Peggy Aldrich Kidwell. Calculating a Natural World: Scientists, Engineers, and Computers during the Rise of U.S. Cold War Research (review). Technology and Culture, 49(1): 249–250, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/230915.

Kielbowicz:2001:RLE

[Kie01] Richard Burket Kielbowicz. Relays: Literature as an Epoch of the Postal System (review). Technology and Culture, 42(3):563– 564, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33773.

Kielbowicz:2002:FNH

[Kie02] Richard Burket Kielbowicz. The Form of News: A History (review). Technology and Culture, 43(2):438–440, April 2002.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33917.

Kiesling:2003:ERF

[Kie03] Lynne L. Kiesling. Energy and the Rise and Fall of Political Economy (review). Technology and Culture, 44(1):218–220, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40165.

Kielbowicz:2009:HBA

[Kie09] Richard B. Kielbowicz. A History of the Book in America, vol. 3: The Industrial Book, 1840–1880 (review). Technology and Culture, 50(2):499–500, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263357.

Kim:2004:OCA

[Kim04] Julie Ju-Youn Kim. The Organizational Complex: Architecture, Media, and Corporate Space (review). Technology and Culture, 45(2):468–470, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55670.

Kinney:2000:CWG

[Kin00] Jeremy R. Kinney. Curtiss–Wright: Greatness and Decline (review). Technology and Culture, 41(1):147–149, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33483.

Kinchy:2009:AAA

[Kin09a] Abby J. Kinchy. African Americans in the Atomic Age: Postwar perspectives on race and the bomb, 1945–1967. Technology and Culture, 50(2):291–315, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263328.

Kinkela:2009:APL

[Kin09b] David Kinkela. American Pests: The Losing War on Insects from Colonial Times to DDT (review). Technology and Culture, 50(4):922–924, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363424.

Kirby:2000:FNA

[Kir00] Jack Temple Kirby. The Fruits of Natural Advantage: Making the Industrial Countryside in California (review). Technology and Culture, 41(3):613–614, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33570.

Kirsch:2008:NPU

[Kir08] Stuart Kirsch. The Navajo People and Uranium Mining (review). Technology and Culture, 49(1):247–249, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230916.

Kleiman:2003:EAL

[Kle03a] Jordan Benson Kleiman. Edward Abbey: A Life (review). Technology and Culture, 44(1):167–168, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40166.

Klein:2003:IIN

[Kle03b] Maury Klein. Internal Improvement: National Public Works and the Promise of Popular Government in the Early United States (review). Technology and Culture, 44(1):175–176, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40167.

Kleinman:2004:EEE

[Kle04] Daniel Lee Kleinman. Everyday Engineering: An Ethnography of Design and Innovation (review). Technology and Culture, 45(4):849–850, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175757.

Kleiman:2006:RUR

[Kle06a] Jordan Benson Kleiman. Resources under Regimes: Technology, Environment, and the State (review). Technology and Culture, 47(4):819–821, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/204986.

Klein:2006:RSW

[Kle06b] Maury Klein. The Railroad and the State: War, Politics, and Technology in Nineteenth-Century America (review). Technology and Culture, 47(1):220–221, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196048.

Kleinman:2006:SIN

[Kle06c] Daniel Lee Kleinman. The Science-Industry Nexus: History, Policy, Implications (review). Technology and Culture, 47(2): 467–468, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200858.

Klein:2008:EAD

[Kle08] Maury Klein. Evolution of the American Diesel Locomotive (review). Technology and Culture, 49(3):803–805, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243097.

Kleiman:2009:LFP

[Kle09] Jordan Kleiman. Local food and the problem of public authority. Technology and Culture, 50(2):399–417, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263337.

Kline:2000:ELI

[Kli00] Ronald R. Kline. Edison: A Life of Invention (review). Technology and Culture, 41(3):605–606, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33571.

Klingle:2004:NSE

[Kli04] Matthew W. Klingle. Natural States: The Environmental Imagination in Maine, Oregon, and the Nation (review). Technology and Culture, 45(4):862–864, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175758.

Kline:2006:CMS

[Kli06] Ronald R. Kline. Cybernetics, management science, and technology policy: The emergence of “Information technology” as a keyword, 1948–1985. Technology and Culture, 47(3):513– 535, July 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/201462.

Kline:2008:PES

[Kli08] Ronald R. Kline. From progressivism to engineering studies: Edwin T. Layton’s The Revolt of the Engineers. Technology and Culture, 49(4):1018–1024, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255448.

Kargon:2004:CCN

[KM04] Robert Kargon and Arthur P. Molella. The city as communications net: Norbert Wiener, the atomic bomb, and urban dispersal. Technology and Culture, 45(4):764–777, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175766.

Knowles:2001:CET

[Kno01a] Anne Kelly Knowles. Confederate Engineer: Training and Campaigning with John Morris Wampler (review). Technology and Culture, 42(3):590–592, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33774.

Knowles:2001:LRT

[Kno01b] Anne Kelly Knowles. Labor, race, and technology in the Confederate iron industry. Technology and Culture, 42(1):1–26, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33675.

Knowles:2003:WWA

[Kno03] Anne Kelly Knowles. Wealth, Waste, and Alienation: Growth and Decline in the Connellsville Coke Industry (review). Technology and Culture, 44(1):220–221, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40168.

Knowles:2005:KCA

[Kno05a] Anne Kelly Knowles. In the Kingdom of Coal: An American Family and the Rock that Changed the World (review). Technology and Culture, 46(4):847–849, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192302.

Knowles:2005:CSR

[Kno05b] Scott Gabriel Knowles. City in the Sky (review). Technology and Culture, 46(3):670–672, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186623.

Knowles:2006:CUV

[Kno06a] Anne Kelly Knowles. Capital’s Utopia: Vandergrift, Pennsylvania, 1855–1916 (review). Technology and Culture, 47(1): 211–213, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196049.

Knowles:2006:IRI

[Kno06b] Anne Kelly Knowles. The Industrial Revolution in Iron: The Impact of British Coal Technology in Nineteenth-Century Europe (review). Technology and Culture, 47(3):668–669, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201463.

Knowles:2007:NMH

[Kno07a] Anne Kelly Knowles. Networked Machinists: High-Technology Industries in Antebellum America (review). Technology and Culture, 48(4):859–861, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223367.

Knowles:2007:AFT

[Kno07b] Scott Gabriel Knowles. Asbestos and Fire: Technological Trade-Offs and the Body at Risk (review). Technology and Culture, 48(1):175–177, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209844.

Koloski-Ostrow:2002:FLE

[KO02] Ann Olga Koloski-Ostrow. Frontinus’ Legacy: Essays on Frontinus’ “de aquis urbis Romae” (review). Technology and Culture, 43(3):585–587, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33972.

Koloski-Ostrow:2007:HFP

[KO07] Ann Olga Koloski-Ostrow. A Handbook of Food Processing in Classical Rome: For Her Bounty No Winter (review). Technology and Culture, 48(4):844–846, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223373.

Koloski-Ostrow:2008:NPA

[KO08] Ann Olga Koloski-Ostrow. New perspectives on ancient technology and engineering in Greece and Rome. Technology and Culture, 49(4):1025–1030, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255449.

Koloski-Ostrow:2009:OOH

[KO09] Ann Olga Koloski-Ostrow. An overture to the Oxford handbook: Serafina cuomo, technology and culture in Greek and Roman antiquity. Technology and Culture, 50(2):450–452, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263332.

Koizumi:2002:SWC

[Koi02] Kenkichiro Koizumi. In search of wakon: The cultural dynamics of manufacturing technology in postwar Japan. Technology and Culture, 43(1):29–49, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33874.

Kolb:2006:CCP

[Kol06] Carolyn Kolb. Crescent city, post-apocalypse. Technology and Culture, 47(1):108–111, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196050.

Kolb:2007:BFH

[Kol07] Carolyn Kolb. Breach of Faith: Hurricane Katrina and the Near Death of a Great American City (review). Technology and Culture, 48(4):870–871, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223368.

Kolb:2008:EME

[Kol08] Carolyn G. Kolb. Energy Metropolis: An Environmental History of Houston and the Gulf Coast (review). Technology and Culture, 49(4):1060–1062, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255490.

Konvitz:2005:MQD

[Kon05] Josef W. Konvitz. The Mapmaker’s Quest: Depicting New Worlds in Renaissance Europe (review). Technology and Culture, 46(1):246–247, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/179765.

Konvitz:2006:UCO

[Kon06] Josef W. Konvitz. Unbuilding Cities: Obduracy in Urban Sociotechnical Change (review). Technology and Culture, 47(4): 853–855, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204987.

Konvitz:2009:MFO

[Kon09] Josef Konvitz. Maps: Finding Our Place in the World (review). Technology and Culture, 50(1):254–256, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258552.

Koukal:2001:TGL

[Kou01a] D. R. Koukal. Technology and the Good Life? and: Holding On to Reality: The Nature of Information at the Turn of the Millennium (review). Technology and Culture, 42(4):776–778, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33822.

Koukal:2001:RGT

[Kou01b] D. R. Koukal. The Robot in the Garden: Telerobotics and Telepistemology in the Age of the Internet (review). Technology and Culture, 42(3):618–620, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33775.

Koukal:2002:BTR

[Kou02] D. R. Koukal. Bodies in Technology (review). Technology and Culture, 43(4):837–838, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34024.

Koukal:2003:APT

[Kou03] D. R. Koukal. American Philosophy of Technology: The Empirical Turn (review). Technology and Culture, 44(1):225–226, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40169.

Koukal:2006:LLT

[Kou06] D. R. Koukal. Living in the Labyrinth of Technology (review). Technology and Culture, 47(4):814–815, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204988.

Koukal:2007:EST

[Kou07] D. R. Koukal. Encyclopedia of Science, Technology, and Ethics (review). Technology and Culture, 48(3):614–617, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218710.

Kraemer:2000:NSI

[Kra00] Sylvia K. Kraemer. NASA and the Space Industry (review). Technology and Culture, 41(4):849–851, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33622.

Kraemer:2001:HES

[Kra01] Sylvia K. Kraemer. A History of the European Space Agency, 1958–1987 (review). Technology and Culture, 42(4):820–822, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33823.

Kranakis:2004:FBO

[Kra04] Eda Kranakis. Fixing the blame: Organizational culture and the Quebec bridge collapse. Technology and Culture, 45(3):487– 518, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/172990.

Kraft:2005:CSH

[Kra05a] James P. Kraft. Capturing Sound: How Technology Has Changed Music (review). Technology and Culture, 46(2):429– 430, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184324.

Kranakis:2005:STH

[Kra05b] Eda Kranakis. Surveying technology and history: Essential tensions and postmodern possibilities. Technology and Culture, 46(4):805–812, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192303.

Kraft:2007:MEL

[Kra07a] James P. Kraft. Making Easy Listening: Material Culture and Postwar American Recording (review). Technology and Culture, 48(2):438–440, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215400.

Kranakis:2007:PPE

[Kra07b] Eda Kranakis. Patents and power: European patent-system integration in the context of globalization. Technology and Culture, 48(4):689–728, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/223369.

Kranakis:2009:HPL

[Kra09] Eda Kranakis. Heat, Power, and Light: Revolutions in Energy Services (review). Technology and Culture, 50(3):699–701, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270567.

Krige:2000:AAH

[Kri00a] John Krige. Air Apparent: How Meteorologists Learned to Map, Predict, and Dramatize Weather (review). Technology and Culture, 41(4):830–831, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33624.

Krige:2000:CIR

[Kri00b] John Krige. Crossing the interface from R&D to operational use: The case of the European meteorological satellite. Technology and Culture, 41(1):27–50, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33484.

Krige:2003:CAE

[Kri03] John Krige. Changing the Atmosphere: Expert Knowledge and Environmental Governance (review). Technology and Culture, 44(2):428–430, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42023.

Krupar:2002:TFD

[Kru02] Jason N. Krupar. Technical Fouls: Democratic Dilemmas and Technological Change (review). Technology and Culture, 43(1): 214–215, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33875.

Krupar:2004:SCA

[Kru04] Jason N. Krupar. Survival City: Adventures Among the Ruins of Atomic America (review). Technology and Culture, 45 (3):673–675, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172991.

Krupar:2008:WHK

[Kru08] Jason N. Krupar. The Worlds of Herman Kahn: The Intuitive Science of Thermonuclear War (review). Technology and Culture, 49(1):244–245, January 2008. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230917.

Krupar:2009:HWB

[Kru09] Jason Krupar. Hiroshima: The World’s Bomb (review). Technology and Culture, 50(3):720–721, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270569.

Krysko:2004:FFS

[Kry04] Michael A. Krysko. Forbidden frequencies: Sino–American relations and Chinese broadcasting during the Interwar Era. Technology and Culture, 45(4):712–739, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175759.

Kulik:2003:RLD

[Kul03] Gary Kulik. Raymond Loewy: Designs for a consumer culture at the Hagley Museum and Library. Technology and Culture, 44(3):566–573, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46098.

Kuwahara:2000:MIS

[Kuw00] Tetsuya Kuwahara. Manufacturing Ideology: Scientific Management in Twentieth-Century Japan (review). Technology and Culture, 41(4):807–808, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33625.

Lackner:2003:GDH

[Lac03] Helmut Lackner. Die Geschichte der Dampfmaschine: Historische Entwicklung, Industriegeschichte, Technische Denkmale (review). Technology and Culture, 44(2):386–387, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42024.

Laird:2003:CFA

[Lai03a] Frank N. Laird. Constructing the future: Advocating energy technologies in the Cold War. Technology and Culture, 44(1): 27–49, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40170.

Laird:2003:BYP

[Lai03b] Pamela Walker Laird. Brought to You By: Postwar Television Advertising and the American Dream (review). Technology and Culture, 44(2):417–419, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42025.

Laird:2007:MBT

[Lai07] Pamela Walker Laird. Made to Break: Technology and Obsolescence in America (review). Technology and Culture, 48 (2):428–429, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215401.

Lakwete:2001:SGE

[Lak01] Angela Lakwete. The Second Great Emancipation: The Mechanical Cotton Picker, Black Migration, and How They Shaped the Modern South (review). Technology and Culture, 42 (2):350–352, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33721.

Lakwete:2004:SNW

[Lak04] Angela Lakwete. The South, the Nation, and the World: Perspectives on Southern Economic Development (review). Technology and Culture, 45(2):434–436, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/55671.

Lakwete:2005:MML

[Lak05] Angela Lakwete. Minding the Machine: Languages of Class in Early Industrial America (review). Technology and Culture, 46 (2):441–442, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184325.

Lankton:2000:IMS

[Lan00] Larry D. Lankton. Iron-Making Societies: Early Industrial Development in Sweden and Russia, 1600–1900 (review). Technology and Culture, 41(3):582–583, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33572.

Landis:2001:RH

[Lan01a] Leo Landis. Reaping the harvest. Technology and Culture, 42(4):750–753, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33824.

Langins:2001:PAP

[Lan01b] Janis Langins. Production d’armes et puissance des nations (review). Technology and Culture, 42(2):388–390, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33722.

Landels:2002:SIG

[Lan02a] John G. Landels. Surveying Instruments of Greece and Rome (review). Technology and Culture, 43(2):406–407, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33918.

Landis:2002:DSL

[Lan02b] Leo Landis. Delta Sugar: Louisiana’s Vanishing Plantation Landscape (review). Technology and Culture, 43(3):619–620, July 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33973.

Landis:2003:CWV

[Lan03a] Leo Landis. Changing Works: Visions of a Lost Agriculture (review). Technology and Culture, 44(4):838–839, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49704.

Lang:2003:FNS

[Lan03b] James Lang. Food Nations: Selling Taste in Consumer Societies (review). Technology and Culture, 44(2):419–421, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
42026.

Langley:2003:MUS

[Lan03c] Harold D. Langley. Medicine under Sail (review). Technology and Culture, 44(3):609–610, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46099.

Lang:2004:ETT

[Lan04a] James Lang. Exploring the Tomato: Transformations of Nature, Society, and Economy (review). Technology and Culture, 45(1):222–224, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51756.

Langins:2004:VGD

[Lan04b] Janis Langins. Vauban: De la gloire du roi au service de l’etat (review). Technology and Culture, 45(4):874–876, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175760.

Lang:2007:BCA

[Lan07a] James Lang. Banana Cultures: Agriculture, Consumption, and Environmental Change in Honduras and the United States (review). Technology and Culture, 48(2):429–431, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215402.

Lang:2007:SRP

[Lan07b] James Lang. The Sausage Rebellion: Public Health, Private Enterprise, and Meat in Mexico City, 1890–1917 (review). Technology and Culture, 48(3):624–625, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218711.

Landecker:2008:DMR

[Lan08] Hannah Landecker. Dolly Mixtures: The Remaking of Genealogy (review). Technology and Culture, 49(3):826–827, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243101.

Lankton:2009:MPD

[Lan09] Larry Lankton. Morgan Park: Duluth, U.S. Steel, and the Forging of a Company Town (review). Technology and Culture, 50(2):484–486, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/263354.

Lapini:2006:PDE

[Lap06] Gianluca Lapini. La Passione dell’invenzione: Enrico Forlanini, ingegnere e aeronauta (review). Technology and Culture, 47(2):458–460, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200859.

Laprise:2009:SSI

[Lap09] John Laprise. Seduced by Secrets: Inside the Stasi’s Spy-Tech World (review). Technology and Culture, 50(3):723–725, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270571.

Larson:2003:LNC

[Lar03] John Lauritz Larson. Listening to Nineteenth-Century America (review). Technology and Culture, 44(1):173–175, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40171.

Larson:2004:DIS

[Lar04] John Lauritz Larson. Demythologizing innovation: Steven usselman, regulating railroad innovation. Technology and Culture, 45(1):159–161, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51757.

Larson:2006:WWD

[Lar06] John Lauritz Larson. What are we doing wrong? Technology and Culture, 47(4):803–807, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/204989.

Lara:2008:MMA

[Lar08] Fernando Luiz Lara. Mexican Modernity: The Avant-Garde and the Technological Revolution (review). Technology and Culture, 49(2):500–501, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236990.

Lassman:2003:IRT

[Las03] Thomas C. Lassman. Industrial research transformed: Edward Condon at the Westinghouse Electric and Manufacturing Company, 1935–1942. Technology and Culture, 44(2):306–339, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
42027.

Lassman:2006:ERC

[Las06] Thomas Charles Lassman. Establishing Research Corporation: A Case Study of Patents, Philanthropy, and Organized Research in Early Twentieth-Century America (review). Technology and Culture, 47(1):227–229, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196051.

Laux:2002:OAD

[Lau02] James Michael Laux. Organise a l’aube du taylorisme: La pratique d’Ernest Mattern chez Peugeot, 1906–1919 (review). Technology and Culture, 43(4):812–814, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34025.

Lauer:2008:RWR

[Lau08] Josh Lauer. From rumor to written record: Credit reporting and the invention of financial identity in Nineteenth-Century America. Technology and Culture, 49(2):301–324, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236961.

Lawrence:2000:ASI

[Law00] Ghislaine Lawrence. American Surgical Instruments: The History of Their Manufacture and a Directory of Instrument Makers to 1900 (review). Technology and Culture, 41(4):814–815, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33627.

Lawrence:2003:MOH

[Law03] Christopher Lawrence. Machines in Our Hearts: The Cardiac Pacemaker, the Implantable Defibrillator, and American Health Care (review). Technology and Culture, 44(1):163–165, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40172.

Liel:2008:EIB

[LB08] Abbie B. Liel and David P. Billington. Engineering innovation at Bonneville Dam. Technology and Culture, 49(3):727– 751, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243086.

LeCoze:2005:FAL

[Le 05] Jean Le Coze. De la forge au laboratoire: Naissance de la metallurgie physique, 1860–1914 (review). Technology and Culture, 46(4):840–841, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192287.

LeRoux:2007:HCO

[Le 07] Muriel Le Roux. Henri Le Châtelier (1850–1936): ou la science appliquée à l’industrie (review). Technology and Culture, 48(1): 222–224, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/211297.

Leonard:2001:MSI

[Leo01] Jane Kate Leonard. Maritime Sector, Institutions, and Sea Power of Premodern China (review). Technology and Culture, 42(3):609–610, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33776.

Leslie:2001:RDR

[Les01] Stuart W. Leslie. Regional disadvantage: Replicating Silicon Valley in New York’s capital region. Technology and Culture, 42 (2):236–264, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33723.

Leslie:2004:ASC

[Les04a] Stuart W. Leslie. After the Smoke Clears: Struggling to Get By in Rustbelt America (review). Technology and Culture, 45(4): 864–866, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175761.

Lessoff:2004:MCC

[Les04b] Alan Lessoff. The Middle-Class City: Transforming Space and Time in Philadelphia, 1876–1926 (review). Technology and Culture, 45(4):842–844, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175762.

Leslie:2007:LB

[Les07] Stuart W. Leslie. Lynwood Bryant, 1908–2005. Technology and Culture, 48(1):236–244, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209845.

Lesko:2008:DDM

[Les08a] Jason Lesko. Devices and Designs: Medical Technologies in Historical Perspective (review). Technology and Culture, 49 (2):516–517, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236991.

Leslie:2008:LSC

[Les08b] Thomas Leslie. “As large as the situation of the columns would allow”: Building cladding and plate glass in the Chicago skyscraper, 1885–1905. Technology and Culture, 49(2):399– 419, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236965.

Levinson:2000:ESB

[Lev00] Mark Levinson. From Engineering Science to Big Science: The NACA and NASA Collier Trophy Research Project Winners (review). Technology and Culture, 41(2):378–380, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33530.

Levin:2001:DMA

[Lev01] Miriam R. Levin. La democratie et les médias au 20e siècle (review). Technology and Culture, 42(3):592–593, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33777.

Levin:2003:LDP

[Lev03a] Miriam R. Levin. L’age du plip: Chroniques de l’innovation technique (review). Technology and Culture, 44(2):409–410, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42028.

Levinson:2003:JCH

[Lev03b] Mark Levinson. Jerome C. Hunsaker and the Rise of American Aeronautics (review). Technology and Culture, 44(2):441– 442, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42029.

Levinson:2005:BWA

[Lev05] Mark Levinson. The Bird is on the Wing: Aerodynamics and the Progress of the American Airplane (review). Technology and Culture, 46(1):231–232, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179766.

Levine:2008:FSA

[Lev08] Deborah Levine. The First Scientific American: Benjamin Franklin and the Pursuit of Genius (review). Technology and Culture, 49(3):794–796, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/243107.

Lewis:2000:IAT

[Lew00] Jeffrey Lewis. The Intellectual Appropriation of Technology: Discourses on Modernity, 1900–1939 (review). Technology and Culture, 41(3):595–597, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33573.

Lewis:2001:RWN

[Lew01] Robert Lewis. Redesigning the workplace: The North American factory in the interwar period. Technology and Culture, 42(4):665–684, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33825.

Licht:2001:RFM

[Lic01] Walter Licht. The Rise and Fall of Mass Production (review). Technology and Culture, 42(2):347–349, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33724.

Lichtenstein:2002:CPA

[Lic02] Nelson Lichtenstein. Corporate Power, American Democracy, and the Automobile Industry (review). Technology and Culture, 43(1):179–181, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33876.

Lienhard:2001:HSM

[Lie01] John H. Lienhard. Hysplex: The Starting Mechanism in Ancient Stadia (review). Technology and Culture, 42(3):567– 568, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33778.

Light:2002:SUN

[Lig02] Jennifer S. Light. Splintering Urbanism: Networked Infrastructures, Technological Mobilities, and the Urban Condition (review). Technology and Culture, 43(3):614–615, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33974.

Light:2006:CCS

[Lig06] Jennifer S. Light. Computers and Commerce: A Study of Technology and Management at Eckert-Mauchly Computer Company, Engineering Research Associates, and Remington Rand, 1946–1957 (review). Technology and Culture, 47(3):682–683, July 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/201464.

Light:2008:TGS

[Lig08] Jennifer Light. Taking games seriously. Technology and Culture, 49(2):347–375, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236963.

Light:2009:IET

[Lig09] Jennifer Light. Invented Edens: Techno-Cities of the Twentieth Century (review). Technology and Culture, 50(2):481– 483, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263355.

Lindstrom:2000:TAB

[Lin00] Richard Lindstrom. “They all believe they are undiscovered Mary Pickfords”: Workers, photography, and scientific management. Technology and Culture, 41(4):725–751, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33629.

Lindeberg:2002:WWS

[Lin02a] J. D. Lindeberg. Waste and Want: A Social History of Trash (review). Technology and Culture, 43(3):615–617, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33975.

Lintsen:2002:TCC

[Lin02b] Harry Lintsen. Two centuries of central water management in the Netherlands. Technology and Culture, 43(3):549–568, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33976.

Lincoln:2005:BAA

[Lin05] Evelyn Lincoln. The Body of the Artisan: Art and Experience in the Scientific Revolution (review). Technology and Culture, 46(3):644–646, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186624.

Lincoln:2006:WVP

[Lin06a] Evelyn Lincoln. The World in Venice: Print, the City, and Early Modern Identity (review). Technology and Culture, 47(4): 842–844, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204990.

Linde:2006:SRC

[Lin06b] C. Linde. Shouldering Risks: The Culture of Control in the Nuclear Power Industry (review). Technology and Culture, 47 (1):226–227, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196052.

Lipartito:2000:ENS

[Lip00a] Kenneth Lipartito. Endless Novelty: Specialty Production and American Industrialization, 1865–1925 (review). Technology and Culture, 41(1):141–143, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33485.

Lipartito:2000:HRG

[Lip00b] Kenneth Lipartito. The historian in the Rose garden? Technology and Culture, 41(3):537–548, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33574.

Lipartito:2002:RUR

[Lip02] Kenneth Lipartito. Risk, Uncertainty, and Rational Action (review). Technology and Culture, 43(2):453–455, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33919.

Lipartito:2003:PIA

[Lip03] Kenneth Lipartito. Picturephone and the Information Age: The social meaning of failure. Technology and Culture, 44(1): 50–81, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40173.

Lipartito:2006:SIA

[Lip06] Kenneth Lipartito. Structuring the Information Age: Life Insurance and Technology in the Twentieth Century (review). Technology and Culture, 47(4):872–873, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204991.

Littmann:2003:AUS

[Lit03] William Littmann. Architecture in the United States, 1800– 1850 (review). Technology and Culture, 44(4):831–833, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
49705.

Levinson:2004:C

[LL04] Mark Levinson and Sarah Lowengard. Communications. Technology and Culture, 45(3):677–678, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/172992.

Logsdon:2004:SSO

[Log04] John M. Logsdon. Single Stage to Orbit: Politics, Space Technology, and the Quest for Reusable Rocketry (review). Technology and Culture, 45(3):633–635, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172993.

Lonnquest:2000:UIW

[Lon00] John C. Lonnquest. Under Ice: Waldo Lyon and the Development of the Arctic Submarine (review). Technology and Culture, 41(2):381–382, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33531.

Long:2001:IIR

[Lon01a] Pamela O. Long. The Invention of the Italian Renaissance Printmaker (review). Technology and Culture, 42(3):559–561, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33779.

Long:2001:CNC

[Lon01b] Priscilla Long. Caverns of Night: Coal Mines in Art, Literature, and Film (review). Technology and Culture, 42(4): 789–790, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33826.

Long:2002:SID

[Lon02a] Pamela O. Long. La seta in Italia dal Medioevo al Seicento: Dal baco al drappo (review). Technology and Culture, 43(4): 784–785, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34026.

Long:2002:ODC

[Lon02b] Pamela O. Long. Origins of Democratic Culture: Printing, Petitions and the Public Sphere in Early Modern England (review). Technology and Culture, 43(2):418–419, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33920.

Long:2003:MSP

[Lon03a] Pamela O. Long. Of mining, smelting, and printing: Agricola’s De re metallica. Technology and Culture, 44(1):97–101, January 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/40174.

Long:2003:WGA

[Lon03b] Priscilla Long. Working the Garden: American Writers and the Industrialization of Agriculture (review). Technology and Culture, 44(2):421–422, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42030.

Long:2004:ELG

[Lon04a] Pamela O. Long. The Eye of the Lynx: Galileo, His Friends, and the Beginnings of Modern Natural History (review). Technology and Culture, 45(1):171–173, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51759.

Longmore:2004:APP

[Lon04b] Paul K. Longmore. Artificial Parts, Practical Lives: Modern Histories of Prosthetics (review). Technology and Culture, 45 (4):899–901, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175763.

Long:2005:AHT

[Lon05a] Pamela O. Long. The annales and the history of technology. Technology and Culture, 46(1):177–186, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179767.

Long:2005:ASD

[Lon05b] Pamela O. Long. Arte e scienza delle acque nel Rinascimento (review). Technology and Culture, 46(3):638–640, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186625.

Long:2006:LR

[Lon06] Pamela O. Long. Leonardo (review). Technology and Culture, 47(1):184–186, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196053.

Long:2007:FDP

[Lon07] Priscilla Long. The Face of Decline: The Pennsylvania Anthracite Region in the Twentieth Century (review). Technology and Culture, 48(1):205–206, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209846.

Long:2009:MRF

[Lon09] Pamela O. Long. Michael of Rhodes: A Fifteenth-Century mariner and his book. Technology and Culture, 50(1):103–109, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258560.

Lowengard:2000:TGT

[Low00a] Sarah Lowengard. Die Tuchlebarone: Zur Geschichte der Textildruck- und Textilfarbeindustrie in Hard vom spaten 18. bis sum fruhen 20. Jahrhundert (review). Technology and Culture, 41(4):801–802, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33601.

Lowood:2000:CBH

[Low00b] Henry Lowood. Current bibliography in the history of technology (1998). Technology and Culture, 41(Supplement):1–4, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33653.

Lowen:2001:BWM

[Low01a] Rebecca S. Lowen. For Better or for Worse: The Marriage of Science and Government in the United States (review). Technology and Culture, 42(4):825–826, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33827.

Lowood:2001:STC

[Low01b] Henry Lowood. Sorting Things Out: Classification and Its Consequences (review). Technology and Culture, 42(2):392– 394, April 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33725.

Lowengard:2002:IEH

[Low02a] Sarah Lowengard. Instruments and Experimentation in the History of Chemistry (review). Technology and Culture, 43 (3):600–602, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33977.

Lowood:2002:CBH

[Low02b] Henry Lowood. Current bibliography in the history of technology (1999). Technology and Culture, 43(Supplement):1–4, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34002.

Lowen:2004:CWL

[Low04a] Rebecca S. Lowen. Cold War Laboratory: RAND, the Air Force, and the American State, 1945–1950 (review). Technology and Culture, 45(4):886–888, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175764.

Lowengard:2004:DCS

[Low04b] Sarah Lowengard. Discussing Chemistry and Steam: The Minutes of a Coffee House Philosophical Society, 1780–1787 (review). Technology and Culture, 45(3):647–648, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172994.

Lowengard:2004:EIR

[Low04c] Sarah Lowengard. From Energy to Information: Representation in Science and Technology, Art, and Literature (review). Technology and Culture, 45(1):227–229, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51760.

Lowen:2005:FTS

[Low05a] Rebecca S. Lowen. Fred Terman at Stanford: Building a Discipline, a University, and Silicon Valley (review). Technology and Culture, 46(4):854–855, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192305.

Lowengard:2005:SVS

[Low05b] Sarah Lowengard. Supple as voile, sturdy as canvas: The Cambridge History of Western Textiles. Technology and Culture, 46(1):192–197, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/179768.

Lowengard:2008:DR

[Low08a] Sarah Lowengard. The digital remains. Technology and Culture, 49(4):1031–1035, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255450.

Lowood:2008:GMU

[Low08b] Henry Lowood. Google and the Myth of Universal Knowledge (review). Technology and Culture, 49(1):298–300, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230918.

Law:2000:PTS

[LS00] John Law and Vicky Singleton. Performing Technology’s stories: On social constructivism, performance, and performativity. Technology and Culture, 41(4):765–775, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33626.

Luckin:2002:EAC

[Luc02] Bill Luckin. Effluent America: Cities, Industry, Energy, and the Environment (review). Technology and Culture, 43(4): 800–801, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34027.

Luckin:2003:IDL

[Luc03] Bill Luckin. Inviting Disaster: Lessons from the Edge of Technology (review). Technology and Culture, 44(4):848–850, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49706.

Luckin:2004:CPB

[Luc04] Bill Luckin. Constructing a Pipe-Bound City: A History of Water Supply, Sewerage, and Excreta Removal in Norrkoping and Linkoping, Sweden, 1860–1910 (review). Technology and Culture, 45(2):428–430, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55672.

Lucas:2005:IMA

[Luc05a] Adam Robert Lucas. Industrial milling in the ancient and medieval worlds: a survey of the evidence for an industrial revolution in Medieval Europe. Technology and Culture, 46(1):1–30, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179769.

Luckin:2005:NMC

[Luc05b] Bill Luckin. Nuclear meltdown and the culture of risk. Technology and Culture, 46(2):393–399, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184326.

Luckin:2006:LTR

[Luc06] Bill Luckin. London’s Thames: The River That Shaped a City and Its History (review). Technology and Culture, 47(4): 846–847, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204992.

Luckin:2007:DRR

[Luc07] Bill Luckin. Death Rode the Rails: American Railroad Accidents and Safety, 1828–1965 (review). Technology and Culture, 48(2):448–449, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215403.

Lucsko:2009:JBR

[Luc09] David N. Lucsko. John Bell Rae and the automobile 1959, 1965, 1971, 1984. Technology and Culture, 50(4):894–914, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363420.

Luey:2005:JHR

[Lue05] Beth Luey. JSTOR: A History (review). Technology and Culture, 46(3):693–695, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186626.

Luger:2003:MSC

[Lug03] Stan Luger. Making and Selling Cars: Innovation and Change in the U.S. Automotive Industry (review). Technology and Culture, 44(4):830–831, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49707.

Lundin:2009:CCL

[Lun09] Per Lundin. Cars for Comrades: The Life of the Soviet Automobile (review). Technology and Culture, 50(4):940–942, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363435.

Langdon:2005:TWM

[LW05] John Langdon and Martin Watts. Tower windmills in Medieval England: a case of arrested development? Technology and Culture, 46(4):697–718, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192304.

Lyman:2003:DTC

[Lym03] Fred Lyman. Deconstructing technology and culture. Technology and Culture, 44(1):227–229, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40175.

Lynn:2009:LDR

[Lyn09] Michael R. Lynn. L’Enquête du Régent 1716–1718: Sciences, techniques et politique dans la France pré-industrielle (review). Technology and Culture, 50(4):936–937, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363432.

Meyer-Arendt:2009:CLH

[MA09] Klaus J. Meyer-Arendt. Coast Lines: How Mapmakers Frame the World and Chart Environmental Change (review). Technology and Culture, 50(3):703–704, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270555.

MacLeod:2000:MMI

[Mac00] Christine MacLeod. The Making of Modern Intellectual Property Law: The British Experience, 1760–1911 (review). Technology and Culture, 41(4):808–810, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33609.

MacLeod:2001:BPI

[Mac01] Christine MacLeod. British Patents of Invention, 1617–1977: A Guide for Researchers (review). Technology and Culture, 42 (2):340–341, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33726.

MacLeod:2002:IEL

[Mac02] Christine MacLeod. Industrializing English Law: Entrepreneurship and Business Organization, 1720–1844 (review). Technology and Culture, 43(3):597–599, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33978.

Macnaughtan:2004:SWK

[Mac04] Helen Macnaughtan. The Silk Weavers of Kyoto: Family and Work in a Changing Traditional Industry (review). Technology and Culture, 45(1):177–178, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51761.

Mackowski:2005:SAS

[Mac05a] Maura Phillips Mackowski. The Secret of Apollo: Systems Management in American and European Space Programs (review). Technology and Culture, 46(1):234–236, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
179770.

MacLeod:2005:TSI

[Mac05b] Christine MacLeod. Trade Secrets: Intellectual Piracy and the Origins of American Industrial Power (review). Technology and Culture, 46(2):443–444, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184327.

Mackowski:2006:MSW

[Mac06] Maura Phillips Mackowski. “Manned” spaceflight for women? Technology and Culture, 47(1):170–174, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196054.

MacDougall:2009:GIA

[Mac09] Robert MacDougall. The Genesis of Industrial America, 1870– 1920 (review). Technology and Culture, 50(1):222–223, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258573.

Madison:2006:TPN

[Mad06] Mark Glen Madison. Tapping the Pines: The Naval Stores Industry in the American South (review). Technology and Culture, 47(1):206–208, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196055.

Maerker:2009:VIA

[Mae09] Anna Maerker. Visions of the Industrial Age, 1830–1914: Modernity and the Anxiety of Representation in Europe (review). Technology and Culture, 50(4):934–935, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363431.

Magoun:2000:RTC

[Mag00] Alexander B. Magoun. Off the Record: The Technology and Culture of Sound Recording in America (review). Technology and Culture, 41(4):818–819, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33623.

Magoun:2001:GFT

[Mag01] Alexander B. Magoun. Gramophone, Film, Typewriter (review). Technology and Culture, 42(1):174–176, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33676.

Magoun:2005:PVM

[Mag05a] Alexander B. Magoun. Playback: From the Victrola to MP3, 100 Years of Music, Machines, and Money (review). Technology and Culture, 46(2):427–429, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184328.

Magoun:2005:FND

[Mag05b] Alexander B. Magoun. The Forgotten Network: Du Mont and the Birth of American Television (review). Technology and Culture, 46(4):830–831, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192306.

Magliari:2009:WRM

[Mag09] Michael F. Magliari. The West the Railroads Made (review). Technology and Culture, 50(1):220–221, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258574.

Maher:2002:HCC

[Mah02] Neil M. Maher. Hoover, Conservation, and Consumerism: Engineering the Good Life (review). Technology and Culture, 43 (2):435–436, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33921.

Maines:2004:BEH

[Mai04] Rachel Maines. The Body Electric: How Strange Machines Built the Modern American (review). Technology and Culture, 45(3):653–654, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172995.

Maines:2007:WGT

[Mai07] Rachel Maines. Women, Gender, and Technology (review). Technology and Culture, 48(4):902–904, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223370.

Maines:2009:FSO

[Mai09] Rachel Maines. Falling for Science: Objects in Mind (review). Technology and Culture, 50(2):502–503, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263360.

Malcolmson:2001:SLG

[Mal01] Patricia E. Malcolmson. Steam Laundries: Gender, Technology, and Work in the United States and Great Britain, 1880– 1940 (review). Technology and Culture, 42(2):356–357, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33727.

Malin:2009:MET

[Mal09a] Brenton J. Malin. Mediating emotion: Technology, social science, and emotion in the Payne fund motion-picture studies. Technology and Culture, 50(2):366–390, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263327.

Mallea:2009:RRU

[Mal09b] Amahia Mallea. Rivertown: Rethinking Urban Rivers (review). Technology and Culture, 50(1):217–218, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258580.

Marks:2000:SSI

[Mar00a] Gregory A. Marks. The Social Shaping of Information Superhighways: European and American Roads to the Information Society (review). Technology and Culture, 41(3):621–623, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33575.

Marovich:2000:MPH

[Mar00b] Lisa A. Marovich. Made to Play House: Dolls and the Commercialization of American Girlhood, 1830–1930 (review). Technology and Culture, 41(1):145–147, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33486.

Marsh:2000:DC

[Mar00c] Joseph O. Marsh. Donald Cardwell (1919–1998). Technology and Culture, 41(1):184–186, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33487.

Marche:2001:RSF

[Mar01] Jordan D. Marche. Reconsidering Sputnik: Forty Years since the Soviet Satellite (review). Technology and Culture, 42(4): 818–820, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33828.

Martin:2002:PIA

[Mar02] Patrick Edward Martin. Perspectives on Industrial Archaeology, and: Twentieth Century Industrial Archaeology (review). Technology and Culture, 43(1):157–160, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33877.

Marschall:2003:BWW

[Mar03] Luitgard Marschall. Bier, Wissenschaft und Wirtschaft in Deutschland, 1800–1914: Ein Beitrag zur deutschen Industrialisierungsgeschichte (review). Technology and Culture, 44 (3):626–627, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46100.

Marche:2005:VGT

[Mar05a] Jordan D. Marche. Voyager’s Grand Tour: To the Outer Planets and Beyond (review). Technology and Culture, 46(1): 236–237, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179771.

Martschukat:2005:EEC

[Mar05b] Jürgen Martschukat. Edison and the Electric Chair (review). Technology and Culture, 46(4):870–871, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192307.

Martello:2006:ASC

[Mar06a] Robert Martello. America as Second Creation: Technology and Narratives of New Beginnings (review). Technology and Culture, 47(2):419–421, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/200860.

Martello:2006:IER

[Mar06b] Robert Martello. Inventing for the Environment (review). Technology and Culture, 47(2):446–447, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/200861.

Marrs:2009:ESU

[Mar09] Aaron W. Marrs. Entrepreneurs in the Southern Upcountry: Commercial Culture in Spartanburg, South Carolina, 1845– 1880 (review). Technology and Culture, 50(4):929–930, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363428.

Matthews:2002:GCL

[Mat02] Jeffrey J. Matthews. The Generations of Corning: The Life and Times of a Global Corporation (review). Technology and Culture, 43(2):445–446, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33922.

Mayring:2000:MTC

[May00] Eva Alexandra Mayring. Munich’s technology collections. Technology and Culture, 41(3):521–524, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33576.

Mayr:2009:EIR

[May09] Otto Mayr. Engines of the Imagination: Renaissance Culture and the Rise of the Machine (review). Technology and Culture, 50(3):683–685, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270565.

Mazzaoui:2001:SIR

[Maz01] Maureen Fennell Mazzaoui. The Silk Industry of Renaissance Venice (review). Technology and Culture, 42(2):338–340, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33728.

Mazumdar:2003:GDC

[Maz03] Sucheta Mazumdar. The Great Divergence: China, Europe, and the Making of the Modern World Economy (review). Technology and Culture, 44(3):604–606, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46101.

Mazzotti:2004:VSC

[Maz04a] Massimo Mazzotti. Volta: Science and Culture in the Age of Enlightenment (review). Technology and Culture, 45(2):420– 421, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55674.

Mazzotti:2004:EMM

[Maz04b] Massimo Mazzotti. Enlightened mills: Mechanizing olive oil production in Mediterranean Europe. Technology and Culture, 45(2):277–304, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55704.

Mazzotti:2007:WWW

[Maz07] Massimo Mazzotti. Wind, Water, Work: Ancient and Medieval Milling Technology (review). Technology and Culture, 48(4): 848–850, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223371.

Mitcham:2006:CTS

[MB06] Carl Mitcham and Juan Bautista Bengoetxea. Culture and technology in Spain: From philosophical analysis to STS. Technology and Culture, 47(3):607–622, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/201467.

McBride:2000:MSL

[McB00] William M. McBride. Monitor: The Story of the Legendary Civil War Ironclad and the Man Whose Invention Changed the Course of History, and: War, Technology, and Experience aboard the USS Monitor (review). Technology and Culture, 41 (4):838–841, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33628.

McBride:2001:PNY

[McB01] William M. McBride. The Philadelphia Navy Yard: From the Birth of the U.S. Navy to the Nuclear Age (review). Technology and Culture, 42(4):810–812, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33829.

McBride:2002:NWR

[McB02] William M. McBride. Naval Warfare, 1815–1914 (review). Technology and Culture, 43(2):457–458, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33923.

McBride:2004:CWI

[McB04] William M. McBride. Civil War Ironclads: The U.S. Navy and Industrial Mobilization (review). Technology and Culture, 45(1):188–189, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51762.

McBride:2005:RSB

[McB05] William M. McBride. Resurrection: Salvaging the Battle Fleet at Pearl Harbor (review). Technology and Culture, 46(3):654– 655, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186627.

McBride:2006:MWP

[McB06] William M. McBride. Making Waves: Politics, Propaganda, and the Emergence of the Imperial Japanese Navy, 1868–1922 (review). Technology and Culture, 47(4):833–834, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
204994.

McBride:2007:DGB

[McB07] William M. McBride. Dreadnought Gunnery and the Battle of Jutland: The Question of Fire Control (review). Technology and Culture, 48(3):649–651, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218712.

McCraw:2000:RPR

[McC00a] Thomas K. McCraw. Rescuing Prometheus (review). Technology and Culture, 41(3):616–619, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33577.

McCray:2000:TPU

[McC00b] Patrick McCray. Traces of the Past: Unraveling the Secrets of Archaeology Through Chemistry (review). Technology and Culture, 41(1):127–128, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33488.

McCray:2001:WMF

[McC01a] Patrick McCray. What makes a failure? Designing a new national telescope, 1975–1984. Technology and Culture, 42(2): 265–291, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33729.

McCurdy:2001:DSE

[McC01b] Howard E. McCurdy. The Dream of Spaceflight: Essays on the Near Edge of Infinity (review). Technology and Culture, 42(4): 815–817, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33830.

McCray:2003:HGR

[McC03] Patrick McCray. A History of Glassforming (review). Technology and Culture, 44(2):383–384, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42031.

McCray:2006:SLT

[McC06] Patrick McCray. Stargazer: The Life and Times of the Telescope (review). Technology and Culture, 47(3):641–643, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201465.

McCaskey:2007:RFF

[McC07a] John McCaskey. The Religious Foundations of Francis Bacon’s Thought (review). Technology and Culture, 48(3):618– 620, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218713.

McCray:2007:HVH

[McC07b] Patrick McCray. Holographic Visions: A History of New Science (review). Technology and Culture, 48(3):675–677, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218714.

McCarthy:2008:MSU

[McC08a] John P. McCarthy. Metropolis on the Styx: The Underworlds of Modern Urban Culture, 1800—2001 (review). Technology and Culture, 49(4):1057–1058, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255491.

McCray:2008:THT

[McC08b] W. Patrick McCray. The Telescope: Its History, Technology, and Future (review). Technology and Culture, 49(3):789– 791, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243093.

McCaskey:2009:TST

[McC09a] John P. McCaskey. Theology and Science in the Thought of Francis Bacon (review). Technology and Culture, 50(3):685– 686, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270558.

McCray:2009:LIS

[McC09b] W. Patrick McCray. From lab to iPod: A story of discovery and commercialization in the Post-Cold War Era. Technology and Culture, 50(1):58–81, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258568.

McCrossen:2009:OTF

[McC09c] Alexis McCrossen. One Time Fits All: The Campaigns for Global Uniformity (review). Technology and Culture, 50(1): 252–254, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258548.

McElheny:2003:CCI

[McE03] Victor K. McElheny. Corning and the Craft of Innovation (review). Technology and Culture, 44(2):384–386, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42032.

McElheny:2004:TPW

[McE04] Victor K. McElheny. Tuxedo Park: A Wall Street Tycoon and the Secret Palace of Science that Changed the Course of World War II (review). Technology and Culture, 45(2):456–457, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55675.

McElheny:2006:PTH

[McE06] Victor K. McElheny. The People’s Tycoon: Henry Ford and the American Century (review). Technology and Culture, 47 (2):433–435, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200862.

McElheny:2007:ERA

[McE07] Victor K. McElheny. Eddie Rickenbacker: An American Hero in the Twentieth Century (review). Technology and Culture, 48 (2):465–466, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215404.

McGee:2000:HWA

[McG00a] David McGee. History of Warships: From Ancient Times to the Twenty-First Century (review). Technology and Culture, 41 (3):632–633, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33578.

McGee:2000:LPV

[McG00b] David McGee. On Line and On Paper: Visual Representations, Visual Culture, and Computer Graphics in Design Engineering (review). Technology and Culture, 41(2):388–390, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33532.

McGlade:2000:PRB

[McG00c] J. M. McGlade. The Productivity Race: British Manufacturing in International Perspective 1850–1990, and: 0The Conservatives and Industrial Efficiency, 1951–64: Thirteen Wasted Years? (review). Technology and Culture, 41(2):363–366, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33533.

McGee:2002:AAV

[McG02] David McGee. The Art and Archaeology of Venetian Ships and Boats (review). Technology and Culture, 43(2):411–412, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33924.

McGee:2003:AAP

[McG03] David McGee. Architecture in the Age of Printing: Orality, Writing, Typography, and Printed Images in the History of Architectural Theory (review). Technology and Culture, 44(2): 377–379, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42033.

McGaw:2004:MBH

[McG04] Judith A. McGaw. Memorial: Brooke hindle, 1918–2001. Technology and Culture, 45(1):241–246, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51763.

McGee:2005:HWS

[McG05] David McGee. To Harness the Wind: A Short History of the Development of Sails (review). Technology and Culture, 46(1): 245–246, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179772.

McGee:2006:PSC

[McG06] David McGee. The Philosophy of Shipbuilding: Conceptual Approaches to the Study of Wooden Ships (review). Technology and Culture, 47(2):426–427, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200863.

McIntyre:2000:FFR

[McI00] Stephen L. McIntyre. The failure of fordism: Reform of the automobile repair industry, 1913–1940. Technology and Culture, 41(2):269–299, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33534.

McManamon:2000:SSS

[McM00] John M. McManamon. Seagoing Ships and Seamanship in the Bronze Age Levant (review). Technology and Culture, 41(1): 129–130, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33489.

McMath:2006:ELG

[McM06] Robert C. McMath. Engineering in a Land-Grant Context: The Past, Present, and Future of an Idea (review). Technology and Culture, 47(2):428–429, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200864.

McNally:2002:MMM

[McN02] Larry McNally. Manufacturing Montreal: The Making of an Industrial Landscape, 1850–1930 (review). Technology and Culture, 43(2):423–424, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33925.

McSwain:2000:IDE

[McS00] Rebecca McSwain. Instruments of Desire: The Electric Guitar and the Shaping of Musical Experience (review). Technology and Culture, 41(4):821–822, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33630.

McShane:2002:AAH

[McS02] Clay McShane. The Automobile in American History and Culture: A Reference Guide (review). Technology and Culture, 43 (3):637–639, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33979.

Meade:2005:PBP

[Mea05] Roger A. Meade. People of the Bomb: Portraits of America’s Nuclear Complex (review). Technology and Culture, 46(2):455– 456, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184329.

Meilinger:2000:IDF

[Mei00a] Phillip S. Meilinger. Innovation and the Development of Flight (review). Technology and Culture, 41(4):843–845, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33631.

Meilinger:2000:FGE

[Mei00b] Phillip S. Meilinger. To Fool a Glass Eye: Camouflage Versus Photoreconnaissance in World War II (review). Technology and Culture, 41(2):385–387, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33535.

Meilinger:2002:WMT

[Mei02] Phillip S. Meilinger. War Machines: Transforming Technologies in the U.S. Military, 1920–1940 (review). Technology and Culture, 43(3):643–644, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33980.

Meikle:2003:LMM

[Mei03] Jeffrey L. Meikle. Leo Marx’s The Machine in the Garden. Technology and Culture, 44(1):147–159, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/40176.

Meikle:2005:GMW

[Mei05] Jeffrey L. Meikle. Ghosts in the machine: Why it’s hard to write about design. Technology and Culture, 46(2):385– 392, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184330.

Meikle:2007:SSS

[Mei07] Jeffrey L. Meikle. From Submarines to Suburbs: Selling a Better America, 1939–1959 (review). Technology and Culture, 48 (3):661–663, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218715.

Melosh:2001:DDG

[Mel01] Barbara Melosh. Devices and Desires: Gender, Technology, and American Nursing (review). Technology and Culture, 42(4): 787–788, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33831.

Melosi:2006:DAW

[Mel06] Martin V. Melosi. Designing America’s Waste Landscapes (review). Technology and Culture, 47(1):224–226, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
196057.

Mende:2007:OMB

[Men07] Michael Mende. Oskar von Miller, 1855–1934: Eine Biographie (review). Technology and Culture, 48(1):219–220, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209848.

Mentzel:2009:CHF

[Men09] Peter Mentzel. On Chariots with Horses of Fire and Iron: The Excursionists and the Narrow Gauge Railroad from Jaffa to Jerusalem (review). Technology and Culture, 50(3):710– 711, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270564.

Mernitz:2005:GIG

[Mer05] Kenneth S. Mernitz. German Industry and Global Enterprise: BASF — The History of a Company (review). Technology and Culture, 46(2):446–447, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184331.

Michel:2008:IMA

[Mic08] Alain P. Michel. An image mise en abyme. Technology and Culture, 49(4):967–973, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255443.

Middleton:2003:WTS

[Mid03] William D. Middleton. In the Wake of Tacoma: Suspension Bridges and the Quest for Aerodynamic Stability (review). Technology and Culture, 44(3):618–620, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46102.

Middleton:2005:ERR

[Mid05a] William D. Middleton. Electric Railways, 1880–1990 (review). Technology and Culture, 46(1):211–213, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179774.

Middleton:2005:NYS

[Mid05b] William D. Middleton. New York Subways: An Illustrated History of New York City’s Transit Cars (review). Technology and Culture, 46(3):668–669, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186628.

Middleton:2006:BDS

[Mid06] William D. Middleton. Bridging Deep South Rivers: The Life and Legend of Horace King (review). Technology and Culture, 47(3):695–696, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201466.

Mierzejewski:2004:BGS

[Mie04] Alfred C. Mierzejewski. The Business of Genocide: The SS, Slave Labor, and the Concentration Camps (review). Technology and Culture, 45(1):199–201, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51765.

Mierzejewski:2009:ABF

[Mie09] Alfred C. Mierzejewski. Across the Borders: Financing the World’s Railways in the Nineteenth and Twentieth Centuries (review). Technology and Culture, 50(3):708–710, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270563.

Miller:2001:STH

[Mil01] Carol Poh Miller. Spans in Time: A History of Nebraska Bridges (review). Technology and Culture, 42(1):158–159, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33677.

Miller:2002:GCT

[Mil02] Carol Poh Miller. Grand Central Terminal: Railroads, Engineering, and Architecture in New York City (review). Technology and Culture, 43(4):805–807, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34028.

Miller:2003:RAM

[Mil03] Carol Poh Miller. Rails across the Mississippi: A History of the St. Louis Bridge, and: Elegant Arches, Soaring Spans: C. B. McCullough, Oregon’s Master Bridge Builder (review). Technology and Culture, 44(3):616–618, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46103.

Millard:2004:HBH

[Mil04a] A. J. Millard. How Breakthroughs Happen: The Surprising Truth about How Companies Innovate (review). Technology and Culture, 45(2):464–465, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55676.

Miller:2004:IGE

[Mil04b] Carol Poh Miller. Invisible Giants: The Empires of Cleveland’s Van Sweringen Brothers (review). Technology and Culture, 45 (2):444–446, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55677.

Miller:2004:IBS

[Mil04c] Carol Poh Miller. The Iron Bridge: Symbol of the Industrial Revolution, 2nd ed. (review). Technology and Culture, 45(3):649–650, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/172996.

Miller:2006:AIC

[Mil06] Marla R. Miller. The Artisan of Ipswich: Craftsmanship and Community in Colonial New England (review). Technology and Culture, 47(2):417–419, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200865.

Miller:2008:JMF

[Mil08] Carol Poh Miller. James Marston Fitch: Selected Writings on Architecture, Preservation, and the Built Environment (review). Technology and Culture, 49(1):289–291, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230922.

Miller:2009:GSP

[Mil09] Jeanine Head Miller. Goods for Sale: Products and Advertising in the Massachusetts Industrial Age (review). Technology and Culture, 50(1):225–226, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258566.

Mindell:2000:OBB

[Min00] David A. Mindell. Opening Black’s box: Rethinking feedback’s myth of origin. Technology and Culture, 41(3):405–434, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33579.

Mindell:2003:OCA

[Min03] David A. Mindell. An Ocean in Common: American Naval Officers, Scientists, and the Ocean Environment (review). Technology and Culture, 44(2):437–439, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/42034.

Mirza:2008:ISM

[Mir08] Mahan Mirza. Islamic Science and the Making of the European Renaissance (review). Technology and Culture, 49(3):786– 787, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243109.

Misa:2002:BSF

[Mis02] Thomas J. Misa. Big Steel: The First Century of the United States Steel Corporation, 1901–2001 (review). Technology and Culture, 43(2):443–444, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33926.

Misa:2007:TMQ

[Mis07] Thomas J. Misa. Technology Matters: Questions to Live With (review). Technology and Culture, 48(2):411–413, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215405.

Misa:2008:TIS

[Mis08] Thomas J. Misa. Transforming the international system: Geoffrey L. Herrera’s Technology and International Transformation. Technology and Culture, 49(1):230–233, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230920.

Mom:2001:TTH

[MK01] Gijs Mom and David A. Kirsch. Technologies in tension: Horses, electric trucks, and the motorization of American cities, 1900–1925. Technology and Culture, 42(3):489–518, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33781.

MacLeod:2006:PPI

[MN06] Christine MacLeod and Alessandro Nuvolari. The pitfalls of prosopography: Inventors in the dictionary of national biography. Technology and Culture, 47(4):757–776, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
204993.

Mnookin:2000:TPN

[Mno00] Jennifer L. Mnookin. Technology and Privacy: The New Landscape (review). Technology and Culture, 41(4):828–830, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33632.

Mody:2006:TIE

[Mod06a] Cyrus C. M. Mody. Technology, Institutions, and Economic Growth (review). Technology and Culture, 47(4):817–819, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204995.

Mody:2006:CUI

[Mod06b] Cyrus C. M. Mody. Corporations, universities, and instrumental communities: Commercializing probe microscopy, 1981–1996. Technology and Culture, 47(1):56–80, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
196058.

Mody:2008:MPB

[Mod08] Cyrus C. M. Mody. Managing Path-Breaking Innovations: CERN-ATLAS, Airbus, and Stem Cell Research (review). Technology and Culture, 49(2):514–515, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/236993.

Mohanty:2004:WCC

[Moh04] Gail Fowler Mohanty. The Weaver’s Craft: Cloth, Commerce, and Industry in Early Pennsylvania (review). Technology and Culture, 45(4):836–837, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175765.

Mohanty:2008:FLP

[Moh08] Gail Fowler Mohanty. Female Labour Power: Women Workers’ Influence on Business Practices in the British and American Cotton Industries, 1780–1860 (review). Technology and Culture, 49(4):1094–1095, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255537.

Mokyr:2001:HTL

[Mok01] Joel Mokyr. High technology in the Low Countries: Geschiedenis van de techniek in Nederland, 1800–1890. Technology and Culture, 42(1):133–137, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33678.

Mokyr:2005:HSE

[Mok05] Joel Mokyr. Hockey stick economics. Technology and Culture, 46(3):613–617, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186629.

Mokyr:2009:CMC

[Mok09] Joel Mokyr. The culture of modern capitalism. Technology and Culture, 50(2):441–448, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263333.

Molella:2000:CFB

[Mol00a] Arthur P. Molella. Crystal Fire: The Birth of the Information Age (review). Technology and Culture, 41(3):623–625, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33580.

Molloy:2000:SRG

[Mol00b] Scott Molloy. Streets, Railroads, and the Great Strike of 1877 (review). Technology and Culture, 41(3):636–638, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33581.

Molella:2002:SMS

[Mol02] Arthur P. Molella. Science moderne: Sigfried Giedion’s space, time and architecture and mechanization takes command. Technology and Culture, 43(2):374–389, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33927.

Molella:2003:TVI

[Mol03] Arthur P. Molella. Technomanifestos: Visions from the Information Revolutionaries (review). Technology and Culture, 44 (3):639–640, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46104.

Molella:2004:DLB

[Mol04] Arthur P. Molella. Draw the Lightning Down: Benjamin Franklin and Electrical Technology in the Age of Enlightenment (review). Technology and Culture, 45(4):840–842, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175767.

Molella:2005:LDA

[Mol05] Arthur P. Molella. The longue durée of Abbott Payson Usher. Technology and Culture, 46(4):779–796, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192308.

Molella:2006:WAC

[Mol06] Arthur P. Molella. At Work in the Atomic City: A Labor and Social History of Oak Ridge, Tennessee (review). Technology and Culture, 47(2):447–448, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200866.

Molella:2007:HIB

[Mol07] Arthur P. Molella. How Invention Begins: Echoes of Old Voices in the Rise of New Machines (review). Technology and Culture, 48(3):637–638, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218716.

Molella:2008:WDT

[Mol08] Arthur P. Molella. Walt Disney: The Triumph of the American Imagination (review). Technology and Culture, 49(3):823– 824, July 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/243088.

Mom:2002:FLA

[Mom02] Gijs Mom. Von Ford lernen? Automobilbau und Motorisierung in Deutschland bis 1933 (review). Technology and Culture, 43 (4):814–815, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34029.

Mom:2004:HSA

[Mom04] Gijs Mom. Die holprige Siegeszug des Automobils, 1895–1930, and: Geschichte des Autos (review). Technology and Culture, 45(1):195–197, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51766.

Mom:2005:RRE

[Mom05] Gijs Mom. Roads without rails: European highway-network building and the desire for long-range motorized mobility. Technology and Culture, 46(4):745–772, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192309.

Monivas:2007:CTQ

[Moñ07] Jesús Romero Moñivas. Cultures of Technology and the Quest for Innovation (review). Technology and Culture, 48(3):641– 643, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218717.

Moon:2000:GGS

[Moo00] Suzanne Moon. Guns, Germs, and Steel: The Fates of Human Societies (review). Technology and Culture, 41(3):570– 571, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33582.

Moon:2002:ZSF

[Moo02a] Suzanne Moon. Zapotec Science: Farming and Food in the Northern Sierra of Oaxaca (review). Technology and Culture, 43(3):623–625, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33981.

Moon:2002:IAU

[Moo02b] Suzanne M. Moon. Insatiable Appetite: The United States and Ecological Degradation of the Tropical World (review). Technology and Culture, 43(1):212–213, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33878.

Moon:2004:TKS

[Moo04a] Suzanne Moon. Tracy Kidder, the soul of a new machine. Technology and Culture, 45(3):597–602, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/172997.

Moon:2004:RST

[Moo04b] Suzanne M. Moon. Research in Science and Technology Studies: Knowledge and Technology Transfer (review). Technology and Culture, 45(1):229–230, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51767.

Morton:2000:CEA

[Mor00] David Morton. Children of Ezekiel: Aliens, UFOs, the Crisis of Race, and the Advent of End Time (review). Technology and Culture, 41(1):170–172, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33490.

Morrison:2001:STT

[Mor01] James C. Morrison. Seeing Things: Television in the Age of Uncertainty, and: TV Living: Television, Culture, and Everyday Life (review). Technology and Culture, 42(1):176–178, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33679.

Morrow:2003:KAT

[Mor03] John Howard Morrow. The Kaiser’s Army, 1870–1918: Technological, Tactical, and Operational Dilemmas in Germany during the Machine Age (review). Technology and Culture, 44 (3):642–643, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46105.

Morrison:2005:RCP

[Mor05a] James C. Morrison. The Rise of Cable Programming in the United States: Revolution or Evolution? (review). Technology and Culture, 46(2):432–434, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184332.

Morrow:2005:RBA

[Mor05b] John Howard Morrow. Hat in the Ring: The Birth of American Air Power in the Great War (review). Technology and Culture, 46(1):223–225, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/179775.

Morse:2006:KWN

[Mor06a] Kathryn Taylor Morse. Knock on Wood: Nature as Commodity in Douglas Fir Country (review). Technology and Culture, 47 (1):208–209, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196059.

Morton:2006:CSR

[Mor06b] David Morton. The Coming of Sound (review). Technology and Culture, 47(3):653–655, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201468.

Morrison:2007:TPH

[Mor07a] James C. Morrison. Transmitting the Past: Historical and Cultural Perspectives on Broadcasting (review). Technology and Culture, 48(1):227–228, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209849.

Morrow:2007:DDT

[Mor07b] John Howard Morrow. Dominance by Design: Technological Imperatives and America’s Civilizing Mission (review). Technology and Culture, 48(2):413–414, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215406.

Morrison:2008:LWJ

[Mor08a] Alan Morrison. The Lost World of James Smithson: Science, Revolution, and the Birth of the Smithsonian (review). Technology and Culture, 49(1):240–242, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230921.

Morton:2008:EST

[Mor08b] David L. Morton. Electric Sounds: Technological Change and the Rise of Corporate Mass Media (review). Technology and Culture, 49(2):505–506, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/236994.

Mosco:2002:TNT

[Mos02] Vincent Mosco. Telecom Nation: Telecommunications, Computers, and Governments in Canada (review). Technology and Culture, 43(3):627–628, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33982.

Moser:2006:AKD

[Mos06] Kurt Moser. Automobilitat: Karl Drais und die unglaublichen Anfange (review). Technology and Culture, 47(3):667–668, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201469.

Moudry:2008:NTH

[Mou08] Roberta Moudry. Nearest Thing to Heaven: The Empire State Building and American Dreams (review). Technology and Culture, 49(4):1067–1069, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/255506.

Moy:2002:BAS

[Moy02] Timothy Moy. The bombing of Auschwitz: Should it have been attempted? Technology and Culture, 43(3):641–643, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33983.

Moy:2005:SA

[Moy05] Timothy Moy. Structure ascendant. Technology and Culture, 46(4):797–804, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/192310.

Moy:2006:FTH

[Moy06] Timothy Moy. Fast Tanks and Heavy Bombers: Innovation and the U.S. Army, 1917–1945 (review). Technology and Culture, 47(1):242–244, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196060.

Merwood-Salisbury:2008:BCP

[MS08] Joanna Merwood-Salisbury. Building a Century of Progress: The Architecture of Chicago’s 1933–34 World’s Fair (review).

Technology and Culture, 49(2):476–477, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/236992.

Muir:2003:RSE

[Mui03] Malcolm Muir. Rickover: The Struggle for Excellence (review). Technology and Culture, 44(1):169–170, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/40177.

Mukerji:2003:IUE

[Muk03] Chandra Mukerji. Intelligent uses of engineering and the legitimacy of state power. Technology and Culture, 44(4):655–676, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49708.

Mukerji:2006:TKC

[Muk06] Chandra Mukerji. Tacit knowledge and classical technique in Seventeenth-Century France: Hydraulic cement as a living practice among masons and military engineers. Technology and Culture, 47(4):713–733, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/204996.

Mullen:2001:CAB

[Mul01a] Megan Gwynne Mullen. La cite au bout du fil (review). Technology and Culture, 42(4):784–785, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33832.

Multhauf:2001:CCT

[Mul01b] Robert P. Multhauf. Communicating Chemistry: Textbooks and Their Audiences, 1789–1939 (review). Technology and Culture, 42(3):564–565, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33782.

Mullen:2004:CWC

[Mul04] Megan Gwynne Mullen. Cold War, Cool Medium: Television, McCarthyism, and American Culture (review). Technology and Culture, 45(4):888–889, October 2004. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175768.

Mullen:2005:TTC

[Mul05] Megan Gwynne Mullen. Technology, Television, and Competition: The Politics of Digital TV (review). Technology and Culture, 46(4):833–835, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192311.

Mullen:2006:CTF

[Mul06] Megan Gwynne Mullen. Coming to terms with the future He foresaw. Technology and Culture, 47(2):373–380, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200867.

Mullen:2007:RMC

[Mul07] Megan Gwynne Mullen. The Republic of Mass Culture: Journalism, Filmmaking, and Broadcasting in America since 1941 (review). Technology and Culture, 48(1):228–230, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209850.

Mullen:2008:RMR

[Mul08] Megan Mullen. Residual Media (review). Technology and Culture, 49(2):506–508, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236995.

Mullen:2009:SBT

[Mul09] Megan Mullen. Space bias/time bias: Harold Innis, empire and communications. Technology and Culture, 50(1):175–186, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258583.

Muller-Wille:2007:WBP

[MW07] Staffan Müller-Wille. “Wir beherrschen den pflanzlichen Organismus besser… ”: Wissenschaftliche Pflanzenzüchtung in Deutschland 1889–1945 (review). Technology and Culture, 48 (1):188–190, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209851.

Nadis:2004:MEC

[Nad04a] Fred Nadis. Modern Enchantments: The Cultural Power of Secular Magic (review). Technology and Culture, 45(4):895– 897, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175769.

Nadis:2004:USS

[Nad04b] Fred Nadis. Undead Science: Science Studies and the Afterlife of Cold Fusion (review). Technology and Culture, 45(2):462– 463, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55678.

Nadis:2006:ETR

[Nad06] Fred Nadis. The Enchantments of Technology (review). Technology and Culture, 47(4):816–817, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/204997.

Nadis:2007:NAW

[Nad07] Fred Nadis. Nature at Aichi World’s Expo 2005. Technology and Culture, 48(3):575–581, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218718.

Nadis:2008:RGW

[Nad08] Fred Nadis. Robot Ghosts and Wired Dreams: Japanese Science Fiction from Origins to Anime (review). Technology and Culture, 49(4):1087–1089, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255507.

Nagel:2008:DTA

[Nag08] Chris Nagel. Democratizing Technology: Andrew Feenberg’s Critical Theory of Technology (review). Technology and Culture, 49(2):519–521, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236996.

Nakazawa:2007:BWV

[Nak07] Satoshi Nakazawa. Building on Water: Venice, Holland, and the Construction of the European Landscape in Early Modern Times (review). Technology and Culture, 48(3):629–630, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218719.

Norton:2009:FT

[NE09] Peter D. Norton and Bruce Epperson. Fighting traffic. Technology and Culture, 50(4):982–986, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363459.

Nebeker:2001:EER

[Neb01] Frederik Nebeker. Exposing Electronics (review). Technology and Culture, 42(4):798–800, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33833.

Neiberg:2002:ATW

[Nei02] Michael S. Neiberg. Anticipating Total War: The German and American Experiences, 1871–1914, and: Great War, Total War: Combat and Mobilization on the Western Front, 1914– 1918 (review). Technology and Culture, 43(1):205–207, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33879.

Neiberg:2005:DCW

[Nei05] Michael S. Neiberg. The Dream of Civilized Warfare: World War I Flying Aces and the American Imagination (review). Technology and Culture, 46(1):225–226, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179776.

Neiberg:2007:WAF

[Nei07] Michael S. Neiberg. Why Air Forces Fail: The Anatomy of Defeat (review). Technology and Culture, 48(1):234–235, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209852.

Nelson:2004:RSE

[Nel04] Andrew Nelson. Robert Stephenson: The Eminent Engineer (review). Technology and Culture, 45(3):650–652, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
172998.

Neri:2008:RVS

[Ner08] Janice Neri. Renaissance Vision from Spectacles to Telescopes (review). Technology and Culture, 49(3):788–789, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243094.

Neushul:2001:CAC

[Neu01] Peter Neushul. Contraception Across Cultures: Technologies, Choices, Constraints (review). Technology and Culture, 42 (3):575–576, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33783.

Neufeld:2002:SVT

[Neu02] Michael J. Neufeld. Stalin’s V-2: Technologietransfer der deutschen Fernlenkwaffentechnik in die USSR und der Aufbau der sowjetischen Raketenindustrie 1945 bis 1959 (review). Technology and Culture, 43(4):820–822, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34020.

Neufeld:2006:FML

[Neu06a] Michael J. Neufeld. First Man: The Life of Neil A. Armstrong (review). Technology and Culture, 47(2):457–458, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200868.

Neushul:2006:MPB

[Neu06b] Peter Neushul. The Male Pill: A Biography of a Technology in the Making (review). Technology and Culture, 47(3):688– 690, July 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/201470.

Newman:2000:VWI

[New00] Karen Newman. The Visible Woman: Imaging Technologies, Gender, and Science (review). Technology and Culture, 41 (3):599–600, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33583.

Newman:2006:ABR

[New06] Kathy M. Newman. American Babel: Rogue Radio Broadcasters of the Jazz Age (review). Technology and Culture, 47 (3):655–656, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201471.

Nickles:2002:PWR

[Nic02] Shelley Nickles. “Preserving women”: Refrigerator design as social process in the 1930s. Technology and Culture, 43(4): 693–727, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34030.

Nielsen:2002:WPV

[Nie02] Kristian H. Nielsen. Wind Power in View: Energy Landscapes in a Crowded World (review). Technology and Culture, 43(4): 825–826, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34031.

Niemann:2007:SDM

[Nie07a] Linda Niemann. Steel Drivin’ Man: John Henry, the Untold Story of an American Legend (review). Technology and Culture, 48(3):656–658, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218720.

Niemann:2007:WSV

[Nie07b] Linda Niemann. Wives of Steel: Voices of Women from the Sparrows Point Steelmaking Communities (review). Technology and Culture, 48(2):446–447, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/215407.

Nishiyama:2007:WPN

[Nis07] Takashi Nishiyama. War, peace, and nonweapons technology: The Japanese national railways and products of defeat, 1880s-1950s. Technology and Culture, 48(2):286–302, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215408.

Nishiyama:2008:BEU

[Nis08] Takashi Nishiyama. Bankrupting the Enemy: The U.S. Financial Siege of Japan before Pearl Harbor (review). Technology and Culture, 49(3):810–811, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243121.

Nixon:2009:CSR

[Nix09] Denver Nixon. Cycling and Society (review). Technology and Culture, 50(1):232–233, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258569.

Norton:2007:SRJ

[Nor07] Peter D. Norton. Street rivals: Jaywalking and the invention of the Motor Age street. Technology and Culture, 48(2):331– 359, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215409.

Norton:2009:WBW

[Nor09] Peter D. Norton. The World beyond the Windshield: Roads and Landscapes in the United States and Europe (review). Technology and Culture, 50(2):489–491, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263353.

Newman:2007:EUE

[NV07] William M. Newman and Walter G. Vincenti. On an engineering use of engineering history. Technology and Culture, 48(1): 245–247, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209853.

Nye:2004:REF

[Nye04] David E. Nye. Reinventing Eden: The Fate of Nature in Western Culture (review). Technology and Culture, 45(2):413– 414, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55679.

Nye:2005:HBW

[Nye05] David E. Nye. Human-Built World: How to Think about Technology and Culture (review). Technology and Culture, 46(3): 626–627, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186630.

Nye:2006:WSP

[Nye06] David E. Nye. Wonder Shows: Performing Science, Magic, and Religion in America (review). Technology and Culture, 47 (2):421–423, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200869.

Ochs:2001:DDC

[Och01] Kathleen Ochs. Dreams of Development: Colombia’s National School of Mines and Its Engineers, 1887–1970 (review). Technology and Culture, 42(1):163–165, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33680.

Ochs:2002:CBB

[Och02] Kathleen Ochs. Crossing Boundaries, Building Bridges: Comparing the History of Women Engineers, 1870s-1990s (review). Technology and Culture, 43(1):175–177, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33880.

ODonnell:2003:TSR

[O’D03] Brian O’Donnell. Technology and the Spirit (review). Technology and Culture, 44(4):857–858, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49709.

OGorman:2003:CDP

[O’G03] Marcel O’Gorman. Cybering Democracy: Public Space and the Internet (review). Technology and Culture, 44(3):637–638, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
46106.

OGorman:2004:NMR

[O’G04a] Marcel O’Gorman. New Media, 1740–1915 (review). Technology and Culture, 45(1):209–211, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51768.

OGorman:2004:DRC

[O’G04b] Marcel O’Gorman. The Digital Revolution and the Coming of the Postmodern University (review). Technology and Culture, 45(2):473–475, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55680.

OGorman:2005:DPB

[O’G05] Marcel O’Gorman. Digital People: From Bionic Humans to Androids (review). Technology and Culture, 46(2):437–439, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184333.

OGorman:2006:SO

[O’G06] Marcel O’Gorman. Swimming to obsolescence. Technology and Culture, 47(1):164–169, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196061.

OGorman:2008:IMH

[O’G08] Marcel O’Gorman. Imitating machines: Humanities research for a culture of data. Technology and Culture, 49(2):459– 461, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236971.

OHar:2000:MMA

[O’H00] George M. O’Har. Magic in the Machine Age. Technology and Culture, 41(4):862–864, October 2000. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33633.

OHar:2001:NML

[O’H01a] George M. O’Har. The Nuclear Muse: Literature, Physics, and the First Atomic Bombs (review). Technology and Culture, 42 (2):377–378, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33730.

OHar:2001:TCH

[O’H01b] George M. O’Har. Time, Conflict, and Human Values (review). Technology and Culture, 42(1):185–187, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33681.

OHar:2003:VPH

[O’H03] George O’Har. Veblen in Perspective: His Life and Thought (review). Technology and Culture, 44(1):165–167, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/40178.

OHar:2004:TD

[O’H04] George O’Har. Technology and its discontents. Technology and Culture, 45(2):479–485, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55681.

OHar:2005:HGS

[O’H05] George O’Har. The Holodeck in the Garden: Science and Technology in Contemporary American Fiction (review). Technology and Culture, 46(4):864–865, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192312.

OHar:2006:WBR

[O’H06] George O’Har. Where the buffalo roam. Technology and Culture, 47(1):156–163, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/196062.

OHar:2007:WHT

[O’H07] George O’Har. The Worst Hard Time: The Untold Story of Those Who Survived the Great American Dust Bowl (review). Technology and Culture, 48(4):872–873, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223372.

OHar:2009:TLN

[O’H09] George O’Har. Trying Leviathan: The Nineteenth-Century New York Court Case that Put the Whale on Trial and Challenged the Order of Nature (review). Technology and Culture, 50(1): 251–252, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/258588.

Oldenziel:2006:SSH

[Old06] Ruth Oldenziel. Signifying semantics for a history of technology. Technology and Culture, 47(3):477–485, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201472.

Olivier:2007:GEM

[Oli07] M. Olivier. George Eastman’s modern stone-age family: Snapshot photography and the Brownie. Technology and Culture, 48 (1):1–19, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209854.

Olsen:2007:HGW

[Ols07] Jonathan Olsen. How Green Were the Nazis? Nature, Environment, and Nation in the Third Reich (review). Technology and Culture, 48(1):207–208, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209855.

Olwell:2001:SBB

[Olw01] Russell B. Olwell. In the Shadow of the Bomb: Bethe, Oppenheimer, and the Moral Responsibility of the Scientist (review). Technology and Culture, 42(4):823–825, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL http://muse.jhu.edu/journals/technology_

and_culture/v042/42.4olwell.html;   https://muse.jhu.
edu/pub/1/article/33834.

Olwell:2004:TNA

[Olw04] Russell B. Olwell. Toward Nuclear Abolition: A History of the World Nuclear Disarmament Movement, 1971–Present (review). Technology and Culture, 45(4):868–869, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175770.

Olwell:2009:PWR

[Olw09] Russell Olwell. Poison in the Well: Radioactive Waste in the Oceans at the Dawn of the Nuclear Age (review). Technology and Culture, 50(1):246–247, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258577.

Orland:2002:GKR

[Orl02] Barbara Orland. Geschichte der Konsumgesellschaft (review). Technology and Culture, 43(1):165–167, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33882.

Orr:2002:DFL

[Orr02] Patricia Orr. Double Fold: Libraries and the Assault on Paper (review). Technology and Culture, 43(4):835–837, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34032.

Orvell:2009:FRG

[Orv09] Miles Orvell. Faux Real: Genuine Leather and 200 Years of Inspired Fakes (review). Technology and Culture, 50(1):260–261, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258585.

Olwell:2005:KSD

[OS05] Russell B. Olwell and Maureen Sullivan. Knocking at the schoolhouse door: The SHOT/AHA booklet series historical perspectives on technology, culture, and society. Technology and Culture, 46(2):400–404, April 2005. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184334.

Osietzki:2000:EGE

[Osi00] Maria Osietzki. Elektrizitat in der Geistesgeschichte, and: Elektrifizierung als Vision: Zur Symbolgeschichte einer Technik im Alltag (review). Technology and Culture, 41(3):603– 604, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33585.

Oslund:2004:UFE

[Osl04] Karen Oslund. The Unending Frontier: An Environmental History of the Early Modern World (review). Technology and Culture, 45(4):861–862, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175771.

Oslund:2007:BMM

[Osl07] Karen Oslund. Beyond men and machines: New contributions to the history of polar exploration. Technology and Culture, 48 (3):594–597, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218721.

Otsubo:2003:RMC

[Ots03] Sumiko Otsubo. Reconfiguring Modernity: Concepts of Nature in Japanese Political Ideology (review). Technology and Culture, 44(2):432–433, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42035.

Otter:2008:ENA

[Ott08] A. A. Den Otter. Encyclopedia of North American Railroads (review). Technology and Culture, 49(2):472–474, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
236997.

Owens:2003:BHM

[Owe03] Larry Owens. Between Human and Machine: Feedback, Control, and Computing before Cybernetics (review). Technology and Culture, 44(4):843–844, October 2003. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49710.

Owens:2004:EPC

[Owe04] Larry Owens. Engineering the perfect cup of coffee: Samuel prescott and the sanitary vision at MIT. Technology and Culture, 45(4):795–807, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/175772.

Pacey:2002:WKN

[Pac02] Arnold Pacey. Ways of Knowing: A New History of Science, Technology and Medicine (review). Technology and Culture, 43 (3):653–654, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33984.

Palmer:2000:PPS

[Pal00] Scott W. Palmer. Peasants into pilots: Soviet air-mindedness as an ideology of dominance. Technology and Culture, 41(1):1– 26, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33491.

Palmer:2002:ZGA

[Pal02] Scott W. Palmer. Zeppelin! Germany and the Airship, 1900– 1939 (review). Technology and Culture, 43(4):816–817, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34033.

Pannabecker:2000:MMN

[Pan00] John R. Pannabecker. Mine et metal, 1780–1880: Les nonferreux et l’industrialisation (review). Technology and Culture, 41(3):585–586, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33586.

Pannabecker:2002:SIL

[Pan02a] John R. Pannabecker. School for industry: L’École d’Arts et Metiers of Chalons-sur-Marne under Napoleon and the Restoration. Technology and Culture, 43(2):254–290, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33928.

Panzer:2002:TVD

[Pan02b] Mary Panzer. Tender Violence: Domestic Visions in an Age of U.S. Imperialism (review). Technology and Culture, 43(1): 210–212, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33883.

Pannabecker:2003:LAS

[Pan03a] John R. Pannabecker. L’Etat, l’armèe, la science: L’Invention de la recherche publique en France (1763–1830) (review). Technology and Culture, 44(4):809–810, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49711.

Pannabecker:2003:LFA

[Pan03b] John R. Pannabecker. L’industrie française au milieu du 19e siècle: Les enquêtes de la Statistique générale de la France (review). Technology and Culture, 44(2):392–394, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42036.

Pannabecker:2005:MEP

[Pan05a] John R. Pannabecker. The myth of the ecole polytechnique. Technology and Culture, 46(3):618–622, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186631.

Panzer:2005:IHC

[Pan05b] Mary Panzer. Invention of Hysteria: Charcot and the Photographic Iconography of the Salpetriere (review). Technology and Culture, 46(3):688–690, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186632.

Panzer:2006:PPP

[Pan06] Mary Panzer. Picturing Poverty: Print Culture and FSA Photographs (review). Technology and Culture, 47(1):203–205, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196063.

Pannabecker:2007:PTF

[Pan07] John R. Pannabecker. The Path Not Taken: French Industrialization in the Age of Revolution, 1750–1830 (review). Technology and Culture, 48(4):852–854, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223374.

Pannabecker:2009:CRI

[Pan09] John R. Pannabecker. Le choix de la révolution industrielle: Les enterprises de Marc Seguin et ses frères (1815– 1835) (review). Technology and Culture, 50(2):459–461, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263362.

Parr:2002:PEE

[Par02a] Joy Parr. The Perception of the Environment: Essays in Livelihood, Dwelling, and Skill (review). Technology and Culture, 43 (2):401–402, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33929.

Parr:2002:EIM

[Par02b] Joy Parr. Editor’s introduction: Modern kitchen, good home, strong nation. Technology and Culture, 43(4):657–667, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34034.

Parr:2005:RSC

[Par05] Joy Parr. Ruth Schwartz Cowan, more work for mother. Technology and Culture, 46(3):604–612, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186633.

Parthasarathy:2008:STP

[Par08] Shobita Parthasarathy. Science and Technology Policy in the United States: Open Systems in Action (review). Technology and Culture, 49(3):819–820, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243090.

Parr:2009:RER

[Par09] Joy Parr. Rethinking Expertise (review). Technology and Culture, 50(2):507–508, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/263366.

Pasqualetti:2002:TLR

[Pas02] Martin J. Pasqualetti. Technologies of Landscape: From Reaping to Recycling (review). Technology and Culture, 43(1): 161–163, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33884.

Pasqualetti:2006:CSH

[Pas06] Martin J. Pasqualetti. Children of the Sun: A History of Humanity’s Unappeasable Appetite for Energy (review). Technology and Culture, 47(3):638–639, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201473.

Patnode:2003:WPN

[Pat03] Randall Patnode. “What these people need is radio”: New technology, the Press, and otherness in 1920s America. Technology and Culture, 44(2):285–305, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42037.

Pavri:2009:WYS

[Pav09] Vera Pavri. What you say is what you get: Policy discourse and the regulation of Canada’s first domestic communications satellite system. Technology and Culture, 50(3):569–585, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270570.

Payne:2005:BRM

[Pay05] Phillip Payne. Beyond the Ruins: The Meanings of Deindustrialization (review). Technology and Culture, 46(2):444– 446, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184336.

Pinch:2003:SOA

[PB03] T. J. Pinch and Karin Bijsterveld. “Should one applaud?” breaches and boundaries in the reception of new technology in music. Technology and Culture, 44(3):536–559, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46107.

Perkins:2000:SSH

[Per00] Edwin J. Perkins. The Speed of Sound: Hollywood and the Talkie Revolution, 1926–1930 (review). Technology and Culture, 41(2):384–385, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33536.

Perkins:2002:EEE

[Per02] Edwin J. Perkins. Engines of Enterprise: An Economic History of New England (review). Technology and Culture, 43(2):426– 428, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33930.

Perdue:2006:JNP

[Per06a] Peter C. Perdue. Joseph Needham’s problematic legacy. Technology and Culture, 47(1):175–178, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196065.

Perkins:2006:FAW

[Per06b] John H. Perkins. The Fire Ant Wars: Nature, Science, and Public Policy in Twentieth-Century America (review). Technology and Culture, 47(2):451–452, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/200870.

Perdue:2007:MZI

[Per07] Peter C. Perdue. The Merchants of Zigong: Industrial Entrepreneurship in Early Modern China (review). Technology and Culture, 48(3):626–627, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218723.

Petzold:2000:TTK

[Pet00] Hartmut Petzold. Der “Transistor” als technisches und kulturelles Phanomen: Die Transistorisierung der Radio- und Fernsehempfanger in der deutschen Rundfunkindustrie 1955 bis 1965 (review). Technology and Culture, 41(2):387–388, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33537.

Peters:2001:KCG

[Pet01a] Tom F. Peters. Karl Culmann und die graphische Statik: Anhang mit umfangreichen Culmann-Texten (review). Technology and Culture, 42(1):162–163, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33683.

Petroski:2001:SF

[Pet01b] Henry Petroski. The success of failure. Technology and Culture, 42(2):321–328, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33731.

Petzold:2001:TFG

[Pet01c] Hartmut Petzold. Die “Tonenden Funken”: Geschichte eines fruhen drahtlosen Kommunikationssystems, 1905–1914 (review). Technology and Culture, 42(3):595–596, July 2001. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33785.

Peters:2003:BCH

[Pet03] Tom F. Peters. Building in China: Henry K. Murphy’s “Adaptive Architecture,” 1914–1935 (review). Technology and Culture, 44(1):207–209, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/40179.

Peters:2004:GBR

[Pet04a] Tom F. Peters. Geschichte der Baustatik (review). Technology and Culture, 45(3):664–666, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/172999.

Peters:2004:ASD

[Pet04b] Tom F. Peters. The Art of Structural Design: A Swiss Legacy (review). Technology and Culture, 45(4):851–852, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175773.

Peters:2004:WTB

[Pet04c] Tom F. Peters. Wilhelm von Traitteur: Ein badischer Baumeister als Neuerer in der russischen Architektur 1814–1832 (review). Technology and Culture, 45(2):426–428, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/55682.

Petrina:2004:SPA

[Pet04d] Stephen Petrina. Sidney Pressey and the automation of education, 1924–1934. Technology and Culture, 45(2):305–330, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55683.

Peters:2005:EHM

[Pet05] Tom F. Peters. Essays on the History of Mechanics: In Memory of Clifford Ambrose Truesdell and Edoardo Benvenuto (review). Technology and Culture, 46(4):875–876, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/192313.

Peters:2007:HCA

[Pet07a] Tom F. Peters. Hardy Cross: American Engineer (review). Technology and Culture, 48(4):882–884, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223375.

Petrie:2007:CTC

[Pet07b] Ian Petrie. Creating the Twentieth Century: Technical Innovations of 1867–1914, and: Culture and Technology: A Primer (review). Technology and Culture, 48(3):639–641, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218724.

Peters:2009:ABM

[Pet09] Tom F. Peters. The art of building in its many forms. Technology and Culture, 50(3):669–672, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270531.

Pfaffenberger:2004:GCI

[Pfa04] Bryan Pfaffenberger. Global Control: Information Technology and Globalization since 1845 (review). Technology and Culture, 45(1):214–216, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51769.

Pfaffenberger:2006:EDC

[Pfa06] Bryan Pfaffenberger. Electric Dreams: Computers in American Culture (review). Technology and Culture, 47(4):874–875, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/204998.

Phillips:2003:GMM

[Phi03] Gervase Phillips. Guns and Men in Medieval Europe, 1200– 1500 (review). Technology and Culture, 44(4):803–805, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
49712.

Phillips:2004:WEP

[Phi04] Jonathan F. Phillips. Warfare as enterprise: Peter W. Singer, Corporate Warriors. Technology and Culture, 45(3):608–612, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/173000.

Pickering:2000:HWB

[Pic00] Andrew Pickering. How We Became Posthuman: Virtual Bodies in Cybernetics, Literature, and Informatics, and: Silicon Second Nature: Culturing Artificial Life in a Digital World (review). Technology and Culture, 41(2):392–395, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33538.

Picon:2001:RFN

[Pic01] Antoine Picon. The Radiance of France: Nuclear Power and National Identity after World War II (review). Technology and Culture, 42(1):140–141, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33684.

Pickering:2006:AAA

[Pic06] Andrew Pickering. Archigram: Architecture without Architecture (review). Technology and Culture, 47(3):661–663, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201474.

Pietrykowski:2006:RMA

[Pie06] Bruce Pietrykowski. Recasting the Machine Age: Henry Ford’s Village Industries (review). Technology and Culture, 47(4): 855–857, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/204999.

Pinkus:2008:CCL

[Pin08] Karen Pinkus. On climate, cars, and literary theory. Technology and Culture, 49(4):1002–1009, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255446.

Platt:2006:CCE

[Pla06] Harold L. Platt. City, Country, Empire: Landscapes in Environmental History (review). Technology and Culture, 47(4): 862–864, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/205000.

Pohlad:2000:PCC

[Poh00] Mark B. Pohlad. Photography and Its Critics: A Cultural History, 1839–1900 (review). Technology and Culture, 41(2):382– 384, April 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33539.

Pohlad:2007:CSA

[Poh07] Mark B. Pohlad. Charles Sheeler: Across Media (review). Technology and Culture, 48(2):370–376, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215410.

Poh:2008:CGG

[Poh08] Carol Poh. Conquering Gotham: A Gilded Age Epic: The Construction of Penn Station and Its Tunnels (review). Technology and Culture, 49(4):1065–1067, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255508.

Poitras:2008:QES

[Poi08] Claire Poitras. Questions d’égouts: Santé publique, infrastructures et urbanisation à Montréal au XIXe siècle (review). Technology and Culture, 49(1):282–284, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230923.

Poos:2003:DMB

[Poo03] Lawrence R. Poos. At the Dawn of Modernity: Biology, Culture, and Material Life in Europe after the Year 1000 (review). Technology and Culture, 44(3):597–599, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46108.

Pope:2000:LKN

[Pop00] Daniel Pope. Licensed to Kill?: The Nuclear Regulatory Commission and the Shoreham Power Plant (review). Technology and Culture, 41(1):156–158, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33492.

Pope:2001:GEN

[Pop01a] Norris Pope. The Great Exhibition of 1851: A Nation on Display (review). Technology and Culture, 42(2):341–343, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33732.

Pope:2001:DSI

[Pop01b] Norris Pope. Dickens’s “The Signalman” and information problems in the Railway Age. Technology and Culture, 42(3):436– 461, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33786.

Pope:2002:UKV

[Pop02] Norris Pope. Useful Knowledge: The Victorians, Morality, and the March of Intellect (review). Technology and Culture, 43(4): 789–791, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34035.

Pope:2007:EAS

[Pop07] Norris Pope. The Economics of Attention: Style and Substance in the Age of Information (review). Technology and Culture, 48 (3):673–675, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218725.

Portuondo:2003:PFS

[Por03] Maria M. Portuondo. Plantation factories: Science and technology in Late-Eighteenth-Century Cuba. Technology and Culture, 44(2):231–257, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42038.

Porter:2004:OMP

[Por04] Glenn Porter. Ornament: A Modern Perspective (review). Technology and Culture, 45(2):418–420, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/55684.

Portuondo:2005:OMA

[Por05] Maria M. Portuondo. Una Obra maestra: El acueducto Albear de La Habana (review). Technology and Culture, 46(1):203–205, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179777.

Post:2001:VSR

[Pos01a] Robert C. Post. “A very special relationship”: SHOT and the Smithsonian’s Museum of History and Technology. Technology and Culture, 42(3):401–435, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33787.

Post:2001:CNA

[Pos01b] Robert C. Post. A corner of the Nation’s attic. Technology and Culture, 42(3):519–522, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33788.

Post:2003:LSR

[Pos03] Robert C. Post. “The Last Steam Railroad in America”: Shaffers Crossing, Roanoke, Virginia, 1958. Technology and Culture, 44(3):560–565, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46109.

Post:2004:NOT

[Pos04] Robert C. Post. A narrative for our time: The enola gay “and after that, period”. Technology and Culture, 45(2):373– 395, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55685.

Post:2005:HKT

[Pos05] Robert C. Post. Henry Kaiser, troy ruttman, and madman muntz: Three originals. Technology and Culture, 46(4):773– 778, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192314.

Post:2006:MGD

[Pos06] Robert C. Post. The machine in the garden district. Technology and Culture, 47(1):91–94, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196066.

Post:2007:TE

[Pos07] Robert C. Post. Two enthusiasts. Technology and Culture, 48(4):804–809, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/223376.

Post:2008:BMS

[Pos08a] Robert C. Post. The bridge at Mackinac Straits: Another fiftieth anniversary. Technology and Culture, 49(3):752–763, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243103.

Post:2008:SB

[Pos08b] Robert C. Post. Silvio A. Bedini, 1917–2007. Technology and Culture, 49(2):522–529, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/237006.

Post:2009:CCP

[Pos09a] Robert C. Post. Chance and contingency: Putting Mel Kranzberg in context. Technology and Culture, 50(4):839–872, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363416.

Post:2009:LSR

[Pos09b] Robert C. Post. The land speed record and the last green monster. Technology and Culture, 50(3):586–593, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270544.

Powaski:2002:PAB

[Pow02] Ronald E. Powaski. Planning Armageddon: Britain, the United States, and the Command of Western Nuclear Forces, 1945– 1964 (review). Technology and Culture, 43(1):207–209, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33885.

Powers:2004:EIB

[Pow04] Patricia E. Powers. Ethical Issues in Biotechnology (review). Technology and Culture, 45(1):221–222, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51770.

Powers:2009:ALT

[Pow09] John C. Powers. Andreas Libavius and the Transformation of Alchemy: Separating Chemical Cultures with Polemical Fire (review). Technology and Culture, 50(2):461–462, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263369.

Puig-Pla:2003:OEN

[PP03] Carles Puig-Pla. Orígenes de la enseñanza técnica en Alcoy (review). Technology and Culture, 44(4):814–815, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49714.

Puig-Pla:2009:EEI

[PP09] Carles Puig-Pla. La Escuela Especial de Ingenieros Industriales de Bilbao, 1897–1936: Educación y Tecnología en el Primer Tercio del Siglo XX (review). Technology and Culture, 50(3):691–692, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270542.

Pratt:2001:PLO

[Pra01] Joseph A. Pratt. Power Loss: The Origins of Deregulation and Restructuring in the American Electric Utility System (review). Technology and Culture, 42(3):598–600, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33789.

Pratt:2003:REP

[Pra03] Joseph A. Pratt. Redefining Efficiency: Pollution Concerns, Regulatory Mechanisms, and Technological Change in the U.S. Petroleum Industry (review). Technology and Culture, 44(1): 221–223, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40180.

Pretzer:2000:MCE

[Pre00] William S. Pretzer. Museums in Central Europe: A traveler’s introduction. Technology and Culture, 41(3):516, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33587.

Pretzer:2001:LEE

[Pre01] William S. Pretzer. The Languages of Edison’s Electric Light (review). Technology and Culture, 42(1):171–173, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33685.

Pretzer:2002:SGW

[Pre02] William S. Pretzer. Scripts, Grooves, and Writing Machines: Representing Technology in the Edison Era (review). Technology and Culture, 43(1):191–193, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33886.

Pretzer:2003:IMV

[Pre03] William S. Pretzer. Ideas, Machines and Values: An Introduction to Science, Technology, and Society Studies, and: Visions of STS: Counterpoints in Science, Technology, and Society Studies (review). Technology and Culture, 44(4):799–802, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49713.

Prentice:2004:FCB

[Pre04a] Rachel Prentice. Figurations: Child, Bodies, Worlds (review). Technology and Culture, 45(1):224–225, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51771.

Pretzer:2004:CRP

[Pre04b] William S. Pretzer. A Consumers’ Republic: The Politics of Mass Consumption (review). Technology and Culture, 45(2): 454–455, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55686.

Pretzer:2004:BIA

[Pre04c] William S. Pretzer. Black Inventors in the Age of Segregation: Granville T. Woods, Lewis H. Latimer, and Shelby J. Davidson (review). Technology and Culture, 45(4):848–849, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175774.

Pretzer:2004:ELB

[Pre04d] William S. Pretzer. Edison’s last breath. Technology and Culture, 45(3):679–682, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/173001.

Pretzer:2005:PTE

[Pre05] William S. Pretzer. The Papers of Thomas A. Edison. Vol. 5: Research to Development at Menlo Park, January 1879– March 1881 (review). Technology and Culture, 46(4):868–870, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192315.

Pretzer:2006:RCS

[Pre06] William S. Pretzer. Recasting a Craft: St. Louis Typefounders Respond to Industrialization (review). Technology and Culture, 47(4):844–845, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/205001.

Pretzer:2008:PTE

[Pre08a] William S. Pretzer. The Papers of Thomas A. Edison, vol. 6: Electrifying New York and Abroad, April 1881–March 1883 (review). Technology and Culture, 49(4):1064–1065, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
255492.

Pretzer:2008:WMP

[Pre08b] William S. Pretzer. The Wizard of Menlo Park: How Thomas Alva Edison Invented the Modern World (review). Technology and Culture, 49(2):478–479, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/236998.

Pritchard:2000:EEB

[Pri00] Michael S. Pritchard. Engineering Ethics: Balancing Cost, Schedule, and Risk–Lessons Learned from the Space Shuttle (review). Technology and Culture, 41(1):164–166, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33493.

Pritchard:2002:MEF

[Pri02] Sara B. Pritchard. The Modernist Enterprise: French Elites and the Threat of Modernity, 1900–1940 (review). Technology and Culture, 43(2):421–423, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33931.

Priest:2003:HAE

[Pri03] Tyler Priest. High-Altitude Energy: A History of Fossil Fuels in Colorado (review). Technology and Culture, 44(3):622– 623, July 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/46110.

Pritchard:2004:TDF

[Pri04] Sara B. Pritchard. La trace du fleuve: La Seine et Paris, 1750–1850 (review). Technology and Culture, 45(4):876–877, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175775.

Pursell:2008:SOT

[Pur08] Carroll W. Pursell. The Shock of the Old: Technology and Global History since 1900 (review). Technology and Culture, 49(1):237–238, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/230924.

Pyenson:2001:QGH

[Pye01] Lewis Pyenson. Quantum Generations: A History of Physics in the Twentieth Century (review). Technology and Culture, 42 (2):373–375, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33733.

Pyne:2009:AOW

[Pyn09] Stephen J. Pyne. America’s Ocean Wilderness: A Cultural History of Twentieth-Century Exploration (review). Technology and Culture, 50(3):707–708, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270539.

Quivik:2001:SSN

[Qui01] Fred Quivik. Smelter Smoke in North America: The Politics of Transborder Pollution (review). Technology and Culture, 42(1): 151–152, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33686.

Quivik:2003:CWH

[Qui03a] Fred Quivik. The Chimney of the World: A History of Smoke Pollution in Victorian and Edwardian Manchester (review). Technology and Culture, 44(3):620–621, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46111.

Quivik:2003:SFE

[Qui03b] Fred Quivik. The Subterranean Forest: Energy Systems and the Industrial Revolution (review). Technology and Culture, 44(1): 216–218, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40181.

Quivik:2005:CHH

[Qui05] Fred Quivik. Coal: A Human History (review). Technology and Culture, 46(4):846–847, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192316.

Quirke:2008:PTT

[Qui08] Viviane Quirke. Penicillin: Triumph and Tragedy (review). Technology and Culture, 49(3):829–831, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/243111.

Raber:2004:RAR

[Rab04] Michael Scott Raber. Rum and Axes: The Rise of a Connecticut Merchant Family, 1795–1850 (review). Technology and Culture, 45(1):185–186, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51772.

Ramaswamy:2005:RAM

[Ram05] Vijaya Ramaswamy. Reproductive Agency, Medicine, and the State (review). Technology and Culture, 46(3):692–693, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186634.

Randall:2001:MMM

[Ran01] Adrian Randall. Merchants, Markets, and Manufacture: The English Wool Textile Industry in the Eighteenth Century (review). Technology and Culture, 42(1):194–195, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33687.

Raphals:2001:TTT

[Rap01] Lisa Ann Raphals. Ten Thousand Things: Module and Mass Production in Chinese Art (review). Technology and Culture, 42(4):793–795, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33835.

Rasmussen:2001:CRM

[Ras01] Chris Rasmussen. A Century of Recorded Music: Listening to Musical History (review). Technology and Culture, 42(4): 803–805, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33836.

Rau:2001:SCW

[Rau01] Erik P. Rau. Science, Cold War, and the American State: Lloyd V. Berkner and the Balance of Professional Ideals (review). Technology and Culture, 42(4):826–828, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33837.

Rau:2003:SMP

[Rau03] Erik P. Rau. Science, Money, and Politics: Political Triumph and Ethical Erosion (review). Technology and Culture, 44(1): 199–201, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40182.

Rau:2005:BPW

[Rau05] Erik P. Rau. Blackett: Physics, War, and Politics in the Twentieth Century (review). Technology and Culture, 46(3):655– 657, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186635.

Ravina:2001:JPE

[Rav01] Mark Ravina. Japan’s Protoindustrial Elite: The Economic Foundation of the Gono (review). Technology and Culture, 42 (4):834–836, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33838.

Rees:2005:LHT

[Ree05a] Jonathan Rees. Linde: History of a Technology Corporation, 1879–2004 (review). Technology and Culture, 46(4):849–850, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192317.

Rees:2005:DKL

[Ree05b] Jonathan Rees. “I did not know … any danger was attached”: Safety consciousness in the early American ice and refrigeration industries. Technology and Culture, 46(3):541–560, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186599.

Reinhardt:2005:GSC

[Rei05] Anne Reinhardt. Gutenberg in Shanghai: Chinese Print Capitalism, 1876–1937 (review). Technology and Culture, 46(2): 411–413, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184337.

Rentetzi:2009:TMC

[Ren09] Maria Rentetzi. The Tobacco Museum of the city of Kavala. Technology and Culture, 50(3):649–657, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270533.

Reuss:2002:LDT

[Reu02] Martin Reuss. Learning from the Dutch: Technology, management, and water resources development. Technology and Culture, 43(3):465–472, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33985.

Reuss:2005:ICR

[Reu05] Martin Reuss. Inventing the Charles River (review). Technology and Culture, 46(1):202–203, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179778.

Reuss:2006:SSB

[Reu06] Martin Reuss. Searching for Sophocles on Bourbon Street. Technology and Culture, 47(2):349–356, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/200871.

Reuss:2008:SLE

[Reu08] Martin Reuss. Seeing like an engineer: Water projects and the mediation of the incommensurable. Technology and Culture, 49 (3):531–546, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243125.

Reynolds:2001:GKN

[Rey01a] Terry S. Reynolds. The Great Kanawha Navigation (review). Technology and Culture, 42(2):364–366, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33734.

Reynolds:2001:BBV

[Rey01b] Terry S. Reynolds. On not burning bridges: Valuing the passe. Technology and Culture, 42(3):523–530, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33790.

Reynard:2002:MMC

[Rey02a] Pierre-Claude Reynard. Merchants and Marvels: Commerce, Science, and Art in Early Modern Europe (review). Technology and Culture, 43(3):591–593, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33986.

Reynard:2002:POP

[Rey02b] Pierre-Claude Reynard. Public order and privilege: Eighteenth-Century French roots of environmental regulation. Technology and Culture, 43(1):1–28, January 2002. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33887.

Reynard:2005:CEF

[Rey05] Pierre-Claude Reynard. Conserving the Enlightenment: French Military Engineering from Vauban to the Revolution (review). Technology and Culture, 46(1):218–220, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179779.

Reynolds:2007:IFG

[Rey07] Terry S. Reynolds. Infrastructure: A Field Guide to the Industrial Landscape (review). Technology and Culture, 48(3):633– 635, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218726.

Richards:2008:SGB

[Ric08a] Stephen Richards. The SS Great Britain (review). Technology and Culture, 49(1):127–132, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230925.

Richelson:2008:SSW

[Ric08b] Jeffrey T. Richelson. Spying on Science: Western Intelligence in Divided Germany, 1945–1961 (review). Technology and Culture, 49(2):486–487, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/236999.

Rieger:2005:LKH

[Rie05] Bernhard Rieger. Luxus und Konsum: Eine historische Annaherung (review). Technology and Culture, 46(4):817–818, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192318.

Riis:2003:MME

[Rii03] Thomas Laurence Riis. Manufacturing the Muse: Estey Organs and Consumer Culture in Victorian America (review). Technology and Culture, 44(4):828–830, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49715.

Ritvo:2007:OHI

[Rit07] Harriet Ritvo. The Ocean at Home: An Illustrated History of the Aquarium (review). Technology and Culture, 48(1):186–188, January 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/209856.

Russell:2008:MLA

[RK08] Edmund Russell and Jennifer Kane. The missing link: Assessing the reliability of Internet citations in history journals. Technology and Culture, 49(2):420–429, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/236966.

Robb:2000:MIB

[Rob00] Francie Robb. Making Iron on the Bald Eagle: Roland Curtin’s Ironworks and Workers’ Community (review). Technology and Culture, 41(4):798–799, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33634.

Robertson:2002:LTI

[Rob02] David Robertson. A Landscape Transformed: The Ironmaking District of Salisbury, Connecticut (review). Technology and Culture, 43(3):607–609, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33987.

Roberts:2004:AAI

[Rob04] Lissa Louise Roberts. An arcadian apparatus: The introduction of the steam engine into the Dutch landscape. Technology and Culture, 45(2):251–276, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55687.

Robertson:2005:AAP

[Rob05a] Frances Robertson. The aesthetics of authenticity: Printed banknotes as industrial currency. Technology and Culture, 46 (1):31–50, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179780.

Robinson:2005:SSR

[Rob05b] Michael F. Robinson. Science Serialized: Representations of the Sciences in Nineteenth-Century Periodicals (review). Technology and Culture, 46(2):413–415, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184338.

Roberts:2006:MTD

[Rob06a] Lissa Louise Roberts. Models: The Third Dimension of Science (review). Technology and Culture, 47(2):468–470, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200872.

Robertson:2006:NEP

[Rob06b] Frances Robertson. Nature Exposed: Photography as Eyewitness in Victorian Science (review). Technology and Culture, 47(4):837–839, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/205002.

Robertson:2006:WBA

[Rob06c] Frances Robertson. Walter Benjamin and Art (review). Technology and Culture, 47(3):698–699, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/201475.

Roberts:2007:CFS

[Rob07] Lissa Louise Roberts. The content of the form: Sarah Lowengard’s The Creation of Color in Eighteenth-Century Europe. Technology and Culture, 48(4):831–835, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/223377.

Roberts:2008:NBD

[Rob08a] Jody Roberts. Nylon and Bombs: DuPont and the March of Modern America (review). Technology and Culture, 49(2):493– 494, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/237000.

Robinson:2008:TYW

[Rob08b] Eric H. Robinson. A thousand years of Western technology in one volume: Is all-inclusive ever conclusive? Technology and Culture, 49(1):215–229, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230926.

Rockman:2002:MHC

[Roc02] Seth Rockman. Making Houses, Crafting Capitalism: Builders in Philadelphia, 1790–1850 (review). Technology and Culture, 43(3):609–610, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33988.

Roland:2000:HGF

[Rol00a] Alex Roland. A History of Greek Fire and Gunpowder (review). Technology and Culture, 41(1):131–132, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33494.

Roland:2000:IIR

[Rol00b] Alex Roland. Inventing the Internet (review). Technology and Culture, 41(4):826–828, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33635.

Roland:2001:SSD

[Rol01] Alex Roland. The Space Shuttle Decision: NASA’s Search for a Reusable Space Vehicle (review). Technology and Culture, 42 (2):386–388, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33735.

Roland:2003:OMS

[Rol03] Alex Roland. Once more into the stirrups: Lynn White, Jr., medieval technology and social change. Technology and Culture, 44(3):574–585, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/46112.

Roland:2005:GFP

[Rol05] Alex Roland. Greek Fire, Poison Arrows, and Scorpion Bombs: Biological and Chemical Warfare in the Ancient World (review). Technology and Culture, 46(4):878–879, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
192319.

Roland:2006:DD

[Rol06a] Alex Roland. David dibner, 1927–2005. Technology and Culture, 47(2):472–475, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/200873.

Roland:2006:WCE

[Rol06b] Alex Roland. The worst century ever. Technology and Culture, 47(3):597–606, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201476.

Roland:2007:CC

[Rol07] Alex Roland. Containers and causality. Technology and Culture, 48(2):386–392, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/215411.

Rome:2003:CLI

[Rom03] Adam Ward Rome. City Lights: Illuminating the American Night (review). Technology and Culture, 44(4):827–828, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49716.

Rombes:2008:SCO

[Rom08] Nick Rombes. Stereoscopic Cinema and the Origins of 3-D Film, 1838–1952 (review). Technology and Culture, 49(4):1084– 1085, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255510.

Rosenband:2003:JUN

[Ros03] Leonard N. Rosenband. John U. Nef, the conquest of the material world. Technology and Culture, 44(2):364–370, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42039.

Rosenband:2005:NJB

[Ros05a] Leonard N. Rosenband. Never just business. Technology and Culture, 46(1):168–176, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179781.

Rosner:2005:TVH

[Ros05b] Lisa Rosner. Technological Visions: The Hopes and Fears that Shape New Technologies (review). Technology and Culture, 46 (3):625–626, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186636.

Rosenthal:2006:BMA

[Ros06] Joel Thomas Rosenthal. Beer in the Middle Ages and the Renaissance (review). Technology and Culture, 47(1):182–184, January 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/196067.

Rosen:2007:CYT

[Ros07] Christine Meisner Rosen. Cash for Your Trash: Scrap Recycling in America (review). Technology and Culture, 48(2):426– 427, April 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/215412.

Rothenberg:2001:PTE

[Rot01] Marc Rothenberg. The Papers of Thomas A. Edison, vol. 4, The Wizard of Menlo Park, 1878 (review). Technology and Culture, 42(3):600–601, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33791.

Roth:2002:MJT

[Rot02] Matthew W. Roth. The Museum of Jurassic Technology, Culver City, California. Technology and Culture, 43(1):102–109, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33888.

Roth:2004:EAC

[Rot04a] Matthew W. Roth. Everyday America: Cultural Landscape Studies after J. B. Jackson (review). Technology and Culture, 45(4):859–860, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175776.

Rothenberg:2004:LHA

[Rot04b] Marc Rothenberg. Latitude: How American Astronomers Solved the Mystery of Variation (review). Technology and Culture, 45(2):439–440, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/55688.

Rothschild:2004:GTR

[Rot04c] Joan Rothschild. Gender and Technology: A Reader (review). Technology and Culture, 45(4):908–910, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175777.

Rothenberg:2005:NR

[Rot05] Marc Rothenberg. Nathan Reingold, 1927–2004. Technology and Culture, 46(2):479–483, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184339.

Roth:2008:BTS

[Rot08] Matthew W. Roth. The Best Transportation System in the World: Railroads, Trucks, Airlines, and American Public Policy in the Twentieth Century (review). Technology and Culture, 49(3):799–801, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243099.

Roush:2002:LHR

[Rou02] Wade Roush. Lords of the Harvest (review). Technology and Culture, 43(2):450–451, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33932.

Roush:2005:CMP

[Rou05] Wade Roush. The Creation of the Media: Political Origins of Modern Communications (review). Technology and Culture, 46 (2):417–418, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184340.

Rozwadowski:2009:UHS

[Roz09] Helen Rozwadowski. The Unnatural History of the Sea (review). Technology and Culture, 50(1):247–249, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258591.

Rymsza-Pawlowska:2009:SYS

[RP09] Malgorzata Rymsza-Pawlowska. Shivers Down Your Spine: Cinema, Museums, and the Immersive View (review). Technology and Culture, 50(4):951–952, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363441.

Rudolph:2009:RBE

[Rud09] John Rudolph. The Race between Education and Technology (review). Technology and Culture, 50(4):970–971, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363452.

Russell:2000:ASA

[Rus00] Edmund Russell. American Science in an Age of Anxiety: Scientists, Anticommunism, and the Cold War (review). Technology and Culture, 41(1):161–162, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33495.

Russell:2007:HHH

[Rus07] Edmund Russell. Hunters, Herders, and Hamburgers: The Past and Future of Human–Animal Relationships (review). Technology and Culture, 48(1):182–184, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/209857.

Russo:2008:RAG

[Rus08] Alexander Russo. Radio’s America: The Great Depression and the Rise of Modern Mass Culture (review). Technology and Culture, 49(4):1079–1080, October 2008. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255511.

Russell:2009:TBC

[Rus09] Andrew L. Russell. Two Bits: The Cultural Significance of Free Software (review). Technology and Culture, 50(4):964–966, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363449.

Raven:2004:DSL

[RV04] Rob Raven and Geert Verbong. Dung, sludge, and landfill: Biogas technology in the Netherlands, 1970–2000. Technology and Culture, 45(3):519–539, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/173002.

Ryden:2001:CAI

[Ryd01] Goran Ryden. Connecticut’s Ames Iron Works: Family, Community, Nature, and Innovation in an Enterprise of the Early American Republic (review). Technology and Culture, 42(2): 349–350, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33736.

Rydell:2003:DCH

[Ryd03] Robert W. Rydell. Designing the Centennial: A History of the 1876 International Exhibition in Philadelphia (review). Technology and Culture, 44(2):407–409, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/42040.

Sadowski:2006:BDB

[Sad06] Thomas J. Sadowski. Big Dish: Building America’s Deep Space Connection to the Planets (review). Technology and Culture, 47(3):643–644, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201477.

Saliba:2001:SMI

[Sal01] George Saliba. Studies in Medieval Islamic Technology (review). Technology and Culture, 42(1):189–190, January 2001.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33688.

Saliba:2005:ESI

[Sal05] George Saliba. The Enterprise of Science in Islam: New Perspectives (review). Technology and Culture, 46(1):251–252, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179782.

Samson:2005:ZAM

[Sam05] M. David Samson. Zoomscape: Architecture in Motion and Media (review). Technology and Culture, 46(4):862–864, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
192320.

Samuels:2007:SWW

[Sam07] Richard J. Samuels. Secret Weapons and World War II: Japan in the Shadow of Big Science (review). Technology and Culture, 48(2):460–462, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/215413.

Sanabria:2000:SCC

[San00] Sergio Sanabria. The Sun in the Church: Cathedrals as Solar Observatories (review). Technology and Culture, 41(4):832–833, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33636.

Sanabria:2002:AGA

[San02] Sergio Sanabria. Architecture and Geometry in the Age of the Baroque (review). Technology and Culture, 43(4):785–787, October 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/34036.

Sandman:2003:INJ

[San03] Alison Sandman. Un inventor navarro: Jeronimo de Ayanz y Beaumont, 1553–1613 (review). Technology and Culture, 44 (2):379–381, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/42041.

Sanabria:2006:AAD

[San06] Sergio Sanabria. Architecture in the Age of Divided Representation: The Question of Creativity in the Shadow of Production (review). Technology and Culture, 47(3):659–661, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201478.

Satya:2004:WMP

[Sat04] Laxman D. Satya. Western Medicine and Public Health in Colonial Bombay, 1845–1895 (review). Technology and Culture, 45(1):179–180, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51773.

Sato:2005:LES

[Sat05] Yasushi Sato. Local engineering and systems engineering: Cultural conflict at NASA’s Marshall space flight center, 1960–1966. Technology and Culture, 46(3):561–583, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186637.

Sato:2007:BJA

[Sat07] Yasushi Sato. Into the Black: JPL and the American Space Program, 1976–2004 (review). Technology and Culture, 48(4): 893–895, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223378.

Sato:2008:DPT

[Sat08] Yasushi Sato. The Development of Propulsion Technology for U.S. Space-Launch Vehicles, 1926–1991 (review). Technology and Culture, 49(4):1099–1100, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255538.

Sauer:2005:TWF

[Sau05] Beverly J. Sauer. True Warnings and False Alarms about Technology, 1948–1971 (review). Technology and Culture, 46(2): 458–459, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/184341.

Scanlon:2005:KKH

[Sca05] Jennifer Scanlon. Kotex, Kleenex, and Huggies: Kimberly-Clark and the Consumer Revolution (review). Technology and Culture, 46(4):824–825, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192321.

Scanlon:2008:PIM

[Sca08] Jennifer Scanlon. Popular Ideologies: Mass Culture at Mid-Century (review). Technology and Culture, 49(4):1077–1079, October 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/255512.

Schiffer:2000:MCR

[Sch00a] Michael B. Schiffer. Material Culture (review). Technology and Culture, 41(4):791–793, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33637.

Schrag:2000:CTI

[Sch00b] Zachary M. Schrag. Chicago Transit: An Illustrated History (review). Technology and Culture, 41(3):638–640, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33589.

Schrag:2000:BYH

[Sch00c] Zachary M. Schrag. “The bus is young and honest”: Transportation politics, technical choice, and the motorization of Manhattan surface transit, 1919–1936. Technology and Culture, 41(1):51–79, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33496.

Schwarzkopf:2000:WWI

[Sch00d] Jutta Schwarzkopf. Women, Writing, and the Industrial Revolution (review). Technology and Culture, 41(2):359–360, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33540.

Schneller:2001:TCU

[Sch01a] Robert John Schneller. Technological Change and the United States Navy, 1865–1945 (review). Technology and Culture, 42 (4):812–814, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33839.

Schramm:2001:DFD

[Sch01b] Jeff Schramm. Dropping the Fire: The Decline and Fall of the Steam Locomotive (review). Technology and Culture, 42 (2):368–369, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33737.

Schwarz:2001:FFE

[Sch01c] Frederic D. Schwarz. Facts On File Encyclopedia of Science, Technology, and Society (review). Technology and Culture, 42 (2):397–399, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33738.

Schwarz:2001:SAA

[Sch01d] Frederic D. Schwarz. Social Authorship and the Advent of Print (review). Technology and Culture, 42(3):561–562, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33792.

Schatzberg:2002:NCC

[Sch02a] Eric Schatzberg. Natural Capitalism: Creating the Next Industrial Revolution (review). Technology and Culture, 43(1): 218–221, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33889.

Schmalzer:2002:SCC

[Sch02b] Sigrid Schmalzer. Science and Civilisation in China. Vol. 6: Biology and Biological Technology. Pt. 5: Fermentations and Food Science (review). Technology and Culture, 43(2):407– 409, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33933.

Schneller:2002:UNO

[Sch02c] Robert John Schneller. The U.S. Navy and the Origins of the Military-Industrial Complex, 1847–1883 (review). Technology and Culture, 43(2):458–460, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33934.

Schrag:2002:BCS

[Sch02d] Zachary M. Schrag. The Bulldozer in the Countryside: Suburban Sprawl and the Rise of American Environmentalism (review). Technology and Culture, 43(4):802–803, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/34037.

Schwantes:2002:RPC

[Sch02e] Carlos A. Schwantes. The Railroad Passenger Car: An Illustrated History of the First Hundred Years, with Accounts by Contemporary Passengers (review). Technology and Culture, 43(1):170–172, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33890.

Schrag:2003:CCR

[Sch03] Zachary M. Schrag. Concrete and Clay: Reworking Nature in New York City (review). Technology and Culture, 44(1): 211–212, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40183.

Schatzberg:2004:ACA

[Sch04a] Eric Schatzberg. On attempting to construct alternative narratives. Technology and Culture, 45(2):406–412, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/55689.

Schiffer:2004:BFB

[Sch04b] Michael B. Schiffer. Bolt of Fate: Benjamin Franklin and His Electric Kite Hoax (review). Technology and Culture, 45(4): 839–840, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175778.

Schrag:2004:WWD

[Sch04c] Zachary M. Schrag. From Warfare to Welfare: Defense Intellectuals and Urban Problems in Cold War America (review). Technology and Culture, 45(4):885–886, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/175779.

Schramm:2004:UPM

[Sch04d] Jeff Schramm. The Union Pacific Museum and Archives (review). Technology and Culture, 45(4):824–829, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
175780.

Schiffer:2005:ELN

[Sch05a] Michael B. Schiffer. The electric lighthouse in the Nineteenth Century: Aid to navigation and political technology. Technology and Culture, 46(2):275–305, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184342.

Schmucki:2005:SBP

[Sch05b] Barbara Schmucki. Strasse, Bahn, Panorama: Verkehrswege und Landschaftsveranderung in Deutschland von 1930 bis 1990 (review). Technology and Culture, 46(1):215–216, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
179783.

Schneider:2005:LML

[Sch05c] Wendie Ellen Schneider. Laws of Men and Laws of Nature: The History of Scientific Expert Testimony in England and America (review). Technology and Culture, 46(3):647–649, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186638.

Schramm:2005:PAS

[Sch05d] Jeff Schramm. Perfecting the American Steam Locomotive (review). Technology and Culture, 46(1):209–211, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
179784.

Schwantes:2005:BCN

[Sch05e] Carlos A. Schwantes. The Birth of California Narrow Gauge: A Regional Study of the Technology of Thomas and Martin Carter (review). Technology and Culture, 46(1):208–209, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179785.

Schatzberg:2006:TCA

[Sch06a] Eric Schatzberg. Technik comes to America: Changing meanings of technology before 1930. Technology and Culture, 47 (3):486–512, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201479.

Schipper:2006:SPC

[Sch06b] Frank Schipper. Second plenary conference of the tensions of Europe network: Lappeenranta, Finland, 24-28 May 2006. Technology and Culture, 47(4):799–802, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/205003.

Schmucki:2006:TCM

[Sch06c] Barbara Schmucki. Technology and the Culture of Modernity in Britain and Germany, 1890–1945 (review). Technology and Culture, 47(4):849–850, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/205004.

Schramm:2006:FFH

[Sch06d] Jeff Schramm. “Follow the Flag”: A History of the Wabash Railroad Company (review). Technology and Culture, 47(1): 217–218, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196068.

Schwantes:2006:TLH

[Sch06e] Carlos A. Schwantes. The Tootin’ Louie: A History of the Minneapolis and St. Louis Railway (review). Technology and Culture, 47(2):430–431, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/200874.

Schwarz:2006:STS

[Sch06f] Frederic D. Schwarz. Science, Technology, and Society: An Encyclopedia (review). Technology and Culture, 47(4):879–881, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/205005.

Schmitt:2007:HCD

[Sch07a] Stefan Schmitt. Half a Century of Digital Gaming: Game On, at the Science Museum, London, 21 October 2006–25 February 2007. Technology and Culture, 48(3):582–588, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218727.

Schrag:2007:MPS

[Sch07b] Zachary M. Schrag. The Merchant of Power: Sam Insull, Thomas Edison, and the Creation of the Modern Metropolis (review). Technology and Culture, 48(1):218–219, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209858.

Schwantes:2007:IGW

[Sch07c] Benjamin Schwantes. Industrial Genius: The Working Life of Charles Michael Schwab (review). Technology and Culture, 48(4):886–887, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/223379.

Schwarz:2007:TWH

[Sch07d] Frederic D. Schwarz. Technology in World History (review). Technology and Culture, 48(3):612–614, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218728.

Schwarz:2007:WSA

[Sch07e] Frederic D. Schwarz. We should all be friends. Technology and Culture, 48(2):407–410, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/215414.

Schatzberg:2008:MEC

[Sch08a] Eric Schatzberg. Means without End: A Critical Survey of the Ideological Geneaology of Technology without Limits, from Apollonian Techne to Postmodern Technoculture (review). Technology and Culture, 49(3):781–782, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/243126.

Schiffer:2008:CAC

[Sch08b] Michael Brian Schiffer. A cognitive analysis of componentstimulated invention: Electromagnet, telegraph, and the Capitol Dome’s electric gas-lighter. Technology and Culture, 49(2): 376–398, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236964.

Schramm:2008:TTR

[Sch08c] Jeff Schramm. The Train of Tomorrow (review). Technology and Culture, 49(2):474–476, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/237001.

Schiffer:2009:MCM

[Sch09a] Michael Brian Schiffer. Material Cultures, Material Minds: The Impact of Things on Human Thought, Society, and Evolution (review). Technology and Culture, 50(3):677–679, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270546.

Schmid:2009:NLC

[Sch09b] Sonja D. Schmid. Nuclear Legacies: Communication, Controversy, and the U.S. Nuclear Weapons Complex (review). Technology and Culture, 50(1):244–245, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/258558.

Scranton:2000:MTC

[Scr00] Philip Scranton. Missing the target? A comment on Edward Constant’s “Reliable Knowledge and Unreliable Stuff”. Technology and Culture, 41(4):752–764, October 2000. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33638.

See [Con99].

Scranton:2005:HUM

[Scr05] Philip Scranton. How Users Matter: The Co-Construction of Users and Technology (review). Technology and Culture, 46(4): 818–820, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192322.

Scranton:2009:CTU

[Scr09] Philip Scranton. The challenge of technological uncertainty. Technology and Culture, 50(2):513–518, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263336.

Smith-Divita:2001:TSS

[SD01] Stephanie Smith-Divita. Twenty-sixth symposium of the International Committee for the History of Technology. Technology and Culture, 42(1):138–139, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33692.

Simms:2009:AS

[SD09] Dennis L. Simms and Stephanie Dalley. The Archimedean screw. Technology and Culture, 50(3):730–735, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270538.

Seely:2001:YSS

[See01] Bruce Edsall Seely. Yellow Steel: The Story of the Earthmoving Equipment Industry (review). Technology and Culture, 42 (2):363–364, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33739.

Seely:2003:RCM

[See03] Bruce Edsall Seely. Reinforced Concrete and the Modernization of American Building, 1900–1930 (review). Technology and Culture, 44(1):209–211, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40184.

Seely:2007:EHT

[See07] Bruce Edsall Seely. Economic history as technological history: George Rogers Taylor’s The Transportation Revolution, 1815–1860. Technology and Culture, 48(4):824–830, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
223380.

Seely:2008:BPD

[See08] Bruce Edsall Seely. Big promises: David Berube’s Nano-Hype. Technology and Culture, 49(1):234–236, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230927.

Segal:2000:SAI

[Seg00] Howard P. Segal. Space and the American Imagination (review). Technology and Culture, 41(1):168–170, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33497.

Segal:2008:DPF

[Seg08] Howard P. Segal. Dreams of Peace and Freedom: Utopian Moments in the Twentieth Century (review). Technology and Culture, 49(1):259–261, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230928.

Segal:2009:PHC

[Seg09] Howard P. Segal. Pioneer in the hardware of culture: Roger Burlingame’s March of the Iron Men and Engines of Democracy. Technology and Culture, 50(2):426–440, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263335.

Seidel:2003:SCD

[Sei03] Robert W. Seidel. Strategic Computing: DARPA and the Quest for Machine Intelligence, 1983–1993 (review). Technology and Culture, 44(4):844–846, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49717.

Self:2009:DDR

[Sel09] David R. Self. To a Distant Day: The Rocket Pioneers (review). Technology and Culture, 50(3):719–720, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270560.

Serchuk:2002:SES

[Ser02a] Adam Serchuk. From Space to Earth: The Story of Solar Electricity (review). Technology and Culture, 43(2):460–461, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33935.

Serchuk:2002:SET

[Ser02b] Adam Serchuk. Solar Energy, Technology Policy, and Institutional Values (review). Technology and Culture, 43(1):184–186, January 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33891.

Servos:2002:WNF

[Ser02c] John W. Servos. War and Nature: Fighting Humans and Insects with Chemicals from World War I to Silent Spring (review). Technology and Culture, 43(2):448–450, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33936.

Servos:2005:ICU

[Ser05] John W. Servos. Impure Cultures: University Biology and the World of Commerce (review). Technology and Culture, 46(1): 255–257, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179786.

Sewell:2006:THA

[Sew06] Jessica Sewell. Town House: Architecture and Material Life in the Early American City, 1780–1830 (review). Technology and Culture, 47(4):826–827, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/205006.

Shackelford:2000:WCT

[Sha00a] Jole Shackelford. Women in Chemistry: Their Changing Roles from Alchemical Times to the Mid-Twentieth Century (review). Technology and Culture, 41(1):134–136, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33498.

Shallat:2000:WCT

[Sha00b] Todd A. Shallat. The Welland Canals and Their Communities: Engineering, Industrial, and Urban Transformation (review). Technology and Culture, 41(2):371–373, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33541.

Shapiro:2000:TLH

[Sha00c] Stuart S. Shapiro. Turing’s Legacy: A History of Computing at the National Physical Laboratory, 1945–1995 (review). Technology and Culture, 41(1):172–174, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33499.

Shallat:2001:BCW

[Sha01a] Todd A. Shallat. Bay Cities and Water Politics: The Battle for Resources in Boston and Oakland (review). Technology and Culture, 42(1):154–155, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33689.

Shayt:2001:OGT

[Sha01b] David H. Shayt. One Good Turn: A Natural History of the Screwdriver and the Screw (review). Technology and Culture, 42(4):807–808, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33840.

Shapiro:2004:HTT

[Sha04] Helen Shapiro. High-Tech Trade Wars: U.S.-Brazilian Conflicts in the Global Economy (review). Technology and Culture, 45(1):217–219, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51774.

Shapiro:2005:SCS

[Sha05a] Stuart S. Shapiro. The Soft Cage: Surveillance in America, from Slave Passes to the Patriot Act (review). Technology and Culture, 46(1):254–255, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179787.

Sharrer:2005:SFB

[Sha05b] G. Terry Sharrer. Safe Food: Bacteria, Biotechnology, and Bioterrorism (review). Technology and Culture, 46(1):257–259, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179788.

Shackleford:2006:TFH

[Sha06a] Ben Shackleford. The Technological Fix: How People Use Technology to Create and Solve Problems (review). Technology and Culture, 47(1):246–248, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196069.

Shallat:2006:HL

[Sha06b] Todd A. Shallat. Holding Louisiana. Technology and Culture, 47(1):102–107, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/196070.

Shapiro:2006:ITP

[Sha06c] Stuart S. Shapiro. Information Technology Policy: An International History (review). Technology and Culture, 47(2):465– 467, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200875.

Shayt:2006:ADC

[Sha06d] David H. Shayt. Artifacts of disaster: Creating the Smithsonian’s Katrina Collection. Technology and Culture, 47(2):357– 368, April 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/200876.

Shayt:2006:CCC

[Sha06e] David H. Shayt. Carriages and Clocks, Corsets and Locks: The Rise and Fall of an Industrial City–New Haven, Connecticut (review). Technology and Culture, 47(1):213–215, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
196071.

Shallat:2007:PPP

[Sha07] Todd A. Shallat. Public Power, Private Dams: The Hells Canyon High Dam Controversy (review). Technology and Culture, 48(4):873–875, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/223381.

Shah:2008:TOH

[Sha08] Esha Shah. Telling otherwise: A historical anthropology of tank irrigation technology in South India. Technology and Culture, 49(3):652–674, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243113.

Shallat:2009:TWS

[Sha09a] Todd Shallat. Troubled Waters: Steamboat Disasters, River Improvements, and American Public Policy, 1821–1860 (review). Technology and Culture, 50(1):219–220, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258579.

Sharpe:2009:BWF

[Sha09b] Pamela Sharpe. The Business of Women: Female Enterprise and Urban Development in Northern England, 1760– 1830 (review). Technology and Culture, 50(1):215, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258576.

Showalter:2002:WAT

[Sho02] Dennis E. Showalter. War in the Age of Technology: Myriad Faces of Modern Armed Conflict (review). Technology and Culture, 43(2):455–457, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33937.

Showalter:2005:GWI

[Sho05] Dennis E. Showalter. The Great War: An Imperial History (review). Technology and Culture, 46(1):222–223, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
179789.

Showalter:2008:CUC

[Sho08] Dennis E. Showalter. Considering the unthinkable: Chemical weapons in modern warfare. Technology and Culture, 49 (2):456–458, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/236970.

Shulman:2002:FTC

[Shu02] Bonnie J. Shulman. Feminism in Twentieth-Century Science, Technology, and Medicine (review). Technology and Culture, 43(4):782–784, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34038.

Shulman:2008:CVO

[Shu08] Peter A. Shulman. Canada’s Victorian Oil Town: The Transformation of Petrolia from Resource Town into a Victorian Community (review). Technology and Culture, 49(2):496–498, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/237002.

Shulman:2009:SBF

[Shu09a] Bonnie Shulman. Sciences from Below: Feminisms, Postcolonialities, and Modernities (review). Technology and Culture, 50 (3):682–683, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270566.

Shulman:2009:EEE

[Shu09b] Peter A. Shulman. Energy and everything else: Vaclav Smil’s Energy in Nature and Society. Technology and Culture, 50(4): 915–918, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363421.

Smith:2003:AEE

[SHW03] Crosbie Smith, Ian Higginson, and Phillip Wolstenholme. “Avoiding equally extravagance and parsimony”: The moral economy of the ocean steamship. Technology and Culture, 44 (3):443–469, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46115.

Siddiqi:2002:RSN

[Sid02] Asif A. Siddiqi. Russian Strategic Nuclear Forces (review). Technology and Culture, 43(4):822–823, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34039.

Siddiqi:2003:RRG

[Sid03] Asif A. Siddiqi. The rockets’ red glare: Technology, conflict, and terror in the Soviet Union. Technology and Culture, 44 (3):470–501, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/46114.

Siddiqi:2004:SMB

[Sid04] Asif A. Siddiqi. Science for the Masses: The Bolshevik State, Public Science, and the Popular Imagination in Soviet Russia, 1917–1934 (review). Technology and Culture, 45(1):206–207, January 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/51775.

Siddiqi:2006:SGS

[Sid06a] Asif A. Siddiqi. Stalin’s Great Science: The Times and Adventures of Soviet Physicists (review). Technology and Culture, 47 (3):674–676, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201480.

Siddiqi:2006:FSR

[Sid06b] Asif A. Siddiqi. The First Space Race: Launching the World’s First Satellites (review). Technology and Culture, 47(1):244– 246, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196072.

Siddiqi:2007:STE

[Sid07] Asif A. Siddiqi. Spacecraft Technology: The Early Years (review). Technology and Culture, 48(4):895–897, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
223382.

Siddiqi:2008:DAA

[Sid08] Asif A. Siddiqi. Dictatorship of the Air: Aviation Culture and the Fate of Modern Russia (review). Technology and Culture, 49 (4):1101–1102, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255539.

Siegelbaum:2009:SCC

[Sie09] Lewis H. Siegelbaum. On the side: Car culture in the USSR, 1960s–1980s. Technology and Culture, 50(1):1–22, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258581.

Silverman:2000:THD

[Sil00] M. Joshua Silverman. Target Hiroshima: Deak Parsons and the Creation of the Atomic Bomb, and: Fallout: A Historian Reflects on America’s Half-Century Encounter with Nuclear Weapons (review). Technology and Culture, 41(1):158–161, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33500.

Simon:2003:CBH

[Sim03] Roger D. Simon. From Cottage to Bungalow: Houses and the Working Class in Metropolitan Chicago, 1869–1929 (review). Technology and Culture, 44(2):402–404, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/42042.

Simon:2006:CBE

[Sim06] Roger D. Simon. City Building on the Eastern Frontier: Sorting the New Nineteenth-Century City (review). Technology and Culture, 47(1):210–211, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/196073.

Simon:2009:CMR

[Sim09] Roger D. Simon. Cities in Modernity: Representations and Productions of Metropolitan Space, 1840–1930 (review). Technology and Culture, 50(3):693–694, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270541.

Singer:2000:AKG

[Sin00] Bayla Singer. Artificial Knowing: Gender and the Thinking Machine (review). Technology and Culture, 41(1):174–175, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33501.

Singer:2005:AR

[Sin05a] Bayla Singer. Aircraft (review). Technology and Culture, 46(1): 230–231, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179790.

Singer:2005:BIM

[Sin05b] Bayla Singer. Beyond the Image Machine: A History of Visual Technologies (review). Technology and Culture, 46(2):424– 426, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184343.

Sinclair:2006:BHH

[Sin06] Bruce Sinclair. The Battle over Hetch Hetchy: America’s Most Controversial Dam and the Birth of Modern Environmentalism (review). Technology and Culture, 47(2):444–445, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200877.

Sinclair:2009:WBR

[Sin09] Bruce Sinclair. William Barton Rogers and the Idea of MIT (review). Technology and Culture, 50(4):968–970, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363451.

Skramstad:2000:EDI

[Skr00] Harold Skramstad. Exhibiting Dilemmas: Issues of Representation at the Smithsonian (review). Technology and Culture, 41 (2):353–355, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33542.

Skramstad:2004:DJB

[Skr04] Harold Skramstad. Daniel J. Boorstin, 1914–2004. Technology and Culture, 45(4):922–930, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/175781.

Skramstad:2007:ED

[Skr07] Harold Skramstad. The exhibiting dilemma. Technology and Culture, 48(3):603–611, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/218729.

Slaton:2001:NPP

[Sla01] Amy E. Slaton. “As near as practicable”: Precision, ambiguity, and the social features of industrial quality control. Technology and Culture, 42(1):51–80, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33690.

Slater:2002:CNE

[Sla02a] Ian J. Slater. Canadian Nuclear Energy Policy: Changing Ideas, Institutions, and Interests (review). Technology and Culture, 43(4):823–825, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/34040.

Slaton:2002:SFF

[Sla02b] Amy E. Slaton. Suitable for framing: Fingerprints and the rise of criminal identification. Technology and Culture, 43(4): 777–781, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34041.

Slade:2004:CFR

[Sla04] Joseph W. Slade. Cognitive Fictions (review). Technology and Culture, 45(1):225–227, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51776.

Slaton:2009:MMT

[Sla09] Amy Slaton. The Measure of Merit: Talents, Intelligence, and Inequality in the French and American Republics, 1750– 1940 (review). Technology and Culture, 50(1):209–211, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258555.

Slinn:2001:PIR

[Sli01] Judy Slinn. Pharmaceutical Innovation: Revolutionizing Human Health (review). Technology and Culture, 42(1):202–203, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33691.

Slinn:2009:MMA

[Sli09] Judy Slinn. Medicating Modern America: Prescription Drugs in History (review). Technology and Culture, 50(2):476–478, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263371.

Slotten:2002:FPR

[Slo02a] Hugh Richard Slotten. Fireside Politics: Radio and Political Culture in the United States, 1920–1940 (review). Technology and Culture, 43(1):196–198, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33892.

Slotten:2002:SCG

[Slo02b] Hugh Richard Slotten. Satellite communications, globalization, and the Cold War. Technology and Culture, 43(2):315– 350, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33938.

Slotten:2004:AAE

[Slo04] Hugh Richard Slotten. Anytime, Anywhere: Entrepreneurship and the Creation of a Wireless World (review). Technology and Culture, 45(1):213–214, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51777.

Slotten:2008:RHV

[Slo08] Hugh Richard Slotten. Radio’s hidden voice: Noncommercial broadcasting, extension education, and state universities during the 1920s. Technology and Culture, 49(1):1–20, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230929.

Slotten:2009:TER

[Slo09] Hugh R. Slotten. Telecommunications and Empire (review). Technology and Culture, 50(3):725–726, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270537.

Smart:2005:TDM

[Sma05] James Smart. Transport of Delight: The Mythical Conception of Rail Transit in Los Angeles (review). Technology and Culture, 46(3):661–662, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186639.

Smith:2002:BTT

[Smi02] Stephanie Smith. Boys and Their Toys?: Masculinity, Class, and Technology in America (review). Technology and Culture, 43(3):634–635, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33989.

Smith:2003:CWH

[Smi03] Robert W. Smith. Cold War, Hot Science: Applied Research in Britain’s Defence Laboratories, 1945–1990 (review). Technology and Culture, 44(3):646–647, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46116.

Smith:2005:FGH

[Smi05a] Robert Smith. Firearms: A Global History to 1700 (review). Technology and Culture, 46(2):407–408, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/184344.

Smith:2005:SOJ

[Smi05b] Robert W. Smith. Sky and Ocean Joined: The U.S. Naval Observatory, 1830–2000 (review). Technology and Culture, 46 (1):240–241, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/179791.

Smith:2006:CWT

[Smi06] Merritt Roe Smith. Civil War Time: Temporality and Identity in America, 1861–1865 (review). Technology and Culture, 47 (2):424–426, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200878.

Smith:2007:SCA

[Smi07] Merritt Roe Smith. Samuel Colt: Arms, Art, and Invention (review). Technology and Culture, 48(2):455–457, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
215415.

Smith:2009:ACA

[Smi09] Merritt Roe Smith. America’s coming of age: Daniel Walker Howe’s What Hath God Wrought. Technology and Culture, 50 (1):187–192, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/258582.

Smulyan:2005:RAA

[Smu05] Susan Smulyan. Radio Active: Advertising and Consumer Activism, 1935–1947 (review). Technology and Culture, 46(4): 825–827, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192323.

Smulyan:2007:RIP

[Smu07] Susan Smulyan. Radio’s Intimate Public: Network Broadcasting and Mass-Mediated Democracy (review). Technology and Culture, 48(1):230–231, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209859.

Socolow:2008:WEN

[Soc08] Michael J. Socolow. A wavelength for every network: Synchronous broadcasting and national radio in the United States, 1926–1932. Technology and Culture, 49(1):89–113, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
230930.

Soluri:2003:SNS

[Sol03] John Soluri. States of Nature: Science, Agriculture, and Environment in the Spanish Caribbean, 1760–1940 (review). Technology and Culture, 44(3):593–595, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46117.

Soluri:2004:TPG

[Sol04] John Soluri. Trouble in Paradise: Globalization and Environmental Crises in Latin America (review). Technology and Culture, 45(3):636–638, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/173003.

Sparke:2001:ICD

[Spa01] Penny Sparke. Imagining Consumers: Design and Innovation from Wedgwood to Corning (review). Technology and Culture, 42(2):345–346, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33740.

Sparke:2002:CWA

[Spa02] Penny Sparke. Candace Wheeler: The Art and Enterprise of American Design, 1875–1900 (review). Technology and Culture, 43(4):797–798, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/34042.

Spector:2004:MMM

[Spe04] Tami I. Spector. Magick, Mayhem, and Mavericks: The Spirited History of Physical Chemistry (review). Technology and Culture, 45(2):422–423, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/55690.

Spicer:2000:ESS

[Spi00a] Dag Spicer. Eye in the Sky: The Story of the Corona Spy Satellites (review). Technology and Culture, 41(3):633–634, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33590.

Spickermann:2000:TCI

[Spi00b] Roland Spickermann. The Traditional Chinese Iron Industry and its Modern Fate (review). Technology and Culture, 41 (3):583–584, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33591.

Spicer:2001:FCH

[Spi01] Dag Spicer. The First Computers: History and Architectures (review). Technology and Culture, 42(3):615–617, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33793.

Spiller:2004:RCC

[Spi04] James Spiller. Radiant cuisine: The commercial fate of food irradiation in the United States. Technology and Culture, 45(4): 740–763, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175782.

Spooner:2002:WMR

[Spo02] Denise S. Spooner. William Mulholland and the Rise of Los Angeles (review). Technology and Culture, 43(2):433–434, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33940.

Squatriti:2000:AAV

[Squ00] Paolo Squatriti. Acque e agricoltura: Venezia, l’Olanda, e la bonifica europea in eta moderna (review). Technology and Culture, 41(2):368–370, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33543.

Savage-Smith:2000:WAR

[SS00] Emilie Savage-Smith. Western Astrolabes (review). Technology and Culture, 41(3):578–579, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33588.

Savage-Smith:2002:GGC

[SS02] Emilie Savage-Smith. Globes at Greenwich: A Catalogue of the Globes and Armillary Spheres in the National Maritime Museum (review). Technology and Culture, 43(2):413–414, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33939.

Savage-Smith:2003:PST

[SS03] Emilie Savage-Smith. Persian Steel: The Tanavoli Collection (review). Technology and Culture, 44(3):601–603, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46113.

Smith:2007:TPB

[SS07] Crosbie Smith and Anne Scott. “Trust in Providence”: Building confidence into the Cunard line of steamers. Technology and Culture, 48(3):471–496, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218730.

Stamper:2000:WFR

[Sta00] John W. Stamper. World’s Fairs (review). Technology and Culture, 41(3):626–627, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33592.

Stapleton:2001:SCU

[Sta01] Darwin H. Stapleton. The Sanitary City: Urban Infrastructure in America from Colonial Times to the Present (review).

Technology and Culture, 42(2):357–359, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33741.

Staudenmaier:2002:EN

[Sta02] John M. Staudenmaier. Editor’s note. Technology and Culture, 43(Supplement):iv, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/34001.

Stapleton:2005:LCE

[Sta05] Darwin H. Stapleton. A lost chapter in the early history of DDT: The development of anti-typhus technologies by the Rockefeller Foundation’s Louse Laboratory, 1942–1944. Technology and Culture, 46(3):513–540, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186640.

Stapleton:2006:GKM

[Sta06a] Darwin H. Stapleton. Genghis Khan and the Making of the Modern World (review). Technology and Culture, 47(4):830– 831, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/205007.

Stapleton:2006:MWH

[Sta06b] Darwin H. Stapleton. The Mosquito Wars: A History of Mosquito Control in Florida (review). Technology and Culture, 47(2):449–450, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200879.

Staudenmaier:2009:SF

[Sta09] John M. Staudenmaier, S.J. SHOT at fifty. Technology and Culture, 50(3):623–630, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/270572.

Steinberg:2001:LER

[Ste01a] Theodore Steinberg. Lake Erie Rehabilitated: Controlling Cultural Eutrophication, 1960s-1990s (review). Technology and Culture, 42(1):152–154, January 2001. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33693.

Stephan:2001:EPM

[Ste01b] Karl David Stephan. Experts at play: Magnetron research at Westinghouse, 1930–1934. Technology and Culture, 42(4): 737–749, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33841.

Stephens:2001:MTG

[Ste01c] Carlene E. Stephens. Manufacturing Time: Global Competition in the Watch Industry, 1795–2000 (review). Technology and Culture, 42(4):791–793, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33843.

Steinberg:2004:BUC

[Ste04] Theodore Steinberg. Big is ugly: Corporate enclosure and the global water supply. Technology and Culture, 45(3):618– 623, July 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/173004.

Stephan:2005:MTS

[Ste05a] Karl David Stephan. Masterworks of Technology: The Story of Creative Engineering, Architecture, and Design (review). Technology and Culture, 46(3):632–633, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186642.

Stephens:2005:ECP

[Ste05b] Carlene E. Stephens. Einstein’s Clocks, Poincare’s Maps: Empires of Time (review). Technology and Culture, 46(1):241–243, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179792.

Steele:2006:IWS

[Ste06] Brett D. Steele. Instrumental in War: Science, Research, and Instruments between Knowledge and the World (review). Technology and Culture, 47(2):403–405, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/200880.

Steen:2007:PNS

[Ste07a] Kathryn Steen. Pull: Networking and Success since Benjamin Franklin (review). Technology and Culture, 48(3):670–671, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218731.

Stephenson:2007:ENS

[Ste07b] Marcia Stephenson. Experiencing Nature: The Spanish American Empire and the Early Scientific Revolution (review). Technology and Culture, 48(3):622–624, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/218732.

Stephenson:2008:SBP

[Ste08a] Anne Stephenson. The Seattle Bungalow: People & Houses, 1900–1940 (review). Technology and Culture, 49(1):281–282, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230931.

Stern:2008:GTI

[Ste08b] Marc Jeffrey Stern. Glass Towns: Industry, Labor, and Political Economy in Appalachia, 1890–1930s (review). Technology and Culture, 49(1):287–289, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230932.

Stine:2003:MMI

[Sti03] Jeffrey K. Stine. Me, myself and infrastructure: Private lives and public works in America: The National Building Museum, Washington, D.C. Technology and Culture, 44(4):778–785, October 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/49718.

Stilgoe:2004:MNO

[Sti04a] John R. Stilgoe. The Mighty Niagara: One River, Two Frontiers (review). Technology and Culture, 45(2):430–431, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55691.

Stine:2004:TPP

[Sti04b] Jeffrey K. Stine. Technology, policy, and a passion for shad: John McPhee, the founding fish. Technology and Culture, 45(4): 830–833, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175783.

Stine:2007:SPD

[Sti07] Jeffrey K. Stine. A sense of place: Donald Worster’s Dust Bowl. Technology and Culture, 48(2):377–385, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215416.

Stine:2008:SSP

[Sti08] Jeffrey K. Stine. Science for Sale: The Perils, Rewards, and Delusions of Campus Capitalism (review). Technology and Culture, 49(3):821–822, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/243089.

Stine:2009:EPS

[Sti09] Jeffrey K. Stine. Environment, Power, and Society for the Twenty-First Century: The Hierarchy of Energy (review). Technology and Culture, 50(2):505–506, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/263361.

Stokes:2001:ALR

[Sto01a] Raymond G. Stokes. Americanization and Its Limits: Reworking U.S. Technology and Management in Postwar Europe and Japan (review). Technology and Culture, 42(4):828–830, October 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33844.

Stokes:2001:EZU

[Sto01b] Raymond G. Stokes. Elektrizitatswirtschaft zwischen Umwelt, Technik und Politik: Aspekte aus 100 Jahren RWE-Geschichte, 1898–1998 (review). Technology and Culture, 42(3):594–595, July 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33794.

Storey:2001:STM

[Sto01c] William Kelleher Storey. Science, Technology and Medicine in Colonial India (review). Technology and Culture, 42(1): 195–197, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33694.

Storey:2002:STC

[Sto02] William Kelleher Storey. Space in the Tropics: From Convicts to Rockets in French Guiana (review). Technology and Culture, 43(2):462–463, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33941.

Stoll:2003:GPL

[Sto03a] Steven Stoll. A Green and Permanent Land: Ecology and Agriculture in the Twentieth Century (review). Technology and Culture, 44(2):426–428, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42043.

Storey:2003:MSU

[Sto03b] William Kelleher Storey. Missions for Science: U.S. Technology and Medicine in America’s African World (review). Technology and Culture, 44(3):608–609, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46118.

Stover:2003:MLA

[Sto03c] John F. Stover. Merging Lines: American Railroads, 1900– 1970 (review). Technology and Culture, 44(1):177–178, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40185.

Storey:2004:ICH

[Sto04a] William Kelleher Storey. Images and Contexts: The Historiography of Science and Modernity in India (review). Technology and Culture, 45(2):467–468, April 2004. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55692.

Storey:2004:GRS

[Sto04b] William Kelleher Storey. Guns, race, and skill in Nineteenth-Century Southern Africa. Technology and Culture, 45(4):687– 711, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175784.

Stokes:2005:PSK

[Sto05] Raymond G. Stokes. Die PAL-SECAM-Kontroverse in der DDR: Die politisch-ideologische Instrumentalisierung der Farbfernsehfrage durch den ostdeutschen Staat zwischen 1965 und 1969 (review). Technology and Culture, 46(4):832–833, October 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/192324.

Stokes:2006:KCA

[Sto06a] Raymond G. Stokes. Knowledge and Competitive Advantage: The Coevolution of Firms, Technology, and National Institutions (review). Technology and Culture, 47(1):229–231, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
196074.

Stoll:2006:SD

[Sto06b] Steven Stoll. The smallholder’s dilemma. Technology and Culture, 47(4):808–813, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/205008.

Stokes:2007:KFI

[Sto07a] Raymond G. Stokes. Kampfstoff-Forschung im Nationalsozialismus: Zur Kooperation von Kaiser-Wilhelm-Instituten, Militär und Industrie (review). Technology and Culture, 48(1): 208–210, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209860.

Storey:2007:SCM

[Sto07b] William Kelleher Storey. Science, Culture, and Modern State Formation (review). Technology and Culture, 48(3):627–629, July 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/218733.

Stoneman:2009:BNV

[Sto09a] Timothy Stoneman. A bold new vision: The VOA Radio Ring Plan and global broadcasting in the Early Cold War. Technology and Culture, 50(2):316–344, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/263331.

Storey:2009:TEI

[Sto09b] William Kelleher Storey. Technology and Ethical Idealism: A History of Development in the Netherlands East Indies (review). Technology and Culture, 50(1):206–207, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258564.

Strasser:2001:APA

[Str01] Susan Strasser. Advertising Progress: American Business and the Rise of Consumer Marketing (review). Technology and Culture, 42(1):181–182, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33695.

Strasser:2002:MCC

[Str02] Susan Strasser. Making consumption conspicuous: Transgressive topics go mainstream. Technology and Culture, 43(4): 755–770, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34043.

Stranges:2005:FHC

[Str05] Anthony N. Stranges. Fritz Haber: Chemist, Nobel Laureate, German, Jew (review). Technology and Culture, 46(2):448– 449, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184345.

Stradling:2009:SPR

[Str09a] David Stradling. Sprawling Places (review). Technology and Culture, 50(3):696–698, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/270548.

Stranges:2009:IIF

[Str09b] Anthony Stranges. Inside IG Farben: Hoechst during the Third Reich (review). Technology and Culture, 50(3):712–713, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270562.

Sturdy:2007:LS

[Stu07] Steve Sturdy. So long, succor. Technology and Culture, 48(1): 169–174, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209861.

Stephan:2001:C

[STW01] Karl David Stephan, Andrew D. Todd, and Rosalind H. Williams. Communications. Technology and Culture, 42(4): 839–852, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33842.

Sudrow:2008:PPK

[Sud08] Anne Sudrow. Plastik und politische Kultur in Westdeutschland (review). Technology and Culture, 49(2):482–484, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
237003.

Sverrisson:2002:SBL

[Sve02] Arni Sverrisson. Small boats and large ships: Social continuity and technical change in the Icelandic fisheries, 1800–1960. Technology and Culture, 43(2):227–253, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33942.

Sverrisson:2003:FMT

[Sve03] Arni Sverrisson. A Fishery for Modern Times: The State and the Industrialization of the Newfoundland Fishery, 1934–1968 (review). Technology and Culture, 44(3):595–597, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46119.

Storm:2007:TTE

[SW07] Anna Storm and Nina Wormbs. Tom Tits Experiment, Södertälje, Sweden (review). Technology and Culture, 48(4): 810–815, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/223383.

Swanson:2009:EPP

[Swa09] Kara W. Swanson. The emergence of the professional patent practitioner. Technology and Culture, 50(3):519–548, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
270553.

Swedberg:2008:PIJ

[Swe08] Richard Swedberg. Prophet of Innovation: Joseph Schumpeter and Creative Destruction (review). Technology and Culture, 49(1):239–240, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/230933.

Tagliacozzo:2005:LAC

[Tag05] Eric Tagliacozzo. The lit archipelago: Coast lighting and the Imperial optic in Insular Southeast Asia, 1860–1910. Technology and Culture, 46(2):306–328, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/184346.

Takahashi:2000:NTA

[Tak00] Yuzo Takahashi. A network of tinkerers: The advent of the radio and television receiver industry in Japan. Technology and Culture, 41(3):460–484, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33594.

Takahashi:2001:AJE

[Tak01] Yuzo Takahashi. Assembled in Japan: Electrical Goods and the Making of the Japanese Consumer (review). Technology and Culture, 42(1):180–181, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33696.

Tapdrup:2002:AFS

[Tap02] Jan Tapdrup. Adams of Fleet Street: Instrument Makers to King George III (review). Technology and Culture, 43(3):599– 600, July 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33990.

Tatarewicz:2005:GTA

[Tat05] Joseph N. Tatarewicz. Giant Telescopes: Astronomical Ambition and the Promise of Technology (review). Technology and Culture, 46(1):238–240, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/179793.

Thompson:2007:C

[TCOA07] Steven L. Thompson, Edward W. Constant, Jeffrey A. Oaks, and Mark Aldrich. Communications. Technology and Culture, 48(3):680–688, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218734.

Teaford:2005:TCS

[Tea05] Jon C. Teaford. Twentieth-Century Sprawl: Highways and the Reshaping of the American Landscape (review). Technology and Culture, 46(3):676–677, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186643.

Teaford:2008:CVM

[Tea08] Jon C. Teaford. Caro versus Moses, round two: Robert Caro’s The Power Broker. Technology and Culture, 49(2):442–448, April 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/236968.

Teaford:2009:PWN

[Tea09] Jon C. Teaford. Paving the Way: New York Road Building and the American State, 1880–1956 (review). Technology and Culture, 50(2):486–487, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/263372.

TeBrake:2002:TWH

[TeB02] William H. TeBrake. Taming the waterwolf: Hydraulic engineering and water management in the Netherlands during the Middle Ages. Technology and Culture, 43(3):475–499, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33991.

Teich:2001:FSA

[Tei01] Albert H. Teich. Funding Science in America: Congress, Universities, and the Politics of the Academic Pork Barrel (review). Technology and Culture, 42(3):622–623, July 2001. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33795.

Teich:2005:IIB

[Tei05] Albert H. Teich. Ivy and Industry: Business and the Making of the American University, 1880–1980 (review). Technology and Culture, 46(4):852–854, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192325.

Tenner:2000:AHI

[Ten00] Edward Tenner. Accidents in History: Injuries, Fatalities, and Social Relations (review). Technology and Culture, 41(1): 117–120, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33502.

Tenner:2005:BPS

[Ten05a] Edward Tenner. Bred for Perfection: Shorthorn Cattle, Collies, and Arabian Horses since 1800 (review). Technology and Culture, 46(2):468–469, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/184347.

Tenner:2005:RYE

[Ten05b] Edward Tenner. Replaceable You: Engineering the Body in Postwar America (review). Technology and Culture, 46(4): 871–873, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192326.

Terino:2000:HBH

[Ter00] John Terino. Helicopters before Helicopters (review). Technology and Culture, 41(4):846–847, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33639.

Tester:2009:EAS

[Tes09] Jeff Tester. Energy and American Society — Thirteen Myths (review). Technology and Culture, 50(2):504–505, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263364.

Thayer:2004:RCN

[Tha04] Ralph E. Thayer. A River and Its City: The Nature of Landscape in New Orleans (review). Technology and Culture, 45 (2):433–434, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55693.

Theriot:2008:SPR

[The08] Jason P. Theriot. The Strategic Petroleum Reserve: U.S. Energy Security and Oil Politics, 1975–2005 (review). Technology and Culture, 49(4):1103–1104, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/255540.

Thiesen:2006:SFS

[Thi06] William H. Thiesen. Ships’ Fastenings: From Sewn Boat to Steamship (review). Technology and Culture, 47(4):871–872, October 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/205010.

Thompson:2000:AMB

[Tho00] Steven L. Thompson. The arts of the motorcycle: Biology, culture, and aesthetics in technological choice. Technology and Culture, 41(1):99–115, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33503.

Thompson:2001:SBE

[Tho01] Emily Ann Thompson. “I Sing the Body Electric”: Music and Technology in the Twentieth Century (review). Technology and Culture, 42(4):800–801, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33845.

Thompson:2005:SMP

[Tho05] Gregory Lee Thompson. Suburbanizing the Masses: Public Transport and Urban Development in Historical Perspective (review). Technology and Culture, 46(3):662–664, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186644.

Thomas:2006:HSB

[Tho06] William Thomas. Herbert A. Simon: The Bounds of Reason in Modern America (review). Technology and Culture, 47(4): 877–879, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/205012.

Thompson:2007:GSS

[Tho07] Gregory Lee Thompson. The Great Society Subway: A History of the Washington Metro (review). Technology and Culture, 48(1):216–218, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/209863.

Thompson:2008:WTG

[Tho08] Gregory Lee Thompson. Watching the Traffic Go By: Transportation and Isolation in Urban America (review). Technology and Culture, 49(1):291–293, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230935.

Thompson:2009:CCC

[Tho09a] Gregory Thompson. Conspiracy or consumer choice? Technology and Culture, 50(3):673–676, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270530.

Thompson:2009:ALH

[Tho09b] Steven L. Thompson. Autophobia: Love and Hate in the Automotive Age (review). Technology and Culture, 50(4):942–944, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363436.

Thompson:2009:MCH

[Tho09c] Steven L. Thompson. The Model T: A Centennial History (review). Technology and Culture, 50(1):237–238, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258584.

Timmermann:2009:MKE

[Tim09] Anke Timmermann. Making Knowledge in Early Modern Europe: Practices, Objects, and Texts, 1400–1800 (review). Technology and Culture, 50(3):687–688, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/270557.

Todd:2002:ICB

[Tod02] Edmund N. Todd. Industrial Culture and Bourgeois Society: Business, Labor, and Bureaucracy in Modern Germany, 1800– 1918 (review). Technology and Culture, 43(3):605–607, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33992.

Todd:2008:MMA

[Tod08] Edmund N. Todd. The Morals of Measurement: Accuracy, Irony, and Trust in Late Victorian Electrical Practice (review). Technology and Culture, 49(4):1052–1053, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
255493.

Todd:2009:CNW

[Tod09] Edmund N. Todd. The Conquest of Nature: Water, Landscape, and the Making of Modern Germany (review). Technology and Culture, 50(1):216–217, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258575.

Tomayko:2003:BDB

[Tom03] J. E. Tomayko. Behind Deep Blue: Building the Computer that Defeated the World Chess Champion (review). Technology and Culture, 44(3):634–635, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/46121.

Tomayko:2004:ARS

[Tom04] J. E. Tomayko. From Airline Reservations to Sonic the Hedgehog: A History of the Software Industry (review). Technology and Culture, 45(2):447–448, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55694.

Tomayko:2005:EEF

[Tom05] J. E. Tomayko. Electronics in the Evolution of Flight (review). Technology and Culture, 46(3):681–682, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186645.

Townsend:2002:THB

[Tow02] Camilla Townsend. Time, History, and Belief in Aztec and Colonial Mexico (review). Technology and Culture, 43(2):409– 410, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33943.

Travis:2002:CTH

[Tra02] Anthony S. Travis. Colouring Textiles: A History of Natural Dyestuffs in Industrial Europe (review). Technology and Culture, 43(3):602–603, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33993.

Travis:2005:BEA

[Tra05] Anthony S. Travis. Bright Earth: Art and the Invention of Color (review). Technology and Culture, 46(1):259–260, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179794.

Trinder:2003:MST

[Tri03] Barrie Stuart Trinder. The Millers: A Story of Technological Endeavor and Industrial Success, 1870–2001 (review). Technology and Culture, 44(2):391–392, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42044.

Trimble:2004:TFI

[Tri04] William F. Trimble. Taking Flight: Inventing the Aerial Age, from Antiquity through the First World War (review). Technology and Culture, 45(3):626–627, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/173005.

Tribbia:2006:CSW

[Tri06a] John Tribbia. Cellphone: The Story of the World’s Most Mobile Medium and How It Has Transformed Everything (review). Technology and Culture, 47(3):687–688, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/201482.

Trimble:2006:IFW

[Tri06b] William F. Trimble. Inventing Flight: The Wright Brothers and Their Predecessors (review). Technology and Culture, 47(1): 241–242, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196076.

Tribbia:2009:MPC

[Tri09] John Tribbia. Mobile Phone Cultures (review). Technology and Culture, 50(3):728–729, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/270573.

Turner:2005:WCM

[Tur05] Fred Turner. Where the counterculture met the new economy: The WELL and the origins of virtual community. Technology and Culture, 46(3):485–512, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/186646.

Turner:2006:CCO

[Tur06] Fred Turner. Code: Collaborative Ownership and the Digital Economy (review). Technology and Culture, 47(3):685– 686, July 2006. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/201483.

Turner:2008:WCI

[Tur08] Fred Turner. Who Controls the Internet? Illusions of a Borderless World (review). Technology and Culture, 49(1):296–297, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230936.

Turner:2009:ZCB

[Tur09a] Fred Turner. Zero Comments: Blogging and Critical Internet Culture (review). Technology and Culture, 50(2):508–509, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263368.

Turner:2009:WNG

[Tur09b] Roger Turner. Weather by the Numbers: The Genesis of Modern Meteorology (review). Technology and Culture, 50(4): 966–968, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363450.

Twardy:2000:AML

[Twa00] Charles R. Twardy. Antonio Meucci: L’Inventore e il suo tempo (review). Technology and Culture, 41(1):125–126, January 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33504.

Tweedale:2005:LTT

[Twe05] Geoffrey Tweedale. Landscapes of Technology Transfer: Swedish Ironmakers in India, 1860–1864 (review). Technology and Culture, 46(4):844–846, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192328.

Tyabji:2004:GTK

[Tya04] Nasir Tyabji. Gaining technical know-how in an unequal world: Penicillin manufacture in Nehru’s India. Technology and Culture, 45(2):331–349, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/55695.

Tympas:2000:FCE

[Tym00] Aristotle Tympas. Frankenstein’s Children: Electricity, Exhibition, and Experiment in Early-Nineteenth-Century London (review). Technology and Culture, 41(4):810–812, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33640.

Uekotter:2007:GNC

[Uek07] Frank Uekötter. Germany’s Nature: Cultural Landscapes and Environmental History (review). Technology and Culture, 48 (3):632–633, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218735.

Unger:2005:MME

[Ung05] Richard W. Unger. Mills in the Medieval Economy: England, 1300–1540 (review). Technology and Culture, 46(3):636– 638, July 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/186647.

Usselman:2003:RAL

[Uss03] Steven W. Usselman. Railroads and American Law (review). Technology and Culture, 44(4):824–825, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/49719.

Usselman:2004:MPC

[Uss04] Steven W. Usselman. Mechanizing Proof: Computing, Risk, and Trust (review). Technology and Culture, 45(4):893–895, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175785.

Usselman:2006:SV

[Uss06] Steven W. Usselman. Still visible. Technology and Culture, 47 (3):584–596, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201484.

Usselman:2009:HTH

[Uss09] Steven W. Usselman. The historian of technology and her true country. Technology and Culture, 50(1):110–126, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
258562.

Vaillant:2000:SAH

[Vai00] Derek Vaillant. Speaking into the Air: A History of the Idea of Communication (review). Technology and Culture, 41(4): 795–796, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33641.

VanderLinden:2000:FWN

[Van00a] F. Robert Van der Linden. Flying Without Wings: NASA Lifting Bodies and the Birth of the Space Shuttle (review). Technology and Culture, 41(4):848–849, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33642.

VanderMeulen:2000:FKM

[Van00b] Jacob A. Vander Meulen. Fundable Knowledge: The Marketing of Defense Technology (review). Technology and Culture, 41(1): 162–164, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33505.

VanSlyck:2002:KTM

[Van02] Abigail Ayres Van Slyck. Kitchen technologies and mealtime rituals: Interpreting the food axis at American summer camps, 1890–1950. Technology and Culture, 43(4):668–692, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
34044.

VanOosterhout:2008:CPI

[Van08] Dianne Van Oosterhout. From colonial to postcolonial irrigation technology: Technological romanticism and the revival of colonial water tanks in Java, Indonesia. Technology and Culture, 49(3):701–726, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/243116.

Vastokas:2001:CSS

[Vas01] Joan M. Vastokas. The Culture of Stone: Sacred and Profane Uses of Stone Among the Dani (review). Technology and Culture, 42(1):187–189, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33697.

Vasta:2008:FSO

[Vas08] Michelangelo Vasta. Fabbriche, sistemi, organizzazioni: Storia dell’ingegneria industriale (review). Technology and Culture, 49(1):265–266, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/230937.

Valderrama:2009:EEI

[VCM^{+}09] Andrés Valderrama, Juan Camargo, Idelman Mejía, Antonio Mejía, Ernesto Lleras, and Antonio García. Engineering education and the identities of engineers in Colombia, 1887–1972. Technology and Culture, 50(4):811–838, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363415.

vanDam:2002:ECT

[vD02] Petra J. E. M. van Dam. Ecological challenges, technological innovations: The modernization of sluice building in Holland, 1300–1600. Technology and Culture, 43(3):500–520, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33994.

vanDam:2003:WTM

[vD03] Petra J. E. M. van Dam. Water Technology in the Middle Ages: Cities, Monasteries, and Waterworks after the Roman Empire (review). Technology and Culture, 44(2):376–377, April 2003.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42045.

vanDam:2009:WCB

[vD09] Petra J. E. M. van Dam. Waterways and Canal-Building in Medieval England (review). Technology and Culture, 50(2):458– 459, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263349.

vanDriel:2005:RIM

[vDS05] H. van Driel and Johan Schot. Radical innovation as a multilevel process: Introducing floating grain elevators in the port of Rotterdam. Technology and Culture, 46(1):51–76, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
179754.

vanderVleuten:2008:TTH

[vdV08] Erik van der Vleuten. Toward a transnational history of technology: Meanings, promises, pitfalls. Technology and Culture, 49(4):974–994, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/255444.

Veenendaal:2000:GBV

[Vee00] A. J. Veenendaal. Die Geschichte der Bahnen in Vorarlberg (review). Technology and Culture, 41(4):803–804, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33643.

Veenendaal:2001:BBV

[Vee01] A. J. Veenendaal. Bahnen und Berge: Verkehrstechnik, Tourismus und Naturschutz in den Schweizer Alpen, 1870–1939 (review). Technology and Culture, 42(3):606–607, July 2001. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33796.

Veenendaal:2003:GTR

[Vee03] A. J. Veenendaal. The Great Train Race: Railways and the Franco-German Rivalry, 1815–1914 (review). Technology and Culture, 44(4):822–823, October 2003. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/49720.

Veenendaal:2005:CRE

[Vee05] A. J. Veenendaal. The City and the Railway in Europe (review). Technology and Culture, 46(3):665–666, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/186641.

Ventre:2000:BES

[Ven00] Francis T. Ventre. Building the Empire State (review). Technology and Culture, 41(1):151–152, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33506.

Verstegen:2000:TGC

[Ver00] Wybren Verstegen. Technology and Global Change (review). Technology and Culture, 41(3):572–573, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33595.

Vermeir:2008:RDM

[Ver08] Koen Vermeir. RoboCop dissected: Man-machine and mindbody in the enlightenment. Technology and Culture, 49(4): 1036–1044, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255451.

vonFalkenhausen:2005:WBS

[vF05] Lothar von Falkenhausen. Written on Bamboo and Silk: The Beginnings of Chinese Books and Inscriptions (review). Technology and Culture, 46(2):410–411, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/184316.

Vietor:2006:USF

[Vie06] Richard H. K. Vietor. United States Foreign Oil Policy since World War I: For Profits and Security (review). Technology and Culture, 47(3):670–671, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201485.

Vietor:2008:OIS

[Vie08] Richard Vietor. The Offshore Imperative: Shell Oil’s Search for Petroleum in Postwar America (review). Technology and Culture, 49(2):494–496, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/237004.

Vining:2003:IWW

[Vin03] Margaret Vining. Improbable Warriors: Women Scientists and the U.S. Navy in World War II (review). Technology and Culture, 44(2):439–440, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42046.

Vining:2006:GHA

[Vin06] Margaret Vining. Grace Hopper: Admiral of the Cyber Sea (review). Technology and Culture, 47(2):462–463, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
200881.

vanLente:2009:ICI

[vL09] Harro van Lente. Insatiable Curiosity: Innovation in a Fragile Future (review). Technology and Culture, 50(4):977–978, October 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/363456.

Vogel:2002:OGR

[Vog02] Robert M. Vogel. Otis: Giving Rise to the Modern City (review). Technology and Culture, 43(2):431–433, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33944.

Vogel:2004:AEE

[Vog04] Robert M. Vogel. Up, down, across: Elevators, escalators, and moving sidewalks at the National Building Museum, Washington, D.C. Technology and Culture, 45(1):154–158, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
51778.

Volti:2001:TTC

[Vol01] Rudi Volti. Tiger Technology: The Creation of a Semiconductor Industry in East Asia (review). Technology and Culture, 42(4):836–838, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33846.

Volti:2003:MWA

[Vol03a] Rudi Volti. Mao’s War against Nature: Politics and the Environment in Revolutionary China (review). Technology and Culture, 44(2):430–432, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/42047.

Volti:2003:TTI

[Vol03b] Rudi Volti. Technology Transfer and International Production: The Development of the Electronics Industry in Korea (review). Technology and Culture, 44(1):197–199, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/40186.

Volti:2004:CIA

[Vol04a] Rudi Volti. Crisis and Innovation in Asian Technology (review). Technology and Culture, 45(2):465–467, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/55697.

Volti:2004:SCR

[Vol04b] Rudi Volti. Social Change with Respect to Culture and Original Nature (review). Technology and Culture, 45(2):396–405, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
55705.

Volti:2005:BHR

[Vol05a] Rudi Volti. Bicycle: The History (review). Technology and Culture, 46(3):680–681, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/186648.

Volti:2005:PTP

[Vol05b] Rudi Volti. Playing with Trains: A Passion Beyond Scale (review). Technology and Culture, 46(2):473–474, April 2005.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184348.

Volti:2005:EVT

[Vol05c] Rudi Volti. The Electric Vehicle: Technology and Expectations in the Automobile Age (review). Technology and Culture, 46(1): 213–215, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/179795.

Volti:2007:PLN

[Vol07a] Rudi Volti. Pushing the Limits: New Adventures in Engineering (review). Technology and Culture, 48(1):181–182, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
209864.

Volti:2007:RIA

[Vol07b] Rudi Volti. Route 66: Iconography of the American Highway (review). Technology and Culture, 48(3):658–659, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218737.

Volti:2008:MIR

[Vol08a] Rudi Volti. Making IT: The Rise of Asia in High Tech (review). Technology and Culture, 49(2):481–482, April 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/237005.

Volti:2008:CGA

[Vol08b] Rudi Volti. A car for the great Asian multitude. Technology and Culture, 49(4):995–1001, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255445.

Volti:2009:BME

[Vol09a] Rudi Volti. Bodies in Motion: Evolution and Experience in Motorcycling (review). Technology and Culture, 50(2):492– 493, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263367.

Volti:2009:EAI

[Vol09b] Rudi Volti. The Encyclopedia of the Age of the Industrial Revolution, 1700–1920 (review). Technology and Culture, 50(3):689– 690, July 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/270575.

vanRooij:2007:MIV

[vR07] Arjan van Rooij. Methods for innovation: The varying role of industrial research in DSM’s nitrogen fertilizer business, 1925–1970. Technology and Culture, 48(3):550–574, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218736.

Wachelder:2005:TTN

[Wac05] Joseph Wachelder. TIN-20: Techniek in Nederland in de twintigste eeuw. Technology and Culture, 46(1):187–191, January 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/179796.

Wajcman:2000:MTM

[Waj00] Judy Wajcman. Making Technology Masculine: Men, Women and Modern Machines in America, 1870–1945 (review). Technology and Culture, 41(4):788–789, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33644.

Wakefield:2002:PEC

[Wak02] Andre Wakefield. Papermaking in Eighteenth-Century France: Management, Labor, and Revolution at the Montgolfier Mill, 1761–1805 (review). Technology and Culture, 43(2):420–421, April 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33945.

Wakefield:2003:MAT

[Wak03] Andre Wakefield. The Measure of All Things: The Seven-Year Odyssey and Hidden Error that Transformed the World (review). Technology and Culture, 44(4):810–812, October 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/49721.

Wakefield:2004:OSA

[Wak04a] Andre Wakefield. Openness, Secrecy, Authorship: Technical Arts and the Culture of Knowledge from Antiquity to the Renaissance (review). Technology and Culture, 45(2):415–416, April 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/55698.

Wakefield:2004:CCF

[Wak04b] Andre Wakefield. Comment: The California fires. Technology and Culture, 45(1):247–249, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/51779.

Wakeman:2005:LGS

[Wak05] Rosemary Wakeman. The Light-Green Society: Ecology and Technological Modernity in France, 1960–2000 (review). Technology and Culture, 46(1):217–218, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179797.

Wakefield:2007:PSL

[Wak07] Andre Wakefield. Pierre Simon Laplace, 1749–1827: A Determined Scientist (review). Technology and Culture, 48(1): 221–222, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209865.

Walton:2000:MWH

[Wal00] Steven A. Walton. Medieval Warfare: A History (review). Technology and Culture, 41(4):836–837, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33645.

Walker:2001:INB

[Wal01a] J. Samuel Walker. India’s Nuclear Bomb: The Impact on Global Proliferation (review). Technology and Culture, 42(2): 379–380, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33742.

Walker:2001:RAN

[Wal01b] J. Samuel Walker. Regulating against nuclear terrorism: The domestic safeguards issue, 1970–1979. Technology and Culture, 42(1):107–132, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33698.

Walter:2001:ESA

[Wal01c] Maila L. Walter. The Establishment of Science in America: 150 Years of the American Association for the Advancement of Science (review). Technology and Culture, 42(2):394–396, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33743.

Wald:2002:HSR

[Wal02] Stephen E. Wald. The History of Science and Religion in the Western Tradition: An Encyclopedia (review). Technology and Culture, 43(3):654–656, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33995.

Walton:2004:ABD

[Wal04] Steven A. Walton. Ars Belli: Deutsche taktische und kriegstechnische Bilderhandschriften und Traktate im 15. und 16. Jahrhundert (review). Technology and Culture, 45(1):168–170, January 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/51780.

Wald:2005:MMN

[Wal05] Stephen E. Wald. Morality and the Mail in Nineteenth-Century America (review). Technology and Culture, 46(2):419–420, April 2005. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/184349.

Walton:2006:FFP

[Wal06a] Steven A. Walton. The Field and the Forge: Population, Production, and Power in the Pre-industrial West (review). Technology and Culture, 47(1):179–180, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196077.

Walton:2006:HAS

[Wal06b] Steven A. Walton. The Heirs of Archimedes: Science and the Art of War through the Age of Enlightenment (review). Technology and Culture, 47(4):828–830, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/205013.

Walker:2009:NWW

[Wal09] J. Samuel Walker. Nuclear Weapons: What You Need to Know (review). Technology and Culture, 50(2):511–512, April 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
263373.

Wang:2002:ISS

[Wan02] Jennifer Hyland Wang. Invisible Stars: A Social History of Women in American Broadcasting (review). Technology and Culture, 43(3):630–632, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/33996.

Wang:2004:EAP

[Wan04] Jessica Wang. Eisenhower’s Atoms for Peace (review). Technology and Culture, 45(1):202–204, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51781.

Wang:2007:BND

[Wan07] Jessica Wang. Building New Deal Liberalism: The Political Economy of Public Works, 1933–1956 (review). Technology and Culture, 48(3):659–661, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/218738.

Wang:2008:TEI

[Wan08] Mian Rachel Wang. Technological Empowerment: The Internet, State, and Society in China (review). Technology and Culture, 49(4):1089–1090, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255513.

Waring:2001:ASD

[War01a] Stephen P. Waring. Aiming for the Stars: The Dreamers and Doers of the Space Age (review). Technology and Culture, 42 (2):382–384, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33744.

Warner:2001:SBJ

[War01b] Deborah Jean Warner. Spectrum of Belief: Joseph von Fraunhofer and the Craft of Precision Optics (review). Technology and Culture, 42(4):797–798, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33847.

Warner:2002:EIM

[War02] Deborah Jean Warner. Elizabethan Instrument Makers: The Origins of the London Trade in Precision Instrument Makers (review). Technology and Culture, 43(2):412–413, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33946.

Warner:2003:LEP

[War03] Deborah Jean Warner. L’Art d’enseigner la physique: Les appareils de demonstration de Jean-Antoine Nollet, 1700–1770 (review). Technology and Culture, 44(3):603–604, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/46122.

Ward:2004:RRC

[War04] James Arthur Ward. Riding the Roller Coaster: A History of the Chrysler Corporation (review). Technology and Culture, 45(1):194–195, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/51782.

Ware:2005:AWF

[War05a] Susan Ware. American Women and Flight since 1940 (review). Technology and Culture, 46(1):228–229, January 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/179798.

Warner:2005:CGU

[War05b] Deborah Jean Warner. The Claude Glass: Use and Meaning of the Black Mirror in Western Art (review). Technology and Culture, 46(4):865–866, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192329.

Ward:2006:DBM

[War06a] James Arthur Ward. The Dodge Brothers: The Men, the Motor Cars, and the Legacy (review). Technology and Culture, 47 (2):435–436, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200882.

Warner:2006:DWS

[War06b] Deborah Jean Warner. Defining Women’s Scientific Enterprise: Mount Holyoke Faculty and the Rise of American Science (review). Technology and Culture, 47(3):697–698, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
201486.

Warner:2007:AGH

[War07a] Deborah Jean Warner. An Acre of Glass: A History and Forecast of the Telescope (review). Technology and Culture, 48(1): 231–232, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209866.

Warner:2007:HTP

[War07b] Deborah Jean Warner. Harmonious Triads: Physicists, Musicians, and Instrument Makers in Nineteenth-Century Germany (review). Technology and Culture, 48(4):856–857, October 2007. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/223385.

Warner:2008:FCG

[War08] Deborah Jean Warner. Flower Confidential: The Good, the Bad, and the Beautiful in the Business of Flowers (review). Technology and Culture, 49(1):293–294, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230938.

Warner:2009:RCG

[War09a] Deborah Jean Warner. Raising Cane in the ’Glades. The Global Sugar Trade and the Transformation of Florida (review). Technology and Culture, 50(4):926–927, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/363426.

Warner:2009:MEH

[War09b] Deborah Jean Warner. The Microscope and the Eye: A History of Reflections, 1740–1870 (review). Technology and Culture, 50 (3):690–691, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/270574.

Warner:2009:TVE

[War09c] Deborah Jean Warner. The Transit of Venus Enterprise in Victorian Britain (review). Technology and Culture, 50(2):463– 464, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263365.

Waters:2000:SRC

[Wat00] Wendy Waters. Sugar and Railroads: A Cuban History, 1837– 1959 (review). Technology and Culture, 41(1):149–151, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33507.

Watkins:2002:AAR

[Wat02] Nicholas Jay Watkins. Architectural Anthropology (review). Technology and Culture, 43(2):404–406, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33947.

Weatherwax:2002:MCV

[Wea02] Sarah J. Weatherwax. Making Culture Visible: Photography and Its Diplay at Industrial Fairs, International Exhibitions, and Institutional Exhibitions in the U.S., 1847–1900 (review). Technology and Culture, 43(4):833–835, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34045.

Weber:2008:RAE

[Web08] Leonard J. Weber. Real American Ethics: Taking Responsibility for Our Country (review). Technology and Culture, 49 (3):817–818, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/243091.

Weinstein:2001:MMB

[Wei01] Deborah Weinstein. Manifesting Medicine: Bodies and Machines (review). Technology and Culture, 42(1):197–199, January 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33699.

Weir:2003:FAB

[Wei03] Gary E. Weir. Fast attacks and boomers: Submarines in the Cold War, the National Museum of American History. Technology and Culture, 44(2):359–363, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/42048.

Weir:2004:BTN

[Wei04] Gary E. Weir. Building the Trident Network: A Study of the Enrollment of People, Knowledge, and Machines (review). Technology and Culture, 45(2):450–452, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/55699.

Weir:2006:OCW

[Wei06] Gary E. Weir. Oceanographers and the Cold War: Disciples of Marine Science (review). Technology and Culture, 47(1): 239–240, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/196079.

Wells:2002:TFS

[Wel02] Christopher W. Wells. Tunneling to the Future: The Story of the Great Subway Expansion that Saved New York (review). Technology and Culture, 43(4):803–805, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/34046.

Wells:2007:RMC

[Wel07] Christopher W. Wells. The road to the Model T: Culture, road conditions, and innovation at the dawn of the American Motor Age. Technology and Culture, 48(3):497–523, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218739.

Wells:2008:TPC

[Wel08] Christopher W. Wells. Trust and Power: Consumers, the Modern Corporation, and the Making of the United States Automobile Market (review). Technology and Culture, 49(1):275–277, January 2008. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/230939.

Wengenroth:2001:LWS

[Wen01] Ulrich Wengenroth. Laboratories, Workshops, and Sites: Concepts and Practices of Research in Industrial Europe, 1800– 1914 (review). Technology and Culture, 42(2):333–334, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33745.

Wengenroth:2002:STD

[Wen02] Ulrich Wengenroth. Streit um die Technikgeschichte in Deutschland, 1945–1975 (review). Technology and Culture, 43 (3):651–653, July 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/33997.

Wengenroth:2007:WPU

[Wen07] Ulrich Wengenroth. Wankel auf dem Prüfstand: Ursprung, Entwicklung, und Niedergang eines innovativen Motorenkonzeptes (review). Technology and Culture, 48(3):652–653, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
218740.

Wengenroth:2008:TMT

[Wen08] Ulrich Wengenroth. Technografie: Zur Mikrosoziologie der Technik (review). Technology and Culture, 49(3):831–832, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243108.

Wengenroth:2009:VSS

[Wen09] Ulrich Wengenroth. Vernetzte Steuerung: Soziale Prozesse im Zeitalter technischer Netzwerke (review). Technology and Culture, 50(3):680–681, July 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/270820.

Wermiel:2002:LDB

[Wer02] Sara E. Wermiel. L’Invention du beton arme: Hennebique, 1890–1914 (review). Technology and Culture, 43(3):610–612, July 2002. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33998.

Wermiel:2003:NER

[Wer03] Sara E. Wermiel. No exit: The rise and demise of the outside fire escape. Technology and Culture, 44(2):258–284, April 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/42049.

Wermiel:2004:GGH

[Wer04] Sara E. Wermiel. Gaining Ground: A History of Landmaking in Boston (review). Technology and Culture, 45(4):844–846, October 2004. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/175786.

Wermiel:2005:FAI

[Wer05a] Sara E. Wermiel. Forging America: Ironworkers, Adventurers, and the Industrious Revolution (review). Technology and Culture, 46(4):842–843, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/192330.

Wermiel:2005:BCP

[Wer05b] Sara E. Wermiel. The Birth of City Planning in the United States, 1840–1917 (review). Technology and Culture, 46(3): 674–675, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/186649.

Wermiel:2006:ACA

[Wer06] Sara E. Wermiel. Adolf Cluss, architect: From Germany to America — shaping a capital city worthy of a republic. Technology and Culture, 47(3):570–577, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/201487.

Wermiel:2007:BDE

[Wer07a] Sara E. Wermiel. Bracing for Disaster: Earthquake-Resistant Architecture and Engineering in San Francisco, 1838–1933 (review). Technology and Culture, 48(4):877–879, October 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
223386.

Werrett:2007:DTM

[Wer07b] Simon Werrett. Deep Time of the Media: Toward an Archeology of Hearing and Seeing by Technical Means (review). Technology and Culture, 48(2):440–442, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/215418.

Westfall:2000:MPB

[Wes00] Catherine Westfall. Making Physics: A Biography of Brookhaven National Laboratory, 1946–1972 (review). Technology and Culture, 41(2):375–376, April 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33544.

Westfall:2001:LMD

[Wes01] Catherine Westfall. Lise Meitner and the Dawn of the Nuclear Age, by Patricia Rife (review). Technology and Culture, 42(2):375–377, April 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL http:/

/www.jstor.org/stable/25147727;   https://muse.jhu.
edu/pub/1/article/33746.

Westfall:2003:SBS

[Wes03] Catherine Westfall. Scientists, Business, and the State, 1890– 1960 (review). Technology and Culture, 44(1):201–202, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
40188.

Westfall:2004:NLS

[Wes04] Catherine Westfall. The National Labs: Science in an American System, 1947–1974 (review). Technology and Culture, 45 (2):459–460, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/55700.

Westwick:2007:REM

[Wes07] Peter J. Westwick. Reengineering engineers: Management philosophies at JPL in the 1990s. Technology and Culture, 48 (1):67–91, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/209867.

Westbrook:2008:MEC

[Wes08] Jack H. Westbrook. Materials in Eighteenth-Century Science: A Historical Ontology (review). Technology and Culture, 49(4): 1104–1105, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/255541.

Warren-Findley:2003:EEF

[WF03] Jannelle Warren-Findley. Expanding the Envelope: Flight Research at NACA and NASA (review). Technology and Culture, 44(1):183–184, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/40187.

Wiesner-Hanks:2009:WWP

[WH09] Merry Wiesner-Hanks. Women at Work in Preindustrial France (review). Technology and Culture, 50(1):211–213, January 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/258572.

Whalen:2006:CIH

[Wha06] David J. Whalen. Communications: An International History of the Formative Years (review). Technology and Culture, 47 (2):463–465, April 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/200883.

Wheeler:2000:ASW

[Whe00] Everett L. Wheeler. Ancient Siege Warfare (review). Technology and Culture, 41(4):834–835, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33646.

White:2001:LDJ

[Whi01] John H. White. Long Day’s Journey: The Steamboat and Stagecoach Era in the Northern West (review). Technology and Culture, 42(3):607–608, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/33797.

White:2002:RTR

[Whi02] William J. White. Radio and Television Regulation: Broadcast Technology in the United States, 1920–1960 (review). Technology and Culture, 43(1):198–200, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33893.

Whitehouse:2004:GWH

[Whi04] David Whitehouse. Glass: A World History (review). Technology and Culture, 45(2):416–418, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55701.

White:2005:WWC

[Whi05] John H. White. War of the wires: A curious chapter in street railway history. Technology and Culture, 46(2):374–384, April 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
184350.

Williams:2000:EWE

[Wil00a] James C. Williams. Empowering the West: Electrical Politics Before FDR (review). Technology and Culture, 41(3):608– 610, July 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33596.

Williams:2000:HAP

[Wil00b] Michael R. Williams. Howard Aiken: Portrait of a Computer Pioneer, and: Makin’ Numbers: Howard Aiken and the Computer (review). Technology and Culture, 41(4):823–825, October 2000. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33647.

Williams:2000:ASM

[Wil00c] Rosalind H. Williams. “All that is solid melts into air”: Historians of technology in the information revolution. Technology and Culture, 41(4):641–668, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33648.

Williams:2001:SBN

[Wil01] James C. Williams. Strictly business: Notes on deregulating electricity. Technology and Culture, 42(3):626–630, July 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33798.

Williams:2002:LMT

[Wil02a] Rosalind H. Williams. Lewis Mumford’s Technics and Civilization. Technology and Culture, 43(1):139–149, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33894.

Williams:2002:TWW

[Wil02b] Rosalind H. Williams. A technological world we can live in. Technology and Culture, 43(1):222–226, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33895.

Williams:2003:CES

[Wil03a] James C. Williams. California Earthquakes: Science, Risk, and the Politics of Hazard Mitigation (review). Technology and Culture, 44(1):214–216, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/40189.

Williams:2003:DPF

[Wil03b] Rosalind H. Williams. Does progress have a future? Joel Moky’s Gifts of Athena. Technology and Culture, 44(2):371– 375, April 2003. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/42050.

Williams:2004:WIC

[Wil04] Michael R. Williams. Who Invented the Computer?: The Legal Battle that Changed Computing History (review). Technology and Culture, 45(2):449–450, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/55702.

Williams:2006:BD

[Wil06] Rosalind H. Williams. The Big Dig. Technology and Culture, 47(3):707–711, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/201488.

Williams:2007:OBB

[Wil07] Rosalind H. Williams. Opening the big Box. Technology and Culture, 48(1):104–116, January 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/209868.

Wills:2009:MMD

[Wil09] John Wills. The Mouse Machine: Disney and Technology (review). Technology and Culture, 50(4):952–954, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
363442.

Winpenny:2002:BQR

[Win02] Thomas R. Winpenny. The Bridge at Quebec (review). Technology and Culture, 43(1):173–175, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33896.

Winer:2005:ESF

[Win05] Daniel Winer. Eating Smoke: Fire in Urban America, 1800– 1950 (review). Technology and Culture, 46(3):672–674, July 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
186650.

Winling:2008:CSM

[Win08] LaDale Winling. From a Cause to a Style: Modernist Architecture’s Encounter with the American City (review). Technology and Culture, 49(3):798–799, July 2008. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/243105.

Wisnioski:2009:TAM

[Wis09a] Matt Wisnioski. Tools of American Mathematics Teaching, 1800–2000 (review). Technology and Culture, 50(2):500–502, April 2009. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/263370.

Wisnioski:2009:LEF

[Wis09b] Matthew H. Wisnioski. “Liberal education has failed”: Reading like an engineer in 1960s America. Technology and Culture, 50 (4):753–782, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363413.

Wittner:2007:ENC

[Wit07] David G. Wittner. Earthquake Nation: The Cultural Politics of Japanese Seismicity, 1868–1930 (review). Technology and Culture, 48(2):462–465, April 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/215419.

Wittner:2008:TGW

[Wit08] David G. Wittner. Technology Gatekeepers for War and Peace: The British Ship Revolution and Japanese Industrialization (review). Technology and Culture, 49(3):808–810, July 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
243122.

Wolters:2008:TAS

[Wol08] Timothy S. Wolters. Testing American Sea Power: U.S. Navy Strategic Exercises, 1923–1940 (review). Technology and Culture, 49(4):1097–1099, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/255542.

Woods:2005:IOI

[Woo05] Abigail Woods. Industrializing Organisms: Introducing Evolutionary History (review). Technology and Culture, 46(4): 873–875, October 2005. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/192331.

Worthington:2001:IFM

[Wor01] William E. Worthington. Invisible Fuel: Manufactured and Natural Gas in America, 1800–2000 (review). Technology and Culture, 42(1):145–146, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33700.

Wosk:2000:TOH

[Wos00] Julie Wosk. The Technology of Orgasm: Hysteria, the Vibrator, and Women’s Sexual Satisfaction (review). Technology and Culture, 41(3):600–602, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33597.

Wosk:2002:ABR

[Wos02a] Julie Wosk. The Artist and the Bridge, 1700–1920 (review). Technology and Culture, 43(2):414–416, April 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33948.

Wosk:2002:PDT

[Wos02b] Julie Wosk. Photographing desolation: Three New York Exhibits of 11 September 2001. Technology and Culture, 43(4): 771–776, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/34047.

Wosk:2004:EEM

[Wos04] Julie Wosk. Edison’s Eve: A Magical History of the Quest for Mechanical Life (review). Technology and Culture, 45(4): 897–899, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/175787.

Wosk:2006:DS

[Wos06] Julie Wosk. Designing for safety. Technology and Culture, 47(4): 791–798, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/205014.

Wright:2000:RAH

[Wri00] Helena Wright. Revisions: An Alternative History of Photography (review). Technology and Culture, 41(4):816–817, October 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33649.

Wright:2004:TSA

[Wri04] Rita P. Wright. Technology and Social Agency (review). Technology and Culture, 45(1):231–232, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51783.

Wright:2006:VRA

[Wri06] Helena Wright. Victor Regnault and the Advance of Photography: The Art of Avoiding Errors (review). Technology and Culture, 47(3):650–651, July 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/201489.

Wu:2003:RRI

[Wu03] Tim Wu. Ruling the Root: Internet Governance and the Taming of Cyberspace (review). Technology and Culture, 44(1): 191–192, January 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/40190.

Wootton:2000:EAL

[WW00] Charles W. Wootton and Carel M. Wolk. The evolution and acceptance of the loose-leaf accounting system, 1885–1935. Technology and Culture, 41(1):80–98, January 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33508.

Woods:2004:PSB

[WW04] Brian Woods and Nick Watson. In pursuit of standardization: The British Ministry of healthis model 8F wheelchair, 1948–1962. Technology and Culture, 45(3):540–568, July 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
173006.

Wyatt:2004:MHT

[Wya04] Kyle Williams Wyatt. Making Histories in Transport Museums (review). Technology and Culture, 45(1):234–235, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
51784.

Yale:2008:CEM

[Yal08] Elizabeth Yale. Communication and Empire: Media, Markets, and Globalization, 1860–1930 (review). Technology and Culture, 49(4):1071–1072, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/255514.

Yarnell:2007:GGA

[Yar07] Damon Yarnell. Georeferencing: The Geographical Associations of Information (review). Technology and Culture, 48 (3):672–673, July 2007. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/218741.

Yarnell:2008:GLW

[Yar08] Damon Yarnell. The Great Lead Water Pipe Disaster (review). Technology and Culture, 49(1):285–287, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/230940.

Yarnell:2009:CMF

[Yar09a] Damon Yarnell. Chicago Made: Factory Networks in the Industrial Metropolis (review). Technology and Culture, 50(4): 932–934, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/363430.

Yarnell:2009:PHH

[Yar09b] Damon Yarnell. Placing History: How Maps, Spatial Data, and GIS Are Changing Historical Scholarship (review). Technology and Culture, 50(1):256–258, January 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/258557.

Yates:2001:CCR

[Yat01] JoAnne Yates. Cultures of Control (review). Technology and Culture, 42(4):771–772, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33848.

Yeang:2004:SFE

[Yea04] Chen-Pang Yeang. Scientific fact or engineering specification? The U.S. Navy’s experiments on long-range wireless telegraphy circa 1910. Technology and Culture, 45(1):1–29, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/51785.

Young:2001:ARL

[You01] Terence Young. The Los Angeles River: Its Life, Death, and Possible Rebirth (review). Technology and Culture, 42(2):361– 363, April 2001. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/33747.

Youngs:2009:WTY

[You09] Larry R. Youngs. Are We There Yet? The Golden Age of American Family Vacations (review). Technology and Culture, 50(4):957–959, October 2009. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/363445.

Zachmann:2001:CST

[Zac01] Karin Zachmann. Constructing Socialism: Technology and Change in East Germany, 1945–1990 (review). Technology and Culture, 42(4):830–832, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/33849.

Zachmann:2002:SCJ

[Zac02] Karin Zachmann. A socialist consumption junction: Debating the mechanization of housework in East Germany, 1956–1957. Technology and Culture, 43(1):73–99, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33897.

Zachmann:2003:FSD

[Zac03] Karin Zachmann. Forschen fur Stalin: Deutsche Fachleute in der Sowjetischen Rustungsindustrie, 1945–1958 (review). Technology and Culture, 44(3):644–645, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46123.

Zachmann:2006:AEO

[Zac06] Karin Zachmann. The Authority of Everyday Objects: A Cultural History of West German Industrial Design (review). Technology and Culture, 47(1):234–236, January 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/196080.

Zaidi:2008:JFT

[Zai08] S. Waqar H. Zaidi. The Janus-face of techno-nationalism: Barnes Willis and the “Strength of England”. Technology and Culture, 49(1):62–88, January 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https:

//muse.jhu.edu/pub/1/article/230941.

Zeilinger:2001:WNR

[Zei01] Stefan Zeilinger. The Works National Railway Museum, York. Technology and Culture, 42(4):760–763, October 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33850.

Zeilinger:2003:FTF

[Zei03] Stefan Zeilinger. On the Fast Track: French Railway Modernization and the Origins of the TGV, 1944–1983 (review). Technology and Culture, 44(3):613–614, July 2003. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/46124.

Zeller:2001:MVA

[Zel01] Thomas Zeller. The Most Valuable Asset of the Reich: A History of the German National Railway, 1920–1932 (review). Technology and Culture, 42(1):168–169, January 2001. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/33701.

Zeller:2002:MVA

[Zel02] Thomas Zeller. The Most Valuable Asset of the Reich: A History of the German National Railway, vol. 2, 1933–1945 (review). Technology and Culture, 43(4):809–810, October 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/34048.

Zeller:2004:AOF

[Zel04] Thomas Zeller. Agrarmodernisierung und ökologische Folgen: Westfalen vom 18. bis zum 20. Jahrhundert (review). Technology and Culture, 45(1):197–199, January 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/51786.

Zeller:2006:BMK

[Zel06] Thomas Zeller. Begrenzte Mobilität: Eine Kulturgeschichte der Autobahnen in der DDR (review). Technology and Culture, 47(4):858–860, October 2006. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/205015.

Zembala:2002:MAG

[Zem02] Dennis M. Zembala. Making the Amalgamated: Gender, Ethnicity, and Class in the Baltimore Clothing Industry, 1899– 1939 (review). Technology and Culture, 43(1):177–179, January 2002. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/
33898.

Zetti:2008:PGV

[Zet08] Daniela Zetti. “Politique de la grandeur” versus “Made in Germany”: Politische Kulturgeschichte der Technik am Beispiel der PAL-SECAM-Kontroverse (review). Technology and Culture, 49(4):1085–1087, October 2008. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/255515.

Zeilinger:2000:MTG

[ZH00] Stefan Zeilinger and Michael Hascher. Museums of technology in Germany. Technology and Culture, 41(3):525–529, July 2000. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/33598.

Zimmerman:2004:PRG

[Zim04] David Zimmerman. The Precision Revolution: GPS and the Future of Aerial Warfare (review). Technology and Culture, 45(4):883–884, October 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/175788.